A Complete Bibliography of Stanford University Computer Science Technical Reports

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

17 December 2025 Version 1.07

Title word cross-reference

(h, k, l) [741]. 0 [668]. 1 [668]. [g, d] [244]. ^{TM} [1298]. AX + XB = C [728]. Ax = λBx [278, 284, 237]. C^{m} [651]. C_{1,1} [151]. C_{2} [151]. δ^{2} [734]. ℓ_{2} [327]. k [656, 752]. L(4, n) [778]. l_{1} [767]. LR [284]. LZ [374]. m [165]. m × n [332]. N [5, 165, 335, 271, 434]. n^{log n} [704]. N log N [192, 190, 200, 198]. O(n · I log^{2} I) [716]. O(nm log n) [847]. p [587, 399]. QD [9, 56]. QL [649]. QZ [750]. r [499, 963]. s [61]. V log V [206].

-Algorithm [9]. -chromatic [499]. -dimensional [61, 656]. -formula [734]. -method [56]. -Ominoes [335, 271]. -ordered [399]. -Person [5]. -shellsort [741]. -Stirling [963]. -th [752]. -tree [587]. -way [434].

0.5 [120].

1 [730, 1372]. 1.2 [37]. 1975 [582]. 1976 [582]. 1978/79 [886]. 1980/81 [886]. 1981/82 [1082]. 1984/85 [1082].

2 [878, 181]. 2-D [878]. 2-Isomorphism [723]. 21st [1580].

3 [878]. 3-D [878]. 360 [79, 202, 175, 220, 257, 256]. 3D [1734, 1725, 1669].

60 [205, 413]. 6000 [265].

= [645].

A0 [593]. Absence [883, 829]. Abstract [1117, 1531, 993, 1329, 1200, 1013, 206, 622, 900]. Abstraction [1383, 1586, 1635, 1530, 1546]. Abstraction-based [1635]. abstractions [1313]. Academic [1575]. Access [1134, 161, 1026, 1157, 1632, 225, 1329, 1612, 215, 215]. accessing [103]. Accommodation [182]. accumulation [163]. Accuracy [1510, 1701, 670]. Accurate [41, 385, 1473, 121]. Achievements [1580]. Achieving [1510]. ACM [1452]. Acquiring [1087]. Acquisition [1096, 1725, 1214, 1115]. Across [1600, 1675]. Action [1515, 1606, 1162, 1373, 1368, 1255, 1184, 1390, 1519, 1724]. Action-Based [1515]. Actions [357]. Active [1512, 434]. Actively [1549]. Activity [975, 1529]. Acyclic [994, 1247, 917, 870, 400, 865]. ADA [805, 884]. ADAM [884]. Adapt [1720]. Adaptation [635, 396]. Adaptive [1622, 943, 920, 1484, 1245, 1513, 1538, 1685, 1705, 1423, 541]. Added [1599]. Addenda [194, 403]. Addition [1219, 1292, 1293, 553]. Additive [876, 879, 616, 815]. Addressing [1355, 641]. Adel’son [1339]. Adel’son-Vel’ski [1339]. Adequate [1022]. Adjectives [1735]. adjoint [1265]. adjustment [788]. Admissibility [295]. Admission [1565, 1609, 1592]. advantages [1013]. Adverbs [298]. Advice [1214]. Advising [475]. Aesthetics [350]. affects [481]. After [422]. Again [864, 257]. Agent [1439, 1360, 1504]. Agent-Oriented [1360]. AGENT0 [1415]. Agents [1606, 1063, 1100]. Aggregation [529]. aggregations [1407]. Aggressive [1456, 1594]. agreement [1470]. AI [771, 764, 772, 1127, 954, 447]. Aid [1046]. Aided [83, 680]. Aids [1731, 836, 738]. Aitken [734]. AL [467, 733, 906]. Alarms [1636]. Algebra [46, 133, 310, 860]. Algebraic [1522, 24, 311, 879, 2, 628, 828, 210, 792, 483]. Algebras [1335]. ALGOL [12, 19, 99, 111, 102, 118, 316, 73, 205, 167, 385, 235, 258, 20, 35, 413]. ALGOL-based [316]. ALGOL-Like [167]. Algorithm [9, 790, 720, 679, 344, 1060, 456, 493, 550, 765, 972, 1609, 1247, 1273, 1327, 1463, 1453, 1498, 1617, 278, 310, 1490, 192, 190, 71, 613, 366, 374, 548, 614, 860, 306, 1719, 917, 665, 872, 1183, 1422, 841, 847, 462, 259, 249, 27, 521, 1692, 373, 640, 787, 336, 54, 649, 200, 198, 163, 237, 704, 653, 716, 750]. Algorithmic [1141, 542, 1235, 1635, 952]. Algorithms [1020, 1110, 1220, 838, 655, 540, 157, 203, 1585, 1378, 1628, 1497, 1392, 1195, 759, 1228, 1230, 1277, 1339, 1398, 1444, 1455, 569, 671, 1041, 1189, 1044, 1093, 1563, 211, 802, 323, 322, 610, 216, 30, 928, 1244, 1459, 561, 1639, 1400, 767, 876, 899, 1501, 1445, 1475, 756, 426, 445, 1424, 1382, 1144, 1188, 527, 1323, 1474, 60, 1304, 1681, 266, 1071, 212, 43, 478, 1401, 846, 1009, 694, 662, 743, 1570, 560, 624, 749, 810, 677, 763]. Algoritmy [1339]. aliasing [755]. Aligning [1657]. Allocating [1636]. allocation [725]. Almost [366, 59, 1740]. Almost-Optimal [366]. Alpha [452]. alphabet [777]. alter [1077]. Altering [1545]. Alternating [291]. Always [1237, 1288, 571, 572]. AM [584]. ambulatory [947]. Among [878, 1636, 1645, 1063, 577, 1100]. Analogical [1163]. Analogies [1091, 1533]. Analogy [221]. Analysis [1500, 1522, 679, 1359, 172, 170, 615, 717, 398, 1226, 361, 1221, 89, 569, 1532, 408, 613, 113, 1625, 211, 452, 548, 1011, 1473, 1374, 1375, 305, 1324, 756, 49, 616, 279, 448, 426, 250, 940, 841, 673, 460, 420, 58, 1239, 817, 27, 1448, 391, 521, 714, 360, 789, 21, 325, 1677, 6, 32, 122, 390, 389, 11, 368, 146, 307, 315, 560, 741, 725]. analyst [62]. Analytic [356, 36, 137]. Analytical [1374]. Analyzing [1108]. {And/OR} [1487]. angles [230]. Animation [1655]. annotated [254]. annotation [646]. Annotations [1599]. Annual [1177, 1441]. Anomalies [318]. Answering [138]. Antimicrobial [475]. APP [899]. Applicability [1067]. Application [929, 138, 1420, 1572, 1116, 1732, 827, 1016, 1739, 128, 155]. Application-controlled [1420]. Applications [1274, 771, 1074, 349, 764, 772, 578, 1387, 1425, 1698, 1138, 1190, 221, 1691, 643, 400, 1097, 485, 1581, 441, 1501, 523, 1621, 1704, 664, 1585, 1263, 1266, 121, 375, 183, 531]. Applications-Oriented [771, 764, 772]. Applied [830]. Applying [1165]. appraisal [275]. Apprentice [1194]. Apprenticeship [1261]. Approach [1522, 1222, 877, 1298, 1004, 1056, 1052, 545, 465, 770, 354, 1089, 1320, 1216, 1427, 584, 1114, 48, 718, 311, 1202, 49, 1462, 682, 1300, 423, 1390, 1281, 983, 984, 405, 280, 393, 549, 705, 797, 941, 744, 842, 1120, 272]. Approaches [1096, 1611, 1403, 1356]. Approximate [824, 14, 875, 29]. Approximating [1395]. Approximation [1220, 1378, 1628, 1617, 851, 1563, 1244, 1375, 1459, 1445, 689, 490, 815, 72, 1401, 387, 31, 23]. Approximations [1030, 73, 1552, 76]. AppSwitch [1739]. Architectural [1176, 1207]. Architecture [1513, 1545, 1160, 441, 1737, 1599, 1329, 1369, 730, 1049, 1055, 792]. Architecture-Altering [1545]. Architectures [1156, 1378, 1232, 891]. Area [893, 1203]. Areas [449]. Argument [1268]. Arithmetic [1696, 277, 24, 38, 876, 875, 216, 109, 62]. arithmetics [1009]. Arm [1547, 1526, 319, 585]. Arms [1548]. Array [438, 1473]. Arrays [1663, 1664, 562]. Arrivals [1592, 341]. Art [194, 636, 564]. Articulated [940]. Artificial [969, 333, 1004, 830, 757, 1052, 1037, 584, 400, 710, 826, 1315, 833, 1111, 470, 422, 415, 476, 636, 316, 247, 272]. Ascribing [740]. asking [1201]. Aspects [1394, 1080, 127, 516, 555, 592, 542]. Assemblies [827]. Assembling [1413]. Assembly [606, 1607, 366, 1482, 1442, 1466, 1495, 582, 407, 547, 568]. assertions [571, 572]. assessment [1728]. Assignment [1498, 1508, 1476]. assignments [557]. Assistance [1357]. Associated [803, 497, 504, 1742]. Associative [1623, 426, 792]. Assumption [971]. Asymptotic [335, 434, 734, 61]. Asynchronous [1151, 1042]. Attachment [1410]. Attention [1702, 1636]. Augmented [1684, 1686, 1706]. augmenting [1728]. Autistic [359]. Automata [1350, 1249, 504]. Automated [979, 947, 1072, 1467, 1626]. Automatic [1404, 956, 655, 1164, 469, 1710, 1383, 773, 263, 17, 686, 919, 545, 376, 1595, 71, 463, 533, 562, 1158, 30, 887, 1307, 483, 565, 1586, 1682, 484, 600, 233, 174, 633, 482, 946]. Automatically [1676, 1721, 509]. Automating [816, 444]. Automation [827, 467]. Automaton [190, 198]. automorphisms [281]. Autonomous [1606, 1610, 1743]. Average [1253, 223, 434, 752]. average-case [752]. Avoidance [831]. avoiding [148]. Aware [1344]. Axiomatic [1048, 393, 405]. Axiomatization [894, 1275]. Axioms [338, 343].

B [583]. B-splines [583]. Back [898]. Backchaining [1216]. Backtrack [1282, 453]. Backward [1243]. Backward-Chaining [1243]. BAIL [534]. Bairstow [54]. Balanced [267, 446, 523]. Balancing [1532, 1241, 1306]. band [638]. Bandwidth [601]. Banks [1616]. Banyan [1374, 1375]. BAOBAB [683]. Barehands [1646]. Base [816, 472, 1200, 391, 487, 478, 64]. Based [1515, 949, 945, 1613, 1429, 1714, 1648, 1524, 783, 910, 929, 359, 1046, 530, 253, 785, 1428, 370, 1553, 981, 1393, 1509, 770, 1611, 1719, 1514, 1454, 1427, 356, 1499, 1558, 1214, 1206, 627, 1054, 1033, 1669, 1624, 1133, 962, 993, 1357, 839, 475, 1616, 1523, 711, 1261, 60, 1678, 836, 714, 683, 275, 1156, 604, 316, 1115, 1024, 1338, 971, 1694, 884, 662, 1391, 1088, 608, 1546, 738, 1570, 611, 941, 525, 1635]. Bases [1166, 578, 377, 506, 1397, 1201, 754]. Basic [826, 266, 65]. Basis [351, 376, 1268, 611, 882]. Bayesian [1427, 1443]. BB1 [1049, 1055, 1057]. BDDs [1542]. Be [1187]. beginner [1003]. behaved [1335]. Behavior [1620, 719, 567, 593, 398, 1049, 1055]. Behaviors [1667]. Behaviour [1487]. Belief [545, 1544, 1038, 1256, 1417, 1443, 82, 1519, 298]. beliefs [1608]. Benchmarks [1726]. Benefits [1385]. Best [493, 550, 261, 327, 387, 752]. Beta [1149, 1250, 452]. better [511, 571, 572]. Between [439, 527, 230, 1211, 340, 780, 1275, 210, 1255]. Beyond [1661, 1716]. Bi [136]. Bi-Directional [136]. Bibliography [980, 1118, 966, 543, 1525, 1209, 1364, 932, 204, 270, 321, 406, 695, 254, 314]. Bidiagonalization [303]. Biharmonic [850, 630]. Bilinear [1605, 876, 879]. Bin [1220, 677]. Binary [267, 1578, 502, 149, 1065, 1066]. Binary-Search [1065, 1066]. Binding [868, 518]. Biologists [1665]. Biology [1287]. Bipartite [1573]. Birthday [1195]. Bit [1695, 875]. Bit-Vectors [1695]. Bits [1305]. Blackboard [1143, 1319, 1049, 1055]. Blazenet} [1203]. Blendshape [1655]. Block [1161, 1, 1462, 507, 650]. Block-Oriented [1462]. Blocking [1247, 1641]. Blum [363]. Board [1103, 1153]. Bone [1688]. Book [147, 1339]. Boolean [1562, 1600, 1328, 570]. Boor [583]. Bose [177]. Bottleneck [1355]. Bottlenecks [1468]. Bound [1215, 271, 243, 660, 748]. Boundaries [285]. Boundary [4, 635, 824, 14, 907, 579, 151, 497, 735, 541, 492, 592, 18]. boundary-value [492]. Bounded [228, 1373]. Bounds [364, 363, 1614, 141, 365, 1387, 143, 335, 760, 1372, 879, 1353, 828, 803, 66, 490, 685, 121, 558, 1742, 480]. Box [1280, 1661]. Boyer [1318]. Brain [1021]. Branching [1503]. Breaking [844]. breeding [1340]. Bricolage [1719]. Bridging [1582]. Broadcast [1215, 939]. Brook [1660]. Browsers [1704]. Bruijn [436]. Brush [1090]. Buffer [1671]. bug [1711]. Building [857, 909, 786, 1424]. Bulletin [1103, 1153]. Burden [1690]. Burnett [372]. Bursts [154]. Byzantine [864].

C [978]. Cache [1389, 1325, 1420]. Caches [1128, 1419]. Caching [1623, 1195, 922, 1589, 1480, 1290]. Calculating [203, 8, 583]. Calculation [51, 81, 205, 413, 583, 319, 36, 67]. Calculations [14, 993, 124]. Calculus [288, 115, 166, 1019, 91, 797]. Calgen [91]. Calibration [1664, 1682]. Call [1721]. Calls [691]. Camera [1649, 1668, 1663]. Cameras [1650]. Can [334, 1187]. Canary [1737]. Cancellations [1353]. Canonical [489, 1741, 750]. Can’t [333]. CAOS [1145]. capabilities [608]. Capture [1649, 1656, 1067, 1025]. Car [1370]. CAREL [1146]. Carlo [568, 1627]. Cartesian [1060]. cascaded [1733]. Case [1428, 221, 1272, 176, 636, 752]. Case-Based [1428]. Cases [1670]. Caster [1165]. Catalog [885, 416, 215]. categories [307]. Causal [921, 1046, 796, 942]. Causality [1487]. Causally [1502]. CDC [265]. CDE [1721]. Center [538]. central [368]. Centric [1713, 1624]. Century [1580]. Certain [838, 367, 277, 65]. Cervical [1603]. Chain [1375, 892, 778]. Chaining [1243]. Chains [1387, 553]. Challenges [833]. Change [1405]. channel [339]. character [840]. Characteristic [3]. characterization [327]. Characters [935, 866]. Chebyshev [67, 1288, 72, 1296, 109]. Checker [1010, 443, 1023]. Checking [477, 1553, 1568, 1307, 431]. Chemical [400, 398, 402]. Chess [116]. Chi [690]. Children [84, 83, 359, 380]. Chinese [935, 992, 840, 866]. CHISEL [978]. Choice [463, 1058, 64]. Cholesky [1403, 1462, 1483, 1432, 260, 309]. choosing [637]. Christoffel [726]. Chromatic [281, 590, 499]. Chu [1539]. CIL [135]. Cinder [1717]. circle [652]. Circuit [1002, 1439, 1334, 1337, 1303]. Circuits [1229, 1318, 814, 988, 1574, 946]. Circulation [1228]. Circumscription [1035, 804, 1097]. ciruits [431]. Class [304, 306, 689, 331, 694, 653, 31, 708]. Classes [628, 398, 342]. Classification [1053, 1436, 1034, 1086, 498, 1604, 1448, 1418, 396]. classnotes [1127]. Classroom [1706]. Clausal [1019]. Cleanly [429, 428]. Clear [1161]. Client [1623]. Clinical [816, 921, 1345, 1427, 1214, 1142, 833, 1582, 1626, 1546, 738, 942]. clinical-practice [1626]. Clinical-Trial [1214]. Clocked [1583]. Close [59, 1740]. closed [342]. Closest [10]. Closure [1352, 519, 662]. Closures [219]. Cloud [1729, 1736, 1738, 1737]. clusters [952]. Co [1529]. Co-Learning [1529]. Coalesced [817]. Code [1712, 1720, 1670, 509, 265]. Codes [154, 791]. Coding [463]. Coffman [372]. COGENT [37]. Cognition [969, 955, 1169]. Collaboration [1678]. Collaborative [1708].

Collected [1629, 1642]. Collection [1528, 294]. Collections [701]. Collectively [76]. Collectors [1195]. Collision [1549, 1556]. Collision-Free [1549, 1556]. Color [1637, 1664]. Coloring [924, 846]. Com [995]. Com-put-er [995]. comb [389]. Combination [1693, 1428]. Combinations [282]. Combinatorial [1392, 1141, 1228, 1384, 1471, 1617, 301, 928, 1475, 1434, 372, 292, 300, 522, 299, 478, 624, 810]. combinatorics [516, 747]. Combining [1257, 1643, 1543, 946]. comes [77]. Command [758, 1643]. Commentaries [1744]. comments [514]. Committee [187]. Common [883, 1563, 1097, 521, 488, 829]. Communicating [1601, 889, 577, 207, 944]. Communication [1271, 1027, 1232, 1502, 1524, 1716, 1477, 1001, 1430, 842]. Communication-Time [1001]. Communications [1197, 1175, 435, 436]. compact [76]. Comparative [1667, 1331]. Comparing [1358, 409]. Comparison [687, 224, 262, 175]. Comparisons [1124, 447]. competence [1087]. Competitive [1609]. Compilation [814, 290]. Compiler [1446, 990, 38, 912, 820, 135]. Compilers [1492, 245, 254]. Compiling [1079, 57]. Compleat [1099]. complementarity [450]. Complementary [126]. Complements [219]. Complete [1030, 1251, 1620, 723, 499, 1724, 1275]. completeness [941]. Completing [1321]. Completion [1669]. Complex [1710, 1707, 408, 1508, 719, 474, 381, 565, 1653, 1652]. Complexities [876]. Complexity [1333, 1110, 601, 1607, 1031, 1006, 875, 879, 570, 1301, 1025, 1337, 566, 624, 1107, 1278, 1482, 125, 261, 780, 1349, 1303, 810, 644, 752]. Compliant [1394]. Components [460]. Composing [1631]. composition [983]. Compositional [1517]. Comprehensive [692, 1082, 886]. Compress [935]. compressible [70]. compression [523]. Computable [296]. Computation [969, 1659, 274, 1744, 352, 147, 1283, 348, 489, 1741, 152, 1373, 892, 1451, 13, 1308, 1080, 66, 1447, 1231, 70, 796, 103, 280, 726, 843]. Computational [94, 67, 46, 1394, 835, 1021, 408, 465, 30, 719, 1142, 448, 1356, 1234, 1337, 674]. Computations [246, 1296, 753, 700, 1285, 196, 1424, 1278, 266, 241]. Compute [1068, 650]. Computed [153]. Computer [832, 329, 461, 473, 980, 1062, 84, 83, 359, 1046, 417, 551, 1153, 180, 648, 26, 106, 117, 98, 88, 112, 213, 883, 320, 472, 15, 1661, 377, 449, 1421, 1620, 1393, 161, 1082, 168, 1210, 709, 194, 554, 564, 438, 1427, 191, 356, 1198, 703, 1525, 48, 680, 1177, 1441, 1209, 1364, 441, 319, 93, 1673, 224, 289, 49, 430, 932, 302, 130, 193, 371, 236, 675, 475, 673, 429, 460, 891, 392, 204, 321, 406, 695, 182, 82, 127, 1239, 1390, 567, 1703, 582, 160, 1369, 251, 334, 21, 39, 77, 92, 727, 802, 1340, 103, 199, 745]. computer [626, 639, 867, 314, 368, 446, 201, 738, 852, 532, 265, 1118, 178, 543, 798, 1571, 110, 129, 886, 591]. Computer-Aided [83, 680]. Computer-Based [1046, 1393, 1427, 356, 738]. Computer-Oriented [130]. Computer-Support [1390]. Computers [333, 1169, 1041, 440, 432, 1380, 1460, 33, 53]. Computing [877, 344, 73, 1722, 1547, 1738, 1737, 833, 1186, 230, 171, 1585, 603, 787, 156, 519, 178, 175, 566, 538]. Concave [5]. Concept [1299, 903, 718, 412].

Concepts [1534, 739, 273, 65]. Conceptual [472, 545, 430, 130, 307, 380, 1085]. Conceptually [370]. Concerning [596]. Concrete [1254]. Concurrency [1539, 1011, 1135, 1367, 900, 1069]. Concurrent [1381, 854, 1636, 794, 861, 1022, 1249, 1319, 170, 855, 936, 985, 622, 513]. Condition [1512]. Conditional [164]. Conditioned [489, 1741]. Conditions [1690, 907, 7, 497]. Cone [85]. Configurable [1359]. Configuration [1451, 1507, 1537, 1554]. Configuration-Space [1451]. congruence [662]. Congruential [818]. conjecture [308, 184]. Conjugate [544, 561, 599, 653, 552]. conjunctions [566]. Connected [192, 200]. Connection [1437]. Connectionist [1581]. Connections [870]. Consecutive [282]. Consensus [1305]. Considerations [1173, 770, 1571, 711, 67]. Considering [1649]. Consistency [1307, 1325, 941]. constant [791]. Constrained [671, 580]. Constraint [761, 1171, 1045, 1553, 712, 704]. Constraint- [761]. Constraints [1342, 1257, 1208, 459, 142, 781, 800, 1681]. Constructing [540, 581, 1064, 1108, 836, 656]. Construction [655, 706, 758, 578, 1531, 919, 1326, 1393, 30, 1603, 373, 444, 401, 254]. constructions [1069]. Constructive [326]. consultant [714]. Consultation [530, 833, 836, 1014, 608]. Consultations [832, 942]. Contact [1211]. Containing [536]. Content [1685, 430]. Contested [632]. Context [228, 448, 1416, 1678]. Contexts [1425]. Contingency [1494]. Contingency-Tolerant [1494]. Continuations [1452]. Continuous [1541, 72, 389, 396]. contouring [409]. Contract [682]. Contribution [35]. Control [1224, 619, 1494, 1117, 90, 1565, 1609, 1057, 161, 1592, 1317, 191, 317, 1032, 93, 585, 7, 236, 1496, 675, 574, 762, 1200, 1438, 1013, 731, 1049, 1055, 233, 199, 1313, 525]. Controllability [1342, 877, 1370]. Controllability/Observability [877]. Controlled [1613, 1128, 320, 319, 675, 1420]. Controlling [1083, 1130]. Convergence [1387, 527, 651, 649, 629]. Convex [15, 904, 335, 7, 23, 144, 101, 748]. Cook [967]. cooperating [667]. Cooperation [1100]. Cooperative [1673, 1674, 1184, 1390]. Coordinate [540]. Coordination [320, 1474]. COPILOT [423]. Coping [554]. corner [397]. Correct [240, 794]. Correcting [1050]. Correction [1723, 18]. corrections [29]. Correctness [822, 245, 189, 38, 311, 485, 509, 197, 393, 571, 572]. corresponding [650]. cosets [326]. Cost [1246, 559, 1444, 1498, 1493, 1617, 432, 1683, 553]. Counterexample [308]. Counterfactuals [1051]. Coupled [1148]. Coupon [1195]. Course [1421, 52]. Coverage [1710]. Covering [1445, 582]. coverings [837]. Coverity [1711]. Crash [1670]. Create [1721]. creation [1626]. Creative [510]. CREW [1259]. Crisis [1739, 1731]. criterion [55, 867]. Critic [908]. Critical [1456]. criticisms [468]. Critique [333, 1728]. cross [637]. cross-validation [637]. Cryptographic [1469, 1489]. crystallographic [604]. CS [1571]. CS-TR [1571]. CS139 [100]. CS229b [1127]. CS363 [1384]. CS363/OR349 [1384]. Cues [1656]. current [538]. Curve [1681]. Curved [313, 1113, 474]. Curves [1500, 234]. customers [342]. Cut [1197]. Cut-Through [1197]. CW92 [1452]. CWEB [1361]. Cybernetics [833]. Cycle [1353]. Cycles [1238]. Cyclic [400, 154]. cyclically [1265].

D [878]. d.mix [1698]. d.tools [1672]. Daily [975]. Data [94, 1332, 937, 816, 717, 706, 1659, 1045, 819, 180, 759, 1643, 411, 1089, 1640, 701, 463, 1125, 440, 131, 225, 565, 1414, 1329, 1681, 950, 208, 1266, 604, 1612, 103, 693, 622, 900, 754, 1432, 1546, 617, 531]. data-accessing [103]. Data-Fitting [131]. Database [1522, 994, 921, 1179, 1312, 1484, 759, 1405, 1557, 1572, 971, 1089, 1461, 1316, 993, 1638, 839, 1580, 1518, 1480, 869, 911, 986, 813, 1407, 961, 792, 737, 842]. Databases [1512, 1336, 1623, 901, 1429, 1061, 1371, 895, 1408, 1426, 1125, 974, 1412, 962, 1423, 1382, 1582, 1017, 806, 1172, 1418, 1470, 1120]. dataflow [964]. Datagram [1431, 1078]. Datalog [1152]. Deadline [1476, 1508]. Deals [1063]. Death [1676]. debugger [534]. Debugging [581, 1468, 1307, 505, 1199, 1699]. December [582]. Decentralized [1122, 1198]. Decidable [222]. Deciding [1521]. Deciphering [818]. Decision [1696, 1531, 498, 1695, 1084, 1578, 1345, 1577, 1320, 738, 828, 662, 667, 1626, 951]. Decision-Theoretic [1320]. Decisions [881]. Declarative [337, 1685, 238]. Decoded [154]. Decomposing [1631]. Decomposition [877, 73, 603, 1263, 788, 782, 778]. Decompositions [994, 133, 1264, 124]. Dedekind [491]. Deduction [1076, 956, 1038, 923, 1072, 633]. Deductions [1243]. Deductive [1633, 1056, 1028, 872, 1433, 1426, 1125, 1461, 1316, 1412, 1382, 1635, 705]. Deductive-Algorithmic [1635]. Deductive-Tableau [1028]. Deep [472]. Defeasible [1256]. deferred [29]. Definable [1236]. Defined [304, 277, 693]. Defining [1733]. definite [144, 45]. Definition [607, 337, 238, 20, 829, 210]. Definitions [274, 959, 511, 629, 973]. Deformable [1682]. degeneracy [573]. degree [16]. Degrees [276, 961]. Delaunay [1326]. Deletion [502]. Deletions [598]. Delight [1596]. Delta [1512]. DENDRAL [208, 664, 176]. denotational [517]. Dense [1718, 1663, 1438, 515, 1432]. Density [851, 375]. Department [110, 129, 417, 551, 178, 543, 1525, 1209, 1364]. Dependence [506]. Dependencies [933]. Dependency [997, 130, 307]. Depth [949, 1590, 532, 486, 466]. depth-first [466]. Derivation [834]. Derivatives [203, 151]. Deriving [1519, 1137]. Described [1016]. Describing [504]. Description [313, 296, 107, 984, 930, 175, 973]. Descriptions [786, 474]. Design [559, 1359, 717, 1002, 1226, 901, 1446, 1715, 1567, 819, 1056, 1208, 240, 140, 785, 968, 1641, 1405, 1421, 1672, 1689, 1295, 1687, 1719, 1684, 1686, 1705, 1706, 1158, 48, 1131, 974, 1016, 1319, 1674, 1624, 989, 1357, 675, 711, 1358, 1708, 911, 986, 1184, 1185, 1390, 1699, 1013, 983, 984, 1677, 414, 777, 1728, 681, 840, 964, 842, 680, 1380]. designed [103]. Designing [1702, 1575, 1094, 1699, 611]. Designs [1703]. destructively [1077]. detailed [1331]. Detecting [1457, 1150]. Detection [1062, 432, 1692]. determinant [109]. Determination [1332, 658, 25]. Determining [691, 822, 597, 590]. deterministic [796]. developer [1711]. Developing [1699]. Development [1713, 1613, 83, 1673, 35, 576, 380]. developments [316, 810]. Devices [1276, 1717, 504]. Diagnosis [1515, 1558, 1165, 1570]. Diagnostic [1046, 1087, 1014]. Diagonal [41, 1, 433, 59, 1740, 42, 331]. Diagrams [722, 1326, 1578, 1631]. Dialog [960]. Dialogue [783, 442, 487]. Diameter [1562]. Diary [1690]. Dictionary [1735]. Difference [920, 22, 362, 926, 340, 541, 28, 492, 18, 31]. Differences [151, 3]. Different [1358, 342]. Differential [1542, 943, 544, 14, 907, 925, 362, 291, 721, 1194, 490, 599, 670, 28]. differentiation [269]. Difficulties [359]. diffusion [1653, 1652]. digest [340, 339, 368]. Digital [1500, 1723, 988, 1598, 1699, 21]. Digitized [1090, 389]. Dilemma [1047]. Dimension [920, 1487, 664]. Dimensional [679, 1018, 361, 180, 1537, 61, 630, 1603, 656]. dimensions [294]. Dinic [1339]. dinner [845]. Direct [195, 191, 49, 286, 159, 6, 199, 837]. Directed [612, 80, 82, 865, 542, 539]. Directional [136]. Directions [1745, 869, 61]. Directory [1104]. Disambiguation [545]. Discipline [1033]. disciplines [340, 341]. Discontinuities [1590]. Discourse [625]. Discover [381]. Discovery [921, 584]. Discrete [195, 585, 1438, 1541, 159, 568]. discussion [547]. discussions [422]. disjoint [743, 466, 618]. disk [341]. Diskless [991, 1026]. Display [1647, 1646]. Displaying [1729]. Displays [1648, 1645]. dissection [788, 659]. Dissemination [1478]. dissipativity [619]. Distance [508, 1681, 685, 1614]. Distances [528]. Distant [1533]. distinguish [334]. Distributed [1613, 901, 1709, 991, 1154, 1174, 1212, 1146, 1103, 1153, 759, 1389, 1508, 1190, 977, 1116, 1198, 1736, 1140, 1101, 1523, 1591, 682, 715, 1148, 1480, 1458, 1040, 1281, 813, 1325, 1476]. Distributing [1248, 1262, 1243]. Distribution [123, 348, 1469, 974, 1414]. Distribution-Specific [1469]. distributions [262, 1576]. Disturbed [83]. divide [242, 243]. divide-sort-merge [242, 243]. Divisors [521]. DNA [1448]. Do [333, 40, 77]. Doctrine [924]. Document [1236, 1458]. Documentation [998, 1361]. Domain [686, 793, 1543, 836]. Domain-Independent [836]. Domains [1312, 873, 121]. Dominators [665, 466]. Double [1246, 326]. Down [1067, 1025]. downdating [1266]. Drawing [915]. Drawings [1113]. Drazin [751]. Dreams [645]. Dreyfus [333]. Driven [1429, 1655, 1045, 236, 486]. drum [325, 340]. drums [339]. DTEP [1183]. Dual [1400]. Duplex [1730]. duplicate [776]. During [1224, 1457]. Dynamic [937, 877, 857, 1197, 317, 1548, 1182, 1335, 1566, 225, 1414, 1464, 1731, 1608, 518]. Dynamics [1224, 592].

early [576]. Earth [1614]. easier [897]. Easy [1068]. Edge [949, 1062, 466, 328]. Edge-disjoint [466]. Edges [1150]. edition [340, 339, 368]. Editor [1236, 1132, 425]. Editors [914]. Edmonds [292, 264, 336]. Education [764, 106, 1684]. Educational [1579, 39]. Effect [1715]. Effective [1680, 1591]. Effectiveness [1492]. Effects [691]. Efficacy [1715]. Efficiency [1510, 1273, 770, 793, 1371, 453]. Efficient [838, 290, 1149, 456, 1327, 1463, 1498, 1189, 1093, 212, 1578, 306, 322, 1640, 1328, 440, 1308, 1462, 1144, 259, 249, 1304, 286, 336, 1325]. Efficiently [1257, 852]. Egyptian [1276]. Eigenfunctions [14]. eigenproblem [44]. Eigenproblems [497, 507]. Eigensystems [489, 1741]. Eigenvalue [699, 374, 610, 803, 638, 239, 237, 1742]. Eigenvalues [14, 41, 22, 59, 1740, 490, 43, 121, 63]. Eigenvector [803, 1742]. Eigenvectors [34]. EKL [1023]. Elastic [11]. Elastic-plastic [11]. Electron [851]. Electronic [1687]. Element [213, 689]. Elemental [113]. Elementary [526, 150]. elements [42]. Elimination [1589, 849, 1382, 542, 537]. Elliptic [544, 386, 291, 689, 286, 599, 1576]. elucidation [479]. Embedded [1569, 1351]. Embedding [454, 893]. Emotion [1658]. Empirical [1712, 123, 186, 1116, 713, 208]. Emulation [191, 317, 199]. EMYCIN [902]. Enabling [1644]. Encryption [818]. End [116]. Energy [1718, 1717]. Engineer [1201]. Engineering [830, 975, 738, 636]. English [19]. Enhanced [225, 1700]. Enhancements [1577]. Enough [147]. Ensembles [1604]. Ensuring [1224]. Enterprise [1722]. Entropy [713]. Entry [1643, 1694]. Environment [1713, 1002, 823, 968, 1162, 1548, 719, 1140, 510, 715, 964]. Environments [1151, 1714, 1707, 1493, 1725, 1016, 1591, 1464]. Epistemic [1347]. Epistemology [910]. Epsilon [690, 115, 1424]. Epsilon-Calculus [115]. Equal [437, 494]. equality [255]. Equation [9, 850, 195, 74, 259, 433, 27, 151, 67, 128, 155, 40, 159, 630, 6, 29]. Equations [943, 1605, 544, 552, 14, 905, 907, 801, 143, 24, 306, 925, 303, 305, 410, 2, 362, 628, 291, 721, 619, 689, 286, 599, 141, 670, 145, 28, 492, 31, 763, 164]. Equidistributing [781]. Equilibrium [5, 1576]. Equivalence [113, 614, 959, 1275]. Equivalents [19]. EREW [1259]. Errata [194, 885, 1254, 403, 727]. Error [490, 3, 164, 153, 143, 305, 154, 803, 432, 141, 156, 163, 387, 146, 1742]. Errors [883, 1242, 151, 829]. Essays [881]. Essence [1080]. Essential [812]. Estimate [3]. Estimates [713, 387, 164]. Estimating [453]. Estimation [465, 962, 375, 1448, 1414]. Estimator [352, 209]. Ethics [1460]. Euclidean [175]. Euclidian [344]. EULER [20]. Evaluating [361, 1684, 1194, 1087, 360]. Evaluation [1512, 526, 1457, 1217, 1497, 1509, 1625, 875, 914, 1403, 1106, 1331, 792, 109]. even [128, 155]. Event [1438, 1644]. events [512]. eventualities [936]. Evidence [123, 1616]. Evidence-Based [1616]. evidential [1039]. Evolution [1529, 600]. Evolving [1238, 1545]. Examination [1315]. Examinations [692, 1082, 745, 886]. Example [1713, 1719, 1370, 1705, 1669]. Example-Based [1719, 1669]. Example-Centric [1713]. Examples [1720, 169, 1599, 973]. exceptions [1407, 805]. Exchange [439, 1306, 67, 1071]. Excitation [1224]. Exclusion [1332, 173]. EXE [1676]. Execution [1720, 1670, 1676, 1738, 1610, 1743, 1192, 1148, 1289, 209, 1105]. Exercise [1396]. Exhaustive [1334, 400]. exhibiting [487]. Existence [5]. expansions [734]. Expected [365, 550, 614, 340, 363]. Experience [1217, 1575, 990]. Experiences [1572]. Experiential [1543]. Experiment [1156, 785, 1647, 611]. Experimental [1332, 1497, 232, 120]. Experiments [1164, 117, 1206, 867, 275, 789, 501]. Expert [1522, 856, 966, 945, 1095, 1112, 910, 929, 1109, 951, 950, 1013, 334, 1115, 941, 942]. Expertise [783]. explain [1049, 1055]. explaining [208]. Explanation [910, 1391, 1133, 608, 1443]. Explanation-Based [1133]. explanations [1014]. Explanatory [942]. Exploiting [1232, 1208, 1196, 1297, 1483]. Exploration [1714, 931]. Exploratory [582, 1002, 1538]. Exploring [823, 1687, 793, 1440]. exponent [64]. Expression [990, 1182, 834]. Expressions [814, 1362, 38, 58, 1733, 442, 1077, 946]. Expressive [1251]. Expressiveness [1333, 1058]. Extended [1600, 1233, 1196, 829, 206]. Extension [827, 1078]. Extensional [1080]. Extensionality [1479]. External [701, 1420]. Extract [1344, 1594]. Extrapolating [50]. Extrapolation [734]. Extrema [203]. Eye [320, 1683]. EyeExposé [1691]. Eyes [1691].

F [1671]. F-Buffer [1671]. Facial [1655]. Facilities [1148, 1105]. Facility [1122]. Factoring [841, 1009]. Factorization [548, 1403, 1462, 1483, 128, 155, 1311, 1331, 1343, 1432, 260, 309]. Factorizations [459, 330]. Factors [1060]. FAIL [419]. Family [798]. Fast [540, 526, 1041, 1071, 1554, 1401, 665, 1738, 451, 662, 1202, 899, 1445, 1311, 749, 763, 640, 286, 6, 331, 486, 1451, 146]. Father [1073]. Fault [1389, 1447, 1325, 736]. Fault-Tolerance [1389]. Fault-Tolerant [1447, 1325, 736]. Faults [1050]. Faulty [1453]. feature [486]. feature-driven [486]. Features [346, 797]. Feedback [1258, 320, 407, 789, 397]. FETE [209]. Fidelity [1726]. Field [1310, 1668, 1450, 1505, 1665, 79]. Fields [1707, 1576]. File [1196, 1297, 1389, 1026, 1157, 609, 1325]. Files [1196, 1126]. Filling [1653, 1652]. filter [390]. Filtering [1511]. Final [415, 88, 826]. find [1071]. finder [397]. Finding [1246, 9, 203, 1637, 1060, 456, 493, 550, 1247, 904, 665, 520, 563, 55, 12, 748, 763]. Fine [1496]. Finite [920, 689, 3, 213, 190, 351, 22, 338, 362, 926, 933, 151, 198, 541, 343]. Finiteness [554]. Firewall [1575]. Firmware [1202]. First [732, 1238, 295, 739, 1003, 443, 477, 619, 415, 630, 466]. First-Order [295, 443]. Fitting [131, 1266]. Five [964, 470, 846]. five-coloring [846]. Fixed [1695, 137, 1479, 295, 1347, 1278, 524]. Fixed-Point [295, 1347]. Fixed-Width [1695]. fixedpoint [555]. fixedpoints [629]. Fixpoint [280]. Flexibility [1510]. Flexible [1549]. Flexibly [1661]. Floating [385, 64, 163, 216, 62]. Floating-Point [385, 64, 163, 216]. Flores [1169]. Flow [94, 1437, 51, 1540, 1247, 1273, 1277, 1339, 1327, 1398, 1463, 1444, 1617, 665, 1639, 847, 70, 1401, 716, 668, 560]. Flowcharts [697]. FlowMenu [1643]. Flows [1246, 1493, 658]. Floyd [363]. Fluid [873, 592]. fluids [70]. FOL [586, 443, 602]. FOLDS [337]. Font [922, 1276, 992, 903]. For-All [1249]. Foraging [1712]. Force [940, 675]. forest [587]. Fork [1493]. Form [489, 1741, 804, 309, 750]. Formal [1229, 1318, 607, 1588, 337, 835, 1000, 1578, 1595, 113, 169, 485, 107, 1582, 20, 1574, 1618, 108, 984, 702]. Formalization [1425]. Formalizing [1097, 756]. Formally [1307]. Formation [1287, 226, 398]. Forms [142]. formula [734]. Formulae [153, 152, 787]. Formulary [161]. Formulas [1568]. FORTRAN [79, 729, 795, 209, 186, 196]. Fortran-to-Pcode [729]. Fortran-to-Universal [795]. Forum [1177, 1441]. Foundation [1616]. Foundations [112, 1661, 710, 826]. Founded [1252, 936]. four [294]. Fourier [1451, 146, 635, 386, 6, 32, 631]. Fourteen [357]. Fourth [1441]. Fractional [1445, 1475, 1434]. fragment [1335]. Fragments [948]. frame [247]. Framework [910, 1534, 1405, 1299, 1601, 715, 1662, 1106, 1270, 625, 731, 855]. Free [1549, 227, 1136, 1556, 1646, 79]. French [1029, 395]. Friction [1224]. Fujii [308]. Full [1730]. Function [1021, 233, 1136, 1101, 570, 950, 1313]. Function-Free [1136]. Functional [1229, 1318, 1479, 222, 29, 973]. Functions [1257, 203, 526, 274, 732, 1221, 3, 304, 459, 1491, 205, 413, 383, 296, 1621, 58, 36, 144, 72, 1267, 137, 583, 629, 694]. Fundamentals [1433]. Future [1722].

G [1339, 12, 753, 1372]. G-stability [753]. Gaining [1345]. Game [760, 1673, 250, 676, 558, 634]. Games [116, 5]. gap [652]. Gaps [123]. Garbage [1528]. Gauss [81]. Gaussian [849, 537]. Gaze [1700, 1701, 1694]. Gaze-based [1694]. Gaze-enhanced [1700]. GEMS [232]. Gene [1744]. General [1565, 1299, 71, 1592, 306, 1160, 1111, 66, 961]. generality [176]. Generalization [284, 1133, 150, 20, 242]. Generalized [544, 552, 1264, 722, 1228, 637, 87, 374, 491, 659, 1639, 599, 54, 237, 653]. Generals [864]. Generated [1729, 1670, 250, 283, 514]. Generating [1676, 1221]. Generation [1710, 1734, 472, 977, 1328, 400, 1214, 392, 80, 91, 546, 395, 265, 531]. Generative [632, 1535]. Generator [347, 912]. Generators [616]. Genetic [1340, 1545, 1560]. genetically [1340]. Genetics [785, 611]. geodetic [788]. GEOMED [425]. Geometer [1647]. Geometric [473, 1062, 722, 1603, 657, 1442, 425]. Geometry [1367, 1424, 1356, 1234, 1440, 674]. German [356]. GFq [1009]. GFq-polynomials [1009]. Gigapixel [1678]. GLISP [918]. Global [1527, 649, 560]. Globally [1042]. Globally-Asynchronous [1042]. Glue [1461, 1379]. Gluing [1689]. Go [188, 427, 148, 250]. Goal [758, 758]. Goals [913, 379]. Golub [1744]. ’good [78, 888, 1201, 637]. Gordonova [164]. gossip [708]. Gossiping [776]. grade [1003]. Gradient [544, 561, 599, 61, 11, 653, 552]. Grading [17, 356]. Graeffe [63]. Gram [122]. Grammar [1429, 89, 117, 90, 98, 86, 356, 683, 108, 92]. Grammar-Driven [1429]. Grammars [1531, 88, 424, 912]. Grammatical [125, 139]. Graph [1531, 1238, 849, 1044, 1559, 506, 665, 590, 185, 537, 82, 723, 326, 597, 212, 865, 846, 520]. Graphic [970, 1088]. Graphical [232, 1158, 1736]. Graphics [559, 1659, 977, 1101, 821, 678, 314]. Graphs [893, 1388, 1060, 1141, 1506, 1138, 192, 1577, 1054, 1235, 392, 416, 697, 915, 499, 373, 281, 264, 528, 848, 515, 676, 791, 200, 642, 825, 558, 542, 500, 539]. Grassroots [1601]. Greatest [521]. Greedy [1020]. Green [1722]. Grid [285, 735]. Grip [1180]. Group [1557, 1518]. Groups [1421, 1078]. Groupware [1680]. Growing [391]. Growth [304]. Guidable [1650]. Guide [1099, 202, 575, 1692]. guidelines [1626]. GUIDON [1015, 1168, 1088]. GUIDON-WATCH [1088]. Gyroscopes [1723].

H [1744]. Hacking [1689]. Hadamard [315]. Hamilton [1258]. Hamiltonian [264]. Hand [320, 1668]. Hand-Eye [320]. Hand-Held [1668]. Handbook [757, 133]. Handed [1495]. Hands [940]. Haptic [1681]. Hardness [1469, 1489]. Hardware [1659, 981, 1671, 1722, 965, 647]. harmonious [791]. Harper [333]. Hash [378]. Hashing [569, 701, 817, 1059, 522]. Having [907]. HCI [1421]. Head [1658, 252]. healthcare [806]. Heap [1644]. Heart [1596]. Height [223]. Heightened [27]. Held [1668]. Help [1527]. Helping [1720]. helps [1724]. HERACLES [1168]. Hermite [763]. Hessenberg [728]. heterodyne [390]. Heterogeneous [1602, 1600, 1685, 1582]. Heuristic [1171, 454, 1282, 1086, 1004, 354, 1320, 584, 122, 136, 250, 381, 1304, 226, 1137, 476, 208]. Heuristics [1057, 221, 95]. Hierarchical [1707, 1657, 658, 1355, 1403, 1574, 1304, 1039, 792]. Hierarchical-Memory [1403]. Hierarchies [1515]. Hierarchy [1204, 1483, 292]. Hieroglyphs [1276]. High [1710, 1298, 1524, 1134, 972, 1647, 1538, 1399, 1537, 362, 1737, 1319, 1430]. High-Coverage [1710]. High-Dimensional [1537]. High-Performance [1298, 1524, 972, 1399, 1319, 1430]. Higher [1021]. Highlighting [1720]. Highlights [1598]. highly [1369]. Hints [414]. History [1523]. History-Based [1523]. Hoffman [150]. holes [1653, 1652]. homogeneous [398]. Horn [1589]. Hospital [975]. Host [1078, 1399]. ’Hot [1139]. Householder [97]. Hubert [333]. Huffman [860]. Hull [904]. hulls [748]. Human [591, 1421, 1526, 1703, 334, 183]. Human-Computer [1703]. Husbands [1260]. Hy [995]. Hy-phen-a-tion [995]. Hybrid [1567, 1611, 1435, 1679, 561, 492]. Hybrids [1410]. Hyperbolic [943, 920, 706, 824, 907, 619, 492]. Hypercube [1285, 1306]. Hypercubes [1447]. Hypergraphs [917, 870]. hypertext [1407]. Hypotheses [816]. Hypothesis [1287, 1039]. hypothetical [487].

I/O [1154, 339]. IBM [175, 220, 257, 256]. ICCG [1355]. Idea [1684]. Identification [1224, 1500, 329, 168]. Identifying [1718]. Identities [605, 1330]. Identity [1739]. II [155, 1265, 892, 861, 483]. III [479, 450, 484]. Ill [489, 1741]. Ill-Conditioned [489, 1741]. ILLIAC [440]. Image [409, 1714, 1134, 1638, 1634, 1664]. Image-Based [1714]. Images [1500, 878, 1637, 449, 1603]. impact [119, 1432]. Imperfect [180]. Implement [1646]. Implement-Free [1646]. Implementable [1203]. Implementation [1359, 1446, 1149, 213, 1327, 1463, 1617, 135, 1671, 1183, 296, 1324, 1492, 1319, 302, 291, 1017, 457, 111, 522, 336, 119, 376, 808]. Implementing [1437, 1540, 1444, 1407, 725]. Implication [933]. Implications [1231]. Implicit [1344, 1687]. Imprecise [1424]. Improve [1632]. Improvement [1490]. Improvements [841]. Improving [1701, 271, 29, 1343]. incidence [520]. Incomplete [1593, 1172]. Incompletely [579]. Incorrectly [154]. Incremental [1326, 1299, 1470]. Incrementally [786]. Indefinite [410, 1267]. independence [495]. Independent [1470, 294, 385, 1125, 1654, 836, 563]. Index [1478, 1511]. Indian [983]. Indices [1458]. Individual [1248, 1262, 739]. Individuation [1558]. Indoor [1725]. Induction [295, 1368, 754]. Inductive [934, 1115, 248, 1163]. Industrial [827]. Inequalities [720, 529]. Inequality [720]. Inertial [1481]. Inexpensive [1664]. Inference [246, 646, 934, 742, 1589, 139, 400, 1136, 371, 1083, 1130, 125, 261, 394]. Inferences [357, 481]. Infinite [348]. Infinity [1716]. Info [1314]. Informal [786]. Informatics [1616]. Information [1344, 84, 1600, 352, 1615, 685, 1553, 1594, 1601, 960, 1126, 1599, 1632, 166, 1634, 420, 1480, 868, 1172, 1231, 1478, 1511, 215]. Informing [1686, 1706, 1665]. Infrastructure [1644]. inheritance [1349]. Initial [635, 706, 687, 1217, 824, 579, 492, 592]. Initial-Boundary [824, 592]. Inner [658, 175]. Input [1701, 79]. input/output [79]. Inputs [1676]. Inputting [866]. Inquiring [50]. Insertion [86, 471]. Insight [1345]. Instantaneous [1211]. Instantiating [1371]. Instruction [356, 673]. Instrumented [1176, 1207]. Integer [1696, 85, 529]. Integers [282, 338, 496, 343]. Integral [4, 266, 490]. Integrated [814, 1672, 582, 1677]. Integrates [1046]. Integrating [1713, 960]. Integration [1359, 1615, 819, 1410]. Integrity [1553]. Intelligence [969, 1004, 830, 757, 1052, 1248, 1262, 1037, 584, 400, 710, 826, 1315, 1160, 833, 1111, 470, 422, 415, 476, 636, 316, 247, 272]. Intelligent [909, 1170, 853, 1245, 1513, 1636, 1100, 1357, 1194]. Intensional [1080]. Intensionality [648]. Intensity [949]. Intention [193, 201]. intentions [481]. Interacting [1647]. Interaction [1648, 742, 1421, 1672, 1687, 1646, 1680, 1100, 131, 1440, 481]. Interactive [758, 968, 546, 1656, 1644, 1651, 1666, 1010, 1679, 1674, 1662, 423, 1678, 91, 1023, 289]. Interconnection [1308]. Interconnections [439, 399]. interest [961]. Interface [1429, 1154, 1674, 1703, 1088]. Interfaces [1685, 1705, 1678]. Interfacing [1399, 970]. Interference [461]. Interior [1283]. Interior-Point [1283]. Interleaved [461, 434]. Interleaving [1712, 1354, 1610, 1743]. Interlisp [896]. Interlisp-VAX [896]. Intermediate [377]. Intermediate-Level [377]. Intermittent [571, 572]. Internet [1597]. Internetwork [1431]. Internetworking [1298]. Internetworks [1233, 1078]. Interoperability [1598]. interpolants [651, 763]. Interpolating [205, 413, 1068, 583]. Interpolation [1655, 153, 152, 1663, 60]. Interposition [1721]. Interpretation [1045, 180, 950, 604]. Interpreter [1202, 47]. Interpreting [1179, 1113]. Interprocess [1027]. interrupts [512]. Intersection [1235, 696, 1585, 478]. Interval [1611, 987, 109]. Intervals [981]. interviews [334]. Intractable [1032]. Intrinsic [1634]. Introduction [978, 191, 570, 1380, 199]. introductory [547, 747]. Invariance [1614, 1022]. Invariances [861]. Invariant [877, 66]. Inverse [638, 699, 1526, 401, 8, 34, 751]. inversion [593]. Inverted [1458]. Investigation [114, 1712]. Involutions [605]. Involving [353, 509]. irreducibility [1009]. Irregular [195]. Irrelevance [1499]. iSecurity [1662]. Isn’t [147, 613]. Isomers [400]. Isomorphic [15]. Isomorphism [192, 723, 200]. Issues [1674, 294]. item [961]. Iteration [87, 283, 34]. Iterative [97, 1265, 240, 560, 507, 1699, 312, 689]. Itself [1670]. IV [79, 533, 440]. IV-Type [440]. Iverson [47].

J [184]. J. [591]. Jacobi [699, 617]. Janus [1730]. Join [1286]. joins [1377]. Jordan [489, 1741]. Joys [1631]. judges [334]. July [582]. Just [1703]. Justification [1391]. Justification-based [1391]. Justified [1284].

Kaleidoscope [1429]. Karzanov [1339]. Kasami [308]. Kernel [991]. Key [871, 1528]. Keyboard [1645]. Khachian [765]. Khachiyan [790]. Kinematic [1342, 1226, 940, 675]. Kinematics [1526, 93]. KLEE [1710]. knapsack [843]. Knee [867]. Knowledge [832, 809, 1156, 1096, 1734, 1117, 1166, 1201, 1046, 530, 578, 830, 785, 1109, 770, 793, 1544, 1499, 1097, 1543, 1140, 975, 1214, 1206, 1397, 1357, 839, 1256, 1255, 738, 1200, 1261, 836, 714, 1013, 604, 636, 1115, 487, 654, 672, 1088, 1546, 611, 595]. Knowledge-Acquisition [1214]. Knowledge-Based [530, 785, 770, 1206, 1357, 839, 836, 1156, 714, 604, 1088, 1546]. Kowalski [924]. KRL [595]. Kronecker [750]. Krylov [1296].

L [363]. L. [184]. l8r [1690]. labeling [326]. Labelled [1000]. Laboratory [761]. Lagrangian [767]. Lakoff [297]. Lambda [1362]. Lanczos [278, 310, 650, 1296, 507]. Landmark [1454]. Landmark-Based [1454]. Language [913, 953, 774, 275, 758, 83, 359, 1179, 1484, 337, 90, 769, 686, 472, 545, 978, 1010, 1058, 1520, 1028, 220, 257, 967, 48, 1461, 430, 1674, 166, 505, 371, 1504, 58, 821, 807, 632, 1184, 1185, 1390, 78, 442, 316, 119, 414, 884, 256, 840, 964, 307, 120, 394, 447, 33, 53, 96, 595, 187, 135, 575]. Language/Action [1184, 1390]. Languages [1446, 1619, 228, 1000, 1479, 1218, 1302, 311, 954, 1017, 678, 576, 504, 511, 797, 882, 65]. LANs [1233]. Laplace [151, 22]. Laplacian [121]. Lapse [1449]. Large [1729, 921, 1648, 578, 788, 1647, 2, 250, 260, 1591, 1286, 507, 244, 1678, 450, 780, 520, 754, 309]. Large-Scale [1591, 260, 309]. Largest [1563]. Laser [1648]. Laser-Based [1648]. Last [885]. Later [1386]. Lax [619]. Layered [1162, 1054]. layers [541]. Layout [978]. Layouts [888]. LCCD [840]. LCF [457, 458, 345, 338, 343, 531]. LCFsmall [457]. Leaf [1157]. learn [1049, 1055]. Learned [996]. Learners [1569]. Learning [1622, 1606, 1712, 612, 707, 1734, 908, 934, 1715, 1109, 1091, 1421, 1057, 1299, 1351, 1469, 1489, 1728, 1727, 1604, 718, 1674, 1632, 1529, 1586, 95, 1194, 1261, 973, 620, 183]. Leases [1325]. Least [1191, 1269, 73, 133, 410, 445, 97, 712, 12, 134, 269, 573, 788]. Least-Squares [1191, 1269]. Lecture [112, 1384, 1471, 1459, 52, 100]. Lectures [703, 470]. Left [1403]. Left-Looking [1403]. Legal [1052]. Leibniz [422]. lemmas [299]. Length [437, 62]. Lessons [1168, 996]. Letter [811]. Letterform [989]. lettershape [984]. Level [578, 1109, 377, 859, 965, 1290, 103, 574]. Leveraging [1524]. Lexical [86]. Libraries [1598]. Library [1571, 1598, 688]. Light [1707, 1668, 1627]. Like [167, 1077, 1080]. Likelihood [352]. Limitations [1502]. Limited [1240, 1702, 352, 1612]. Line [1134, 180, 366, 1548, 1113]. Linear [790, 720, 838, 1696, 877, 552, 267, 893, 293, 46, 765, 459, 133, 134, 142, 310, 282, 24, 818, 227, 303, 876, 1475, 684, 259, 60, 67, 97, 230, 689, 290, 516, 450, 141, 670, 1265, 712, 12, 241, 283, 753, 331, 846, 217, 704, 18, 31, 260, 309, 618, 779]. linear-time [846]. Linearity [614]. Lines [717, 844]. linguistic [379]. Linguistics [632, 252, 297]. Link [859]. Link-Level [859]. LISP [607, 485, 1146, 1024, 245, 1077, 1202, 1080, 1225, 1289]. Lisp-like [1077, 1080]. List [1111, 1315]. Lists [267, 338, 343]. Literate [999]. Literature [1095, 1112, 1427, 1142]. Liveness [1022]. Load [1532, 1241, 1306, 25]. Loads [1532]. Local [1159, 960, 891, 1102, 852]. Locally [1042]. Locally-Synchronous [1042]. locking [865]. Locomotion [1555]. Log [1196, 1297]. Logarithmic [493, 550, 435, 437]. Logging [1681]. Logic [1178, 1291, 1223, 1251, 1274, 1613, 1132, 854, 295, 1089, 1114, 1147, 703, 766, 889, 1568, 680, 296, 1408, 965, 987, 1410, 1368, 1252, 1417, 1155, 443, 944, 477, 755, 533, 626, 297]. Logic-Based [1613]. Logical [1257, 1089, 876, 1017, 1107, 1106, 376, 483, 882, 431, 1120]. Logics [1333, 1344, 881, 1038, 1347]. logistic [1711]. Logs [1684]. long [515]. Longest [488]. Longitudinal [1686]. Looking [1403]. loop [641]. Loops [658]. Loosely [1148]. Loosely-Coupled [1148]. Low [559, 432]. Low-Cost [432]. Lower [1215, 1614, 760, 387, 879, 828, 243, 660, 748]. Lowering [1690]. LumiPoint [1648].

M [1339, 1372]. MAC [1730]. Machine [1437, 78, 1314, 1620, 1489, 385, 296, 1632, 839, 131, 95, 183, 266, 401, 272]. Machines [246, 1292, 1293, 1587, 1645, 740, 1403, 1187, 618, 946, 660]. macro [514]. MAINSAIL [808, 807]. maintain [618]. Maintaining [1612]. Maintenance [578, 1640, 1414]. Make [1670]. Making [1602, 1245, 738, 951]. Man [131]. Management [1485, 890, 1212, 1405, 1572, 1624, 1717, 119, 1420]. Managing [1270, 1039]. Manipulation [835, 1481, 1547, 1548, 212]. Manipulator [1549, 581, 968, 574, 546]. Manipulators [1226, 938, 93, 675]. Manual [1579, 181, 1361, 1379, 235, 1415, 807, 902, 187, 92, 730, 1023, 746, 733, 906, 918, 37, 1003, 258, 384, 602, 363]. Manufacturing [1543, 568]. Many [1027, 1693]. Many-Sorted [1693]. Mapping [1600, 1133, 1708, 1390]. Mappings [1074, 1593]. Maps [851, 924, 1681]. Marking [1295]. Markov [1387, 348, 1375]. Mashing [1689]. mass [1576]. Massive [1714]. Massively [1241]. Massively-Parallel [1241]. Matches [493, 550]. Matching [935, 449, 1064, 1573, 451, 1558, 1144, 1422, 1188, 1300, 442, 336, 644]. Matchings [1395, 1253, 276]. Math [147]. Mathematical [275, 211, 663, 1213, 189, 197, 227, 162, 96, 187, 575]. Mathematics [771, 554, 1254, 1037, 584, 802]. Matrices [699, 10, 508, 303, 1, 59, 1740, 497, 638, 57, 42, 782]. Matrix [1562, 699, 1744, 459, 8, 73, 124, 348, 41, 1296, 610, 700, 28, 196, 305, 875, 899, 66, 312, 157, 330, 327, 650, 892, 43, 237, 520, 13, 1311, 1331, 1343, 34, 617]. Matroid [355, 478]. Matroids [464, 506]. Max [1491]. Maximal [895, 520]. Maximizing [16]. Maximum [1437, 1540, 352, 1273, 1327, 847, 144, 336, 716, 563]. maximum-flow [716]. MC [1290]. Mean [1353]. Meaning [430]. Means [823]. Measurement [1174, 302]. Measures [1607]. Mechanical [454, 152, 1466]. Mechanism [1054, 1325]. Mechanisms [1686, 1706, 939]. Mechanization [1318, 1396]. Mechanizing [1344]. Mediator [1329]. Medical [1613, 1046, 1581, 1634, 738, 833, 947, 951, 942]. Medicine [772, 1616, 1582]. Meeting [1177, 1441]. Megaprogramming [1366]. member [743]. Memories [1134, 1077, 225, 399]. Memory [461, 890, 1212, 1524, 1468, 1190, 1514, 439, 474, 418, 430, 1355, 1403, 1483, 193, 1050, 1040, 1369, 1420, 1077, 1275, 217, 434, 201, 525, 725]. Memory-Based [1524, 1514]. Mental [740]. Mentally [83]. Mentoring [1421]. Menu [1429]. Merge [293, 684, 242, 243]. Merging [1299, 640, 480]. Meru [1716]. Mesh [943]. Meshes [1237, 1657, 781]. Message [1106, 1102]. Messages [1456, 437]. Messaging [1524, 1514]. Meta [578, 1109, 232, 664, 992, 903]. Meta-DENDRAL [664]. Meta-Font [992, 903]. Meta-Level [1109]. Metacognitive [1201]. METAFONT [777, 996, 1094]. METAFONTware [1279]. Metalinguistic [19]. Metamarks [1280]. metamathematics [477]. Metarules [1117]. Method [1437, 9, 691, 114, 4, 635, 920, 1540, 544, 552, 3, 213, 1657, 284, 1410, 305, 1434, 657, 507, 27, 151, 67, 56, 1137, 128, 155, 599, 522, 61, 262, 650, 637, 728, 1039, 952, 631, 331, 389, 11, 18, 63, 1546]. Methodologies [1409]. Methodology [909, 1488, 911, 986, 484]. Methods [1332, 1062, 454, 687, 386, 46, 330, 1283, 145, 1595, 1190, 925, 248, 1568, 22, 927, 1423, 628, 580, 230, 1418, 619, 286, 670, 1265, 159, 630, 12, 44, 45, 753, 492, 31, 101, 1627]. Metric [1472, 498]. Metrics [1638]. Micro [191, 199]. Micro-Computer [191, 199]. Microprocessor [711]. Microprocessor-Based [711]. microprogram [525]. Microprogrammed [302]. Migration [609]. mild [541]. mildly [689]. Milestones [1744]. Min [1491]. Min-Max [1491]. Minimal [1258, 540, 324]. Minimax [767]. Minimization [679, 253, 601, 24, 527, 233]. Minimizing [190, 1430, 1071, 198, 694].

Minimum [1246, 358, 1444, 1493, 1617, 32, 1353, 144, 656]. Minimum-Cost [1246, 1617]. Minimum-Mean [1353]. MinimumCost [1463]. Mininet [1726]. Mismatched [1312]. MIX [202]. MIX/360 [202]. Mixed [1696, 1650, 342]. Mixed-Integer [1696]. Mixture [1564]. MLISP [105, 179]. MLISP2 [369]. MLP [1318]. Mobile [1342, 1690, 1717, 1450, 1505]. Modal [1123, 766]. ’model [252, 94, 84, 612, 1429, 648, 1589, 377, 1539, 1354, 161, 1038, 1089, 1330, 1558, 1214, 1603, 1070, 1142, 1582, 1511, 593, 160, 1087, 156, 481, 672, 418, 620, 1570, 737, 1120]. Model-Based [1429, 1558, 1214, 1570]. Model-Directed [612]. Model-Matching [1558]. model-theoretic [1120]. Modeling [1224, 473, 1608, 1467, 1367, 279, 1591, 1570, 1165, 1194, 567, 930, 951]. Modelling [1018, 823, 319, 50, 421, 1135, 961]. Models [1606, 878, 707, 1167, 1166, 955, 1472, 1531, 819, 1259, 1345, 1486, 786, 1374, 345, 1581, 1564, 1566, 1574, 1535, 171, 241, 546, 1069, 1626]. Modern [1276, 798, 802, 1311]. Modification [1556, 600, 1338]. Modified [1267, 462, 734, 239]. modifying [330]. Modular [1555, 1631]. Modules [439, 434]. Molecular [479, 785, 1287]. Molecule [381]. molecules [373, 398]. moments [141, 1267]. monadic [222]. Monitoring [1718, 719, 1200]. Monitors [412]. monotone [566]. monotonic [1349, 804]. monotonicity [775]. Monte [568, 1627]. Moore [1318]. Mordell [184]. Moser [294]. Motion [1281, 1211, 1649, 1494, 1650, 1394, 1216, 1317, 1370, 1551, 938, 1550, 1465, 1496, 1464]. Motions [1450, 1505]. Motivation [648]. Mounted [1646]. Mover [1614]. MPL [96, 187]. MRS [1099]. MTPT [340, 341]. Muir [1185]. Multi [1648, 1290, 361, 1024, 1547, 1545, 965, 1396, 668, 1474, 1675, 1369, 103, 884]. Multi-Arm [1547]. multi-computer [1369]. Multi-Level [965, 1290, 103]. Multi-Part [1545]. Multi-Process [1396]. Multi-processing [1024, 884]. Multi-terminal [668]. Multi-User [1648, 1675]. Multiagent [1622]. Multibody [1370]. Multibutterfly [1453]. Multicast [1197, 1175, 1271, 1233, 1431, 1609, 1078]. Multicoloring [596]. Multicommodity [1398, 1617, 1639, 1401]. Multidatabase [1485]. Multidimensional [360, 815]. Multifrontal [1403]. Multigrid [1323]. Multilevel [1419]. Multiple [913, 526, 907, 1645, 1477, 1243, 423, 59, 1740, 339, 1377]. Multiple-precision [526]. Multiplication [1562, 899, 157, 32]. multiplicative [496, 553]. Multiprocessing [891]. Multiprocessor [461, 1173, 1271, 1128, 1217, 1510, 972, 1468, 1514, 1206, 1324, 1355, 1343]. Multiprocessors [1419]. multiprogrammed [867, 368]. multiset [666]. multistep [670, 753, 31]. MUTANT [120]. mutual [173]. MYCIN [475].

NAIL [1205, 1131]. Names [1428]. Naming [1122, 1198]. Natural [913, 774, 1179, 769, 686, 1398, 472, 205, 413, 1520, 430, 166, 371, 1402, 394, 447, 78, 442, 583, 307, 297]. Natural-Language [913, 166]. Navigation [1454, 831, 1638]. NC [1098, 1064]. near [524]. Nearest [456]. Nearly [615]. necessary [62]. Negation [1252]. neglect [42]. neighbor [524]. Neighbors [456]. Nelson [177]. Neomycin [929, 1168]. nested [659]. NESTOR [1046]. Net [682]. Nets [27, 796]. Network [1053, 1174, 1247, 1273, 1277, 1453, 1203, 1393, 1572, 1560, 439, 1183, 1400, 1301, 1624, 437, 847, 1337, 1102, 476, 1071, 1303, 852]. Network-Centric [1624]. Networking [1732]. Networks [1388, 1565, 1341, 1592, 1399, 1477, 1554, 1374, 1375, 1581, 1308, 268, 435, 436, 891, 1443, 342, 1630, 218, 566, 242, 243, 244, 736]. Neumann [262]. Neural [1581]. NEVER [132, 799, 862, 863, 916]. Newton [56, 927, 1434, 628]. Nineteenth [1177]. Ninomiya [308]. nodal [1331]. nodes [1630]. nomenclature [1264]. nominals [307]. Non [1589, 15, 1641, 1019, 312, 398, 689, 1265, 159, 804, 18, 743, 618, 1667]. Non-blocking [1641]. Non-Clausal [1019]. non-disjoint [743]. non-homogeneous [398]. Non-Horn [1589]. Non-Isomorphic [15]. non-linear [689, 18, 618]. non-monotonic [804]. non-rectangular [159]. non-self-adjoint [1265]. non-stationary [312]. non-XP [1667]. Nonclausal [1076, 1147]. nondifferentiable [694]. Nonholonomic [1370]. Nonlinear [1191, 1269, 459, 383, 767, 445, 628, 657, 599, 269, 25, 234, 541]. Nonlinearly [671, 580]. Nonmonotonic [1411, 1397, 1347, 1268]. NonParametric [498]. Nonperiodic [635]. nonseparable [286]. Nonspeaking [83, 359]. Nonstandard [706]. Nonsubsective [1735]. Nonsymmetric [552]. Normal [10, 262]. Normative [1531]. norms [75]. notation [47]. Notations [1295, 227]. Note [372, 751, 184]. Notebook [1686, 1706]. Notes [112, 1384, 1471, 148, 491, 1459, 353, 747, 1665, 100, 52]. Nothing [1480]. Notions [997]. Novel [1730, 441]. NP [723]. NP-Complete [723]. NQTHM [1396]. Number [335, 367, 24, 347, 590, 1253, 1353, 616, 597, 495, 1576, 271, 434, 64]. Numbers [963, 780]. Numerical [1310, 114, 162, 4, 51, 230, 850, 877, 687, 544, 599, 630, 241, 310, 501, 905, 671, 323, 291, 726, 580, 688, 286, 122, 62, 12]. Numerically [153, 617].

O [1154, 339]. O.S [79]. Object [1446, 1619, 1109, 1330, 1523, 1402, 546]. Object-Level [1109]. Object-Oriented [1446, 1619]. Objects [313, 1336, 1528, 1371, 895, 1113, 474, 1270, 970, 1682, 398]. Oblivious [1577]. Observability [877]. Observation [742]. Obstacle [831]. Obstacles [1451]. Obtaining [3]. Occlusion [1649]. odd [128, 155]. odd/even [128, 155]. off [156, 42]. off-diagonal [42]. offs [634]. Ominoes [335, 271]. On-Line [1548]. Once [1196, 1297, 1050]. One [679, 920, 1027, 27, 252]. One-Dimensional [679]. One-to-Many [1027]. Online [1696, 1714, 1609, 1501, 1728]. only [813]. onto [1133]. Open [342]. Operating [1717, 976, 412]. Operation [1149, 1250, 875, 283]. Operational [517, 1275]. Operations [1149, 1312, 24, 1545, 165, 37, 852, 431]. Operator [22, 66, 76]. Operators [4, 1193, 490]. Opportunistic [1712, 1689]. Opportunities [1580]. OPS5 [1192, 1422]. Optical [1314]. Optimal [1515, 1237, 615, 901, 1288, 922, 366, 1334, 887, 7, 1059, 556, 865, 478, 555, 480]. Optimality [684]. Optimization [1392, 1484, 1615, 1384, 1471, 459, 851, 1587, 671, 1538, 1559, 874, 1408, 1426, 767, 1475, 1434, 1286, 580, 735, 287]. Optimize [1512]. Optimized [1524, 509]. Optimizing [1446, 1514, 1136, 1152]. Optimum [149, 1180, 61, 389]. OR349 [1384]. Order [732, 295, 739, 1408, 362, 1654, 443, 477, 266, 619, 670, 1069]. Order-Independent [1654]. order-theoretic [1069]. Ordered [378, 1502, 399]. orderings [666]. Orders [1048, 1135, 775, 837]. Organ [1018]. Organic [381, 373, 402]. Organization [873, 1036, 1005, 103]. Organizing [1195, 1661, 1601]. orientations [528]. Oriented [771, 816, 921, 761, 758, 1446, 764, 772, 1619, 1462, 130, 1360, 1504, 562]. Origins [1065, 1066]. orthogonal [1266, 788, 782]. orthonormalization [122]. other [178, 262]. Outline [1276]. output [79]. overdetermined [67]. Overlapping [1693]. Overview [1485, 535, 1131, 1598, 595, 808]. own [1049, 1055].

P [1030]. P-Complete [1030]. package [79, 583]. Packages [1721]. Packet [1399]. Packet-Switched [1399]. Packing [1220, 1445, 494, 677]. Page [1622, 1420]. page-cache [1420]. Pagination [887]. Paging [196, 462, 593, 339]. Pair [536]. pairwise [787]. palindrome [660]. Palladio [1002]. Pandora [1280]. Paper [3, 1678, 1699]. Paper-Based [1678]. Papers [288, 962, 833, 363, 294]. Parabolic [579, 31]. Paradata [1377]. Paradigm [1065, 1066, 1340, 964, 447, 1369]. Paradox [1195]. Paragraphs [844]. Parallel [1110, 1220, 1156, 1659, 1149, 1250, 1715, 229, 1395, 1247, 1230, 1283, 1453, 1241, 1587, 1041, 1138, 405, 1044, 1093, 323, 306, 322, 1560, 191, 700, 1244, 1285, 1009, 1192, 441, 421, 1070, 1355, 1462, 1483, 1423, 158, 259, 433, 1323, 1107, 697, 1225, 1289, 1231, 172, 160, 171, 199, 897, 1331, 1432, 792, 399, 843, 1377]. Parallelism [1020, 324]. Parallelizing [1492, 1422]. Parameter [1224, 637]. Parameterization [1218]. Parametric [1708]. paranoia [481]. Paranoid [361, 251, 334, 360]. Paraphraser [370]. Parenthesis [227]. Parenthesis-Free [227]. PARRY [468]. parser [683]. Parsing [1520, 107, 697]. Part [892, 1545, 155, 100, 1487, 892, 855, 861]. Partial [943, 288, 966, 274, 544, 14, 1048, 1553, 907, 1625, 925, 189, 1135, 291, 721, 599, 775, 197, 28, 837]. Partially [431]. Partition [1053, 605, 1]. Partitioned [1397, 782]. Partitioning [1012, 972, 355, 301, 974, 1101, 1482]. partitions [500]. parts [568]. Party [1599]. Pascal [746, 458, 265, 829, 1079]. Passing [1106, 1102]. Password [1694]. Path [1310, 1506, 1189, 1537, 1554, 825, 136, 1304, 1440, 523, 539, 749, 744]. Path-regular [825]. Paths [1549, 1258, 1497, 1456, 1455, 1556, 515, 1071, 669]. patients [334]. Pattern [442, 1567, 1520, 451, 888, 375, 644]. Pattern-matching [442]. Patterns [51, 1637, 1328, 762, 70]. PCFORT [729]. PCODE [795, 729]. Pebble [760, 676, 634]. People [1601, 168, 1380, 1724]. Perception [979, 1417, 420, 532]. Perceptual [1036, 1638]. percolation [952]. Perfect [1138, 1064, 438, 158]. Performance [1515, 991, 1217, 1298, 1524, 1655, 1588, 1715, 1269, 339, 759, 972, 1468, 1389, 1726, 1538, 1399, 1577, 1116, 1026, 1374, 1324, 1737, 1319, 673, 460, 1458, 1430, 1343, 1432]. Performing [1312, 438, 852, 431]. period [582, 389]. Periodic [699, 1487, 1532, 497]. peripheral [504]. Permutations [353, 294]. Perseus [976]. Person [5]. Perspective [1625, 1571, 1184]. Pet [967]. PF [1490]. phen [995]. philosophy [422]. photogrammetry [482]. Photographs [1661]. Photography [1668]. Photonic [1203]. Physical [1672, 974, 1467, 911, 986, 1677, 1420, 1728, 681, 973]. Physician [1427]. Physicians [475, 833]. Picture [1321, 91, 289]. Pictures [224, 107]. Piles [1675]. pipeline [1377]. Pipelined [1285]. pitch [389]. Pitfalls [147]. Pixel [1590]. Pixel-To-Pixel [1590]. PL360 [220, 257, 256, 68, 104]. PLACA [1504]. Placement [1649]. placements [1576]. Plan [930, 1521, 1338]. Planar [893, 192, 915, 200, 643, 642, 846]. Planarity [206, 249]. Plane [223]. Planning [913, 1549, 1310, 1281, 1494, 1383, 958, 1615, 1394, 1537, 1554, 1320, 1216, 1317, 1370, 1551, 1548, 1610, 1743, 1550, 1465, 1496, 800, 1450, 1505, 1530, 1442, 1304, 1440]. Plans [853, 1161, 762, 1724]. plastic [11]. Platform [1599]. Play [116]. Pleiades [1629, 1642]. Plenoptic [1668]. plex [119]. Point [1175, 1283, 295, 1611, 385, 1347, 1278, 25, 541, 163, 216, 18, 64, 62]. Point-Based [1611]. Point-to-Point [1175]. Pointer [1645]. Pointer/Keyboard [1645]. pointers [562]. PointRight [1645]. Points [1479, 5, 137, 652]. Poisson [128, 195, 155, 159, 6, 1592, 1630, 27]. Policy [1575]. Poligon [1319]. Polya [294]. Polygon [904, 1180]. polygons [1585]. Polyhedra [536, 15, 1413, 292, 264]. Polyhedral [1031]. polyhedron [328]. Polynomial [9, 720, 358, 1060, 71, 87, 143, 1400, 875, 55, 16, 144, 145, 23, 669]. Polynomials [1288, 36, 1266, 1009]. polyphase [556]. Pomset [1070]. populations [1340]. Portable [1721, 976]. Position [1523]. Positive [9, 144]. Posteriori [143]. PostScript [1276]. Potential [1310, 1576]. Potokovye [1339]. Power [591, 346, 1718, 942]. Practical [615, 607, 275, 592, 751]. Practice [1575, 1573, 1081, 1309, 1639, 1626]. Practicing [1346]. PRAM [1259]. Pratt [363]. Precedence [228, 982, 1033, 231]. Precedence-Based [1033]. precedents [973]. Precisely [1534]. Precision [27, 526]. Predicate [288, 166]. predict [1711]. Prediction [1719, 930]. Predictions [1564]. Preemptable [1105, 1148]. Preference [388]. Preimage [1216]. Preliminary [1280, 1390, 100]. Preparation [1682]. Preprocessing [1507]. Presentation [1685]. Presentations [1158]. Preserve [598]. Pretty [784]. Price [1527]. Primal [1]. Prime [123, 1060]. primer [586]. Primitive [1149, 1250, 273, 357]. Primitives [1234, 1474, 792]. Principles [945, 1022]. PRINTED [132, 799, 862, 863, 916]. Printers [1276]. Printing [1276, 784]. Prints [1678]. Priority [615, 267, 471]. Prisoner [1047]. Privacy [161].

Probabilistic [1378, 1046, 1341, 1393, 1537, 1618, 1535, 1570, 660]. Probabilistic-Network [1393]. Probability [1448, 780, 375]. Problem [1437, 1030, 332, 699, 635, 454, 1376, 1540, 931, 1034, 958, 177, 1228, 1273, 1455, 1498, 278, 1075, 1363, 1008, 366, 374, 1563, 221, 880, 1007, 1092, 1182, 353, 1181, 1294, 682, 715, 372, 866, 1586, 724, 845, 363, 173, 294, 621, 176, 712, 630, 728, 685, 247, 865, 237, 704, 492, 669, 611, 218, 708, 641, 843]. Problem-Solving [931, 1075, 1008, 880, 1007, 1092, 1181, 1586, 621]. Problematic [797]. Problems [838, 4, 1171, 687, 1628, 1392, 824, 14, 386, 1191, 1269, 320, 1444, 1506, 1189, 1559, 1573, 324, 323, 322, 610, 301, 410, 879, 1032, 1445, 136, 1475, 445, 1235, 131, 579, 1300, 1337, 933, 762, 638, 292, 300, 450, 735, 239, 269, 573, 247, 1340, 1401, 541, 653, 592, 18, 101, 539, 749, 744, 656, 12]. Procedural [512, 1252, 594]. Procedure [691, 1696, 1695, 73, 385, 291, 1731, 312, 12, 271]. Procedures [1084, 205, 413, 834, 667]. Proceedings [1452, 1177, 1441]. Process [1472, 1396, 423, 1165, 122, 951]. Processes [361, 854, 1636, 889, 1070, 7, 577, 944, 1535, 251, 334, 513, 1630]. Processing [84, 686, 793, 1550, 430, 715, 158, 420, 1323, 868, 1024, 119, 884, 514, 289]. Processor [1151, 1327, 1363, 1098, 1474, 368]. Processor-Efficient [1327]. Processors [439, 1243, 165]. Producing [1696]. Product [1388, 1060, 1141, 309]. Production [832, 612, 535, 530, 1144, 731, 633, 608]. production-based [608]. Products [875, 892, 175, 837]. Program [857, 255, 349, 530, 116, 770, 354, 1545, 1006, 923, 1433, 786, 485, 1054, 430, 834, 475, 381, 713, 593, 208, 600, 646, 176, 39, 455, 487, 654, 376, 533, 562, 174, 571, 572, 705, 185, 900, 486, 215, 483, 484, 531, 26]. Program-Synthetic [923]. program-understanding [455, 487]. Programmable [606, 1704, 407, 547]. Programmers [1720, 92]. Programming [790, 838, 1223, 1251, 1274, 96, 275, 1712, 1713, 469, 1446, 346, 1376, 1484, 187, 773, 1619, 765, 686, 968, 1075, 1698, 1008, 1538, 1479, 978, 382, 427, 880, 999, 1007, 1545, 1560, 1218, 1028, 1114, 1147, 220, 257, 703, 1365, 1092, 710, 826, 1302, 311, 1016, 1461, 85, 1272, 1181, 1, 1294, 1360, 510, 423, 1313, 657, 954, 1504, 724, 575, 162, 226, 529, 516, 621, 476, 229, 263, 467, 134, 678, 414, 405, 25, 576, 727, 1340, 511, 256, 626, 661, 797, 704, 653, 964, 120, 260, 309, 52, 882, 33, 53, 65, 564, 194]. Programs [1584, 188, 1222, 1710, 732, 581, 240, 17, 873, 883, 919, 113, 186, 248, 645, 766, 861, 894, 1022, 1249, 1439, 50, 1252, 1152, 1307, 429, 428, 565, 574, 1155, 1079, 1107, 404, 1289, 567, 836, 170, 444, 290, 600, 261, 453, 1340, 533, 393, 549, 855, 936, 985, 1077, 1275, 897, 210, 557, 622, 483, 484, 1396]. Progress [1436, 455, 582]. Project [1421, 214, 816, 921, 1629, 1642, 1666, 1571, 1598, 1380]. Projected [767]. projection [11]. Projections [1219]. Prolegomena [702]. PROLOG [1222]. Pronouncing [1428]. Proof [1344, 1696, 758, 1132, 1409, 1010, 967, 1022, 404, 443, 150, 649, 1023, 533, 562, 805, 985, 1488]. Proof-Producing [1696]. Proofs [835, 477, 622]. Propagation [1045, 153, 926]. Properties [1222, 1074, 349, 304, 1481, 248, 982, 1022, 1204, 658, 527, 151, 1519, 222, 1263, 775, 277, 216, 1077]. Property [986]. Proposal [1597, 1264]. Proposed [123]. Propositional [1132, 1019]. Propositions [739]. Protein [1332, 604]. proteins [1137]. Protocol [1197, 1175, 1174, 1597, 1730, 1157, 1523]. Protocols [871, 859, 1599, 476]. Protograms [1348]. Prototypes [832]. Prototyping [1715, 1672, 1698, 1677]. Prover [454, 1579, 1536]. Provers [1589]. Providers [1599]. Providing [1601, 1724]. Proving [1123, 1178, 1222, 1164, 666, 138, 794, 248, 861, 982, 1077, 509, 429, 428, 829, 571, 572, 936, 1313]. Proxies [1704]. Pruning [452]. Pseudo [8, 1334]. Pseudo-Exhaustive [1334]. Pseudo-inverse [8]. pseudoinverse [13]. pseudoinverses [269]. psychiatric [334]. Public [871]. Publications [1557, 1518]. PUFF [950]. pulmonary [950]. Punctuality [1385]. Pure [1368]. Purpose [919]. Push [1437, 1540, 1444]. Push-Relabel [1437, 1540]. put [995]. PYGMALION [510].

QLISP [1192, 1365, 1272, 1422]. Quadratic [74, 142, 527, 40, 144, 712, 134, 653]. Quadrature [153, 152, 81]. Quadri [416]. Quadri/Trivalent [416]. Qualifying [745, 1315]. Qualitative [1167, 1166, 1287, 1165]. Qualities [740]. Quantities [879]. Quasi [628]. Quasi-Newton [628]. Queries [1587, 1286, 779]. Query [1240, 1600, 1457, 1615, 1593, 1509, 793, 874, 1408, 1426, 1550, 1017, 1107, 1106, 961]. Query-Limited [1240]. Question [138]. Question-Answering [138]. questions [1201]. Queue [615, 1024, 471]. Queue-based [1024]. Queueing [437, 1630]. Queues [267, 1372, 342, 218]. Queuing [1033]. Quicksort [503]. Quintic [413].

R [1276]. Radio [1215, 1730]. Radiosurgery [1465]. radius [524, 144]. Random [1074, 1124, 341, 1559, 1064, 1554, 464, 347, 471, 616, 488, 262, 80, 644]. Randomization [1398]. Randomized [1326, 1507, 1550]. Randomness [598]. range [64, 1687]. Rank [573, 13]. Ranking [1703, 936]. Rapid [152]. Rapidly [1698]. Rasterization [1134]. Ratio [142]. Rational [359, 1063, 1100, 763]. ray [604]. Rayleigh [87]. Reachability [1219]. react [1521]. Reactive [1584, 1606, 1517, 1488, 1635, 1535]. Reactivity [1486]. Read [813]. Read-only [813]. Reading [1315, 1111]. Real [1333, 1404, 1515, 1517, 1588, 1532, 1241, 1354, 1409, 1406, 1477, 1508, 1572, 831, 1556, 236, 1530, 1438, 1552, 1108, 1633, 1476, 34]. Real-Time [1333, 1404, 1515, 1588, 1409, 1406, 1477, 1508, 1556, 236, 1438, 1552, 1517, 1572, 1633, 1476]. Realistic [803, 1742]. realization [964]. Really [1291]. Reason [333, 591]. Reasoning [878, 1004, 1052, 1428, 1534, 1162, 1287, 874, 221, 1551, 1499, 1411, 965, 988, 987, 693, 1142, 1163, 1199, 1466, 1541, 517, 1039, 1268, 804, 702]. Reasons [881]. Recall [1656]. Recognition [1567, 1036, 1520, 474, 49, 1402, 396, 127, 442, 930, 375, 660]. Recognizers [888]. Recognizing [1031]. Recommendation [1622]. Reconciling [1538]. Reconfigurable [1555]. Reconfiguring [929, 1453]. Reconstruction [536, 617]. record [947]. records [562]. Recovery [1650, 1640]. rectangular [159]. Recurrence [253, 324, 323, 306, 322]. Recursion [1232, 1125, 661]. Recursive [274, 732, 1012, 411, 1136, 1083, 58, 404, 290, 549, 629]. Recursively [304, 277, 693]. Redirection [1645]. Reduce [181]. reduced [1265]. Reducibilities [1240, 623]. Reducing [1694, 917, 1683]. Reduction [1595, 128, 155, 483]. Redundant [1226, 1457, 1456]. Reference [235, 618, 1023, 258]. References [1473]. Refined [841]. Refinement [943, 1171, 1200, 1137]. Refinements [1217, 97]. Reflective [1677]. Reformulation [1593]. Reformulations [1284]. Regarding [475, 67]. Region [391, 285, 70]. Regions [195, 159]. regression [1711]. Regular [814, 990, 367, 500, 58, 1733, 825, 946]. Rehearse [1720]. Reinforcement [1569]. Relabel [1437, 1540, 1444]. Related [1720, 320, 1398, 294, 247, 753, 656]. Relation [780, 87, 971, 895, 1335, 1255, 527, 959, 519]. Relational [1336, 1061, 1371, 1426, 839, 1582, 869, 1270, 792]. Relations [1512, 253, 948, 923, 1072]. Relationship [1624]. Relationship-Based [1624]. Relationships [921, 1296, 1407, 942]. relative [163]. Relatively [456]. Relaxation [44, 45, 101]. Relaxing [1385]. Release [1579]. Reliability [1588]. Reliable [538, 1564]. remarks [252]. Remixing [1704]. Remote [1591, 1148, 852, 1105]. Removing [557]. Renaming [1151]. Rendering [1709, 1679, 1681]. repair [1570]. Repetition [1725]. Replicated [1153, 759, 1418, 1470]. Report [88, 1571, 896, 582, 417, 551, 415, 455, 214, 826]. Reports [1525, 1209, 1364, 204, 321, 406, 980, 1118, 543, 932, 695, 1711]. Representation [313, 832, 809, 921, 595, 530, 1345, 227, 50, 130, 82, 696, 1281, 328, 395, 434, 942, 531]. representations [496, 64]. Representing [717, 1117, 1126, 166, 1087]. Request [1174]. Request-Response [1174]. require [618]. Requirements [1231, 217]. Research [771, 856, 1095, 1112, 764, 772, 178, 1427, 710, 826, 110, 129, 1580, 954, 1186, 300, 417, 551, 415, 476, 639, 538, 482, 951]. resemblance [251]. RESIDUE [1056]. Resolution [1129, 1647]. Resolving [1739]. Resource [1344, 1093]. Resource-Aware [1344]. Resource-Efficient [1093]. Resources [1373]. Response [1174, 1731]. responses [1711]. Responsibility [1460]. Restoration [577]. restricted [1263]. restyling [678]. Results [1602, 596, 1141, 1693, 601, 1035, 21, 1331]. Rethinking [1732]. Retrieval [1597, 426, 839, 1480, 1458]. Retrieving [1533]. Retrospective [976]. Reuse [1656, 1338, 1432]. reveal [216]. Review [333, 1085, 1169, 187, 1339, 382, 738, 611, 420]. Reviews [591]. Revised [220, 257, 102, 118, 256]. Revisited [351, 379]. rewrite [801]. rhombical [121]. Richardson [312]. ridge [637]. Right [228, 1403]. Right-Looking [1403]. Rights [1624]. Rivest [363]. Roadmaps [1537]. Robert [363]. Robot [1224, 1310, 1281, 1226, 1211, 1494, 1317, 1454, 1551, 1548, 831, 1550, 585, 1450, 1505, 1555, 418]. Robotics [1481]. Robots [1342, 1610, 1743, 1464]. Robust [1649, 1424, 1627]. Role [908, 853, 833, 1596]. roles [1349]. Rolling [1723]. Ronald [363]. Root [358, 55]. Roots [9]. Round [1430, 156]. round-off [156]. Round-Trip [1430]. Roundoff [164, 146]. Routing [1233, 1431, 1565, 1609, 1592, 440]. Rover [831]. Row [333]. RPC [1232]. Rule [1522, 945, 783, 910, 929, 498, 1428, 627, 962, 993, 475, 1199, 683, 398, 1115, 1025, 941]. Rule-Based [945, 783, 910, 929, 1428, 627, 962, 993, 475, 683, 1115, 941]. Rules [832, 612, 530, 153, 152, 1346, 1136, 1067, 442, 646, 81, 801, 562]. Rum [1080]. Run [1029]. Runtime [883, 1307, 829]. Russian [19]. RX [816, 921].

S [730, 811, 1077]. S-1 [730]. S-expressions [1077]. Saccade [1692]. SACON [714]. Safer [1386]. Safety [1436]. Safety-Progress [1436]. SAIL [316, 534, 588, 589, 384]. SAL [1697]. Sample [246, 1698, 787]. Sampling [1282, 1387, 1559, 1663, 262]. Sandwich [1516]. Satisfaction [1171, 997]. Satisfactory [74]. Satisfiability [838, 1328, 1019, 704]. Satisfy [913]. Saulev [3]. scalability [1290]. Scalable [1419, 1514, 1738, 1369]. SCALD [681, 680]. Scale [1359, 1650, 1591, 1703, 788, 260, 309]. Scales [907]. Scaling [1246, 1463, 1444, 1455, 1498, 1716, 1480, 327]. Scan [1134, 1669, 1711]. Scene [1734, 180, 391]. Scenes [329, 536, 1031, 1679]. Schedule [1363, 897]. Scheduling [1628, 848, 1041, 1098, 1138, 366, 1737, 165, 340, 341, 318, 368]. Schema [774]. Schema-Shift [774]. Schemas [349, 222, 255]. scheme [725]. Schemes [701, 440, 926]. Schmidt [122]. Schuffle [439, 438]. Schur [728]. Schwarz [291, 1193, 726]. Science [771, 417, 551, 106, 178, 112, 543, 1082, 554, 1571, 110, 129, 321, 406, 226, 39, 709, 802, 745, 639, 980, 1118, 26, 1525, 1209, 1364, 932, 204, 695, 886]. Scientific [632]. scientist [77]. Screw [827]. Scripting [1732]. scripts [983]. Scrolling [1700]. SDR [367]. Search [1602, 1713, 1195, 757, 15, 149, 584, 1065, 1066, 136, 402, 1561, 762, 1440, 145, 466, 1597]. Searching [722, 1006, 524, 403]. second [670, 16]. second-degree [16]. Sections [100]. Security [871, 1662]. Seeing [831]. Segmentation [1334, 1603, 1634, 1239, 396]. Selected [1744, 300, 1598, 21, 178]. selecting [752]. Selection [364, 365, 1301, 1566, 475, 363, 363]. Selective [939, 1478]. Selectivity [1414]. Self [1446, 1715, 1195, 1016, 1265, 1728, 431, 1538]. self-assessment [1728]. self-checking [431]. Self-Described [1016]. Self-Efficacy [1715]. Self-Organizing [1195]. Semantic [1222, 607, 874, 912, 238, 391, 683, 829, 518, 395, 307]. Semantically [1533]. Semantics [1223, 1274, 1229, 1318, 981, 167, 169, 1727, 895, 485, 1439, 1397, 1252, 1412, 130, 1347, 388, 1503, 458, 517, 594]. semaphores [207]. semi [45]. semi-definite [45]. Semiantichains [837]. Semiconductor [1543]. Semigroups [1352]. Semilinear [1219]. Seminar [1376, 1075, 1008, 880, 1007, 1092, 1181, 1294, 724, 621]. Sense [1602, 1097]. Sensing [1718, 1317]. Sensors [1664]. Sensory [1450, 1505, 407]. Sentence [370]. Sentences [448, 80]. Separability [911, 986]. Separable [386, 445]. separate [269, 103]. Separating [1259]. Separation [358]. Separator [849, 643, 642]. Sequence [1124, 50, 1448, 1377]. Sequences [1607, 488, 652]. Sequencing [94, 332, 938, 1495]. Sequential [894, 1483, 170, 789, 1071]. Sequentialization [1155]. Series [133, 1532, 697]. Server [1623]. Service [1602, 1153, 1196, 1297, 1389, 1104, 1630]. Serving [319]. Set [462, 696, 563]. Sets [1258, 1219, 282, 283, 338, 673, 1199, 277, 343, 743, 618]. Settling [1372]. Several [1139]. Shape [1637, 424, 1159, 1634]. Shapes [1280]. Share [1040]. Shared [1419, 1514, 1599, 1480, 1290, 1369]. Shared-Memory [1514, 1369]. Shared-Nothing [1480]. Sharing [1601, 1190, 1686, 1706]. Sharp [1564]. Shell [1319, 1165]. Shellsort [268, 741]. Shift [774, 87, 143, 56]. shifts [964]. Shoe [1661]. Short [1168, 881, 858]. Shortest [1497, 1455, 1189, 685, 1071]. Should [768]. Shoulder [1694]. Shoulder-surfing [1694]. Shuffle [1048, 1306, 158, 1071]. Shuffle-Exchange [1306, 1071]. shuffle-minimizing [1071]. Shutter [1723]. Side [691, 830]. SIDES [1673]. Signal [1156]. Signals [627]. signature [375, 396, 183]. Signatures [1693]. significance [751]. SIGPLAN [1452]. Silicon [1073, 1079, 1596]. Similarity [1341]. Simple [904, 548, 614, 1037, 803, 231, 1742, 231]. Simplex [114, 1273, 1400]. Simplification [667]. Simplifier [1362]. Simplify [1232]. Simulation [761, 361, 1510, 1176, 1207, 648, 1620, 1736, 1688, 1139, 1591, 82, 1682, 251, 334, 360, 568, 210, 185, 1331, 1627]. Simulator [1225]. Simultaneous [465]. Single [1413]. Singular [8, 73, 133, 603, 1264, 1263, 650, 782]. Singularity [1221]. Sirpent [1298]. situational [797]. situational-calculus [797]. Six [703]. Skew [1506]. Skew-Symmetric [1506]. Skills [1673]. Slicing [1697]. SLTF [340]. Small [123, 432, 163, 252]. Smalltalk [1324]. Smooth [1068]. Smoothing [1692]. Snapshots [1449]. Soar [1160, 1133]. SOARIOPSS [1144]. Social [1673, 1529, 1460, 1667, 1703]. Soft [1359, 1532, 1241, 1508, 1476]. Soft-Real-Time [1241]. Software [1128, 522, 1269, 678, 1721, 1722, 721, 512]. Software-Controlled [1128]. Solution [4, 850, 195, 544, 824, 1191, 164, 905, 143, 24, 324, 306, 322, 303, 410, 2, 362, 259, 507, 27, 1137, 689, 294, 286, 599, 450, 159, 6, 29]. Solutions [133, 301, 151, 67, 97, 31, 708, 12]. Solve [284, 374, 1355, 762, 40, 1340]. Solver [74, 71, 541]. Solvers [433]. Solving [720, 454, 1376, 931, 1034, 958, 1047, 1075, 1008, 1573, 221, 880, 1007, 323, 1092, 305, 1181, 1294, 682, 715, 539, 1586, 724, 621, 176, 12, 145, 331, 217, 653, 611, 749, 763]. Some [266, 1096, 516, 1141, 239, 310, 775, 1425, 299, 1035, 441, 879, 50, 1069, 1475, 841, 623, 428, 628, 1005, 478, 592, 252, 447, 632]. sometime [571, 572]. sometimes [571, 572]. Sooner [1386]. sort [242, 243]. Sorted [717, 1693, 768]. Sorting [1258, 1012, 177, 1285, 268, 218, 480, 242, 243, 244, 736, 779, 556, 403]. sound [1275]. Source [505, 1612]. Sources [1600]. Sowa [1085]. Space [920, 1299, 1451, 1554, 760, 558, 1708, 1675, 1511, 1039, 634, 480, 779]. Spaces [540, 1539, 1537, 718, 656]. Spacial [1171]. Spanning [540, 126, 466, 656]. Sparse [849, 610, 700, 410, 1308, 1403, 1462, 1483, 507, 515, 1311, 1331, 1343, 698]. Spatial [1734, 1656]. Spatiotemporal [1663]. Special [919, 923, 1072, 1070]. Specific [1469]. Specification [1660, 1381, 1406, 1011, 1249, 820, 187, 805]. Specifications [889, 622, 574, 944, 1108]. Specified [1307]. spectral [501, 617]. Spectrum [154]. Specularity [1159]. Speech [49, 127, 389, 315, 396, 183]. Speed [1134]. sphere [16, 144]. Spine [1603]. Spline [205, 413, 383, 815, 583, 234]. Splines [1068, 583]. Split [925]. Splitting [1193]. Spoken [953]. Spot [1139]. SQL [1587]. Square [1237]. Squares [1191, 1269, 73, 133, 410, 445, 97, 494, 712, 12, 134, 269, 573, 788]. Stability [323, 631, 1337, 926, 597, 670, 753, 1303]. Stabilization [1723]. Stable [1388, 1260, 1253, 1300, 480, 617]. stacks [218, 725]. Stage [87, 143]. STAGE2 [514]. Standard [1477]. standardized [1264]. Stanford [1177, 110, 129, 1441, 980, 1579, 417, 551, 415, 26, 178, 1597, 1557, 1666, 709, 746, 1536, 1598, 482, 932, 1518, 204, 695, 1596]. STARTS [1597]. State [1595, 577, 721, 564, 946]. Statements [167, 427, 148]. States [190, 198]. Stationary [142, 348, 312, 587]. Statistical [1520, 993, 124]. Statistically [1532]. Statistics [962]. Steerable [1621]. STeP [1579, 1633, 1536]. Steplength [694]. steps [206]. Stereo [979, 949, 1590, 536, 823, 449, 1561, 486, 532]. Stereotaxic [1465]. Stewart [782]. Sticky [1305]. Stirling [963]. Stokes [630]. stopping [55]. Storage [1196, 1297, 502, 325, 119]. Storing [698]. Strategic [1014, 1186]. Strategies [774, 931, 930, 402]. Strategy [1398, 1087, 242, 243]. Stratified [1535]. Strike [864]. Strikes [898]. String [1229, 1318, 935, 644]. String-Functional [1229, 1318]. Strings [1124, 451]. Stront [113]. Structural [1727, 714, 737]. Structurally [1358]. Structure [1332, 1236, 717, 227, 479, 1338, 471, 652]. Structured [774, 382, 427, 998, 1361, 1719, 661, 474, 450, 511, 897, 680]. Structured-Prediction [1719]. Structures [937, 1322, 545, 411, 191, 317, 701, 463, 565, 1478, 1511, 1681, 1137, 199, 693, 380, 531, 1085]. Structuring [1601, 1054, 412]. Student [1167, 1201, 930]. Studies [1280, 1690, 1686, 1706, 636, 942]. studio [1728]. Study [816, 469, 1689, 139, 221, 186, 1116, 1411, 50, 1272, 236, 1596, 632, 1390, 1667, 582, 226, 176, 234]. studying [952]. Subgoal [1408]. subgraph [495]. subgraphs [499]. Subject [459, 142]. subjects [789]. Sublinear [1230]. Sublinear-Time [1230]. Subproblems [332]. Subrecursive [623]. subroutine [79]. subroutines [175]. Subsequences [353, 488]. subset [245]. Subspaces [66, 230]. Substitution [1574]. subsystem [681]. Subtask [1508]. Subtractive [521]. Subtree [1563, 1382]. Successive [60]. Sufficient [7]. sum [156]. Summary [1557, 1518]. Summation [385]. Sums [491, 163]. Supercomputer [1005]. supernodal [1331]. Supervised [1604, 1564]. Supervisory [279]. Support [1651, 1390, 1665, 688]. Supporting [1705, 1674, 1680, 1731, 1699]. Suppositions [632]. Surf [1622, 1575]. Surfaces [1211, 1653, 1652]. surfing [1694]. Surgery [1688]. Survey [1170, 721, 524, 171, 801, 1127, 447]. Survival [1581]. Survive [1053].

Switched [1399]. Switches [1453]. Switching [1691, 641]. switchyard [218]. Symbolic [878, 1676, 1552]. Symmetric [1506, 278, 41, 507, 43, 42, 778]. Synchronization [1641, 513, 1630]. Synchronous [1229, 1318, 1042, 1574]. syntax [108]. Syntheses [381]. Synthesis [1584, 1056, 1208, 770, 645, 872, 889, 1433, 1316, 574, 1654, 944, 170, 654, 174, 705, 402, 946]. Synthetic [923]. System [877, 606, 1219, 1676, 1154, 1212, 1524, 910, 909, 929, 1176, 1567, 1103, 337, 98, 88, 232, 164, 1641, 1721, 1314, 1620, 1645, 1651, 1601, 1611, 998, 1361, 971, 1218, 1370, 967, 719, 1131, 1206, 1101, 1319, 1717, 1144, 1145, 259, 1165, 1040, 976, 836, 950, 407, 547, 683, 1013, 600, 604, 467, 92, 91, 231, 72, 487, 1014, 412, 1476, 690, 777, 985, 633, 867, 514, 1088, 961, 109, 611, 941, 1105, 68, 104, 1027, 1174, 971]. System/U [971]. Systems [1404, 720, 1622, 1522, 461, 920, 706, 707, 856, 966, 945, 1112, 1271, 1670, 1710, 607, 1381, 1517, 1042, 1378, 1154, 1170, 1605, 552, 535, 1588, 908, 1619, 1109, 853, 350, 138, 1487, 1532, 1241, 1245, 1513, 1409, 1406, 907, 161, 1351, 1190, 794, 367, 24, 977, 1499, 1268, 1435, 1198, 1488, 1583, 1302, 1214, 1467, 1143, 410, 2, 50, 421, 1016, 887, 279, 989, 1357, 833, 460, 628, 579, 423, 82, 1148, 1480, 1458, 1323, 1635, 1194, 1261, 1438, 1552, 1108, 1683, 1618, 67, 582, 1633, 172, 160, 1095, 619, 141, 1265, 69, 1608, 1115, 731, 455]. systems [1049, 1055, 45, 947, 331, 217, 368, 446, 608, 52, 951, 620, 942, 737, 447, 763].

Table [1006, 1674, 845, 698, 396]. Tableau [1028]. Tables [378, 375, 768]. Tabletop [1673, 1680]. TABLOG [1028, 1114]. Tailoring [1704]. Taliesin [1103]. Tape [293]. Tarjan [363]. Task [1117, 574, 402]. Task-level [574]. Tasks [1508, 165, 318, 1724]. Tau [690]. Teaching [1569, 931, 929, 853]. Teams [1667]. Technical [1571, 1525, 1209, 1364, 543, 690, 214]. Techniques [1404, 1602, 1310, 1274, 1045, 1578, 1509, 1700, 887, 1343, 302, 833, 404, 432, 1261, 162, 524, 1290, 805, 1570]. Technology [1665]. teletyped [334]. Template [1193, 933]. Templates [1738]. Temporal [1076, 1178, 1291, 1223, 1251, 1579, 1132, 1322, 854, 981, 1409, 1406, 1350, 889, 861, 894, 967, 982, 1204, 1321, 1488, 1536, 1568, 965, 987, 1070, 944, 855, 985, 1546]. Temporal-Logic [1178]. Ten [468, 415]. Terminal [559, 711, 668]. Terminate [429, 428]. Termination [1222, 527, 666, 533]. Terminology [1613]. Terms [19, 879, 504]. Test [1670, 1119, 1328, 251, 501, 1677, 1043]. Tester [1019]. Testing [822, 88, 1334, 1067, 206, 109]. Tests [1710]. Text [1734, 1643, 914, 1458, 983, 690, 514]. Texts [774]. Texture [1239, 1654]. Textured [329]. TEXware [1121]. th [752]. Their [274, 424, 1000, 317, 357, 664, 1049, 1055, 504, 964, 431]. Themes [636]. Theorem [1123, 1178, 1164, 454, 1589, 138, 351, 150, 294, 643, 642, 1516]. Theorems [3, 849, 338, 28, 343]. Theoretic [113, 1320, 1069, 1120]. Theoretical [1543, 592, 555]. Theories [1693, 295, 739, 1519]. Theory [75, 1344, 732, 1497, 1045, 1567, 1048, 1021, 1573, 1081, 1309, 189, 1161, 227, 827, 1639, 1391, 1032, 1016, 1501, 820, 430, 1284, 1621, 869, 226, 141, 518, 1338, 197, 280, 672, 379, 1080, 537, 702]. Theory-Driven [1045]. therapeutic [233]. Therapy [475]. There [1187]. ThingLab [761].

Thinkers [1210]. Thinking [1187]. Third [977, 1599]. Third-Party [1599]. Thought [361, 360, 273]. Thoughts [428, 1005]. Three [1018, 591, 881, 180, 87, 143, 1603, 962]. Three-Dimensional [1018, 180, 1603]. Three-Stage [87, 143]. Throughput [1609, 1033]. Throughput-Competitive [1609]. thru [100]. Tight [1353]. Tiled [1648]. till [77]. Time [1333, 1404, 1515, 720, 364, 363, 816, 921, 877, 344, 1588, 1449, 1060, 293, 365, 493, 550, 1456, 1230, 1532, 1241, 1354, 1409, 1406, 907, 1477, 1508, 925, 760, 1372, 217, 634, 1556, 236, 1255, 1001, 1530, 1438, 1552, 1503, 1633, 1517, 209, 1476, 318, 1572, 846, 618, 480, 779, 363]. Time-Invariant [877]. Time-Lapse [1449]. Time-Oriented [816, 921]. Time-Space [760, 634, 779]. Time-Split [925]. Timed [1487, 1435]. Timely [1129]. Times [1430]. tion [995]. Today [46]. Toeplitz [386, 763]. Toetjes [1227]. tokens [1733]. TokensRegex [1733]. Tolerance [1389]. toleranced [1585]. tolerancing [568]. Tolerant [1494, 1447, 1325, 736]. Tool [1185, 390, 62]. Toolkit [1699]. Tools [1602, 972, 1346, 1214, 505]. Top [1067, 1025]. Top-Down [1067, 1025]. Topics [1471, 287, 674]. Topological [1356]. Topologies [1173]. Topology [1565, 1592]. Torture [1119, 1043]. Total [261, 393]. Totally [1502]. Tournaments [1258]. TR [1571]. Tracking [1464, 1683]. trade [634]. trade-offs [634]. Tradeoff [1001, 779]. Tradeoffs [760]. Trajectories [1090, 1547]. Trajectory [319]. Transaction [1053, 1485]. Transactional [1732]. Transactions [1316, 813]. transcripts [334]. Transfer [783, 514]. transform [315, 1451, 146]. Transformation [1596, 726]. Transformational [89, 117, 90, 98, 88, 86, 108, 92]. Transformations [1344, 97, 900]. Transforming [1631]. transient [390]. Transition [1583, 57]. Transitive [1493, 219, 1352, 519]. Translation [188, 1614, 272]. Translations [311, 509, 1413]. Translator [795, 69, 729, 231]. Transmissions [1456, 776]. Transonic [51, 70]. Transport [905, 1430, 1627]. Transputers [1560]. Treatment [359]. Tree [1509, 1182, 227, 587, 525]. Tree-Based [1509]. Trees [540, 1282, 267, 508, 502, 149, 250, 828, 1448, 223, 126, 495, 466, 523, 183, 656]. Tri [41, 433, 42, 331]. Tri-Diagonal [41, 433, 42, 331]. Trial [1214, 1616]. Triangular [1355, 285]. Triangularization [1363]. Tridiagonal [259, 649, 43]. trigonometric [651]. Trip [1430]. Triply [192, 200]. Trivalent [416]. Trivial [613, 557]. Truncated [927]. Truncated-Newton [927]. trusses [11]. Tuples [1457]. Turing [246, 401, 660]. Tutorial [783, 1170, 589]. Tutoring [909, 1170, 1194]. Twenty [1168, 1441]. Twenty-Fourth [1441]. Two [720, 363, 288, 1236, 596, 293, 1363, 1098, 43, 1548, 846, 1568, 833, 60, 1495, 165, 294, 488, 630, 541, 245, 704, 18, 669]. two-dimensional [630]. Two-Handed [1495]. two-point [18]. Two-Processor [1363, 1098]. two-variable-per-constraint [704]. Two-View [1236]. txt [1690]. Type [1619, 1302, 440, 619, 983, 984]. Typed [607, 1479, 295, 933]. Typefaces [798, 1094]. types [622, 900]. Typesetting [1745, 887, 821]. Typography [663].

U [971]. U.R. [898]. U.S. [947]. UFORT [795]. UIO [1154]. Ultracomputer [1139]. Ultrasonic [1018]. Unassisted [1710]. Uncertainty [1004, 1216, 1317, 1551, 1496, 1450, 1505, 1199]. unconditional [670]. underlying [273]. Understand [1620, 960]. Understanding [953, 774, 1156, 1502, 769, 1091, 545, 627, 448, 193, 371, 625, 455, 487, 201, 447, 594, 1169]. unichain [837]. Unification [872]. Unified [1212, 1218, 882, 744]. Uniform [1154, 1601, 1059, 1268, 387]. Unifying [1070]. Union [743, 1585]. Union-member [743]. Uniqueness [5]. Unit [1555, 16, 318, 525]. Unit-Modular [1555]. unit-time [318]. units [325]. univariate [241]. Universal [971, 1104, 895, 347, 1410, 959, 795]. Universality [1305]. Universities [1186]. university [39, 26, 178, 110, 129, 1596]. unknown [13]. unscramble [399]. Unstructured [1464]. Untyped [1362]. Unweighted [1573]. Update [759, 1519, 1418]. Updates [1179, 1527, 1470]. Updating [787, 1266, 1061, 1172, 1120]. Upper [760, 271]. URAND [347]. Usage [1686, 1706, 1599]. Use [286, 578, 819, 153, 1314, 1594, 1625, 793, 1037, 1427, 317, 131, 1503, 407, 159, 57, 109, 243, 431, 252]. User [1729, 1579, 1236, 1648, 181, 1361, 733, 1675, 1023, 202, 746, 906, 918, 384, 575]. User-Definable [1236]. User-Generated [1729]. users [602]. Uses [424, 835, 310, 1000]. Using [1515, 1512, 1156, 1676, 1650, 1531, 1721, 935, 1532, 854, 1620, 1578, 1451, 1694, 439, 1328, 875, 1632, 1634, 1523, 1586, 986, 1194, 762, 391, 1711, 1633, 683, 326, 1707, 1655, 334, 141, 1653, 176, 1420, 583, 6, 1656, 1723, 1652, 1564, 260, 218, 375, 396, 1704]. Utilization [1718, 78]. Utterances [913, 430].

V [1339, 164, 562, 991, 1027, 1174, 1212, 1105]. V-System [1027, 1105]. V. [3]. v0.2 [1660]. Vague [1534]. validation [251, 637, 1338]. Valley [1073, 1596]. Value [4, 635, 687, 824, 14, 73, 133, 1599, 579, 1337, 603, 1264, 1263, 735, 1303, 492, 592, 18, 782]. Value-Added [1599]. Values [8, 73, 142, 1330, 151, 1703, 650]. Variability [1725]. Variable [498, 87, 143, 704]. Variable-Shift [87, 143]. Variables [720, 269]. variances [787]. variants [122]. Variational [234]. Variations [676]. Varieties [1416]. Varying [62]. Vaughan [363]. VAX [896]. Vector [1219, 1511]. Vectorcardiographic [21]. Vectors [1695, 650, 399]. Vehicle [823, 236]. Vel’ski [1339]. verbs [273]. Verification [1404, 1562, 606, 1229, 1318, 607, 1517, 1588, 1406, 1578, 1595, 1611, 354, 859, 855, 861, 894, 936, 985, 1249, 1396, 820, 565, 1635, 647, 1574, 1541, 1631, 1618, 1633, 376, 533, 562, 483, 484, 547]. verification-oriented [562]. Verifier [1346, 746]. Verifying [883, 854, 1552, 941, 483, 484]. Version [1600, 1299, 718, 1379, 100]. Version-Space [1299]. versus [1370, 64]. Vertex [392, 542]. Vertex-Graphs [392]. Vertical [974]. via [860, 1370]. Video [559, 1723]. View [1336, 1236, 869]. Viewing [1707, 1166, 1088]. Viewpoint [1210]. Views [536, 1594, 1061, 1371, 1358]. Viron [964]. Virtual [1729, 890, 1709, 1716, 1679]. Visibility [1189]. Visible [1146].

Vision [473, 1062, 606, 823, 377, 1314, 957, 182, 1464, 532, 547]. Visual [329, 320, 168, 1210, 1036, 474, 420, 1678, 397]. Visualization [1603, 1523, 1591]. Visualizing [1720]. Visuohaptic [1688]. VLSI [978, 1358, 946, 842]. VMP [1128, 1217, 1290]. VMP-MC [1290]. VMTP [1174]. Volkov [151]. Volume [1459]. Volumes [1634]. volumetric [1653, 1652]. Voronoi [722, 1326]. vs [1377]. Vulnerability [436].

W [363, 99, 111, 102, 118, 258]. Wafer [1359]. WAITS [858]. Wall [1646]. Wall-Mounted [1646]. Ward [975]. Warehouses [1640, 1612]. Warped [1657]. Waste [1718]. WATCH [1088]. Wave [49, 926, 315]. waveforms [390]. Way [982, 434]. weak [685, 533, 31]. weakly [264]. Web [1622, 1712, 1713, 1698, 1719, 1599, 998, 1727, 1704]. weight [1267, 791]. Weighted [937, 1074, 1573, 1188]. Weizenbaum [591]. Well [897, 1252, 936, 1335]. well-behaved [1335]. well-founded [936]. Wendroff [619]. whether [1521]. which [153]. While [188]. Whiteboard [1687]. Whole [885]. Whose [613, 269]. Wide [1203, 848]. Wide-Area [1203]. Width [1695]. Wielandt [150]. WINDOW [1205]. Winged [328]. Winograd [1169]. Winter [1384, 1127]. wisdom [1728]. Within [606, 547, 1248, 1262, 533, 447]. Without [899, 1561, 203, 776]. Word [545, 78, 995]. Work [966, 1629, 1642, 1184, 1390]. Workbench [1647]. Working [1095, 1112, 462, 43]. Works [1744]. Workshop [1452]. Workspace [1656]. WorkspaceNavigator [1656]. Workspaces [1644, 1651, 1666, 1662]. Workstations [991, 1026, 1311, 1343]. World [831, 1108]. Worlds [1729, 1709, 1716]. Wrappers [1577]. Write [1196, 1297, 1050, 567]. Write-Once [1196, 1297, 1050]. Writing [1712, 88, 1213, 69, 231].

X [604]. X-ray [604]. XII [400]. XP [1667].

YAWN [1205]. years [415]. Yield [1359, 25]. Yield-point [25]. York [333].

Zeros [203, 87, 143, 36].

References

Rosen:1963:PPP

[1] J. B. Rosen. Primal partition programming for block diagonal matrices. Technical Report STAN-CS-63-1 (AD462108), Stanford University, Department of Computer Science, Stanford, CA, USA, November

      1963. 23 pp. URL https://apps.dtic.mil/sti/html/tr/AD0426108/
      index.html.

Pavkovich:1963:SLS

[2] John M. Pavkovich. The solution of large systems of algebraic equations. Technical Report CS-TR-63-2 (AD427753), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1963. 46 pp.

URL http://i.stanford.edu/TR/CS-TR-63-2.html.

Forsythe:1964:TPV

[3] G. E. Forsythe. The theorems in a paper by V. K. Saulev, On an Estimate of the Error in Obtaining Characteristic Functions by the Method of Finite Differences. Technical Report STAN-CS-64-3 (AD430445), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1964. 6 pp. Translation from Russian to English by GEF.

Bergman:1964:NSB

[4] Stefan Bergman and J. G. Herriot. Numerical solution of boundary value problems by the method of integral operators. Technical Report STAN-CS-64-4 (AD434858), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1964. 24 pp.

Rosen:1964:EUE

[5] J. B. Rosen. Existence and uniqueness of equilibrium points for concave N-person games. Technical Report STAN-CS-64-5 (N-6519765), Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1964. 28 pp. URL https://ntrs.nasa.gov/api/citations/
19650010164/downloads/19650010164.pdf.

Hockney:1964:FDS

[6] Roger W. Hockney. A fast direct solution of Poisson’s equation using Fourier analysis. Technical Report STAN-CS-64-6 (AD600164), Stanford University, Department of Computer Science, Stanford, CA, USA, April

14, 1964. ii + 39 pp. URL http://i.stanford.edu/TR/CS-TR-64-6.
html.

Rosen:1964:SCO

[7] J. B. Rosen. Sufficient conditions for optimal control of convex processes. Technical Report STAN-CS-64-7 (PE3176753), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1964. 29 pp.

Golub:1964:CSV

[8] Gene H. Golub and William Kahan. Calculating the singular values and pseudo-inverse of a matrix. Technical Report STAN-CS-64-8 (AD603116), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1964. 33 pp.

Andersen:1964:AMF

[9] Christian Andersen. The QD-algorithm as a method for finding the roots of a polynomial equation when all roots are positive. Technical Report STAN-CS-64-9 (AD604012), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1964. 74 pp. URL http:/

/i.stanford.edu/TR/CS-TR-64-9.html.

Causey:1964:CNM

[10] R. L. Causey. On closest normal matrices. Technical Report STAN-CS-64-10 (AD603163), Stanford University, Department of Computer Science, Stanford, CA, USA, June 30, 1964. v + 131 pp.

Nakamura:1964:EPA

[11] Tsuneyoshi Nakamura and Judah Ben Rosen. Elastic-plastic analysis of trusses by the gradient projection method. Technical Report CS-TR-64-11 (PB176754), Stanford University, Department of Computer Science, Stanford, CA, USA, July 17, 1964. ii + 32 pp. URL http://i.stanford.

edu/TR/CS-TR-64-11.html.

Golub:1964:NMS

[12] Gene H. Golub and Peter A. Businger. Numerical methods for solving linear least squares problems (by G. Golub); An ALGOL procedure for finding linear least squares solutions (by Peter Businger). Technical Report STAN-CS-64-12 (AD608292), Stanford University, Department of Computer Science, Stanford, CA, USA, August 28, 1964. ii + 27 pp.

URL http://i.stanford.edu/TR/CS-TR-64-12.html.

Pereyra:1964:CPM

[13] Victor Pereyra and Judah Ben Rosen. Computation of the pseudoinverse of a matrix of unknown rank. Technical Report CS-TR-64-13 (N6527058), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1964. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-64-13.html.

Efimenko:1964:ACE

[14] V. A. Efimenko. On approximate calculations of the eigenvalues and eigenfunctions of boundary value problems in partial differential equations. Technical Report STAN-CS-64-14 (TT-65-61724), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1964. 20 pp. Translated by G. Reiter and C. Moler.

Grace:1965:CSN

[15] Donald W. Grace. Computer search for non-isomorphic convex polyhedra. Technical Report STAN-CS-65-15 (AD611366), Stanford University, Department of Computer Science, Stanford, CA, USA, January 29, 1965. viii + 137 pp.

Forsythe:1965:MSD

[16] George E. Forsythe and Gene H. Golub. Maximizing a second-degree polynomial on the unit sphere. Technical Report STAN-CS-65-16 (AD611427), Stanford University, Department of Computer Science, Stanford, CA, USA, February 5, 1965. iv + 31 pp. URL http://i.

stanford.edu/TR/CS-TR-65-16.html.

Forsythe:1965:AGP

[17] George E. Forsythe and Niklaus Wirth. Automatic grading programs. Technical Report CS-TR-65-17, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1965. ii + 17 pp.

URL http://i.stanford.edu/pub/cstr/reports/cs/tr/65/17/CS-
TR-65-17.pdf;   http://i.stanford.edu/TR/CS-TR-65-17.html.

Pereyra:1965:DCM

[18] Victor Pereyra. The difference correction method for non-linear twopoint boundary value problems. Technical Report CS-TR-65-18, Stanford University, Department of Computer Science, Stanford, CA, USA,

February 1965. ?? pp. URL http://i.stanford.edu/TR/CS-TR-65-
18.html.

Ageev:1965:EEM

[19] M. I. Ageev and J. Maclaren. English equivalents of metalinguistic terms of Russian ALGOL. Technical Report STAN-CS-65-19 (AD0612686, TT-65-61839), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1965. 115 pp. URL https://apps.dtic.mil/

sti/citations/tr/AD0612686.

Wirth:1965:EGA

[20] Niklaus Wirth and Helmut Weber. EULER: a generalization of ALGOL, and its formal definition. Technical Report CS-TR-65-20, Stanford University, Department of Computer Science, Stanford, CA, USA, April 27,

1965. xv + 115 pp. URL http://i.stanford.edu/pub/cstr/reports/
cs/tr/65/20/CS-TR-65-20.pdf;   http://i.stanford.edu/TR/CS-
TR-65-20.html.

Fisher:1965:VAD

[21] Donald D. Fisher, Jobst von der Groeben, and J. Gerald Toole. Vectorcardiographic analysis by digital computer, selected results. Technical Report CS-TR-65-21, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1965. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-65-21.html.

Moler:1965:FDM

[22] C. B. Moler. Finite difference methods for the eigenvalues of Laplace’s operator. Technical Report STAN-CS-65-22 (AD0616675), Stanford University, Department of Computer Science, Stanford, CA, USA, May

27, 1965. 142 pp. URL https://apps.dtic.mil/sti/citations/
AD0616675.

Rudin:1965:CPA

[23] Bernard D. Rudin. Convex polynomial approximation. Technical Report CS-TR-65-23, Stanford University, Department of Computer Science, Stanford, CA, USA, June 4, 1965. iv + 44

pp. URL http://i.stanford.edu/TR/CS-TR-65-23.html;   https:/
/dl.acm.org/doi/10.5555/891684;   https://www.proquest.com/
pqdtglobal/results/C5640F46CA624452PQ/.

Klyuyev:1965:MNA

[24] V. V. Klyuyev and N. I. Kokovkin Shoherbak. On the minimization of the number of arithmetic operations for the solution of linear algebraic systems of equations. Technical Report STAN-CS-65-24 (AD616611), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1965. 24 pp. Translated by G. J. Tee.

Hodge:1965:YPL

[25] Philip G. Hodge, Jr. Yield-point load determination by nonlinear programming. Technical Report CS-TR-65-25, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1965. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-65-25.html.

Forsythe:1965:SUP

[26] George E. Forsythe. Stanford University’s Program in Computer Science. Technical Report CS-TR-65-26, Stanford University, Department of Computer Science, Stanford, CA, USA, June 25, 1965. i + 27 pp.

URL http://i.stanford.edu/TR/CS-TR-65-26.html.

Volkov:1965:AOA

[27] E. A. Volkov. An analysis of one algorithm of heightened precision of the method of nets for the solution of Poisson’s equation. Technical Report STAN-CS-65-27 (AD618216), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1965. 29 pp. Translated by R. Bartels.

Miller:1965:MTP

[28] John J. H. Miller and Gilbert Strang. Matrix theorems for partial differential and difference equations. Technical Report CS-TR-65-28, Stanford University, Department of Computer Science, Stanford, CA, USA, July

1965. ?? pp. URL http://i.stanford.edu/TR/CS-TR-65-28.html.

Pereyra:1965:IAS

[29] Victor Pereyra. On improving an approximate solution of a functional equation by deferred corrections. Technical Report CS-TR-65-29, Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1965. ?? pp. URL http://i.stanford.edu/TR/CS-TR-65-29.
html.

Marchuk:1965:ACC

[30] S. Marchuk. The automatic construction of computational algorithms. Technical Report STAN-CS-65-30 (SS624-829), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1965. 56 pp. Translated by G. J. Tee.

Raviart:1965:AWS

[31] Pierre Arnaud Raviart. On the approximation of weak solutions of linear parabolic equations by a class of multistep difference methods. Technical Report CS-TR-65-31, Stanford University, Department of Computer Science, Stanford, CA, USA, December 6, 1965. ?? pp. URL http://

i.stanford.edu/TR/CS-TR-65-31.html.

Hockney:1965:MMF

[32] Roger W. Hockney. Minimum multiplication Fourier analysis. Technical Report CS-TR-65-32, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1965. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-65-32.html.

Wirth:1965:PLC

[33] Niklaus Wirth. A programming language for the 360 computers. Technical Report CS-TR-65-33, Stanford University, Department of Computer Science, Stanford, CA, USA, December 24, 1965. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-65-33.html;   https://
bitsavers.org/pdf/stanford/cs_techReports/CS33_Wirth_PL360_
Dec65.pdf.

Varah:1966:ERM

[34] James M. Varah. Eigenvectors of a real matrix by inverse iteration. Technical Report CS-TR-66-34, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1966. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-66-34.html.

Wirth:1966:CDA

[35] N. Wirth and C. Hoare. A contribution to the development of ALGOL. Technical Report STAN-CS-66-35 (PB176758), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1966. 64 pp.

Traub:1966:CZP

[36] J. F. Traub. The calculation of zeros of polynomials and analytic functions. Technical Report STAN-CS-66-36 (PB176759), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1966. iii + 26 pp.

Reynolds:1966:COM

[37] John C. Reynolds. COGENT 1.2 operations manual. Technical Report CS-TR-66-37, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1966. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-66-37.html.

McCarthy:1966:CCA

[38] John McCarthy and James Painter. Correctness of a compiler for arithmetic expressions. Technical Report STAN-CS-66-38 (AIM-40, AD662880), Stanford University, Department of Computer Science, Stanford, CA, USA, April 29, 1966. ii + 13 pp.

Forsythe:1966:UEP

[39] George E. Forsythe. A university’s educational program in computer science. Technical Report CS-TR-66-39, Stanford University, Department of Computer Science, Stanford, CA, USA, May 18, 1966. i + 26 pp. URL

http://i.stanford.edu/TR/CS-TR-66-39.html.

Forsythe:1966:HDY

[40] George E. Forsythe. How do you solve a quadratic equation? Technical Report CS-TR-66-40, Stanford University, Department of Computer Science, Stanford, CA, USA, June 16, 1966. i + 19 pp. URL http://i.

stanford.edu/TR/CS-TR-66-40.html.

Kahan:1966:AES

[41] William M. Kahan. Accurate eigenvalues of a symmetric tri-diagonal matrix. Technical Report CS-TR-66-41, Stanford University, Department of Computer Science, Stanford, CA, USA, July 22, 1966. iv + 53 pp. URL

http://i.stanford.edu/TR/CS-TR-66-41.html. Revised June 1968.

Kahan:1966:WND

[42] William Kahan. When to neglect off-diagonal elements of symmetric tri-diagonal matrices. Technical Report CS-TR-66-42, Stanford University, Department of Computer Science, Stanford, CA, USA, July

25, 1966. i + 10 pp. URL ftp://reports.stanford.edu/pub/cstr/
reports/cs/tr/66/42/CS-TR-66-42.pdf; http://i.stanford.edu/
TR/CS-TR-66-42.html.

Kahan:1966:TWA

[43] William Kahan and James M. Varah. Two working algorithms for the eigenvalues of a symmetric tridiagonal matrix. Technical Report CS-TR-66-43, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1, 1966. iv + 28 pp. URL http://i.stanford.edu/

TR/CS-TR-66-43.html.

Kahan:1966:RME

[44] William Kahan. Relaxation methods for an eigenproblem. Technical Report CS-TR-66-44, Stanford University, Department of Computer Science, Stanford, CA, USA, August 8, 1966. iv + 35 pp. URL http://i.

stanford.edu/TR/CS-TR-66-44.html.

Kahan:1966:RMS

[45] William Kahan. Relaxation methods for semi-definite systems. Technical Report CS-TR-66-45, Stanford University, Department of Computer Science, Stanford, CA, USA, August 9, 1966. iv + 32 pp. URL http:/

/i.stanford.edu/TR/CS-TR-66-45.html.

Forsythe:1966:TCM

[46] G. E. Forsythe. Today’s computational methods of linear algebra. Technical Report STAN-CS-66-46 (SS638-809), Stanford University, Department of Computer Science, Stanford, CA, USA, August 11, 1966. 47 pp.

URL https://apps.dtic.mil/sti/tr/pdf/AD0638809.pdf.

Abrams:1966:IIN

[47] Philip S. Abrams. An interpreter for ‘Iverson notation’. Technical Report CS-TR-66-47, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1966. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-66-47.html.

McKeeman:1966:ACL

[48] W. M. McKeeman. An approach to computer language design. Technical Report STAN-CS-66-48 (AD0639166, SS639-166), Stanford University, Department of Computer Science, Stanford, CA, USA, August

31, 1966. iv + 124 pp. URL https://apps.dtic.mil/sti/citations/
AD0639166.

Reddy:1968:ACS

[49] D. R. Reddy. An approach to computer speech recognition by direct analysis of speech wave. Technical Report STAN-CS-66-49 (AIM-43, SS640-836), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1968. 143 pp.

Persson:1966:SSE

[50] Staffan Persson. Some sequence extrapolating programs: a study of representation and modelling in inquiring systems. Technical Report STAN-CS-66-50 (AIM-46, PB176761), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1966. 176 pp.

Bergman:1966:NCT

[51] Stefan Bergman, John G. Herriot, and Thomas G. Kurtz. Numerical calculation of transonic flow patterns. Technical Report STAN-CS-66-51 (AD648394), Stanford University, Department of Computer Science, Stanford, CA, USA, October 14, 1966. i + 35 pp.

Shaw:1966:LNC

[52] Alan C. Shaw. Lecture notes on a course in systems programming. Technical Report CS-TR-66-52, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1966. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-66-52.html.

Wirth:1966:PLC

[53] Niklaus Wirth. A programming language for the 360 computers. Technical Report CS-TR-66-53, Stanford University, Department of Computer Science, Stanford, CA, USA, December 20, 1966. ii + 81 pp.

URL http://i.stanford.edu/pub/cstr/reports/cs/tr/66/53/CS-
TR-66-53.pdf;   http://i.stanford.edu/TR/CS-TR-66-53.html.

Golub:1967:GBAa

[54] Gene H. Golub and Thomas N. Robertson. A generalized Bairstow algorithm. Technical Report STAN-CS-67-54 (AD662882), Stanford University, Department of Computer Science, Stanford, CA, USA, January 13,

1967. i + 10 pp. URL http://i.stanford.edu/TR/CS-TR-67-54.html.

Adams:1967:SCP

[55] Duane A. Adams. A stopping criterion for polynomial root finding. Technical Report CS-TR-67-55, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1967. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-67-55.html.

Bauer:1967:MNS

[56] Friedrich L. Bauer. QD-method with Newton shift. Technical Report CS-TR-67-56, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1967. ii + 6 pp. URL http://i.stanford.edu/

TR/CS-TR-67-56.html.

Gries:1967:UTM

[57] David Gries. The use of transition matrices in compiling. Technical Report CS-TR-67-57, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1967. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-67-57.html.

Tixier:1967:RFR

[58] Vincent Emile-Louis Tixier. Recursive functions of regular expressions in language analysis. Technical Report STAN-CS-67-58 (PB176766), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1967. 146 pp.

Wilkinson:1967:ADM

[59] James H. Wilkinson. Almost diagonal matrices with multiple or close eigenvalues. Technical Report CS-TR-67-59, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1967. ??

pp. URL http://i.stanford.edu/TR/CS-TR-67-59.html. Published

as [1740].

Wilkinson:1967:TAB

[60] James H. Wilkinson. Two algorithms based on successive linear interpolation. Technical Report CS-TR-67-60, Stanford University, Department of Computer Science, Stanford, CA, USA, April 10, 1967. i + 15 pp.

URL http://i.stanford.edu/TR/CS-TR-67-60.html.

Forsythe:1967:ADD

[61] George E. Forsythe. On the asymptotic directions of the s-dimensional optimum gradient method. Technical Report CS-TR-67-61, Stanford University, Department of Computer Science, Stanford, CA, USA, April

13, 1967. ii + 43 pp. URL http://i.stanford.edu/TR/CS-TR-67-61.
html.

Tienari:1967:VLF

[62] Martti Tienari. Varying length floating point arithmetic: a necessary tool for the numerical analyst. Technical Report CS-TR-67-62, Stanford University, Department of Computer Science, Stanford, CA, USA, April

17, 1967. 38 pp. URL http://i.stanford.edu/TR/CS-TR-67-62.html.

Polya:1967:GME

[63] George Polya. Graeffe’s method for eigenvalues. Technical Report CS-TR-67-63, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1967. ?? pp. URL http://i.stanford.edu/TR/

CS-TR-67-63.html.

Richman:1967:FPN

[64] Paul L. Richman. Floating-point number representations: base choice versus exponent range. Technical Report CS-TR-67-64, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1967.

?? pp. URL http://i.stanford.edu/TR/CS-TR-67-64.html.

Wirth:1967:CBC

[65] Niklaus Wirth. On certain basic concepts of programming languages. Technical Report CS-TR-67-65, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1, 1967. ii + 30 pp.

URL http://i.stanford.edu/pub/cstr/reports/cs/tr/67/65/CS-
TR-67-65.pdf;   http://i.stanford.edu/TR/CS-TR-67-65.html.

Varah:1967:CBI

[66] J. M. Varah. The computation of bounds for the invariant subspaces of a general matrix operator. Technical Report STAN-CS-67-66 (AD652921), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1967. 240 pp.

Bartels:1967:CCR

[67] Richard H. Bartels and Gene H. Golub. Computational considerations regarding the calculation of Chebyshev solutions for overdetermined linear equation systems by the exchange method. Technical Report STAN-CS-67-67 (AD652992), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1967. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-67-67.html.

Wirth:1967:PS

[68] Niklaus Wirth. The PL360 system. Technical Report CS-TR-67-68, Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1967. ?? pp. URL http://i.stanford.edu/TR/CS-TR-67-
68.html.

Feldman:1967:TWS

[69] Jerome A. Feldman and David Gries. Translator writing systems. Technical Report CS-TR-67-69, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1967. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-67-69.html.

Bergman:1967:CFP

[70] Stefan Bergman, John G. Herriot, and Paul L. Richman. On computation of flow patterns of compressible fluids in the transonic region. Technical Report CS-TR-67-70, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1967. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-67-70.html.

Jenkins:1967:AAG

[71] M. A. Jenkins and J. F. Traub. An algorithm for an automatic general polynomial solver. Technical Report STAN-CS-67-71 (AD655230), Stanford University, Department of Computer Science, Stanford, CA, USA, July 21, 1967. iii + 34 pp.

Golub:1967:CAC

[72] Gene H. Golub and Lyle B. Smith. Chebyshev approximation of continuous functions by a Chebyshev system of functions. Technical Report STAN-CS-67-72 (PB175581), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1967. 54 pp. URL http://i.

stanford.edu/TR/CS-TR-67-72.html.

Golub:1967:LSS

[73] Gene H. Golub and Peter Businger. Least squares, singular values and matrix approximations (and an ALGOL procedure for computing the singular value decomposition). Technical Report STAN-CS-67-73 (AD662883), Stanford University, Department of Computer Science, Stanford, CA, USA, July 31, 1967. i + 22 pp. URL https://apps.

dtic.mil/sti/html/tr/AD0662883.

Forsythe:1967:WSQ

[74] George E. Forsythe. What is a satisfactory quadratic equation solver? Technical Report CS-TR-67-74 (AD657639), Stanford University, Department of Computer Science, Stanford, CA, USA, August 7, 1967. ii

+ 9 pp. URL https://apps.dtic.mil/sti/pdfs/AD0657639.pdf.

Bauer:1967:TN

[75] Friedrich L. Bauer. Theory of norms. Technical Report CS-TR-67-75, Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1967. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
67-75.html.

Anselone:1967:CCO

[76] Phillip M. Anselone. Collectively compact operator approximations. Technical Report CS-TR-67-76, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1967. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-67-76.html.

Forsythe:1967:WDT

[77] George E. Forsythe. What to do till the computer scientist comes. Technical Report CS-TR-67-77, Stanford University, Department of Computer Science, Stanford, CA, USA, September 22, 1967. ii + 13 pp. URL

http://i.stanford.edu/TR/CS-TR-67-77.html.

Colby:1967:MUN

[78] Kenneth Mark Colby and Horace J. Enea. Machine utilization of the natural language word ’good’. Technical Report CS-TR-67-78, Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1967. ?? pp. URL http://i.stanford.edu/TR/CS-TR-67-
78.html.

Doran:1967:FIF

[79] Robert W. Doran. 360 O.S. FORTRAN IV free field input/output subroutine package. Technical Report CS-TR-67-79, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1967. ??

pp. URL http://i.stanford.edu/TR/CS-TR-67-79.html.

Friedman:1967:DRG

[80] Joyce Friedman. Directed random generation of sentences. Technical Report CS-TR-67-80, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1967. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-67-80.html.

Golub:1967:CGQ

[81] Gene H. Golub and John H. Welsch. Calculation of Gauss quadrature rules. Technical Report STAN-CS-67-81 (AD661217), Stanford University, Department of Computer Science, Stanford, CA, USA, November

1967. 28 pp. URL http://i.stanford.edu/TR/CS-TR-67-81.html.

Tesler:1967:DGR

[82] Lawrence Tesler, Horace Enea, and Kenneth Mark Colby. A directed graph representation for computer simulation of belief systems. Technical Report STAN-CS-67-82 (PB176775), Stanford University, Department of Computer Science, Stanford, CA, USA, December 29, 1967. iv + 28 pp.

URL https://en.wikipedia.org/wiki/Larry_Tesler.

Colby:1968:CAL

[83] Kenneth Mark Colby. Computer-aided language development in nonspeaking mentally disturbed children. Technical Report STAN-CS-68-85 (PB177426), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1968. 35 pp. URL http://i.stanford.

edu/TR/CS-TR-68-85.html.

Bredt:1968:CMI

[84] Thomas H. Bredt. A computer model of information processing in children. Technical Report STAN-CS-68-100 (PB178877), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1968.

60 pp. URL http://i.stanford.edu/TR/CS-TR-68-100.html.

Pnueli:1968:IPC

[85] Amir Pnueli. Integer programming over a cone. Technical Report STAN-CS-68-102 (AD677982), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1968. 29 pp. URL http://i.

stanford.edu/TR/CS-TR-68-102.html.

Friedman:1968:LIT

[86] Joyce Friedman and Thomas H. Bredt. Lexical insertion in transformational grammar. Technical Report STAN-CS-68-103 (AD692689), Stanford University, Department of Computer Science, Stanford, CA, USA,

June 1968. vi + 47 pp. URL http://i.stanford.edu/TR/CS-TR-68-
103.html.

Jenkins:1968:TSV

[87] M. A. Jenkins and J. F. Traub. A three-stage variable-shift iteration for polynomial zeros and its relation to generalized Rayleigh iteration. Technical Report STAN-CS-68-107 (AD668558, AD0673673), Stanford University, Department of Computer Science, Stanford, CA, USA, August

26, 1968. viii + 40 + 2 pp. URL http://i.stanford.edu/TR/CS-TR-
68-107.html;   https://apps.dtic.mil/sti/html/tr/AD0673673/
index.html.

Friedman:1968:CSW

[88] Joyce Friedman. A computer system for writing and testing transformational grammars — final report. Technical Report STAN-CS-68-109 (AD692690), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1968. 14 pp. URL http://i.stanford.

edu/TR/CS-TR-68-109.html.

Friedman:1968:ATG

[89] Joyce Friedman and Theodore S. Martner. Analysis in transformational grammar. Technical Report STAN-CS-68-111 (AD692691), Stanford University, Department of Computer Science, Stanford, CA, USA, Au-

gust 1968. 18 pp. URL http://i.stanford.edu/TR/CS-TR-68-111.
html.

Friedman:1968:CLT

[90] Joyce Friedman and Bary W. Pollack. A control language for transformational grammar. Technical Report STAN-CS-68-112 (AD692687), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1968. 51 pp. URL http://i.stanford.edu/TR/CS-TR-
68-112.html.

George:1968:CIP

[91] James E. George. Calgen — an interactive picture calculus generation system. Technical Report CS-TR-68-114, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1968. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-68-114.html.

Friedman:1968:PMC

[92] Joyce Friedman, Thomas H. Bredt, Robert W. Doran, Theodore S. Martner, and Bary W. Pollack. Programmers manual for a computer system for transformational grammar. Technical Report CS-TR-68-115, Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1968. ?? pp. URL http://i.stanford.edu/TR/CS-TR-68-115.
html.

Pieper:1968:KMU

[93] Donald Lee Pieper. The kinematics of manipulators under computer control. Technical Report STAN-CS-68-116 (AIM-72, AD680036), Stanford University, Department of Computer Science, Stanford, CA, USA, October 24, 1968. 167 pp.

Adams:1968:CMD

[94] D. Adams. A computational model with data flow sequencing. Technical Report STAN-CS-68-117 (PB182151), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1968. 130 pp.

Waterman:1968:MLH

[95] Donald Arthur Waterman. Machine learning of heuristics. Technical Report STAN-CS-68-118 (AIM-74, AD681027), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1968. 235 pp.

Bayer:1968:MMP

[96] Rudolf Bayer, James H. Bigelow, George B. Dantzig, David J. Gries, Michael B. McGrath, Paul D. Pinsky, Stephen K. Schuck, and Christoph Witzgall. MPL: Mathematical Programming Language. Technical Report CS-TR-68-119, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1968. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-68-119.html.

Bjorck:1968:IRL

[97] Åke Björck and Gene H. Golub. Iterative refinements of linear least squares solutions by Householder transformations. Technical Report STAN-CS-68-83 (AD664237), Stanford University, Department of Computer Science, Stanford, CA, USA, January 19, 1968. i + 28 pp. URL

http://i.stanford.edu/TR/CS-TR-68-83.html.

Friedman:1968:CST

[98] Joyce Friedman. A computer system for transformational grammar. Technical Report CS-TR-68-84 (AD692680), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1968. 31

pp. URL http://i.stanford.edu/TR/CS-TR-68-84.html.

Bauer:1968:AW

[99] Henry R. Bauer, Sheldon I. Becker, and Susan L. Graham. ALGOL W. Technical Report CS-TR-68-86, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1968. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-68-86.html.

Ehrman:1968:CLN

[100] John R. Ehrman. CS139 lecture notes. Part I: Sections 1 thru 21. Preliminary version. Technical Report CS-TR-68-87, Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1968. ??

pp. URL http://i.stanford.edu/TR/CS-TR-68-87.html.

Schechter:1968:RMC

[101] Samuel Schechter. Relaxation methods for convex problems. Technical Report CS-TR-68-88, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1968. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-68-88.html.

Bauer:1968:AWR

[102] Henry R. Bauer, Sheldon I. Becker, Susan L. Graham, George E. Forsythe, and Edwin H. Satterthwaite. ALGOL W (revised). Technical Report CS-TR-68-89, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1968. i + 124 pp. URL http://i.

stanford.edu/TR/CS-TR-68-89.html.

Lesser:1968:MLC

[103] Victor R. Lesser. A multi-level computer organization designed to separate data-accessing from the computation. Technical Report CS-TR-68-90, Stanford University, Department of Computer Science, Stanford,

CA, USA, March 1968. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-68-90.html.

Wirth:1968:PS

[104] Niklaus Wirth, Joseph W. Wells, Jr., and Edwin H. Satterthwaite, Jr. The PL360 system. Technical Report CS-TR-68-91, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1, 1968.

viii + 89 pp. URL http://i.stanford.edu/pub/cstr/reports/cs/
tr/68/91/CS-TR-68-91.pdf;   http://i.stanford.edu/TR/CS-TR-
68-91.html.

Enea:1968:M

[105] Horace J. Enea. MLISP. Technical Report CS-TR-68-92 (PB178078), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1968. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
68-92.html.

Forsythe:1968:CSE

[106] G. E. Forsythe. Computer science and education. Technical Report STAN-CS-68-93 (PB178078), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1968. 50 pp.

Shaw:1968:FDP

[107] A. C. Shaw. The formal description and parsing of pictures. Technical Report STAN-CS-68-94 (SLACR-84), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1968. 205 pp.

Friedman:1968:FST

[108] Joyce Friedman and Robert W. Doran. A formal syntax for transformational grammar. Technical Report CS-TR-68-95, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1968. ??

pp. URL http://i.stanford.edu/TR/CS-TR-68-95.html.

Smith:1968:IAD

[109] Lyle B. Smith. Interval arithmetic determinant evaluation and its use in testing for a Chebyshev system. Technical Report CS-TR-68-96, Stanford University, Department of Computer Science, Stanford, CA, USA, April

1968. ?? pp. URL http://i.stanford.edu/TR/CS-TR-68-96.html.

Miller:1968:RCS

[110] W. F. Miller. Research in the Computer Science Department at Stanford University. Technical Report STAN-CS-68-97, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1968. 49 pp.

Bauer:1968:AWI

[111] Henry R. Bauer, Sheldon I. Becker, and Susan L. Graham. ALGOL W implementation. Technical Report CS-TR-68-98, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1968. ??

pp. URL http://i.stanford.edu/TR/CS-TR-68-98.html.

Friedman:1968:LNF

[112] Joyce Friedman. Lecture notes on foundations for computer science. Technical Report STAN-CS-68-99 (PB179057), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1968. 212 pp.

Kaplan:1968:FTA

[113] D. M. Kaplan. The formal theoretic analysis of stront equivalence for elemental programs. Technical Report STAN-CS-68-101 (AIM-60, AD672923), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1968. 263 pp.

Bartels:1968:NIS

[114] Richard H. Bartels. A numerical investigation of the simplex method. Technical Report STAN-CS-68-104 (AD673010), Stanford University, Department of Computer Science, Stanford, CA, USA, July 31, 1968.

iv + 118 pp. URL https://apps.dtic.mil/sti/pdfs/AD0673010.pdf.

Richman:1968:EC

[115] Paul Lawrence Richman. Epsilon-calculus. Technical Report STAN-CS-68-105 (AD673674), Stanford University, Department of Computer Science, Stanford, CA, USA, August 16, 1968. vii + 138 pp.

Huberman:1968:PPC

[116] Barbara Jane Huberman. A program to play chess end games. Technical Report STAN-CS-68-106 (AIM-65, AD673971), Stanford University, Department of Computer Science, Stanford, CA, USA, August 19, 1968. 168 pp.

Friedman:1968:CET

[117] Joyce Friedman. Computer experiments in transformational grammar. Technical Report STAN-CS-68-108 (AD692681), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1968. 36 pp.

Bauer:1969:AWR

[118] Henry R. Bauer, Sheldon I. Becker, Susan L. Graham, Robert W. Floyd, George E. Forsythe, and Edwin H. Satterthwaite. AL-GOL W (revised). Technical Report CS-TR-68-110, Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1969. ?? pp. URL http://i.stanford.edu/TR/
CS-TR-68-110.html;   https://bitsavers.org/pdf/stanford/cs_
techReports/CS110_ALGOL_W_Revised_Sep69.pdf.

Hansen:1969:ISM

[119] Wildred J. Hansen. The impact of storage management on plex processing language implementation. Technical Report CS-TR-69-113, Stanford University, Department of Computer Science, Stanford, CA, USA, July

1969. ?? pp. URL http://i.stanford.edu/TR/CS-TR-68-113.html.

Satterthwaite:1969:MEP

[120] Edwin H. Satterthwaite. MUTANT 0.5: an experimental programming language. Technical Report CS-TR-69-120, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1969. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-69-120.html.

Moler:1969:ABE

[121] Cleve B. Moler. Accurate bounds for the eigenvalues of the Laplacian and applications to rhombical domains. Technical Report CS-TR-69-121, Stanford University, Department of Computer Science, Stanford, CA, USA, February 19, 1969. i + 17 pp. URL http://i.stanford.

edu/TR/CS-TR-69-121.html.

Mitchell:1969:HAN

[122] William C. Mitchell and Douglas L. McCraith. Heuristic analysis of numerical variants of the Gram–Schmidt orthonormalization process. Technical Report CS-TR-69-122, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1969. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-69-122.html.

Brent:1969:EEP

[123] R. P. Brent. Empirical evidence for a proposed distribution of small prime gaps. Technical Report STAN-CS-69-123 (AD696982), Stanford University, Department of Computer Science, Stanford, CA, USA, February 28, 1969. ii + 18 pp.

Golub:1969:MDS

[124] Gene H. Golub. Matrix decompositions and statistical calculations. Technical Report STAN-CS-69-124 (AD687719), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1969. 52 pp. URL

http://i.stanford.edu/TR/CS-TR-69-124.html.

Feldman:1969:GCI

[125] Jerome A. Feldman, James Gips, James J. Horning, and Stephen Reder. Grammatical complexity and inference. Technical Report CS-TR-69-125, Stanford University, Department of Computer Science, Stanford,

CA, USA, June 1969. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
69-125.html.

Dantzig:1969:CST

[126] George B. Dantzig. Complementary spanning trees. Technical Report CS-TR-69-126, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1969. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-69-126.html.

Vicens:1969:ASR

[127] Pierre Vicens. Aspects of speech recognition by computer. Technical Report STAN-CS-69-127 (AIM-85, AD687720), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1969. 210 pp.

Buzbee:1969:MOE

[128] B. L. Buzbee, Gene H. Golub, and C. W. Nielson. The method of odd/even reduction and factorization with application to Poisson’s equation. Technical Report STAN-CS-69-128 (AD687717), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1969.

iii + 37 pp. URL http://i.stanford.edu/TR/CS-TR-69-128.html.

Miller:1969:RCS

[129] William F. Miller. Research in the Computer Science Department, Stanford University. Technical Report CS-TR-69-129, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1969. ??

pp. URL http://i.stanford.edu/TR/CS-TR-69-129.html.

Schank:1969:CDR

[130] R. C. Schank. A conceptual dependency representation for a computeroriented semantics. Technical Report STAN-CS-69-130 (AIM-83, PB183907), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1969. 201 pp.

Smith:1969:UMM

[131] L. B. Smith. The use of man–machine interaction in data-fitting problems. Technical Report STAN-CS-69-131 (SLAC-96), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1969. 287 pp.

Anonymous:1969:NP

[132] Anonymous. NEVER PRINTED. Technical Report STAN-CS-69-132, Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1969. ?? pp.

Golub:1969:HSL

[133] Gene H. Golub and Christian H. Reinsch. Handbook series linear algebra: Singular value decompositions and least squares solutions. Technical Report STAN-CS-69-133 (AD687718), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1969. ii + 38 pp.

Golub:1969:LLS

[134] Gene H. Golub and Michael A. Saunders. Linear least squares and quadratic programming. Technical Report STAN-CS-69-134 (AD700923), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1969. 35 pp. URL http://i.stanford.edu/

TR/CS-TR-69-134.html.

Gries:1969:CCI

[135] David Gries. CIL: Compiler Implementation Language. Technical Report STAN-CS-69-135 (SLACR-102), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1969. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-69-135.html.

Pohl:1969:BDH

[136] I. Pohl. Bi-directional and heuristic search in path problems. Technical Report STAN-CS-69-136 (SLACR-104), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1969. 157 pp.

Henrici:1969:FPA

[137] Peter Henrici. Fixed points of analytic functions. Technical Report CS-TR-69-137, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1969. iii + 5 pp. URL http://i.stanford.

edu/TR/CS-TR-69-137.html.

Green:1969:ATP

[138] Cordell C. Green. The application of theorem proving to questionanswering systems. Technical Report STAN-CS-69-138 (AIM-96, AD696394), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1969. 162 pp.

Horning:1969:SGI

[139] James Jay Horning. A study of grammatical inference. Technical Report STAN-CS-69-139 (AIM-98, AD695401), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1969. 166 pp.

Forsythe:1969:DTN

[140] George E. Forsythe. Design — then and now. Technical Report STAN-CS-69-140 (AD698799), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1969. ii + 14 pp.

Dahlquist:1969:BEL

[141] Germund Dahlquist, Stanley C. Eisenstat, and Gene H. Golub. Bounds for the error of linear systems of equations using the theory of moments. Technical Report STAN-CS-69-141 (PB188542), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1969. ??

pp. URL http://i.stanford.edu/TR/CS-TR-69-141.html.

Golub:1969:SVR

[142] Gene H. Golub and Richard R. Underwood. Stationary values of the ratio of quadratic forms subject to linear constraints. Technical Report STAN-CS-69-142, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1969. 22 pp. URL http://

i.stanford.edu/TR/CS-TR-69-142.html. Cited in Åke Björck’s bib-

liography on least squares, which is available by anonymous ftp from

math.liu.se in pub/references.

Jenkins:1969:TSV

[143] M. A. Jenkins. Three-stage variable-shift for the solution of polynomial equations with a posteriori error bounds for the zeros. Technical Report STAN-CS-69-143 (AD694464), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1969. 199 pp.

Forsythe:1969:MMP

[144] George E. Forsythe. The maximum and minimum of a positive definite quadratic polynomial on a sphere are convex functions of the radius. Technical Report CS-TR-69-144, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1969. iii + 7 pp. URL

http://i.stanford.edu/TR/CS-TR-69-144.html.

Henrici:1969:MSS

[145] Peter Henrici. Methods of search for solving polynomial equations. Technical Report CS-TR-69-145, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1969. iii + 23 pp. URL

http://i.stanford.edu/TR/CS-TR-69-145.html. Dedicated to D. H.

Lehmer on his 65th birthday.

Ramos:1970:REA

[146] George U. Ramos. Roundoff error analysis of the Fast Fourier Transform. Technical Report STAN-CS-70-146, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1970. ii + 29 pp. URL

http://i.stanford.edu/TR/CS-TR-70-146.html.

Forsythe:1970:PCW

[147] George E. Forsythe. Pitfalls in computation, or why a math book isn’t enough. Technical Report STAN-CS-70-147 (AD699897), Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1970. iv + 43 pp. URL http://i.stanford.edu/pub/cstr/
reports/cs/tr/70/147/CS-TR-70-147.pdf;   http://i.stanford.
edu/TR/CS-TR-70-147.html;   https://apps.dtic.mil/sti/pdfs/
AD0699897.pdf.

Knuth:1970:NAG

[148] Donald E. Knuth and Robert W. Floyd. Notes on avoiding ‘go to’ statements. Technical Report STAN-CS-70-148 (PB188749), Stanford University, Department of Computer Science, Stanford, CA, USA, January

1970. ii + 15 pp. URL https://bitsavers.org/pdf/stanford/cs_
techReports/STAN-CS-70-148_Notes_on_Avoiding_GO_TO_Statements_
Jan1970.pdf.

Knuth:1970:OBS

[149] Donald E. Knuth. Optimum binary search trees. Technical Report STAN-CS-70-149 (PB188748), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1970. 19 pp.

Wilkinson:1970:EPW

[150] James H. Wilkinson. Elementary proof of the Wielandt–Hoffman theorem and of its generalization. Technical Report STAN-CS-70-150 (AD699898), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1970. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-70-150.html.

Volkov:1970:PDS

[151] E. A. Volkov and George E. Forsythe. ‘On the Properties of the Derivatives of the Solutions of Laplace’s Equation and the Errors of the Method of Finite Differences for Boundary Values in C_{2} and C_{1,1}’ by E. A. Volkov. Technical Report STAN-CS-70-151, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1970. ??

pp. URL http://i.stanford.edu/TR/CS-TR-70-151.html.

Gustafson:1970:RCI

[152] Sven-Åke Gustafson. Rapid computation of interpolation formulae and mechanical quadrature rules. Technical Report STAN-CS-70-152, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1970. 23 pp.

Gustafson:1970:EPU

[153] Sven-Åke Gustafson. Error propagation by use of interpolation formulae and quadrature rules which are computed numerically. Technical Report STAN-CS-70-153 (AD701358), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1970. ii + 17 pp.

Stone:1970:SID

[154] H. S. Stone. The spectrum of incorrectly decoded bursts for cyclic error codes. Technical Report STAN-CS-70-154, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1970. 24 pp.

Buzbee:1970:MOE

[155] B. L. Buzbee, Gene H. Golub, and C. W. Nielson. The method of odd/ even reduction and factorization with application to Poisson’s equation, part II. Technical Report STAN-CS-70-155 (AD705508), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1970. 36 pp. URL http://i.stanford.edu/TR/CS-TR-70-155.html.

Dantzig:1970:MCR

[156] George B. Dantzig. On a model for computing round-off error of a sum. Technical Report STAN-CS-70-156 (AD713972), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1970.

?? pp. URL http://i.stanford.edu/TR/CS-TR-70-156.html.

Brent:1970:AMM

[157] Richard P. Brent. Algorithms for matrix multiplication. Technical Report STAN-CS-70-157 (AD705509), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1970. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-70-157.html.

Stone:1970:PPP

[158] H. Stone. Parallel processing with the perfect shuffle. Technical Report STAN-CS-70-158, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1970. 36 pp.

George:1970:UDM

[159] John Alan George. The use of direct methods for the solution of the discrete Poisson equation on non-rectangular regions. Technical Report STAN-CS-70-159 (AD708690), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1970. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-70-159.html.

Bredt:1970:MPC

[160] Thomas H. Bredt and Edward J. McCluskey. A model for parallel computer systems. Technical Report STAN-CS-70-160 (CSL-TR-5, AD707762), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1970. ?? pp. URL http://i.stanford.edu/

TR/CS-TR-70-160.html.

Hoffman:1970:FMA

[161] L. J. Hoffman. The formulary model for access control and privacy in computer systems. Technical Report STAN-CS-70-161 (SLACR-117, UC-32), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1970. 81 pp.

Bartels:1970:NTMa

[162] Richard H. Bartels, Gene H. Golub, and Michael A. Saunders. Numerical techniques in mathematical programming. Technical Report STAN-CS-70-162 (SLACP-760, AD709564), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1970. iv + 61 pp. URL

http://i.stanford.edu/TR/CS-TR-70-162.html.

Malcolm:1970:AFP

[163] Michael A. Malcolm. An algorithm for floating-point accumulation of sums with small relative error. Technical Report STAN-CS-70-163 (AD708691), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1970. 21 pp. URL http://i.stanford.edu/

TR/CS-TR-70-163.html.

Gordonova:1970:ERE

[164] V. I. Gordonova and Linda C. Kaufman. ‘Estimates of the Roundoff Error in the Solution of a System of Conditional Equations’ by V. I. Gordonova. Technical Report STAN-CS-70-164 (AD708692), Stanford University, Department of Computer Science, Stanford, CA, USA, June

1970. ?? pp. URL http://i.stanford.edu/TR/CS-TR-70-164.html.

Bauer:1970:STM

[165] Henry R. Bauer and Harold S. Stone. The scheduling of n tasks with m operations on two processors. Technical Report STAN-CS-70-165, Stanford University, Department of Computer Science, Stanford, CA,

USA, July 1970. ?? pp. URL http://i.stanford.edu/TR/CS-TR-70-
165.html.

Sandewall:1970:RNL

[166] Erik J. Sandewall. Representing natural-language information in predicate calculus. Technical Report STAN-CS-70-166 (AIM-128, AD713841), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1970. iii + 27 pp.

Igarashi:1970:SAL

[167] Shigeru Igarashi. Semantics of ALGOL-like statements. Technical Report STAN-CS-70-167 (AIM-129, AD712460), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1970. iii + 95 pp.

Kelly:1970:VIP

[168] Michael David Kelly. Visual identification of people by computer. Technical Report STAN-CS-70-168 (AIM-130, AD713252), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1970.

Knuth:1970:EFS

[169] Donald E. Knuth. Examples of formal semantics. Technical Report STAN-CS-70-169 (AIM-126, AD711329), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1970. ii + 34 pp.

Bredt:1970:ASC

[170] Thomas H. Bredt. Analysis and synthesis of concurrent sequential programs. Technical Report STAN-CS-70-170 (CSL-TR-6, AD711334), Stanford University, Department of Computer Science, Stanford, CA,

USA, May 1970. ?? pp. URL http://i.stanford.edu/TR/CS-TR-70-
170.html.

Bredt:1970:SMP

[171] Thomas H. Bredt. A survey of models for parallel computing. Technical Report STAN-CS-70-171 (CSL-TR-8, AD714202), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1970. ??

pp. URL http://i.stanford.edu/TR/CS-TR-70-171.html.

Bredt:1970:APS

[172] Thomas H. Bredt. Analysis of parallel systems. Technical Report STAN-CS-70-172 (CSL-TR-7, AD714180), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1970. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-70-172.html.

Bredt:1970:MEP

[173] Thomas H. Bredt. The mutual exclusion problem. Technical Report STAN-CS-70-173 (CSL-TR-9, AD714181), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1970. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-70-173.html.

Manna:1970:TAP

[174] Zohar Manna and Richard J. Waldinger. Towards automatic program synthesis. Technical Report STAN-CS-70-174 (AIM-127, AD711395), Stanford University, Department of Computer Science, Stanford, CA,

USA, July 1970. ?? pp. URL http://i.stanford.edu/TR/CS-TR-70-
174.html.

Malcolm:1970:DCS

[175] Michael A. Malcolm. A description and comparison of subroutines for computing Euclidean inner products on the IBM 360. Technical Report STAN-CS-70-175 (AD713842), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-70-175.html.

Feigenbaum:1970:GPS

[176] Edward A. Feigenbaum, Bruce G. Buchanan, and Joshua Lederberg. On generality and problem solving: a case study using the DENDRAL program. Technical Report STAN-CS-70-176 (AIM-131, AD715128), Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1970. iv + 48 pp. URL http://i.stanford.edu/TR/CS-TR-70-
176.html.

Floyd:1970:BNS

[177] Robert W. Floyd and Donald E. Knuth. The Bose–Nelson sorting problem. Technical Report STAN-CS-70-177 (AD715511), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. ii + 16 pp. Dedicated to R. C. Bose on his seventieth birthday.

Forsythe:1970:RCS

[178] George E. Forsythe and William F. Miller. Research in the Computer Science Department and selected other research in computing at Stanford University. Technical Report STAN-CS-70-178, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. ??

pp. URL http://i.stanford.edu/TR/CS-TR-70-178.html.

Smith:1970:M

[179] David Canfield Smith. MLISP. Technical Report STAN-CS-70-179 (AIM-135, AD716566), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-70-179.html.

Falk:1970:CII

[180] Gilbert Falk. Computer interpretation of imperfect line data as a three-dimensional scene. Technical Report STAN-CS-70-180 (AIM-132, AD715665), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1970. ix + 187 pp.

Hearn:1970:RUM

[181] Anthony C. Hearn. Reduce 2 — user’s manual. Technical Report STAN-CS-70-181 (AIM-133), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. 85 pp.

Tenenbaum:1970:ACV

[182] Jay Martin Tenenbaum. Accommodation in computer vision. Technical Report STAN-CS-70-182 (AIM-134, AD748565), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1970. xiv + 452 pp.

White:1970:MLT

[183] George M. White. Machine learning through signature trees: applications to human speech. Technical Report STAN-CS-70-183 (AIM-136, AD717600), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-70-183.html.

Malcolm:1970:NCJ

[184] Michael A. Malcolm. A note on a conjecture of L. J. Mordell. Technical Report STAN-CS-70-184 (AD715512), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1970. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-70-184.html.

Nelson:1970:GPS

[185] Edward C. Nelson. Graph program simulation. Technical Report STAN-CS-70-185 (TID22593), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-70-185.html.

Knuth:1970:ESF

[186] Donald E. Knuth. An empirical study of Fortran programs. Technical Report STAN-CS-70-186 (AIM-137, AD715513), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1970. ii

+ 42 pp. URL https://apps.dtic.mil/sti/tr/pdf/AD0715513.pdf.

Eisenstat:1970:MMP

[187] Stanley C. Eisenstat, Thomas L. Magnanti, Steven F. Maier, Michael B. McGrath, Vincent J. Nicholson, Christiane Riedl, and George B. Dantzig. MPL, Mathematical Programming Language: specification manual for Committee review. Technical Report STAN-CS-70-187 (AD197154), Stanford University, Department of Computer Science, Stanford, CA,

USA, November 1970. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-70-187.html.

Ashcroft:1971:TGP

[188] Edward A. Ashcroft and Zohar Manna. The translation of “Go To” programs to “While” programs. Technical Report STAN-CS-71-188 (AIM-138, PB197161), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1971. ii + 29 pp. URL http://i.

stanford.edu/TR/CS-TR-71-188.html.

Manna:1970:MTP

[189] Zohar Manna. Mathematical theory of partial correctness. Technical Report STAN-CS-71-189 (AIM-139, AD717601), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1971. ii + 23 pp.

Hopcroft:1970:AMS

[190] John Hopcroft. An N log N algorithm for minimizing states in a finite automaton. Technical Report STAN-CS-70-190 (AD719398), Stanford University, Department of Computer Science, Stanford, CA, USA, December

1970. 12 pp. URL https://apps.dtic.mil/sti/html/tr/AD0719398/
index.html.

Lesser:1970:IDE

[191] V. Lesser. An introduction to the direct emulation of control structures by a parallel micro-computer. Technical Report STAN-CS-70-191 (SLACP-904, PB198494), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1970. 26 pp.

Hopcroft:1970:AIP

[192] John Hopcroft. An N log N algorithm for isomorphism of planar triply connected graphs. Technical Report STAN-CS-70-192 (AD719399), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1970. 6 pp.

Schank:1970:IMC

[193] R. Schank. Intention, memory and computer understanding. Technical Report STAN-CS-70-193 (AIM-140), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1970. 59 pp.

Knuth:1970:ACP

[194] Donald E. Knuth. The Art of Computer Programming — errata et addenda. Technical Report STAN-CS-70-194 (PB198495), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1970. 28 pp.

Buzbee:1970:DSD

[195] B. L. Buzbee, Fred W. Dorr, John Alan George, and Gene H. Golub. The direct solution of the discrete Poisson equation on irregular regions. Technical Report STAN-CS-71-195 (723871), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1970. vi

+ 30 pp. URL http://i.stanford.edu/TR/CS-TR-71-195.html.

Moler:1970:MCF

[196] C. B. Moler. Matrix computations with Fortran and paging. Technical Report STAN-CS-70-196 (AD725167), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1970. 13 pp.

Manna:1971:MTP

[197] Zohar Manna. Mathematical theory of partial correctness. Technical Report STAN-CS-71-189, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1971. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-71-189.html.

Hopcroft:1971:LAM

[198] John E. Hopcroft. An n log n algorithm for minimizing states in a finite automaton. Technical Report STAN-CS-71-190, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1971. ??

pp. URL http://i.stanford.edu/TR/CS-TR-71-190.html.

Lesser:1971:IDE

[199] Victor R. Lesser. An introduction to the direct emulation of control structures by a parallel micro-computer. Technical Report STAN-CS-71-191, Stanford University, Department of Computer Science, Stanford,

CA, USA, January 1971. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-71-191.html.

Hopcroft:1971:LAI

[200] John E. Hopcroft. An n log n algorithm for isomorphism of planar triply connected graphs. Technical Report STAN-CS-71-192, Stanford University, Department of Computer Science, Stanford, CA, USA, January

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-192.html.

Schank:1971:IMC

[201] Roger C. Schank. Intention, memory, and computer understanding. Technical Report STAN-CS-71-193, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1971. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-71-193.html.

Knuth:1971:MUG

[202] Donald E. Knuth and Richard L. Sites. MIX/360 user’s guide. Technical Report STAN-CS-71-197, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1, 1971.

?? pp. URL http://i.stanford.edu/TR/CS-TR-71-197.html;
 http://www-db.stanford.edu/TR/CS-TR-71-197.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-71-197.

Brent:1971:AFZ

[203] Richard P. Brent. Algorithms for finding zeros and extrema of functions without calculating derivatives. Technical Report STAN-CS-71-198 (AD726170), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. + 313 pp.

Staff:1971:BSC

[204] Staff. Bibliography of Stanford Computer Science reports 1963–1971. Technical Report STAN-CS-71-199 (PB198415), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. 28 pp.

Herriot:1971:APC

[205] J. G. Herriot and Christian H. Reinsch. ALGOL 60 procedures for the calculation of interpolating natural spline functions. Technical Report STAN-CS-71-200 (PB198416), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. ii + 30 pp.

Hopcroft:1971:PTV

[206] John E. Hopcroft and Robert Endre Tarjan. Planarity testing in V log V steps: extended abstract. Technical Report STAN-CS-71-201 (AD722434), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-71-201.html.

Saal:1971:CS

[207] Harry J. Saal and William E. Riddle. Communicating semaphores. Technical Report STAN-CS-71-202 (SLAC-117), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-71-202.html.

Buchanan:1971:HDP

[208] Bruce G. Buchanan and Joshua Lederberg. The Heuristic DENDRAL program for explaining empirical data. Technical Report STAN-CS-71-203 (AIM-141, AD730506), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-71-203.html.

Ingalls:1971:FFE

[209] Daniel H. H. Ingalls. FETE: a Fortran execution time estimator. Technical Report STAN-CS-71-204 (PB198510), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-71-204.html.

Milner:1971:ADS

[210] Robin Milner. An algebraic definition of simulation between programs. Technical Report STAN-CS-71-205 (AIM-142, AD731383), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-205.html.

Knuth:1971:MAA

[211] Donald E. Knuth. Mathematical analysis of algorithms. Technical Report STAN-CS-71-206 (AD726158), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1971. iii + 26 pp.

Hopcroft:1971:EAG

[212] John E. Hopcroft and Robert Endre Tarjan. Efficient algorithms for graph manipulation. Technical Report STAN-CS-71-207 (AD726169), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
71-207.html.

George:1971:CIF

[213] John Alan George. Computer implementation of the finite element method. Technical Report STAN-CS-71-208 (AD726171), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. vi + 222 pp.

McCarthy:1971:PTR

[214] John McCarthy, Arthur L. Samuel, Edward A. Feigenbaum, and Joshua Lederberg. Project technical report. Technical Report STAN-CS-71-209 (AIM-143, AD724867), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1971. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-71-209.html.

Purdy:1971:APC

[215] J. Gerry Purdy. ACCESS: a program for the catalog and access of information. Technical Report STAN-CS-71-210 (PB201917), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-210.html.

Malcolm:1971:ARP

[216] Michael A. Malcolm. Algorithms to reveal properties of floating-point arithmetic. Technical Report STAN-CS-71-211 (AD727104), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-211.html.

Morgana:1971:TMR

[217] Maria Aurora Morgana. Time and memory requirements for solving linear systems. Technical Report STAN-CS-71-212 (AD727107), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-212.html.

Tarjan:1971:SPS

[218] Robert Endre Tarjan. The switchyard problem: sorting using networks of queues and stacks. Technical Report STAN-CS-71-213 (PB201629), Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-
213.html.

Graham:1971:CTC

[219] Ronald L. Graham, Donald E. Knuth, and T. S. Motzkin. Complements and transitive closures. Technical Report STAN-CS-71-214 (AD727108), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1971. 13 pp.

Malcolm:1971:PRP

[220] Michael A. Malcolm. PL360 (revised) a programming language for the IBM 360. Technical Report STAN-CS-71-215 (AD727115), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1971.

v + 94 pp. URL http://infolab.stanford.edu/pub/cstr/reports/
cs/tr/71/215/CS-TR-71-215.pdf;   https://bitsavers.org/pdf/
stanford/cs_techReports/STAN-CS-71-215_PL360_Rev_May72.pdf.

Kling:1971:RAA

[221] Robert Elliot Kling. Reasoning by analogy with applications to heuristics problem solving: a case study. Technical Report STAN-CS-71-216 (AIM-147, AD732457), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. ix + 191 pp.

Ashcroft:1971:DPM

[222] Edward A. Ashcroft, Zohar Manna, and Amir Pneuli. Decidable properties of monadic functional schemas. Technical Report STAN-CS-71-217 (AIM-148, AD731730), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-71-217.html.

deBruijn:1971:AHP

[223] N. G. de Bruijn, Donald E. Knuth, and S. O. Rice. The average height of plane trees. Technical Report STAN-CS-71-218 (AD731038), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1971. iv + 7 pp.

Quam:1971:CCP

[224] Lynn H. Quam. Computer comparison of pictures. Technical Report STAN-CS-71-219 (AIM-144), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1971. x + 118 pp. URL https:

//apps.dtic.mil/sti/tr/pdf/AD0785172.pdf.

Stone:1971:DME

[225] Harold S. Stone. Dynamic memories with enhanced data access. Technical Report STAN-CS-71-220 (CSL-14, AD727116), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1971. vii + 29 + 5 pp.

Buchanan:1971:HPS

[226] Bruce G. Buchanan, Edward A. Feigenbaum, and Joshua Lederberg. A heuristic programming study of theory formation in science. Technical Report STAN-CS-71-221 (AIM-145, AD731729), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-71-221.html.

Meyers:1971:LRT

[227] W. J. Meyers. Linear representation of tree structure (a mathematical theory of parenthesis-free notations). Technical Report STAN-CS-71-222 (PB235417/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1971. 245 pp.

Graham:1971:PLB

[228] Susan Graham. Precedence languages and bounded right context languages. Technical Report STAN-CS-71-223 (PB203429), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971. 192 pp.

Ershov:1971:PP

[229] Andrei P. Ershov. Parallel programming. Technical Report STAN-CS-71-224 (AIM-146, PB212183), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971. 15 pp. URL http://i.

stanford.edu/TR/CS-TR-71-224.html.

Bjorck:1971:NMC

[230] Åke Björck and Gene H. Golub. Numerical methods for computing angles between linear subspaces. Technical Report STAN-CS-71-225 (PB203344), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971. 30 pp. URL http://i.stanford.edu/

TR/CS-TR-71-225.html.

George:1971:SSP

[231] James E. George. SIMPLE: a simple precedence translator writing system. Technical Report STAN-CS-71-226 (SLAC-133), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971.

?? pp. URL http://i.stanford.edu/TR/CS-TR-71-226.html.

George:1971:GGE

[232] James E. George. GEMS — a graphical experimental meta system. Technical Report STAN-CS-71-227 (SLAC-134), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971. 184 pp.

Kaufman:1971:FMA

[233] Linda C. Kaufman. Function minimization and automatic therapeutic control. Technical Report STAN-CS-71-228 (PB203343), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1971.

?? pp. URL http://i.stanford.edu/TR/CS-TR-71-228.html.

Lee:1971:VSN

[234] Erastus H. Lee and George E. Forsythe. Variational study of nonlinear spline curves. Technical Report STAN-CS-71-229 (AD732766), Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-229.
html.

Sites:1971:ARM

[235] Richard L. Sites. ALGOL with reference manual. Technical Report STAN-CS-71-230 (PB203601), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. 141 pp.

Schmidt:1971:SRT

[236] Rodney A. Schmidt, Jr. A study of the real-time control of a computer driven vehicle. Technical Report STAN-CS-71-231 (AIM-149, AD732644), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. iv + 164 pp.

Moler:1971:AGM

[237] Cleve B. Moler and Gilbert W. Stewart. An algorithm for the generalized matrix eigenvalue problem Ax = λBx. Technical Report STAN-CS-71-232 (AD733073, CNA-32), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1971. ii + 50 pp. URL https:/

/apps.dtic.mil/sti/tr/pdf/AD0746896.pdf. Issued jointly as report

CNA 32 by the Center for Numerical Analysis, the University of Texas at Austin.

Wilner:1971:DSD

[238] Wayne Wilner. Declarative semantic definition. Technical Report STAN-CS-71-233, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. 211 pp.

Golub:1971:SMEa

[239] Gene H. Golub. Some modified eigenvalue problems. Technical Report STAN-CS-71-234 (SU326 P30-11), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. 38 pp. URL http:

//i.stanford.edu/TR/CS-TR-71-234.html.

Floyd:1971:TID

[240] Robert W. Floyd. Toward iterative design of correct programs. Technical Report STAN-CS-71-235 (AIM-150), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1971. 12 pp.

Golub:1971:NCU

[241] Gene H. Golub and George P. H. Styan. Numerical computations for univariate linear models. Technical Report STAN-CS-71-236 (AD737648), Stanford University, Department of Computer Science, Stanford, CA,

USA, September 1971. 35 pp. URL http://i.stanford.edu/TR/CS-
TR-71-236.html.

VanVoorhis:1971:GDS

[242] David C. Van Voorhis. A generalization of the divide-sort-merge strategy for sorting networks. Technical Report STAN-CS-71-237 (CSL-TR-16, AD737270), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-71-237.html.

VanVoorhis:1971:LBS

[243] David C. Van Voorhis. A lower bound for sorting networks that use the divide-sort-merge strategy. Technical Report STAN-CS-71-238 (CSL-TR-17, AD735901), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-71-238.html.

VanVoorhis:1971:LSN

[244] David C. Van Voorhis. Large [g, d] sorting networks. Technical Report STAN-CS-71-239 (CSL-TR-18, AD736610), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-71-239.html.

London:1971:CTC

[245] Ralph L. London. Correctness of two compilers for a Lisp subset. Technical Report STAN-CS-71-240 (AIM-151, AD738568), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1971. ??

pp. URL http://i.stanford.edu/TR/CS-TR-71-240.html.

Bierman:1971:ITM

[246] Alan W. Biermann. On the inference of Turing machines from sample computations. Technical Report STAN-CS-71-241 (AIM-152, AD732642), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1971. ii + 31 pp.

Hayes:1971:FPR

[247] Patrick J. Hayes. The frame problem and related problems in artificial intelligence. Technical Report STAN-CS-71-242 (AIM-153, AD738569), Stanford University, Department of Computer Science, Stanford, CA,

USA, November 1971. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-71-242.html.

Manna:1971:IMP

[248] Zohar Manna, Stephen Ness, and Jean Vuillemin. Inductive methods for proving properties of programs. Technical Report STAN-CS-71-243 (AIM-154, AD738570), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1971. iv + 24 pp.

Tarjan:1971:EPA

[249] Robert E. Tarjan. An efficient planarity algorithm. Technical Report STAN-CS-71-244 (AD738027), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1971. v + 154 pp.

Ryder:1971:HAL

[250] John Ryder. Heuristic analysis of large trees as generated in the game of Go. Technical Report STAN-CS-71-245 (AIM-155), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1971. 350 pp.

Colby:1971:RTV

[251] Kenneth Mark Colby, Franklin Dennis Hilf, Sylvia Weber, and Helena C. Kraemer. A resemblance test for the validation of a computer simulation of paranoid processes. Technical Report STAN-CS-71-246 (AIM-156, AD740141), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1971. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-71-246.html.

Wilks:1971:OSH

[252] Yorick A. Wilks. One small head — some remarks on the use of ’model’ in linguistics. Technical Report STAN-CS-71-247 (AIM-157), Stanford University, Department of Computer Science, Stanford, CA, USA, December

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-247.html.

Fredman:1971:RRB

[253] Michael L. Fredman and Donald E. Knuth. Recurrence relations based on minimization. Technical Report STAN-CS-71-248 (AD739335), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1971. ii + 35 pp.

Pollack:1971:ABC

[254] Bary W. Pollack. An annotated bibliography on the construction of compilers. Technical Report STAN-CS-71-249, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1971. ??

pp. URL http://i.stanford.edu/TR/CS-TR-71-249.html.

Chandra:1971:PSE

[255] Ashok K. Chandra and Zohar Manna. Program schemas with equality. Technical Report STAN-CS-71-250 (AIM-158, AD740127), Stanford University, Department of Computer Science, Stanford, CA, USA, December

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-250.html.

Malcolm:1972:PRP

[256] Michael A. Malcolm. PL360 (revised): a programming language for the IBM 360. Technical Report STAN-CS-71-215 (AD727115), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972.

vii + 103 pp. URL http://i.stanford.edu/TR/CS-TR-71-215.html.

Malcolm:1972:PRA

[257] Michael A. Malcolm. PL360 (revised again) a programming language for the IBM 360. Technical Report STAN-CS-71-215 (revised again), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972. iv + 103 pp. URL https://bitsavers.org/

pdf/stanford/cs_techReports/STAN-CS-71-215_PL360_Revised_A_
Programming_Language_For_The_IBM_360_May71.pdf. Original ver-

sion May 1971.

Sites:1972:AWR

[258] Richard L. Sites. ALGOL W reference manual. Technical Report STAN-CS-71-230 (PB203601), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. i + 141

pp. URL http://i.stanford.edu/TR/CS-TR-71-230.html;   https:
//bitsavers.org/pdf/stanford/cs_techReports/STAN-CS-71-230_
Algol_W_Reference_Manual_Feb72.pdf.

Stone:1972:EPA

[259] Harold Stone. An efficient parallel algorithm for the solution of a tridiagonal linear system of equation. Technical Report STAN-CS-72-251 (CSL-TR-19, AD736814), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1971. iii + 21 pp.

Saunders:1972:LSL

[260] Michael A. Saunders. Large-scale linear programming using the Cholesky factorization. Technical Report STAN-CS-72-252 (SU326 P30-14), Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1972. 64 pp. URL http://i.stanford.edu/TR/CS-TR-72-
252.html.

Feldman:1972:TCI

[261] Jerome A. Feldman and Paul C. Shields. Total complexity and the inference of best programs. Technical Report STAN-CS-72-253 (AIM-159), Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-
253.html.

Forsythe:1972:NCMa

[262] George E. Forsythe. Von Neumann’s comparison method for random sampling from the normal and other distributions. Technical Report STAN-CS-72-254 (AD740330), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1972. ii + 19

pp. URL http://i.stanford.edu/TR/CS-TR-72-254.html;   https:
//apps.dtic.mil/sti/tr/pdf/AD0740330.pdf.

Feldman:1972:AP

[263] Jerome A. Feldman. Automatic programming. Technical Report STAN-CS-72-255 (AIM-160, AD740140), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-72-255.html.

Chvatal:1972:EPW

[264] Vaclav Chvátal. Edmonds polyhedra and weakly Hamiltonian graphs. Technical Report STAN-CS-72-256 (AD740331), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1972.

?? pp. URL http://i.stanford.edu/TR/CS-TR-72-256.html.

Wirth:1972:PCG

[265] Niklaus Wirth. On ‘PASCAL,’ code generation, and the CDC 6000 computer. Technical Report STAN-CS-72-257 (PB208519), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-257.html;
 https://bitsavers.org/pdf/stanford/cs_techReports/STAN-CS-
72-257_6600_PASCAL_Feb72.pdf.

Brown:1972:SBM

[266] Harold Brown. Some basic machine algorithms for integral order computations. Technical Report STAN-CS-72-258 (AD740332), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-258.html.

Crane:1972:LLP

[267] Clark A. Crane. Linear lists and priority queues as balanced binary trees. Technical Report STAN-CS-72-259 (PB208595), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. 131 pp.

Pratt:1972:SSN

[268] Vaughan R. Pratt. Shellsort and sorting networks. Technical Report STAN-CS-72-260 (AD740110), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. v + 59 pp.

Golub:1972:DPN

[269] Gene H. Golub and Victor Pereyra. The differentiation of pseudoinverses and nonlinear least squares problems whose variables separate. Technical Report STAN-CS-72-261 (SU326 P30 15), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. 35 pp.

URL http://i.stanford.edu/TR/CS-TR-72-261.html.

Staff:1972:B

[270] Staff. Bibliography. Technical Report STAN-CS-72-262 (PB209357), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. 36 pp.

Klarner:1972:PIU

[271] David A. Klarner and Ronald L. Rivest. A procedure for improving the upper bound for the number of n-ominoes. Technical Report STAN-CS-72-263 (AD741189), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-72-263.html.

Wilks:1972:AIA

[272] Yorick A. Wilks. An artificial intelligence approach to machine translation. Technical Report STAN-CS-72-264 (AIM-161, AD741189), Stanford University, Department of Computer Science, Stanford, CA, USA,

February 1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-
264.html.

Schank:1972:PCU

[273] Roger C. Schank, Neil M. Goldman, Charles J. Rieger, and Christopher K. Riesbeck. Primitive concepts underlying verbs of thought. Technical Report STAN-CS-72-265 (AIM-162, AD744634), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1972. iv + 98 pp. URL http://i.stanford.edu/TR/CS-TR-72-265.
html;   https://apps.dtic.mil/sti/tr/pdf/AD0744634.pdf.

Cadiou:1972:RDP

[274] Jean M. Cadiou. Recursive definitions of partial and functions and their computation. Technical Report STAN-CS-72-266 (AIM-163), Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1972. v + 157 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA022048.pdf.

Bonzon:1972:MPL

[275] Pierre E. Bonzon. Mathematical programming language: an appraisal based on practical experiments. Technical Report STAN-CS-72-267 (PB209629), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-267.html.

Chvatal:1972:DM

[276] Vaclav Chvátal. Degrees and matchings. Technical Report STAN-CS-72-268 (AD742348), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-268.html.

Klarner:1972:APC

[277] David A. Klarner and Richard Rado. Arithmetic properties of certain recursively defined sets. Technical Report STAN-CS-72-269 (AD742747), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
72-269.html.

Golub:1972:LAS

[278] Gene H. Golub, Richard R. Underwood, and James H. Wilkinson. The Lanczos algorithm for the symmetric Ax = λBx problem. Technical Report STAN-CS-72-270 (PB209616 SU326P30-16), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. iii

+ 24 pp. URL http://i.stanford.edu/TR/CS-TR-72-270.html.

Riddle:1972:MAS

[279] William E. Riddle. The modeling and analysis of supervisory systems. Technical Report STAN-CS-72-271, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. 174 pp.

Manna:1972:FAT

[280] Zohar Manna and Jean Vuillemin. Fixpoint approach to the theory of computation. Technical Report STAN-CS-72-272 (A7IM-164, AD742748), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-272.html.

Chvatal:1972:CAG

[281] Vaclav Chvátal and Jiri Sichler. Chromatic automorphisms of graphs. Technical Report STAN-CS-72-273 (PB209806), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-72-273.html.

Klarner:1972:LCS

[282] D. Klarner and Richard Rado. Linear combinations of sets of consecutive integers. Technical Report STAN-CS-72-274 (AD742749), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. 12 pp.

Klarner:1972:SGI

[283] David A. Klarner. Sets generated by iteration of a linear operation. Technical Report STAN-CS-72-275 (AD742750), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. 16 pp. URL

http://i.stanford.edu/TR/CS-TR-72-275.html.

Kaufman:1972:GMS

[284] Linda Kaufman. A generalization of the LR method to solve Ax = λBx. Technical Report STAN-CS-72-276 (AD745022), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1972. ii

+ 72 pp. URL https://apps.dtic.mil/sti/tr/pdf/AD0745022.pdf.

Zahn:1972:RBT

[285] C. T. Zahn. Region boundaries on a triangular grid. Technical Report STAN-CS-72-277 (SLAC-149), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1972. 40 pp.

Concus:1972:UFD

[286] Paul Concus and Gene H. Golub. Use of fast direct methods for the efficient numerical solution of nonseparable elliptic equations. Technical Report STAN-CS-72-278 (SU326 P30-17), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1972. ii + 39 pp.

URL http://i.stanford.edu/TR/CS-TR-72-278.html.

Osborne:1972:TO

[287] Michael R. Osborne. Topics in optimization. Technical Report STAN-CS-72-279 (AD744313), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-279.html.

Bochvar:1972:TPP

[288] D. A. Bochvar. Two papers on partial predicate calculus. Technical Report STAN-CS-72-280 (AIM-165, AD742751), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1972. i + 50 pp.

Quam:1972:CIP

[289] Lynn H. Quam, Sidney Liebes, Jr., Robert B. Tucker, Marsha Jo Hannah, and Botond G. Eross. Computer interactive picture processing. Technical Report STAN-CS-72-281 (AIM-166, AD743598), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1972. ??

pp. URL http://i.stanford.edu/TR/CS-TR-72-281.html.

Chandra:1972:ECL

[290] Ashok K. Chandra. Efficient compilation of linear recursive programs. Technical Report STAN-CS-72-282 (AIM-167, AD747254), Stanford University, Department of Computer Science, Stanford, CA, USA, April

1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-282.html.

Stoutemyer:1972:NIS

[291] David R. Stoutemyer. Numerical implementation of the Schwarz alternating procedure for elliptic partial differential equations. Technical Report STAN-CS-72-283, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972. 131 pp.

Chvatal:1972:EPH

[292] Vaclav Chvátal. Edmonds polyhedra and a hierarchy of combinatorial problems. Technical Report STAN-CS-72-284 (O.R. 72-6, AD745778), Stanford University, Department of Computer Science, Stanford, CA,

USA, May 1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-
284.html.

Floyd:1972:LTT

[293] Robert W. Floyd and Alan J. Smith. A linear time two tape merge. Technical Report STAN-CS-72-285 (PB210910), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972. 16 pp.

Chandra:1972:SMP

[294] Ashok K. Chandra. On the solution of Moser’s problem in four dimensions, and related issues. A collection of two papers: On the solution of Moser’s problem in four dimensions and independent permutations as related to a problem of Moser and a theorem of Polya. Technical Report STAN-CS-72-286 (PB211036), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-72-286.html.

Igarashi:1972:AFP

[295] Shigeru Igarashi. Admissibility of fixed-point induction in first-order logic of typed theories. Technical Report STAN-CS-72-287 (AIM-168, AD746146), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972. iii + 41 pp. URL https://dl.acm.org/

doi/10.5555/646795.759643.

Milner:1972:LCF

[296] Robin Milner. Logic for computable functions: Description of a machine implementation. Technical Report STAN-CS-72-288 (AIM-169), Stanford University, Department of Computer Science, Stanford, CA, USA,

May 1972. ii + 36 pp. URL http://i.stanford.edu/TR/CS-TR-72-
288.html;   https://apps.dtic.mil/sti/pdfs/AD0785072.pdf.

Wilks:1972:LLN

[297] Yorick A. Wilks. Lakoff on linguistics and natural logic. Technical Report STAN-CS-72-289 (AIM-170, AD748607), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-72-289.html.

Schank:1972:AB

[298] Roger C. Schank. Adverbs and belief. Technical Report STAN-CS-72-290 (AIM-171, AD746147), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-72-290.html.

Knuth:1972:SCL

[299] Donald E. Knuth. Some combinatorial lemmas. Technical Report STAN-CS-72-291 (AD746189), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1, 1972.

?? pp. URL http://i.stanford.edu/TR/CS-TR-72-291.html;
 http://www-db.stanford.edu/TR/CS-TR-72-291.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-72-291.

Chvatal:1972:SCR

[300] Vaclav Chvátal, David A. Klarner, and Donald E. Knuth. Selected combinatorial research problems. Technical Report STAN-CS-72-292 (AD746150), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1, 1972. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-72-292.html;   http://
www-db.stanford.edu/TR/CS-TR-72-292.html;   http://www.ncstrl.
org:8900/ncstrl/servlet/search?formname=detail&id=oai%3Ancstrlh%
3Astan%3ASTAN%2F%2FCS-TR-72-292.

Lukes:1972:CSP

[301] J. A. Lukes. Combinatorial solutions to partitioning problems. Technical Report STAN-CS-72-293 (CSL-TN-32, PB212234), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. xi + 120 pp.

Saal:1972:MIC

[302] Harry J. Saal and Leonard J. Shustek. Microprogrammed implementation of computer measurement techniques. Technical Report STAN-CS-72-294 (SLACP-1072), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. 20 pp.

Paige:1972:BMS

[303] C. C. Paige. Bidiagonalization of matrices and solution of linear equations. Technical Report STAN-CS-72-295 (PB212130), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. ii + 28 pp.

Fredman:1972:GPC

[304] Michael L. Fredman. Growth properties of a class of recursively defined functions. Technical Report STAN-CS-72-296 (AD748606), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. 84 pp.

Paige:1972:EAM

[305] C. C. Paige. An error analysis of a method for solving matrix equations. Technical Report STAN-CS-72-297 (PB212300), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. 13 pp.

Kogge:1972:PAEa

[306] Peter M. Kogge and Harold S. Stone. A parallel algorithm for the efficient solution of a general class of recurrence equations. Technical Report STAN-CS-72-298 (CSL-TR-25), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1972. 33 pp.

Russell:1972:SCN

[307] Sylvia Weber Russell. semantic categories of nominals for conceptual dependency analysis of natural language. Technical Report STAN-CS-72-299 (AIM-172, AD752801), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1972. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-72-299.html.

Kaufman:1972:CCF

[308] Marc T. Kaufman. Counterexample to a conjecture of Fujii, Kasami and Ninomiya. Technical Report STAN-CS-72-300 (CSL-TN-17, AD749848), Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-
300.html.

Saunders:1972:PFC

[309] Michael A. Saunders. Product form of the Cholesky factorization for large-scale linear programming. Technical Report STAN-CS-72-301 (SU326 P30-21), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-301.html.

Golub:1972:SUL

[310] G. H. Golub. Some uses of the Lanczos algorithm in numerical linear algebra. Technical Report STAN-CS-72-302 (SU326, P30-19), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. 23 pp.

Morris:1972:CTP

[311] Francis Lockwood Morris. Correctness of translations of programming languages — an algebraic approach. Technical Report STAN-CS-72-303 (AIM-174, PB212827, AD-A954771), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. iii + 126 pp.

Anderssen:1972:RNS

[312] Robert S. Anderssen and Gene H. Golub. Richardson’s non-stationary matrix iterative procedure. Technical Report STAN-CS-72-304 (SU326 P30-20), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. 76 pp. URL http://i.stanford.edu/

TR/CS-TR-72-304.html.

Agin:1972:RDC

[313] Gerald Jacob Agin. Representation and description of curved objects. Technical Report STAN-CS-72-305 (AIM-173, AD755139), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1972. iv + 133 pp.

Pollack:1972:BCG

[314] Bary W. Pollack. A bibliography on computer graphics. Technical Report STAN-CS-72-306 (SU326 P23-X-2), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-72-306.html.

Tanaka:1972:HTS

[315] Hozumi Tanaka. Hadamard transform for speech wave analysis. Technical Report STAN-CS-72-307 (AIM-175), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-72-307.html.

Feldman:1972:RDS

[316] Jerome A. Feldman, James R. Low, Daniel C. Swinehart, and Russell H. Taylor. Recent developments in SAIL, an ALGOL-based language for artificial intelligence. Technical Report STAN-CS-72-308 (AIM-176, AD754109), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-308.html.

Lesser:1972:DCS

[317] V. Lesser. Dynamic control structures and their use in emulation. Technical Report STAN-CS-72-309, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. 251 pp.

Kaufman:1972:ASU

[318] Marc T. Kaufman. Anomalies in scheduling unit-time tasks. Technical Report STAN-CS-72-310 (CSL-TR-34, AD750671), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. ??

pp. URL http://i.stanford.edu/TR/CS-TR-72-310.html.

Paul:1972:MTC

[319] Richard Paul. Modelling, trajectory calculation and serving of a computer controlled arm. Technical Report STAN-CS-72-311 (AIM-177, AD785071), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1972. iv + 89 pp. URL https://apps.

dtic.mil/sti/tr/pdf/AD0785071.pdf.

Gill:1972:VFR

[320] Aharon Gill. Visual feedback and related problems in computer controlled hand-eye coordination. Technical Report STAN-CS-72-312 (AIM-178, AD754108), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1972. v + 130 pp.

Staff:1972:BCS

[321] Staff. Bibliography of computer science reports. Technical Report STAN-CS-72-313 (PB218353/1), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1972. 42 pp.

Kogge:1972:PAEb

[322] Peter M. Kogge. Parallel algorithms for the efficient solution of recurrence problems. Technical Report STAN-CS-72-314 (CSL-TR-43, PB212893), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1972. 74 pp.

Kogge:1972:NSP

[323] Peter M. Kogge. The numerical stability of parallel algorithms for solving recurrence problems. Technical Report STAN-CS-72-315 (CSL-TR-44, PB212894), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1972. 49 pp.

Kogge:1972:MPS

[324] Peter M. Kogge. Minimal parallelism in the solution of recurrence problems. Technical Report STAN-CS-72-316 (CSL-TR-45, PB212828), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1972. 45 pp.

Fuller:1972:ADS

[325] Samuel H. Fuller and Forest Baskett. An analysis of drum storage units. Technical Report STAN-CS-72-317 (CSL-TR-26, AD750672), Stanford University, Department of Computer Science, Stanford, CA, USA, Au-

gust 1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-72-317.
html.

Brown:1972:CGL

[326] Harold Brown, Larry M. Masinter, and Larry Hjelmeland. Constructive graph labeling using double cosets. Technical Report STAN-CS-72-318 (AD755140), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-318.html.

Golub:1972:CBS

[327] Gene H. Golub and James M. Varah. On a characterization of the best ℓ_{2} scaling of a matrix. Technical Report STAN-CS-72-319 (SU326 P30-22), Stanford University, Department of Computer Science, Stanford, CA,

USA, October 1972. 14 pp. URL http://i.stanford.edu/TR/CS-TR-
72-319.html.

Baumgart:1972:WEP

[328] Bruce G. Baumgart. Winged edge polyhedron representation. Technical Report STAN-CS-72-320 (AIM-179), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1972. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-72-320.html.

Bajcsy:1972:CIT

[329] Ruzena Bajcsy. Computer identification of textured visual scenes. Technical Report STAN-CS-72-321 (AIM-180, AD759712), Stanford University, Department of Computer Science, Stanford, CA, USA, Octo-

ber 1972. iv + 157 pp. URL https://apps.dtic.mil/sti/tr/pdf/
AD0759712.pdf.

Gill:1972:MMM

[330] Phillip E. Gill, Gene H. Golub, Walter A. Murray, and Michael A. Saunders. Methods for modifying matrix factorizations. Technical Report STAN-CS-72-322 (SU326 P30-23), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1972. 62 pp. URL

http://i.stanford.edu/TR/CS-TR-72-322.html.

Malcolm:1972:FMS

[331] Michael A. Malcolm and John Palmer. A fast method for solving a class of tri-diagonal linear systems. Technical Report STAN-CS-72-323, Stanford University, Department of Computer Science, Stanford, CA,

USA, November 1972. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-72-323.html.

Bauer:1972:SMX

[332] Henry Raymond Bauer, III. Subproblems of the m × n sequencing problem. Technical Report STAN-CS-72-324 (CSL-TR-48, SU-SEL-72-057, PB214612), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1972. viii + 110 pp.

Buchanan:1972:RHD

[333] Bruce G. Buchanan. Review of Hubert Dreyfus’ What Computers Can’t Do: a Critique of Artificial Reason (Harper & Row, New York, 1972).

Technical Report STAN-CS-72-325 (AIM-181), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1972. ??

pp. URL http://i.stanford.edu/TR/CS-TR-72-325.html.

Colby:1972:CEJ

[334] Kenneth Mark Colby and Franklin Dennis Hilf. Can expert judges, using transcripts of teletyped psychiatric interviews, distinguish human paranoid patients from a computer simulation of paranoid processes? Technical Report STAN-CS-72-326 (AIM-182, AD754107), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1972.

?? pp. URL http://i.stanford.edu/TR/CS-TR-72-326.html.

Klarner:1972:ABN

[335] David A. Klarner and Ronald L. Rivest. Asymptotic bounds for the number of convex n-ominoes. Technical Report STAN-CS-72-327 (AD755138), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1972. iii + 14 pp.

Gabow:1972:EIE

[336] Harold N. Gabow. An efficient implementation of Edmonds’ maximum matching algorithm. Technical Report STAN-CS-72-328 (CSL-TR-31, PB218929), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1972. ?? pp. URL http://i.stanford.edu/TR/

CS-TR-72-328.html.

Fang:1972:FDF

[337] Isu Fang. FOLDS, a declarative formal language definition system. Technical Report STAN-CS-72-329 (PB218875), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1972. + 286 pp.

Newey:1972:ATI

[338] Malcolm Newey. Axioms and theorems for integers, lists and finite sets in LCF. Technical Report STAN-CS-73-330 (AIM-184, AD758651), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1972. 53 pp.

Fuller:1972:PCM

[339] Samuel H. Fuller. Performance of an I/O channel with multiple paging drums (digest edition). Technical Report STAN-CS-73-351 (SU-SEL-73-010), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. iii + 6 pp. URL http://i.stanford.edu/TR/

CS-TR-73-351.html.

Fuller:1972:EDB

[340] Samuel H. Fuller. The expected difference between the SLTF and MTPT drum scheduling disciplines (digest edition). Technical Report STAN-CS-73-352 (CSL-TR-28, AD761176), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-73-352.html.

Fuller:1972:RAM

[341] Samuel H. Fuller. Random arrivals and MTPT disk scheduling disciplines. Technical Report STAN-CS-73-353 (CSL-TR-29, AD761185), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1972. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
73-353.html.

Muntz:1972:OCM

[342] Richard R. Muntz and Forest Baskett, III. Open, closed, and mixed networks of queues with different classes of customers. Technical Report STAN-CS-73-360 (CSL-TR-33, AD764014), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1972. iv + 35

pp. URL http://i.stanford.edu/TR/CS-TR-73-360.html.

Newey:1973:ATI

[343] Malcolm C. Newey. Axioms and theorems for integers, lists and finite sets in LCF. Technical Report STAN-CS-73-330 (AIM-184, AD758651), Stanford University, Department of Computer Science, Stanford, CA,

USA, January 1973. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
73-330.html.

Collins:1973:CTE

[344] George E. Collins. The computing time of the Euclidian algorithm. Technical Report STAN-CS-73-331 (AIM-187, AD757364), Stanford University, Department of Computer Science, Stanford, CA, USA, January

1973. 17 pp. URL http://i.stanford.edu/TR/CS-TR-73-331.html.

Milner:1973:ML

[345] Robin Milner. Models of LCF. Technical Report STAN-CS-73-332 (AIM-186, AD758645), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1973. 17 pp. URL http://i.stanford.

edu/TR/CS-TR-73-332.html.

Chandra:1973:PPF

[346] Ashok K. Chandra and Zohar Manna. On the power of programming features. Technical Report STAN-CS-73-333 (AIM-185, AD757367), Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1973. 29 pp. URL http://i.stanford.edu/TR/CS-TR-73-
333.html.

Malcolm:1973:UUR

[347] Michael A. Malcolm and Cleve B. Moler. URAND, a universal random number generator. Technical Report STAN-CS-73-334 (AD757366), Stanford University, Department of Computer Science, Stanford, CA,

USA, January 1973. i + 6 pp. URL http://i.stanford.edu/TR/CS-
TR-73-334.html.

Golub:1973:CSDa

[348] Gene H. Golub and Eugene Seneta. Computation of the stationary distribution of an infinite Markov matrix. Technical Report STAN-CS-73-335 (SU326 P30-24), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1973. 12 pp. URL http://i.stanford.

edu/TR/CS-TR-73-335.html.

Chandra:1973:PAP

[349] Ashok K. Chandra. On the properties and applications of program schemas. Technical Report STAN-CS-73-336 (AIM-188, AD758646), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1973. viii + 231 pp. URL https://apps.dtic.mil/sti/

tr/pdf/AD0758646.pdf.

Gips:1973:AS

[350] James Gips and George Stiny. Aesthetics systems. Technical Report STAN-CS-73-337 (AIM-189, PB218682), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1973. 22 pp.

URL http://i.stanford.edu/TR/CS-TR-73-337.html.

Klarner:1973:FBT

[351] David A. Klarner. A finite basis theorem revisited. Technical Report STAN-CS-73-338 (AD759713), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1973. 10 pp. URL http:/

/i.stanford.edu/TR/CS-TR-73-338.html.

Dent:1973:CLI

[352] Warren T. Dent and Gene H. Golub. Computation of the limited information maximum likelihood estimator. Technical Report STAN-CS-73-339 (SU326 P30-25), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1973. ii + 27 pp. URL http://

i.stanford.edu/TR/CS-TR-73-339.html.

Newey:1973:NPI

[353] Malcolm C. Newey. Notes on a problem involving permutations as subsequences. Technical Report STAN-CS-73-340 (AIM-190, AD759714), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1973. 20 pp. URL http://i.stanford.edu/TR/CS-TR-
73-340.html.

Katz:1973:HAP

[354] Shmuel M. Katz and Zohar Manna. A heuristic approach to program verification. Technical Report STAN-CS-73-341 (AIM-191, AD764272), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1973. 40 pp. URL http://i.stanford.edu/TR/CS-TR-
73-341.html.

Knuth:1973:MP

[355] Donald E. Knuth. Matroid partitioning. Technical Report STAN-CS-73-342 (AD759715), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1, 1973. 12 pp. URL http://i.stanford.

edu/TR/CS-TR-73-342.html;   http://www-db.stanford.edu/TR/CS-
TR-73-342.html;   http://www.ncstrl.org:8900/ncstrl/servlet/
search?formname=detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-
TR-73-342.

Levine:1973:CBA

[356] David R. Levine. Computer-based analytic grading for German grammar instruction. Technical Report STAN-CS-73-343, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1973. 220 pp.

Schank:1973:FPA

[357] Roger C. Schank. The fourteen primitive actions and their inferences. Technical Report STAN-CS-73-344 (AIM-183, AD759716), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1973. 71 pp. URL http://i.stanford.edu/TR/CS-TR-73-344.html.

Collins:1973:MRS

[358] George E. Collins and Ellis Horowitz. The minimum root separation of a polynomial. Technical Report STAN-CS-73-345 (AIM-192), Stanford University, Department of Computer Science, Stanford, CA, USA, April

1973. 71 pp. URL http://i.stanford.edu/TR/CS-TR-73-345.html.

Colby:1973:RCB

[359] Kenneth Mark Colby. The rational for computer based treatment of language difficulties in nonspeaking autistic children. Technical Report STAN-CS-73-346 (AIM-193, AD759717), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1973. 8 pp. URL

http://i.stanford.edu/TR/CS-TR-73-346.html.

Colby:1973:MAE

[360] Kenneth Mark Colby and Franklin Dennis Hilf. Multidimensional analysis in evaluating a simulation of paranoid thought. Technical Report STAN-CS-73-347 (AIM-194, PB221170/4), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-73-347.html.

Colby:1973:MDA

[361] Kenneth M. Colby and Franklin Dennis Hilf. Multi dimensional analysis in evaluating a simulation of paranoid thought processes. Technical Report STAN-CS-73-347 (AIM-194, PB221170/4), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1973. 10 pp.

Pereyra:1973:HOF

[362] Victor Pereyra. High order finite difference solution of differential equations. Technical Report STAN-CS-73-348 (SU326 P30-26, PB222513), Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1973. 86 pp. URL http://i.stanford.edu/TR/CS-TR-73-
348.html.

Blum:1973:TPS

[363] Manual Blum, Robert W. Floyd, Vaughan R. Pratt, Ronald L. Rivest, and Robert Endre Tarjan. Two papers on the selection problem: Time Bounds for Selection [by Manual Blum, Robert W. Floyd, Vaughan Pratt, Ronald L. Rivest, and Robert E. Tarjan] and Expected Time Bounds for Selection [by Robert W. Floyd and Ronald L. Rivest]. Technical Report STAN-CS-73-349 (PB221115), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1973. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-73-349.html.

Blum:1973:TBS

[364] Manuel Blum, Robert Floyd, Vaughn Pratt, Ronald Rivest, and Robert Tarjan. Time bounds for selection. Technical Report STAN-CS-73-349 (PB221115) [part 1], Stanford University, Department of Computer Science, Stanford, CA, USA, April 1973. 51 pp.

Floyd:1973:ETB

[365] Robert Floyd and Ronald Rivest. Expected time bounds for selection. Technical Report STAN-CS-73-349 (PB221115) [part 2], Stanford University, Department of Computer Science, Stanford, CA, USA, April 1973. 51 pp.

Kaufman:1973:AOA

[366] Marc T. Kaufman. An almost-optimal algorithm for the assembly line scheduling problem. Technical Report STAN-CS-73-350 (CSL-TR-53, AD761177, SU-SEL-73-009), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1973. 26 pp. URL http://

i.stanford.edu/TR/CS-TR-73-350.html.

Klarner:1973:NSC

[367] David A. Klarner. The number of SDR’s in certain regular systems. Technical Report STAN-CS-73-354 (PB221165/4), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1973. 7 pp.

URL http://i.stanford.edu/TR/CS-TR-73-354.html.

Price:1973:ACP

[368] Thomas G. Price. An analysis of central processor scheduling in multiprogrammed computer systems (digest edition). Technical Report STAN-CS-73-355 (CSL-TR-57, AD764598, SU-SEL-73-013), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1973.

8 pp. URL http://i.stanford.edu/TR/CS-TR-73-355.html.

Smith:1973:M

[369] David Canfield Smith and Horace J. Enea. MLISP2. Technical Report STAN-CS-73-356 (AIM-195, PB222164), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. 92 pp. URL

http://i.stanford.edu/TR/CS-TR-73-356.html.

Goldman:1973:CBS

[370] Neil M. Goldman and Christopher K. Riesbeck. A conceptually based sentence paraphraser. Technical Report STAN-CS-73-357 (AIM-196, AD762471), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. 88 pp. URL http://i.stanford.edu/

TR/CS-TR-73-357.html.

Schank:1973:ICU

[371] Roger C. Schank and Charles J. Rieger III. Inference and the computer understanding of natural language. Technical Report STAN-CS-73-358 (AIM-197, AD762470), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. 40 pp. URL http://i.

stanford.edu/TR/CS-TR-73-358.html.

Stone:1973:NCP

[372] Harold S. Stone. A note on a combinatorial problem of Burnett and Coffman. Technical Report STAN-CS-73-359 (CSL-TN-25, PB222064), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1973. iv + 7 pp.

Brown:1973:ACG

[373] Harold Brown and Larry M. Masinter. An algorithm for the construction of the graphs of organic molecules. Technical Report STAN-CS-73-361 (Serra, AD764273), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-73-361.html.

Kaufman:1973:ASG

[374] Linda C. Kaufman. The LZ algorithm to solve the generalized eigenvalue problem. Technical Report STAN-CS-73-363 (Serra, PB222099), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. 101 pp.

Thosar:1973:EPD

[375] Ravindra B. Thosar. Estimation of probability density using signature tables for applications to pattern recognition. Technical Report STAN-CS-73-364 (AIM-198, AD763611), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. 36 pp. URL http:/

/i.stanford.edu/TR/CS-TR-73-364.html.

Igarashi:1973:APV

[376] Shigeru Igarashi, Ralph L. London, and David C. Luckham. Automatic program verification I: a logical basis and its implementation. Technical Report STAN-CS-73-365 (AIM-200, AD767331), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. iv + 53

pp. URL http://i.stanford.edu/TR/CS-TR-73-365.html;   https:
//apps.dtic.mil/sti/tr/pdf/AD0767331.pdf.

Grape:1973:MBI

[377] Gunnar Rutger Grape. Model bases (intermediate-level computer vision). Technical Report STAN-CS-73-366 (AIM-201, AD763673), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. 256 pp.

Amble:1973:OHT

[378] Ole Amble and Donald E. Knuth. Ordered hash tables. Technical Report STAN-CS-73-367 (AD763601), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. 34 pp.

Schank:1973:GLT

[379] Roger C. Schank and Yorick A. Wilks. The goals of linguistic theory revisited. Technical Report STAN-CS-73-368 (AIM-202, AD764396), Stanford University, Department of Computer Science, Stanford, CA, USA,

May 1973. 44 pp. URL http://i.stanford.edu/TR/CS-TR-73-368.
html.

Schank:1973:DCS

[380] Roger C. Schank. The development of conceptual structures in children. Technical Report STAN-CS-73-369 (AIM-203, AD764274), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973.

?? pp. URL http://i.stanford.edu/TR/CS-TR-73-369.html.

Sridharan:1973:HPD

[381] N. S. Sridharan, G. Gelernter, A. J. Hart, W. F. Fowler, and H. J. Shue. A heuristic program to discover syntheses for complex organic molecule. Technical Report STAN-CS-73-370 (AIM-205, AD764288), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1973. 30 pp.

Knuth:1973:RSP

[382] Donald E. Knuth. A review of “Structured Programming”. Technical Report STAN-CS-73-371 (AD223572/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1,

1973. 25 pp. URL http://i.stanford.edu/TR/CS-TR-73-371.html;
 http://www-db.stanford.edu/TR/CS-TR-73-371.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-73-371.

Malcolm:1973:NSF

[383] Michael A. Malcolm. Nonlinear spline functions. Technical Report STAN-CS-73-372 (AD767970), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1973. 60 pp. URL https://

apps.dtic.mil/sti/tr/pdf/AD0767970.pdf.

VanLehn:1973:SUM

[384] Kurt A. VanLehn. SAIL user manual. Technical Report STAN-CS-73-373 (AIM-204, AD765353/BWC), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. 124 pp. URL http:

//i.stanford.edu/TR/CS-TR-73-373.html.

Malcolm:1973:MIA

[385] Michael A. Malcolm. A machine independent ALGOL procedure for accurate floating-point summation. Technical Report STAN-CS-73-374 (AD-764275), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1973. iii + 6 pp. URL https://apps.dtic.

mil/sti/tr/pdf/AD0764275.pdf.

Fischer:1973:FTM

[386] D. Fischer, G. H. Golub, O. Hald, C. Levin, and O. Widlund. On Fourier–Toeplitz methods for separable elliptic problems. Technical Report STAN-CS-73-375 (SU-326, P30-27), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. 30 pp.

Meinardus:1973:LEE

[387] Günter Meinardus and Gerald D. Taylor. Lower estimates for the error of best uniform approximation. Technical Report STAN-CS-73-376 (SU326 P30-28), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. 20 pp. URL http://i.stanford.edu/TR/CS-

TR-73-376.html;   https://en.wikipedia.org/wiki/Remez_algorithm.

Wilks:1973:PS

[388] Yorick Wilks. Preference semantics. Technical Report STAN-CS-73-377 (AIM-206, AD764652), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1973. 20 pp.

Moorer:1973:OCM

[389] James Anderson Moorer. The optimum comb method of pitch period analysis of continuous digitized speech. Technical Report STAN-CS-73-378 (AIM-207, AD767333), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. 22 pp. URL http://i.

stanford.edu/TR/CS-TR-73-378.html.

Moorer:1973:HFT

[390] James Anderson Moorer. The heterodyne filter as a tool for analysis of transient waveforms. Technical Report STAN-CS-73-379 (AIM-208, AD767334), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. 27 pp. URL http://i.stanford.edu/

TR/CS-TR-73-379.html.

Yakimovsky:1973:SAU

[391] Yoram Yakimovsky. Scene analysis using a semantic base for region growing. Technical Report STAN-CS-73-380 (AIM-209, AD767695/OWC), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1973. 120 pp.

Sridharan:1973:CGV

[392] N. S. Sridharan. Computer generation of vertex-graphs. Technical Report STAN-CS-73-381 (AD767694), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. 18 pp. See also [416].

Manna:1973:AAT

[393] Zohar Manna and Amir Pnueli. Axiomatic approach to total correctness of programs. Technical Report STAN-CS-73-382 (AIM-210, AD767335), Stanford University, Department of Computer Science, Stanford, CA,

USA, July 1973. ?? pp. URL http://i.stanford.edu/TR/CS-TR-73-
382.html.

Wilks:1973:NLI

[394] Yorick A. Wilks. Natural language inference. Technical Report STAN-CS-73-383 (AIM-211, AD769673), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1973. ii + 24 pp. URL

http://i.stanford.edu/TR/CS-TR-73-383.html.

Herskovits:1973:GFS

[395] Annette Herskovits. The generation of French from a semantic representation. Technical Report STAN-CS-73-384 (AIM-212, AD769379), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1973. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
73-384.html.

Thosar:1973:RCS

[396] Ravindra B. Thosar. Recognition of continuous speech: segmentation and classification using signature table adaptation. Technical Report STAN-CS-73-385 (AIM-213), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1973. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-73-385.html.

Perkins:1973:CFV

[397] W. A. Perkins and Thomas O. Binford. A corner finder for visual feedback. Technical Report STAN-CS-73-386 (AIM-214, AD767332), Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1973. ?? pp. URL http://i.stanford.edu/TR/CS-TR-73-386.
html.

Buchanan:1973:ABC

[398] Bruce G. Buchanan and Natesa S. Sridharan. Analysis of behavior of chemical molecules: rule formation on non-homogeneous classes of objects. Technical Report STAN-CS-73-387 (AIM-215, AD769380), Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1973. ?? pp. URL http://i.stanford.edu/TR/CS-TR-73-387.
html.

Swanson:1973:IPM

[399] Roger C. Swanson. Interconnections for parallel memories to unscramble p-ordered vectors. Technical Report STAN-CS-73-388 (CSL-TR-74, PB226044/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1973. ?? pp. URL http://i.stanford.edu/

TR/CS-TR-73-388.html.

Masinter:1973:AAI

[400] L. Masinter, N. S. Sridharan, J. Lederberg, and D. H. Smit. Applications of artificial intelligence for chemical inference XII. Exhaustive generation of cyclic and acyclic isomers. Technical Report STAN-CS-73-389 (AIM-216, AD771299), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1973. iii + 58 pp.

Gips:1973:CIT

[401] James Gips. A construction for the inverse of a Turing machine. Technical Report STAN-CS-73-390, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1973. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-73-390.html.

Sridharan:1973:SST

[402] Natesa S. Sridharan. Search strategies for the task of organic chemical synthesis. Technical Report STAN-CS-73-391 (AIM-217, AD770610), Stanford University, Department of Computer Science, Stanford, CA,

USA, October 1973. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
73-391.html.

Knuth:1973:SSE

[403] Donald E. Knuth. Sorting and Searching — errata and addenda. Technical Report STAN-CS-73-392, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1, 1973.

?? pp. URL http://i.stanford.edu/TR/CS-TR-73-392.html;

 http://www-db.stanford.edu/TR/CS-TR-73-392.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-73-392.

Vuillemin:1973:PTR

[404] Jean Étienne Vuillemin. Proof techniques for recursive programs. Technical Report STAN-CS-73-393 (AIM-218, AD772063/4WC), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1973. iv + 97 pp.

Hoare:1973:PPA

[405] C. A. R. Hoare. Parallel programming: an axiomatic approach. Technical Report STAN-CS-73-394 (AIM-219, AD769674), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1973. 97

pp. URL http://i.stanford.edu/TR/CS-TR-73-394.html.

Staff:1973:BCS

[406] Staff. Bibliography of computer science reports. Technical Report STAN-CS-73-395, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1973. 48 pp.

Bolles:1973:USF

[407] Robert C. Bolles and Richard P. Paul. The use of sensory feedback in a programmable assembly system. Technical Report STAN-CS-73-396 (AIM-220, AD772064), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1973. 24 pp. URL http://i.

stanford.edu/TR/CS-TR-73-396.html.

Henrici:1973:CCA

[408] Peter Henrici. Computational complex analysis. Technical Report STAN-CS-73-397 (SU326, P30-28A), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1973. 14 pp.

Baumgart:1973:ICC

[409] Bruce G. Baumgart. Image contouring and comparing. Technical Report STAN-CS-73-398 (AIM-199, AD771300) (also mistakenly assigned STAN-CS-73-362), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1973. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-73-398.html.

Paige:1973:SSI

[410] C. C. Paige and M. A. Saunders. Solution of sparse indefinite systems of equations and least squares problems. Technical Report STAN-CS-73-399 (SU326, P30-29), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1973. i + 49 pp.

Hoare:1973:RDS

[411] C. A. R. Hoare. Recursive data structures. Technical Report STAN-CS-73-400 (AIM-223, AD-772509), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1973. iii + 32 pp. URL

https://apps.dtic.mil/sti/tr/pdf/AD0772509.pdf.

Hoare:1973:MOS

[412] C. A. R. Hoare. Monitors: an operating system structuring concept. Technical Report STAN-CS-73-401 (PB226691/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1973.

i + 25 pp. URL http://i.stanford.edu/TR/CS-TR-73-401.html.

Herriot:1974:APC

[413] John G. Herriot and Christian H. Reinsch. Algol 60 procedures for the calculation of interpolating natural quintic spline functions. Technical Report STAN-CS-74-402 (PB229616/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1974. iv + 40 pp.

Hoare:1973:HPL

[414] C. A. R. Hoare. Hints on programming language design. Technical Report STAN-CS-73-403 (AIM-224, AD773391), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1973. ??

pp. URL http://i.stanford.edu/TR/CS-TR-73-403.html.

Earnest:1973:FRF

[415] Lester D. Earnest, John McCarthy, Edward A. Feigenbaum, and Joshua Lederberg. Final report: the first ten years of artificial intelligence research at Stanford. Technical Report STAN-CS-74-409 (AIM-228, AD776233), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-74-409.html.

Sridharan:1974:CQT

[416] Natesa S. Sridharan. A catalog of quadri/trivalent graphs. Technical Report STAN-CS-74-404 (AD775452), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1974. 48 pp. URL

http://i.stanford.edu/TR/CS-TR-74-404.html.

Davis:1974:SCS

[417] Randall Davis and Margaret H. Wright. Stanford Computer Science Department research report. Technical Report STAN-CS-74-405, Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1974. ?? pp. URL http://i.stanford.edu/TR/CS-TR-74-
405.html.

Perkins:1974:MMR

[418] W. A. Perkins. Memory model for a robot. Technical Report STAN-CS-74-406 (AIM-225, AD775645), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1974. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-74-406.html.

Wright:1974:F

[419] F. H. G. Wright, II and Ralph E. Gorin. FAIL. Technical Report STAN-CS-74-407 (AIM-226, AD778310), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1974. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-74-407.html.

Thomas:1974:IPA

[420] Arthur J. Thomas and Thomas O. Binford. Information processing analysis of visual perception: a review. Technical Report STAN-CS-74-408 (AIM-227, AD-A003483), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1974. i + 60 pp.

Peterson:1974:MPS

[421] James L. Peterson. Modelling of parallel systems. Technical Report STAN-CS-74-410 (CSL-TR-46, PB231926/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1974. 241 pp.

Anderson:1974:ALL

[422] D. Bruce Anderson, Thomas O. Binford, Arthur J. Thomas, Richard W. Weyhrauch, and Yorick A. Wilks. After Leibniz …: discussions on philosophy and artificial intelligence. Technical Report STAN-CS-74-411 (AIM-229), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1974. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-74-411.html.

Swinehart:1974:CMP

[423] Daniel Carl Swinehart. COPILOT: a multiple process approach to interactive programming systems. Technical Report STAN-CS-74-412 (AIM-230, AD786721), Stanford University, Department of Computer Science,

Stanford, CA, USA, July 1974. URL https://apps.dtic.mil/sti/tr/
pdf/AD0786721.pdf.

Gips:1974:SGT

[424] James Gips. Shape grammars and their uses. Technical Report STAN-CS-74-413 (AIM-231, AD-A001814), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1974. xi + 233 pp.

Baumgart:1974:GGE

[425] Bruce G. Baumgart. GEOMED — a geometric editor. Technical Report STAN-CS-74-414 (AIM-232, AD780452), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-74-414.html.

Rivest:1974:AAR

[426] Ronald Linn Rivest. Analysis of associative retrieval algorithms. Technical Report STAN-CS-74-415 (PB233065/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. viii +

102 pp. URL https://people.csail.mit.edu/rivest/pubs/rivest-
1974-PhD.pdf.

Knuth:1974:SPG

[427] Donald E. Knuth. Structured programming with Go To statements. Technical Report STAN-CS-74-416 (PB233507/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1974. 100 pp.

Sites:1974:STP

[428] Richard L. Sites. Some thoughts on proving that programs terminate cleanly. Technical Report STAN-CS-74-417 (PB233045/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 68 pp.

Sites:1974:PCP

[429] Richard L. Sites. Proving that computer programs terminate cleanly. Technical Report STAN-CS-74-418 (PB233045/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. vi +

139 pp. URL http://i.stanford.edu/TR/CS-TR-74-417.html.

Rieger:1974:CMT

[430] Charles Rieger III. Conceptual memory: a theory and computer program for processing the meaning content of natural language utterances.

Technical Report STAN-CS-74-419 (AIM-233, AD-A000086), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. xv + 393 pp.

Wakerly:1973:PSC

[431] John F. Wakerly. Partially self-checking ciruits and their use in performing logical operations. Technical Report STAN-CS-74-420 (CSL-TR-50, PB232543/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1973. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-74-420.html.

Wakerly:1974:LCE

[432] John Wakerly. Low-cost error detection techniques for small computers. Technical Report STAN-CS-74-421 (CSL-TR-51, PB232356/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 232 pp.

Stone:1974:PTD

[433] Harold Stone. Parallel tri-diagonal equation solvers. Technical Report STAN-CS-74-422 (CSL-TR-79, NASA-TM-62,370), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 42 pp.

Rao:1974:ARA

[434] Gururaj S. Rao. Asymptotic representation of the average number of active modules in an n-way interleaved memory. Technical Report STAN-CS-74-423 (CSL-TN-41, PB232860/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1974. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-74-423.html.

Schlumberger:1974:LCN

[435] Maurice Schlumberger. Logarithmic communications networks. Technical Report STAN-CS-74-424 (CSL-TR-80, PB232602/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 38 pp.

Schlumberger:1974:VBC

[436] Maurice Schlumberger. Vulnerability of de Bruijn communications networks. Technical Report STAN-CS-74-425 (CSL-TR-81, PB232598/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1974. iii + 65 pp.

Schlumberger:1974:QEL

[437] Maurice Schlumberger. Queueing equal length messages in a logarithmic network. Technical Report STAN-CS-74-426 (CSL-TR-82, PB232597), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 75 pp.

Lang:1974:PPS

[438] Tomas Lang. Performing the perfect schuffle in an array computer. Technical Report STAN-CS-74-427 (CSL-TN-36, PB232624/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 18 pp.

Lang:1974:IBP

[439] Tomas Lang. Interconnections between processors and memory modules using the schuffle–exchange network. Technical Report STAN-CS-74-428 (CSL-TR-76, PB232633/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 32 pp.

Orcutt:1974:EDR

[440] Samuel E. Orcutt. Efficient data routing schemes for ILLIAC IVtype computers. Technical Report STAN-CS-74-429 (CSL-TR-70, PB232623/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. + 26 pp.

Orcutt:1974:NPC

[441] Samuel E. Orcutt. A novel parallel computer architecture and some applications. Technical Report STAN-CS-74-430 (CSL-TR-71, PB234513/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. vi + 41 pp.

Colby:1974:PMR

[442] Kenneth Mark Colby, Roger C. Parkison, and William S. Faught. Pattern-matching rules for the recognition of natural language dialogue expressions. Technical Report STAN-CS-74-431 (AIM-234), Stanford University, Department of Computer Science, Stanford, CA, USA, June

1974. 23 pp. URL http://i.stanford.edu/TR/CS-TR-74-431.html.

Weyhrauch:1974:FPC

[443] Richard W. Weyhrauch and Arthur J. Thomas. FOL: a proof checker for first-order logic. Technical Report STAN-CS-74-432 (AIM-235, AD-A006898), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. vi + 56 pp.

Buchanan:1974:ACP

[444] Jack R. Buchanan and David C. Luckham. On automating the construction of programs. Technical Report STAN-CS-74-433 (AIM-236, AD784513), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 65 pp. URL http://i.stanford.edu/

TR/CS-TR-74-433.html.

Ruhe:1974:ASN

[445] Axel Ruhe and Per Åke Wedin. Algorithms for separable nonlinear least squares problems. Technical Report STAN-CS-74-434 (SU326, P30-31), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1974. 50 pp.

Price:1974:BCS

[446] Thomas G. Price. Balanced computer systems. Technical Report STAN-CS-74-435 (CSL-TR-88, A001-071, SU-SEL-74-037), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1974. 56

pp. URL http://i.stanford.edu/TR/CS-TR-74-435.html.

Wilks:1974:NLU

[447] Yorick A. Wilks. Natural language understanding systems within the AI paradigm: a survey and some comparisons. Technical Report STAN-CS-74-436 (AIM-237, AD-A012477), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1974. ii + 40 pp. URL

http://i.stanford.edu/TR/CS-TR-74-436.html.

Riesbeck:1974:CUA

[448] Christopher Kevin Riesbeck. Computational understanding: Analysis of sentences and context. Technical Report STAN-CS-74-437 (AIM-238, AD-A005040), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. vi + 242 pp.

Hanna:1974:CMA

[449] Marsha Jo Hanna. Computer matching of areas in stereo images. Technical Report STAN-CS-74-438 (AIM-239, AD786720), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1974. v + 94 pp.

Cottle:1974:SLS

[450] Richard W. Cottle, Gene H. Golub, and Richard S. Sacher. On the solution of large, structured linear complementarity problems: III. Technical Report STAN-CS-74-439 (OR-74-7, SU326 P30-32), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1974. 87

pp. URL http://i.stanford.edu/TR/CS-TR-74-439.html.

Morris:1974:FPM

[451] Donald E. Knuth, James H. Morris, Jr., and Vaughn R. Pratt. Fast pattern matching in strings. Technical Report STAN-CS-74-440 (PB237360/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1974. iii + 32 pp.

Knuth:1974:AAB

[452] Donald E. Knuth and Ronald W. Moore. An analysis of alpha–beta pruning. Technical Report STAN-CS-74-441 (AD-A000284), Stanford University, Department of Computer Science, Stanford, CA, USA, August

1974. ii + 64 pp. URL https://apps.dtic.mil/sti/citations/tr/
ADA000284;   https://kodu.ut.ee/~ahto/eio/2011.07.11/ab.pdf.

Knuth:1974:EEB

[453] Donald E. Knuth. Estimating the efficiency of backtrack programs. Technical Report STAN-CS-74-442 (AD-A004208), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1974.

i + 29 pp. URL http://i.stanford.edu/TR/CS-TR-74-442.html;
 http://www-db.stanford.edu/TR/CS-TR-74-442.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-74-442. Dedicated to

Derrich H. Lehmer on his 70th birthday, February 32, 1975.

Brotz:1974:EHP

[454] Douglas K. Brotz. Embedding heuristic problem solving methods in a mechanical theorem prover. Technical Report STAN-CS-74-443 (PB-236471/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1974. 107 pp.

Green:1974:PRP

[455] C. Cordell Green, Richard J. Waldinger, David R. Barstow, Robert A. Elschlager, Douglas B. Lenat, Brian P. McCune, David E. Shaw, and Louis I. Steinberg. Progress report on program-understanding systems. Technical Report STAN-CS-74-444 (AIM-240, AD787035), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1974.

50 pp. URL http://i.stanford.edu/TR/CS-TR-74-444.html.

Friedman:1974:REA

[456] J. H. Friedman, F. Baskett, and L. J. Shustek. A relatively efficient algorithm for finding nearest neighbors. Technical Report STAN-CS-74-445 (SLACP-1448), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. 21 pp.

Aiello:1974:LIL

[457] Luigia Aiello and Richard W. Weyhrauch. LCFsmall: an implementation of LCF. Technical Report STAN-CS-74-446 (AIM-241, AD786723), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1974. iv + 45 pp. URL http://i.stanford.edu/TR/CS-
TR-74-446.html.

Aiello:1974:SPL

[458] Luigia Aiello, Mario Aiello, and Richard W. Weyhrauch. The semantics of PASCAL in LCF. Technical Report STAN-CS-74-447 (AIM-221, AD787631), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1974. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-74-447.html.

Goldfarb:1974:MFO

[459] D. Goldfarb. Matrix factorizations in optimization of nonlinear functions subject to linear constraints. Technical Report STAN-CS-74-448 (SU326, P30-33), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. 45 pp.

Smith:1974:PAC

[460] Alan Jay Smith. Performance analysis of computer systems components. Technical Report STAN-CS-74-449 (CSL-TR-89, AD785027), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. 323 pp. URL https://apps.dtic.mil/sti/

html/tr/AD0785027/index.html.

Baskett:1974:IMC

[461] Forest Baskett and Alan Jay Smith. Interference in multiprocessor computer systems with interleaved memory. Technical Report STAN-CS-74-450 (CSL-TR-90, AD787008), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1974. iii + 41 pp.

Smith:1974:MWS

[462] Alan Jay Smith. A modified working set paging algorithm. Technical Report STAN-CS-74-451 (CSL-TR-91, AD786999), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1974. 43 pp.

Low:1974:ACC

[463] James Richard Low. Automatic coding: Choice of data structures. Technical Report STAN-CS-74-452 (AIM-242, AD-A000-500), Stanford University, Department of Computer Science, Stanford, CA, USA, Au-

gust 1974. v + 103 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA000500.pdf.

Knuth:1974:RM

[464] Donald E. Knuth. Random matroids. Technical Report STAN-CS-74-453 (AD-A000034), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. 30 pp. URL https://apps.dtic.

mil/sti/html/tr/ADA000084/.

Jennings:1974:CAS

[465] L. S. Jennings. A computational approach to simultaneous estimation. Technical Report STAN-CS-74-454 (SU326, P30-35), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. 15 pp.

Tarjan:1974:EDS

[466] Robert Endre Tarjan. Edge-disjoint spanning trees, dominators, and depth-first search. Technical Report STAN-CS-74-455 (AD-A000083), Stanford University, Department of Computer Science, Stanford, CA,

USA, September 1974. 40 pp. URL http://i.stanford.edu/TR/CS-
TR-74-455.html.

Finkel:1974:APS

[467] Raphael A. Finkel, Russell H. Taylor, Robert C. Bolles, Richard P. Paul, and Jerome A. Feldman. AL, a programming system for automation. Technical Report STAN-CS-74-456 (AIM-243, AD-A003815), Stanford University, Department of Computer Science, Stanford, CA, USA,

November 1974. ?? pp. URL http://i.stanford.edu/TR/CS-TR-74-
456.html.

Colby:1974:TCP

[468] Kenneth Mark Colby. Ten criticisms of PARRY. Technical Report STAN-CS-74-457 (AIM-244), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. ?? pp. URL http://

i.stanford.edu/TR/CS-TR-74-457.html.

Buchanan:1974:SAP

[469] Jack R. Buchanan. A study in automatic programming. Technical Report STAN-CS-74-458 (AIM-245, AD784816), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. vii + 148 pp.

Winograd:1974:FLA

[470] Terry A. Winograd. Five lectures on artificial intelligence. Technical Report STAN-CS-74-459 (AIM-246, AD-A000085), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1974. viii + 95 pp.

Porter:1974:RIP

[471] Thomas Porter and Istvan Simon. Random insertion into a priority queue structure. Technical Report STAN-CS-74-460 (PB238148/AS), Stanford University, Department of Computer Science, Stanford, CA,

USA, October 1974. 25 pp. URL http://i.stanford.edu/TR/CS-TR-
74-460.html.

Goldman:1974:CGN

[472] Neil Murray Goldman. Computer generation of natural language from a deep conceptual base. Technical Report STAN-CS-74-461 (AIM-247, AD-A005041), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1974. ix + 309 pp. URL https://apps.

dtic.mil/sti/tr/pdf/ADA005041.pdf.

Baumgart:1974:GMC

[473] Bruce Guenther Baumgart. Geometric modeling for computer vision. Technical Report STAN-CS-74-463 (AIM-249, AD-A002261), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1974. v + 136 pp.

Nevatia:1974:SDC

[474] Ramakant Nevatia. Structured descriptions of complex curved objects for recognition and visual memory. Technical Report STAN-CS-74-464 (AIM-250, AD-A003486), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1974. v + 120 pp.

Shortliffe:1974:MRB

[475] E. H. Shortliffe. MYCIN: a rule-based computer program for advising physicians regarding antimicrobial therapy selection. Technical Report STAN-CS-74-465 (AIM-251, AD-A001373), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1974. 409 pp.

Earnest:1974:RRA

[476] Lester D. Earnest, John McCarthy, Edward A. Feigenbaum, Joshua Lederberg, and Vinton G. Cerf. Recent research in artificial intelligence, heuristic programming, and network protocols. Technical Report STAN-CS-74-466 (AIM-252, AD-A002246), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1974. 79 pp. URL http:/

/i.stanford.edu/TR/CS-TR-74-466.html.

Aiello:1974:CPM

[477] Mario Aiello and Richard W. Weyhrauch. Checking proofs in the metamathematics of first order logic. Technical Report STAN-CS-74-467 (AIM-222, AD-A007562), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1974. viii + 51 pp. URL http://

i.stanford.edu/TR/CS-TR-74-467.html.

Krogdahl:1974:CBS

[478] Stein Krogdahl. A combinatorial base for some optimal matroid intersection algorithms. Technical Report STAN-CS-74-468 (AD-A003832), Stanford University, Department of Computer Science, Stanford, CA,

USA, November 1974. 55 pp. URL http://i.stanford.edu/TR/CS-
TR-74-468.html.

Brown:1974:MSE

[479] Harold Brown. Molecular structure elucidation III. Technical Report STAN-CS-74-469, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1974. 38 pp. URL http://i.stanford.

edu/TR/CS-TR-74-469.html.

TrabbPardo:1974:SSM

[480] Luis Trabb Pardo. Stable sorting and merging with optimal space and time bounds. Technical Report STAN-CS-74-470, Stanford University, Department of Computer Science, Stanford, CA, USA, De-

cember 1974. iv + 74 pp. URL http://i.stanford.edu/pub/cstr/
reports/cs/tr/74/470/CS-TR-74-470.pdf;   http://i.stanford.
edu/TR/CS-TR-74-470.html.

Faught:1974:IIA

[481] William S. Faught, Kenneth Mark Colby, and Roger C. Parkison. The interaction of inferences, affects, and intentions, in a model of paranoia. Technical Report STAN-CS-74-471 (AIM-253, AD-A003487), Stanford University, Department of Computer Science, Stanford, CA, USA, De-

cember 1974. 38 pp. URL http://i.stanford.edu/TR/CS-TR-74-471.
html.

Quam:1974:SAP

[482] Lynn H. Quam and Marsha Jo Hannah. Stanford automatic photogrammetry research. Technical Report STAN-CS-74-472 (AIM-254, AD-A005407), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1974. 15 pp. URL http://i.stanford.

edu/TR/CS-TR-74-472.html.

Suzuki:1974:APV

[483] Norihisa Suzuki. Automatic program verification II: verifying programs by algebraic and logical reduction. Technical Report STAN-CS-74-473 (AIM-255, AD-A005412), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1974. 28 pp. URL http://i.

stanford.edu/TR/CS-TR-74-473.html.

vonHenke:1974:APV

[484] Friedrich W. von Henke and David C. Luckham. Automatic program verification III: a methodology for verifying programs. Technical Report STAN-CS-74-474 (AIM-256, AD-A007563), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1974. 45 pp.

URL http://i.stanford.edu/TR/CS-TR-74-474.html.

Newey:1975:FSL

[485] M. C. Newey. Formal semantics of LISP with applications to program correctness. Technical Report STAN-CS-75-475 (AIM-257, AD-A005413), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1975. 184 pp.

Pingle:1975:FFD

[486] Karl K. Pingle and Arthur J. Thomas. A fast, feature-driven stereo depth program. Technical Report STAN-CS-74-462 (AIM-248), Stanford University, Department of Computer Science, Stanford, CA, USA, May

1975. ii + 17 pp. URL http://i.stanford.edu/TR/CS-TR-74-462.
html.

Green:1975:HDE

[487] C. Cordell Green and David R. Barstow. A hypothetical dialogue exhibiting a knowledge base for a program-understanding system. Technical Report STAN-CS-75-476 (AIM-258, AD-A006294), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1975. 45

pp. URL http://i.stanford.edu/TR/CS-TR-75-476.html.

Chvatal:1975:LCS

[488] Vaclav Chvátal and David Sankoff. Longest common subsequences of two random sequences. Technical Report STAN-CS-75-477, Stanford University, Department of Computer Science, Stanford, CA, USA, January

1975. ?? pp. URL http://i.stanford.edu/TR/CS-TR-75-477.html.

Golub:1975:ICE

[489] Gene H. Golub and James H. Wilkinson. Ill-conditioned eigensystems and the computation of the Jordan canonical form. Technical Report STAN-CS-75-478 (SU326 P30-36)8, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1975. 66 pp. URL

http://i.stanford.edu/TR/CS-TR-75-478.html. Published as [1741].

Chatelin:1975:EBA

[490] François Chatelin and J. Lemordant. Error bounds in the approximation of eigenvalues of differential and integral operators. Technical Report STAN-CS-75-479 (SU326 P30-38), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1975. 24 pp. URL

http://i.stanford.edu/TR/CS-TR-75-479.html.

Knuth:1975:NGD

[491] Donald E. Knuth. Notes on generalized Dedekind sums. Technical Report STAN-CS-75-480 (AD-A008804), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1975. 45 pp. URL http://matwbn.icm.edu.pl/ksiazki/aa/aa33/
aa3341.pdf;   https://apps.dtic.mil/sti/pdfs/ADA008804.pdf;
 https://apps.dtic.mil/sti/tr/pdf/ADA008804.pdf.

Oliger:1975:HDM

[492] Joseph E. Oliger. Hybrid difference methods for the initial boundaryvalue problem for hyperbolic equations. Technical Report STAN-CS-75-481 (SU326 P30-39), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1975. 31 pp. URL http://i.

stanford.edu/TR/CS-TR-75-481.html.

Friedman:1975:AFB

[493] J. A. Friedman, J. L. Bentley, and R. A. Finkel. An algorithm for finding best matches in logarithmic time. Technical Report STAN-CS-75-482 (SLACP-1549), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1975. 31 pp.

Erdos:1975:PSE

[494] Paul Erdős and Ronald L. Graham. On packing squares with equal squares. Technical Report STAN-CS-75-483 (AD-A011835), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1975. 8 pp. URL http://i.stanford.edu/TR/CS-TR-75-483.html.

Graham:1975:SNI

[495] Ronald L. Graham and Endre Szemerédi. On subgraph number independence in trees. Technical Report STAN-CS-75-484 (AD-A011832), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1975. 18 pp. URL http://i.stanford.edu/TR/CS-TR-
75-484.html.

Erdos:1975:MRI

[496] Paul Erdős and Endre Szemerédi. On multiplicative representations of integers. Technical Report STAN-CS-75-485 (AD-A011834), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1975. 18 pp. URL http://i.stanford.edu/TR/CS-TR-75-485.html.

Bjorck:1975:EMA

[497] Åke Björck and Gene H. Golub. Eigenproblems for matrices associated with periodic boundary conditions. Technical Report STAN-CS-75-486 (SU326 P30-37), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1975. 19 pp. URL http://i.stanford.edu/

TR/CS-TR-75-486.html.

Friedman:1975:VMD

[498] J. H. Friedman. A variable metric decision rule for NonParametric classification. Technical Report STAN-CS-75-487 (SLACP-1573), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1975. 34 pp.

Bollabas:1975:CSC

[499] Bela Bollabas, Paul Erdős, and Endre Szemerédi. On complete subgraphs of r-chromatic graphs. Technical Report STAN-CS-75-488 (AD-A011445), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1975. 16 pp. URL http://i.stanford.edu/TR/

CS-TR-75-488.html.

Szemeredi:1975:RPG

[500] Endre Szemerédi. Regular partitions of graphs. Technical Report STAN-CS-75-489 (AD-A011833), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1975. 8 pp. URL http://

i.stanford.edu/TR/CS-TR-75-489.html.

Gosper:1975:NES

[501] R. William Gosper. Numerical experiments with the spectral test. Technical Report STAN-CS-75-490 (AD-A014429), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. 31 pp.

URL http://i.stanford.edu/TR/CS-TR-75-490.html.

Knott:1975:DBS

[502] G. D. Knott. Deletion in binary storage trees. Technical Report STAN-CS-75-491, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. 93 pp.

Sedgewick:1975:Q

[503] Robert Sedgewick. Quicksort. Technical Report STAN-CS-75-492, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. 352 pp.

Kurki-Suonio:1975:DAT

[504] Reino Kurki-Suonio. Describing automata in terms of languages associated with their peripheral devices. Technical Report STAN-CS-75-493 (PB244421/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. 37 pp. URL http://i.stanford.edu/

TR/CS-TR-75-493.html.

Satterthwaite:1975:SLD

[505] E. H. Satterthwaite, Jr. Source language debugging tools. Technical Report STAN-CS-75-494, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. 345 pp.

Krogdahl:1975:DGB

[506] Stein Krogdahl. The dependence graph for bases in matroids. Technical Report STAN-CS-75-495 (AD-A014424), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. iv + 28 pp.

Underwood:1975:IBL

[507] Richard Ray Underwood. An iterative block Lanczos method for the solution of large sparse symmetric eigenproblems. Technical Report STAN-CS-75-496 (SU326 P30-41), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. v + 133 pp. URL

https://www.proquest.com/pqdtglobal/docview/302727419.

Graham:1975:DMT

[508] Ronald L. Graham and L. Lovász. Distance matrices of trees. Technical Report STAN-CS-75-497 (AD-A016825), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1975. 48 pp.

Samet:1975:APC

[509] Hanan Samet. Automatically proving the correctness of translations involving optimized code. Technical Report STAN-CS-75-498 (AIM-259, AD-A017025), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1975. vi + 214 pp. URL http://i.stanford.

edu/TR/CS-TR-75-498.html.

Smith:1975:PCP

[510] David Canfield Smith. PYGMALION: a creative programming environment. Technical Report STAN-CS-75-499 (AIM-260), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1975. 187 pp.

Kurki-Suonio:1975:TBS

[511] Reino Kurki-Suonio. Towards better structured definitions of programming languages. Technical Report STAN-CS-75-500 (PB246708/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975. i + 29 pp. URL http://i.stanford.edu/TR/

CS-TR-75-500.html.

Pettersen:1975:PES

[512] Odd Pettersen. Procedural events as software interrupts. Technical Report STAN-CS-75-501 (AIM-261, AD-A016810), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1975. ??

pp. URL http://i.stanford.edu/TR/CS-TR-75-501.html.

Pettersen:1975:SCP

[513] Odd Pettersen. Synchronization of concurrent processes. Technical Report STAN-CS-75-502 (AIM-262, AD-A016808), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1975. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-75-502.html.

Pettersen:1975:MPS

[514] Odd Pettersen. The macro processing system STAGE2: transfer of comments to the generated text. Technical Report STAN-CS-75-503 (AIM-263, AD-A016807), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1975. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-75-503.html.

Erdos:1975:SGD

[515] Paul Erdős, Ronald L. Graham, and Endre Szemerédi. On sparse graphs with dense long paths. Technical Report STAN-CS-75-504 (AD-A017370), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975. ?? pp. URL http://i.stanford.edu/

TR/CS-TR-75-504.html.

Chvatal:1975:SLP

[516] Vaclav Chvátal. Some linear programming aspects of combinatorics. Technical Report STAN-CS-75-505 (AD-A017053), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975.

?? pp. URL http://i.stanford.edu/TR/CS-TR-75-505.html.

Gordon:1975:ORD

[517] Michael J. C. Gordon. Operational reasoning and denotational semantics. Technical Report STAN-CS-75-506 (AIM-264, AD-A017176), Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1975. ?? pp. URL http://i.stanford.edu/TR/CS-TR-75-506.
html.

Gordon:1975:TST

[518] Michael J. C. Gordon. Towards a semantic theory of dynamic binding. Technical Report STAN-CS-75-507 (AIM-265), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1975. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-75-507.html.

Eve:1975:CTC

[519] James Eve. On computing the transitive closure of a relation. Technical Report STAN-CS-75-508, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-75-508.html.

Overton:1975:FMI

[520] Michael L. Overton and Andrzej Proskurowski. Finding the maximal incidence matrix of a large graph. Technical Report STAN-CS-75-509 (AD-A017331), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-75-509.html.

Yao:1975:ASA

[521] Andrew C. Yao and Donald E. Knuth. Analysis of the subtractive algorithm for greatest common divisors. Technical Report STAN-CS-75-510 (AD-A017054), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975. iv + 11 pp. URL https://www. jstor.org/stable/65273. To the memory of Hans A. Heilbronn, 1908–

Dubost:1975:SIN

[522] Pierre Dubost and Jean-Michel Trousse. Software implementation of a new method of combinatorial hashing. Technical Report STAN-CS-75-511 (AD-A017294), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975. ?? pp. URL http://

i.stanford.edu/TR/CS-TR-75-511.html.

Tarjan:1975:APC

[523] Robert Endre Tarjan. Applications of path compression on balanced trees. Technical Report STAN-CS-75-512 (PB247895/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, August

1975. ?? pp. URL http://i.stanford.edu/TR/CS-TR-75-512.html.

Bentley:1975:STF

[524] Jon Louis Bentley. A survey of techniques for fixed radius near neighbor searching. Technical Report STAN-CS-75-513 (SLACR-186), Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1975. ?? pp. URL http://i.stanford.edu/TR/CS-TR-75-513.
html.

Tokura:1975:MCU

[525] Nobuki Tokura. A microprogram control unit based on a tree memory. Technical Report STAN-CS-75-514 (PB247561/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1975. ??

pp. URL http://i.stanford.edu/TR/CS-TR-75-514.html.

Brent:1975:FMP

[526] Richard P. Brent. Fast multiple-precision evaluation of elementary functions. Technical Report STAN-CS-75-515, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1975. i + 22

pp. URL https://bitsavers.org/pdf/stanford/cs_techReports/
STAN-CS-75_515_Brent_Fast_Multiple-Precision_Evaluation_Of_
Elementary_Functions_Aug75.pdf.

Stoer:1975:RBQ

[527] J. Stoer. On the relation between quadratic termination and convergence properties of minimization algorithms. Technical Report STAN-CS-75-516 (SU326, P30-42), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. 103 pp.

Chvatal:1975:DOG

[528] Vaclav Chvátal and Carsten Thomassen. Distances in orientations of graphs. Technical Report STAN-CS-75-517, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1975. ??

pp. URL http://i.stanford.edu/TR/CS-TR-75-517.html.

Chvatal:1975:AII

[529] Vaclav Chvátal and Peter L. Hammer. Aggregation of inequalities in integer programming. Technical Report STAN-CS-75-518 (AD-A018461), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1975. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
75-518.html.

Davis:1975:PRR

[530] Randall Davis, Bruce Buchanan, and Edward Shortliffe. Production rules as a representation for a knowledge-based consultation program. Technical Report STAN-CS-75-519 (AIM-266, AD-A019641), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. iii + 37 pp.

vonHenke:1975:RDS

[531] Friedrich W. von Henke. On the representation of data structures in LCF with applications to program generation. Technical Report STAN-CS-75-520 (AIM-267, AD-A019664), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1975. 41 pp. URL

http://i.stanford.edu/TR/CS-TR-75-520.html.

Thompson:1975:DPS

[532] Clark Thompson. Depth perception in stereo computer vision. Technical Report STAN-CS-75-521 (AIM-268), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-75-521.html.

Luckham:1975:APV

[533] David C. Luckham and Norihisa Suzuki. Automatic program verification IV: proof of termination within a weak logic of programs. Technical Report STAN-CS-75-522 (AIM-269, AD-A019569), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. ??

pp. URL http://i.stanford.edu/TR/CS-TR-75-522.html.

Reiser:1975:BDS

[534] John F. Reiser. BAIL: a debugger for SAIL. Technical Report STAN-CS-75-523 (AIM-270, AD-A019467), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-75-523.html.

Davis:1975:OPS

[535] Randall Davis and Jonathan King. An overview of production systems. Technical Report STAN-CS-75-524 (AIM-271, AD-A019702), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. v + 38 pp.

Ganapathy:1975:RSC

[536] S. Ganapathy. Reconstruction of scenes containing polyhedra from stereo pair of views. Technical Report STAN-CS-75-525 (AIM-272), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1975. 204 pp.

Tarjan:1975:GTG

[537] Robert Endre Tarjan. Graph theory and Gaussian elimination. Technical Report STAN-CS-75-526 (AD-A020848), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1975. 23

pp. URL http://i.stanford.edu/TR/CS-TR-75-526.html.

McCluskey:1975:CRC

[538] Edward J. McCluskey, John F. Wakerly, and Roy C. Ogus. Center for Reliable Computing: current research. Technical Report STAN-CS-75-527 (CSL-TR-100, SU-SEL-75-044), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. v + 99 pp. URL

http://i.stanford.edu/TR/CS-TR-75-527.html.

Tarjan:1975:SPP

[539] Robert Endre Tarjan. Solving path problems on directed graphs. Technical Report STAN-CS-75-528 (AD-A020597), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. 45

pp. URL http://i.stanford.edu/TR/CS-TR-75-528.html.

Bentley:1975:FAC

[540] J. L. Bentley and J. H. Friedman. Fast algorithms for constructing minimal spanning trees in coordinate spaces. Technical Report STAN-CS-75-529 (SLACP-1665), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1975. 29 pp.

Lentini:1975:AFD

[541] M. Lentini and Victor Pereyra. An adaptive finite difference solver for nonlinear two point boundary problems with mild boundary layers. Technical Report STAN-CS-75-530 (SU326 P30-40), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1975. 42

pp. URL http://i.stanford.edu/TR/CS-TR-75-530.html.

Rose:1975:AAV

[542] Donald J. Rose and Robert Endre Tarjan. Algorithmic aspects of vertex elimination on directed graphs. Technical Report STAN-CS-75-531 (AD-A020847), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1975. 45 pp. URL http://i.stanford.

edu/TR/CS-TR-75-531.html.

Jacobs:1975:BCS

[543] Patricia E. Jacobs. Bibliography of Computer Science Department technical reports. Technical Report STAN-CS-75-532, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1975.

?? pp. URL http://i.stanford.edu/TR/CS-TR-75-532.html.

Concus:1976:GCG

[544] Paul Concus, Gene H. Golub, and Dianne Prost O’Leary. A generalized conjugate gradient method for the numerical solution of elliptic partial differential equations. Technical Report STAN-CS-76-533 (LBL-4604, SU326 P30-44), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1976. 24 pp. URL http://i.stanford.

edu/TR/CS-TR-76-533.html.

Hemphill:1976:CAA

[545] Linda G. Hemphill. A conceptual approach to automatic language understanding and belief structures: With disambiguation of the word “For”. Technical Report STAN-CS-76-534 (AIM-273), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1976. 254 pp.

Grossman:1975:IGO

[546] David D. Grossman and Russell H. Taylor. Interactive generation of object models with a manipulator. Technical Report STAN-CS-75-536, Stanford University, Department of Computer Science, Stanford, CA,

USA, December 1975. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-75-536.html.

Bolles:1975:VVW

[547] Robert C. Bolles. Verification Vision within a programmable assembly system: an introductory discussion. Technical Report STAN-CS-75-537, Stanford University, Department of Computer Science, Stanford, CA,

USA, December 1975. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-75-537.html.

Knuth:1976:ASF

[548] Donald E. Knuth and L. Trabb Pardo. Analysis of a simple factorization algorithm. Technical Report STAN-CS-76-538 (AD-A024416), Stanford University, Department of Computer Science, Stanford, CA,

USA, January 1976. 43 pp. URL https://apps.dtic.mil/sti/tr/
pdf/ADA024416.pdf.

Manna:1975:NAR

[549] Zohar Manna and Adi Shamir. A new approach to recursive programs. Technical Report STAN-CS-75-539 (AIM-276, AD-A021055/9WC), Stanford University, Department of Computer Science, Stanford, CA,

USA, December 1975. 30 pp. URL http://i.stanford.edu/TR/CS-
TR-75-539.html.

Friedman:1976:AFB

[550] Jerome Friedman, Jon Louis Bentley, and Raphael Ari Finkel. An algorithm for finding best matches in logarithmic expected time. Technical Report STAN-CS-75-482 (SLACP-1549), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1976. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-75-482.html. Original ver-

sion February 1975. Revised December 1975 and July 1976.

Davis:1976:SCS

[551] Randall Davis and Margaret H. Wright. Stanford Computer Science Department research report. Technical Report STAN-CS-76-405, Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1976. ?? pp. URL http://i.stanford.edu/TR/CS-TR-76-
405.html.

Concus:1976:GCGb

[552] Paul Concus and Gene H. Golub. A generalized Conjugate Gradient method for nonsymmetric systems of linear equations. Technical Report STAN-CS-76-535, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1976. 12 pp. URL http://i.stanford.

edu/TR/CS-TR-76-535.html.

Graham:1976:ACM

[553] Ronald L. Graham, Andrew Chi-Chih Yao, and F. Frances Yao. Addition chains with multiplicative cost. Technical Report STAN-CS-76-540 (AD-A021587), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1976. 7 pp. URL http://i.stanford.

edu/TR/CS-TR-76-540.html.

Knuth:1976:MCS

[554] Donald E. Knuth. Mathematics and computer science: Coping with finiteness. Technical Report STAN-CS-76-541, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1976. 30 pp.

Manna:1976:TAO

[555] Zohar Manna and Adi Shamir. The theoretical aspects of the optimal fixedpoint. Technical Report STAN-CS-76-542 (AIM-277, AD-A027454), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1976. iv + 22 pp. URL http://i.stanford.edu/TR/CS-
TR-76-542.html.

Zave:1976:OPS

[556] Derek A. Zave. Optimal polyphase sorting. Technical Report STAN-CS-76-543, Stanford University, Department of Computer Science, Stanford,

CA, USA, March 1976. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-76-543.html.

Mont-Reynaud:1976:RTA

[557] Bernard Mont-Reynaud. Removing trivial assignments from programs. Technical Report STAN-CS-76-544, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1976. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-76-544.html.

Paul:1976:SBG

[558] Wolfgang J. Paul, Robert Endre Tarjan, and James R. Celoni. Space bounds for a game on graphs. Technical Report STAN-CS-76-545, Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1976. i + 21 pp. URL http://i.stanford.edu/pub/cstr/
reports/cs/tr/76/545/CS-TR-76-545.pdf;   http://i.stanford.
edu/TR/CS-TR-76-545.html.

Baskett:1976:DLC

[559] F. Baskett and L. Sustek. The design of a low cost video graphics terminal. Technical Report STAN-CS-76-546 (SLACP-1715), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1976. 25 pp.

Tarjan:1976:IAG

[560] Robert Endre Tarjan. Iterative algorithms for global flow analysis. Technical Report STAN-CS-76-547, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1976. ?? pp. URL http://

i.stanford.edu/TR/CS-TR-76-547.html.

OLeary:1976:HCG

[561] Dianne Prost O’Leary. Hybrid conjugate gradient algorithms. Technical Report STAN-CS-76-548, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1976. 120 pp.

Luckham:1976:APV

[562] David C. Luckham and Norihisa Suzuki. Automatic program verification v: verification-oriented proof rules for arrays, records and pointers. Technical Report STAN-CS-76-549 (AIM-278, AD-A027455), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1976. ?? pp. URL http://i.stanford.edu/TR/CS-TR-76-549.html.

Tarjan:1976:FMI

[563] Robert Endre Tarjan and Anthony E. Trojanowski. Finding a maximum independent set. Technical Report STAN-CS-76-550, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1976.

?? pp. URL http://i.stanford.edu/TR/CS-TR-76-550.html.

Knuth:1976:SACb

[564] Donald E. Knuth. The state of the Art of Computer Programming. Technical Report STAN-CS-76-551 (AD-A032347), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1,

1976. ?? pp. URL http://i.stanford.edu/TR/CS-TR-76-551.html;
 http://www-db.stanford.edu/TR/CS-TR-76-551.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-76-551.

Suzuki:1976:AVP

[565] Norihsa Suzuki. Automatic verification of programs with complex data structures. Technical Report STAN-CS-76-552 (AIM-279), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1976. 194 pp.

Tarjan:1976:CMN

[566] Robert Endre Tarjan. Complexity of monotone networks for computing conjunctions. Technical Report STAN-CS-76-553 (AD-A032772), Stanford University, Department of Computer Science, Stanford, CA, USA,

June 1976. 21 pp. URL http://i.stanford.edu/TR/CS-TR-76-553.
html.

Yu:1976:MWB

[567] F. S. Yu. Modeling the write behavior of computer programs. Technical Report STAN-CS-76-554, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1976. 185 pp.

Grossman:1976:MCS

[568] David D. Grossman. Monte Carlo simulation of tolerancing in discrete parts manufacturing and assembly. Technical Report STAN-CS-76-555 (AIM-280), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1976. 25 pp. URL http://i.stanford.edu/TR/

CS-TR-76-555.html.

Guibas:1976:AHA

[569] Leonidas John Guibas. The analysis of hashing algorithms. Technical Report STAN-CS-76-556, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. 136 pp.

URL https://doc.lagout.org/science/0_Computer%20Science/0_
Computer%20History/old-hardware/xerox/parc/techReports/CSL-
76-3_The_Analysis_of_Hashing_Algorithms.pdf; https://en.wikipedia.
org/wiki/Leonidas_J._Guibas;   https://www.proquest.com/pqdtglobal/
docview/302823644.

Paterson:1976:IBF

[570] Michael S. Paterson. An introduction to Boolean function complexity. Technical Report STAN-CS-76-557 (AD-A032122), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. iv + 19 pp.

Manna:1976:SSB

[571] Zohar Manna and Richard J. Waldinger. Is “sometime” sometimes better than “always”? Intermittent assertions in proving program correctness. Technical Report STAN-CS-76-558 (AIM-281, AD-A031406), Stanford University, Department of Computer Science, Stanford, CA, USA,

June 16, 1976. v + 39 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA031406.pdf.

Manna:1977:SSB

[572] Zohar Manna and Richard J. Waldinger. Is “sometime” sometimes better than “always”? Intermittent assertions in proving program correctness. Technical Report STAN-CS-76-558 (AIM-281.1, AD-A042507), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1977. v + 38 pp. URL http://i.stanford.edu/TR/CS-TR-76-558.

html. See also original version [571].

Golub:1976:RDL

[573] Gene H. Golub, Virginia C. Klema, and Gilbert W. Stewart. Rank degeneracy and least squares problems. Technical Report STAN-CS-76-559 (AD-A032348), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. 38 pp. URL http://i.stanford.

edu/TR/CS-TR-76-559.html.

Taylor:1976:SMC

[574] Russell Taylor. Synthesis of manipulator control programs from task-level specifications. Technical Report STAN-CS-76-560 (AIM-282), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1976. 229 pp.

Woods:1976:MPL

[575] Donald R. Woods. Mathematical Programming Language — user’s guide. Technical Report STAN-CS-76-561, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-76-561.html.

Knuth:1976:EDP

[576] Donald E. Knuth and L. Trabb Pardo. The early development of programming languages. Technical Report STAN-CS-76-562 (AD-A032123), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. i + 109 pp. URL https:

//bitsavers.org/pdf/stanford/cs_techReports/STAN-CS-76-562_
EarlyDevelPgmgLang_Aug76.pdf.

Russell:1976:SRA

[577] D. L. Russell. State restoration among communicating processes. Technical Report STAN-CS-76-563, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. 173 pp.

Davis:1976:AML

[578] Randall Davis. Applications of meta level knowledge to the construction, maintenance and use of large knowledge bases. Technical Report STAN-CS-76-564 (AIM-283, HPP-76-112), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1976. 304 pp.

Strikwerda:1976:IBV

[579] John Charles Strikwerda. Initial boundary value problems for incompletely parabolic systems. Technical Report STAN-CS-76-565 (AD-A032802), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1976. iv + 107 pp. URL https://apps.dtic.

mil/sti/tr/pdf/ADA032802.pdf.

Wright:1976:NMN

[580] Margaret Wright. Numerical methods for nonlinearly constrained optimization. Technical Report STAN-CS-76-566, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1976. 262 pp.

Finkel:1976:CDM

[581] Rafael Finkel. Constructing and debugging manipulator programs. Technical Report STAN-CS-76-567 (AIM-284), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. 171 pp.

Binford:1976:ESC

[582] Thomas O. Binford, David D. Grossman, C. Richard Liu, Robert C. Bolles, Raphael A. Finkel, M. Shahid Mujtaba, Michael D. Roderick, Bruce E. Shimano, Russell H. Taylor, Ronald H. Goldman, J. Pitts Jarvis, III, Victor D. Scheinman, and Thomas A. Gafford. Exploratory study of computer integrated assembly systems. Progress report 3, covering the period December 1, 1975 to July 31, 1976. Technical Report STAN-CS-76-568 (AIM-285, PB-259 130/2WC), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976.

356 pp. URL http://i.stanford.edu/TR/CS-TR-76-568.html.

Herriot:1976:CIN

[583] John G. Herriot. Calculation of interpolating natural spline functions using de Boor’s package for calculating with B-splines. Technical Report STAN-CS-76-569 (P261 814/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1976. i + 44 pp. URL

http://i.stanford.edu/TR/CS-TR-76-569.html.

Lenat:1976:AIA

[584] Douglas Lenat. AM: an artificial intelligence approach to discovery in mathematics as heuristic search. Technical Report STAN-CS-76-570 (AIM-286), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1976. viii + 345 pp.

Roderick:1976:DCR

[585] Michael Roderick. Discrete control of a robot arm. Technical Report STAN-CS-76-571 (AIM-287), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1976. 98 pp.

Filman:1976:FP

[586] Robert E. Filman and Richard W. Weyhrauch. An FOL primer. Technical Report STAN-CS-76-572 (AIM-288), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1976. 36 pp.

URL http://i.stanford.edu/TR/CS-TR-76-572.html.

Jonassen:1976:STF

[587] Arne T. Jonassen. The stationary p-tree forest. Technical Report STAN-CS-76-573 (AD-A032945), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1976. iii + 88 pp. URL http://

i.stanford.edu/TR/CS-TR-76-573.html.

Reiser:1976:S

[588] John F. Reiser. SAIL. Technical Report STAN-CS-76-574 (AIM-289), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1976. 178 pp. URL http://i.stanford.edu/pub/cstr/
reports/cs/tr/76/574/CS-TR-76-574.pdf;   http://i.stanford.
edu/TR/CS-TR-76-574.html.

Smith:1976:ST

[589] Nancy W. Smith. SAIL tutorial. Technical Report STAN-CS-76-575 (AIM-290, AD-A042494), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1976. 54 pp. URL http://i.

stanford.edu/TR/CS-TR-76-575.html.

McDiarmid:1976:DCN

[590] Colin McDiarmid. Determining the chromatic number of a graph. Technical Report STAN-CS-76-576 (AD-A035350), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1976. 61 pp.

Buchanan:1976:TRJ

[591] Bruce Buchanan, Joshua Lederberg, and John McCarthy. Three reviews of J. Weizenbaum’s Computer Power and Human Reason. Technical Report STAN-CS-76-577 (AIM-291, A044713), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1976. 28 pp.

Oliger:1976:TPA

[592] Joseph Oliger and Arne Sundström. Theoretical and practical aspects of some initial-boundary value problems in fluid dynamics. Technical Report STAN-CS-76-578 (AD-A035219), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1976. iv + 61 pp.

URL http://i.stanford.edu/TR/CS-TR-76-578.html.

Baskett:1976:AIM

[593] Forest Baskett and Abbas Rafii. The A0 inversion model of program paging behavior. Technical Report STAN-CS-76-579 (SLACP-1826), Stanford University, Department of Computer Science, Stanford, CA, USA,

November 1976. 30 pp. URL http://i.stanford.edu/TR/CS-TR-76-
579.html.

Winograd:1976:TPU

[594] Terry A. Winograd. Towards a procedural understanding of semantics. Technical Report STAN-CS-76-580 (AIM-292), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1976. 30

pp. URL http://i.stanford.edu/TR/CS-TR-76-580.html.

Bobrow:1976:OKK

[595] Daniel G. Bobrow and Terry A. Winograd. An overview of KRL, a Knowledge Representation Language. Technical Report STAN-CS-76-581 (AIM-293, AD-A042508), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1976. 40 pp. URL http:

//i.stanford.edu/TR/CS-TR-76-581.html.

Chvatal:1976:TRC

[596] Vaclav Chvátal, M. R. Garey, and D. S. Johnson. Two results concerning multicoloring. Technical Report STAN-CS-76-582 (AD-A038863), Stanford University, Department of Computer Science, Stanford, CA, USA,

December 1976. iv + 8 pp. URL https://apps.dtic.mil/sti/pdfs/
ADA038863.pdf.

Chvatal:1976:DSN

[597] Vaclav Chvátal. Determining the stability number of a graph. Technical Report STAN-CS-76-583 (AD-A038864), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1976. ii + 39

pp. URL http://i.stanford.edu/TR/CS-TR-76-583.html;   https:
//apps.dtic.mil/sti/pdfs/ADA038864.pdf.

Knuth:1976:DPR

[598] Donald E. Knuth. Deletions that preserve randomness. Technical Report STAN-CS-77-584 (AD-A038865), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1976. v + 32 pp.

URL https://apps.dtic.mil/sti/tr/pdf/ADA038865.pdf.

Concus:1976:NSN

[599] Paul Concus, Gene H. Golub, and Dianne Prost O’Leary. Numerical solution of nonlinear elliptic partial differential equations by a generalized conjugate gradient method. Technical Report STAN-CS-76-585 (SU326 P30-50), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1976. vi + 43 pp. URL http://i.stanford.

edu/TR/CS-TR-76-585.html.

Dershowitz:1976:EPS

[600] Nachum Dershowitz and Zohar Manna. The evolution of programs: a system for automatic program modification. Technical Report STAN-CS-76-586 (AIM-294, AD-A042516), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1976. 45 pp. URL

http://i.stanford.edu/TR/CS-TR-76-586.html.

Garey:1977:CRB

[601] Michael R. Garey, Ronald L. Graham, David S. Johnson, and Donald E. Knuth. Complexity results for bandwidth minimization. Technical Report STAN-CS-77-587 (AD-A038867), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1977. v +

36 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA038867.pdf;
https://www.jstor.org/stable/2100947.

Weyhrauch:1977:UMF

[602] Richard W. Weyhrauch. A users manual for FOL. Technical Report STAN-CS-77-432, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1977. iv + 68 pp. URL http://i.stanford.

edu/TR/CS-TR-77-432.html.

Chan:1977:CSV

[603] Tony Fan C. Chan. On computing the singular value decomposition. Technical Report STAN-CS-77-588 (PB-266 381/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1977. ?? pp. URL http://i.stanford.edu/TR/CS-TR-77-588.html.

Engelmore:1977:KBS

[604] Robert S. Engelmore and H. Penny Nii. A knowledge-based system for the interpretation of protein X-ray crystallographic data. Technical Report STAN-CS-77-589 (HPP-77-2, AD-A038866), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1977.

?? pp. URL http://i.stanford.edu/TR/CS-TR-77-589.html.

Knuth:1977:IPI

[605] Donald E. Knuth and Michael S. Paterson. Identities from partition involutions. Technical Report STAN-CS-77-590 (AD-A038868), Stanford University, Department of Computer Science, Stanford, CA, USA,

February 1977. v + 22 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA038868.pdf.

Bolles:1976:VVW

[606] Robert C. Bolles. Verification vision within a programmable assembly system. Technical Report STAN-CS-77-591 (AIM-295, AD-A045723), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1976. vii + 239 pp.

Cartwright:1976:PFS

[607] Robert Cartwright, Jr. A practical formal semantic definition and verification systems for typed LISP. Technical Report STAN-CS-77-592 (AIM-296), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1976. v + 153 pp.

Scott:1977:ECP

[608] A. Carlisle Scott, William J. Clancey, Randall Davis, and Edward H. Shortliffe. Explanation capabilities of production-based consultation systems. Technical Report STAN-CS-77-593 (HPP-77-1), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1977. 30 pp. URL http://i.stanford.edu/TR/CS-TR-77-593.html.

Stritter:1977:FM

[609] Edwin P. Stritter. File migration. Technical Report STAN-CS-77-594 (SLAC-200), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1977. 112 pp.

Lewis:1977:ASM

[610] John Gregg Lewis. Algorithms for sparse matrix eigenvalue problems. Technical Report STAN-CS-77-595, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1977. 212 pp.

Stefik:1977:RKB

[611] Mark J. Stefik and Nancy Martin. A review of knowledge based problem solving as a basis for a genetics experiment designing system. Technical Report STAN-CS-77-596 (HPP-77-5), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1977. 97 pp. URL

http://i.stanford.edu/TR/CS-TR-77-596.html.

Buchanan:1977:MDL

[612] Bruce G. Buchanan and Tom M. Mitchell. Model-directed learning of production rules. Technical Report STAN-CS-77-597 (HPP-77-6), Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1977. 26 pp. URL http://i.stanford.edu/TR/CS-TR-77-597.
html.

Jonassen:1977:TAW

[613] Arne T. Jonassen and Donald E. Knuth. A trivial algorithm whose analysis isn’t. Technical Report STAN-CS-77-598 (AD-A040486), Stanford University, Department of Computer Science, Stanford, CA, USA,

April 1977. iv + 32 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA040486.pdf.

Knuth:1977:ELS

[614] Donald E. Knuth and Arnold Schönhage. The expected linearity of a simple equivalence algorithm. Technical Report STAN-CS-77-599 (AD-A040441), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1977. iv + 56 pp. URL https://apps.dtic.

mil/sti/tr/pdf/ADA040441.pdf.

Brown:1977:APN

[615] Mark R. Brown. The analysis of a practical and nearly optimal priority queue. Technical Report STAN-CS-77-600 (AD-A040538), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1977. iii + 99 pp.

Reiser:1977:AAR

[616] John Frederick Reiser. Analysis of additive random number generators. Technical Report STAN-CS-77-601 (AD-A045652), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1977. vi +

34 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA045652.pdf.

deBoor:1977:NSR

[617] Carl de Boor and Gene H. Golub. The numerically stable reconstruction of a Jacobi matrix from spectral data. Technical Report STAN-CS-77-602, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1977. vi + 18 pp. URL http://i.stanford.edu/TR/

CS-TR-77-602.html.

Tarjan:1977:RMR

[618] Robert Endre Tarjan. Reference machines require non-linear time to maintain disjoint sets. Technical Report STAN-CS-77-603 (AD-A041292), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1977. 45 pp. URL http://i.stanford.edu/TR/

CS-TR-77-603.html.

Chan:1977:CDL

[619] Tony Fan C. Chan and Joseph Oliger. Control of the dissipativity of Lax– Wendroff type methods for first order systems or hyperbolic equations. Technical Report STAN-CS-77-604 (AD-A040399), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1977. 45

pp. URL http://i.stanford.edu/TR/CS-TR-77-604.html.

Smith:1977:MLS

[620] Reid G. Smith, Tom M. Mitchell, Richard A. Chestek, and Bruce G. Buchanan. A model for learning systems. Technical Report STAN-CS-77-605 (HPP-77-14, AD-A042834), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1977. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-77-605.html.

Clancy:1977:PPS

[621] Michael J. Clancy and Donald E. Knuth. A programming and problemsolving seminar. Technical Report STAN-CS-77-606, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1,

1977. ?? pp. URL http://i.stanford.edu/TR/CS-TR-77-606.html;
 http://www-db.stanford.edu/TR/CS-TR-77-606.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-77-606.

Owicki:1977:SPA

[622] Susan S. Owicki. Specifications and proofs for abstract data types in concurrent programs. Technical Report STAN-CS-77-607 (CSL-TR-133), Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1977. ?? pp. URL http://i.stanford.edu/TR/CS-TR-77-
607.html.

Simon:1977:SSR

[623] Istvan Simon. On some subrecursive reducibilities. Technical Report STAN-CS-77-608, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1977. 102 pp.

Tarjan:1977:CCA

[624] Robert Endre Tarjan. Complexity of combinatorial algorithms. Technical Report STAN-CS-77-609 (AD-A043362), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1977. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-77-609.html.

Winograd:1977:FUD

[625] Terry A. Winograd. A framework for understanding discourse. Technical Report STAN-CS-77-610 (AIM-297), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1977. 97 pp.

Manna:1977:LCP

[626] Zohar Manna and Richard J. Waldinger. The logic of computer programming. Technical Report STAN-CS-77-611 (AIM-298, AD-A046703), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1977. 90 pp. URL http://i.stanford.edu/TR/CS-TR-
77-611.html.

Nii:1977:RBU

[627] H. Penny Nii and Edward A. Feigenbaum. Rule-based understanding of signals. Technical Report STAN-CS-77-612 (HPP-77-7, AD-A042756), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1977. iv + 23 pp.

Spedicato:1977:SCQ

[628] E. Spedicato. On some classes of quasi-Newton methods for systems of nonlinear algebraic equations. Technical Report STAN-CS-77-613, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1977. 23 pp.

Manna:1977:CFF

[629] Zohar Manna and Adi Shamir. The convergence of functions to fixedpoints of recursive definitions. Technical Report STAN-CS-77-614 (AIM-299), Stanford University, Department of Computer Science, Stanford,

CA, USA, May 1977. 45 pp. URL http://i.stanford.edu/TR/CS-TR-
77-614.html.

Glowinski:1977:NMF

[630] Roland Glowinski and Olivier Pironneau. Numerical methods for the first biharmonic equation and for the two-dimensional Stokes problem. Technical Report STAN-CS-77-615, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1977. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-77-615.html.

Kreiss:1977:SFM

[631] Heinz-Otto Kreiss and Joseph Oliger. Stability of the Fourier method. Technical Report STAN-CS-77-616 (AD-A046311), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1977. 25

pp. URL http://i.stanford.edu/TR/CS-TR-77-616.html.

Winograd:1977:SCS

[632] Terry A. Winograd. On some contested suppositions of generative linguistics about the scientific study of language. Technical Report STAN-CS-77-617 (AIM-300), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1977. 25 pp.

Nilsson:1977:PSA

[633] Nils J. Nilsson. A production system for automatic deduction. Technical Report STAN-CS-77-618 (HPP-77-28, AD-A045948), Stanford University, Department of Computer Science, Stanford, CA, USA, August

1977. 42 pp. URL http://i.stanford.edu/TR/CS-TR-77-618.html.

Paul:1977:TST

[634] Wolfgang J. Paul and Robert Endre Tarjan. Time-space trade-offs in a pebble game. Technical Report STAN-CS-77-619 (AD-A046481), Stanford University, Department of Computer Science, Stanford, CA, USA,

July 1977. 8 pp. URL http://i.stanford.edu/TR/CS-TR-77-619.
html.

Bolstad:1977:AFM

[635] J. Bolstad and J. Oliger. Adaptation of the Fourier method to the nonperiodic initial boundary value problem. Technical Report STAN-CS-77-620, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1977. 80 pp.

Feigenbaum:1977:AAI

[636] Edward A. Feigenbaum. The art of artificial intelligence: I. Themes and case studies of knowledge engineering. Technical Report STAN-CS-77-621 (HPP-77-25, AD-A046289), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1977. 80 pp. URL http:

//i.stanford.edu/TR/CS-TR-77-621.html.

Golub:1977:GCV

[637] Gene H. Golub, Michael Heath, and Grace Wahba. Generalized crossvalidation as a method for choosing a good ridge parameter. Technical Report STAN-CS-77-622, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1977. 25 pp. URL http://

www.stat.yale.edu/~jtc5/312_612/readings/generalized-cross-
validation-and-ridge_Golub-Heath-Wahba_79.pdf;   https://www.
jstor.org/stable/10.2307/1268518. Also issued as Department of

Statistics Technical Report number 491, University of Wisconsin, Madison, WI.

Boley:1977:IEPa

[638] D. Boley and Gene H. Golub. Inverse eigenvalue problems for band matrices. Technical Report STAN-CS-77-623, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1977. 9 pp.

McCarthy:1977:RRC

[639] John McCarthy, Thomas O. Binford, Cordell C. Green, David C. Luckham, Zohar Manna, Terry A. Winograd, and Lester D. Earnest. Recent research in computer science. Technical Report STAN-CS-77-624 (AIM-301, AD-A044231), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1977. 118 pp. URL http://i.

stanford.edu/TR/CS-TR-77-624.html;   https://apps.dtic.mil/
sti/tr/pdf/ADA044231.pdf.

Brown:1977:FMA

[640] Mark R. Brown and Robert Endre Tarjan. A fast merging algorithm. Technical Report STAN-CS-77-625 (AD-A046090), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1977.

iv + 42 pp. URL http://i.stanford.edu/TR/CS-TR-77-625.html;
https://apps.dtic.mil/sti/tr/pdf/ADA046090.pdf.

Yao:1977:LSA

[641] Andrew Chi-Chih Yao. On the loop switching addressing problem. Technical Report STAN-CS-77-626, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1977. 21 pp. URL http://

i.stanford.edu/TR/CS-TR-77-626.html.

Lipton:1977:STP

[642] Richard J. Lipton and Robert Endre Tarjan. A separator theorem for planar graphs. Technical Report STAN-CS-77-627 (AD-A048786), Stanford University, Department of Computer Science, Stanford, CA, USA,

October 1977. 31 pp. URL http://i.stanford.edu/TR/CS-TR-77-
627.html.

Lipton:1977:APS

[643] Richard J. Lipton and Robert Endre Tarjan. Applications of a planar separator theorem. Technical Report STAN-CS-77-628 (AD-A048787), Stanford University, Department of Computer Science, Stanford, CA,

USA, October 1977. 34 pp. URL http://i.stanford.edu/TR/CS-TR-
77-628.html.

Yao:1977:CPM

[644] Andrew Chi-Chih Yao. The complexity of pattern matching for a random string. Technical Report STAN-CS-77-629, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1977. 43 pp.

URL http://i.stanford.edu/TR/CS-TR-77-629.html.

Manna:1977:SDP

[645] Zohar Manna and Richard Waldinger. Synthesis: Dreams =¿ programs. Technical Report STAN-CS-77-630 (AIM-302, AD-A049761), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1977. 119 pp.

Dershowitz:1977:IRP

[646] Nachum Dershowitz and Zohar Manna. Inference rules for program annotation. Technical Report STAN-CS-77-631 (AIM-303, AD-A050806), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1977. 46 pp. URL http://i.

stanford.edu/TR/CS-TR-77-631.html;   https://apps.dtic.mil/
sti/tr/pdf/ADA050806.pdf.

Wagner:1977:HV

[647] Todd Jeffery Wagner. Hardware verification. Technical Report STAN-CS-77-632 (AIM-304, AD-A048684), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1977. iv + 99 pp.

Faught:1977:MIC

[648] William S. Faught. Motivation and intensionality in a computer simulation model. Technical Report STAN-CS-77-633 (AIM-305, AD-A048660), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1977. v + 99 pp.

Hoffmann:1977:NPGb

[649] Walter Hoffmann and Beresford N. Parlett. A new proof of global convergence for the tridiagonal QL algorithm. Technical Report STAN-CS-77-634, Stanford University, Department of Computer Science, Stanford,

CA, USA, October 1977. 18 pp. URL http://i.stanford.edu/TR/CS-
TR-77-634.html.

Golub:1977:BLMa

[650] Gene H. Golub, Franklin T. Luk, and Michael L. Overton. A block Lanczos method to compute the singular values and corresponding singular vectors of a matrix. Technical Report STAN-CS-77-635, Stanford University, Department of Computer Science, Stanford, CA, USA, October

1977. 80 pp. URL http://i.stanford.edu/TR/CS-TR-77-635.html.

Bube:1977:CTI

[651] Kenneth P. Bube. C^{m} convergence of trigonometric interpolants. Technical Report STAN-CS-77-636 (AD-A048788), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1977. 25

pp. URL http://i.stanford.edu/TR/CS-TR-77-636.html.

Ramshaw:1977:GSS

[652] Lyle H. Ramshaw. On the gap structure of sequences of points on a circle. Technical Report STAN-CS-77-637, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1977. 26 pp. URL

http://i.stanford.edu/TR/CS-TR-77-637.html.

OLeary:1977:GCG

[653] Dianne Prost O’Leary. A generalized conjugate gradient algorithm for solving a class of quadratic programming problems. Technical Report STAN-CS-77-638 (SU326 P30-57), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1977. 49 pp. URL

http://i.stanford.edu/TR/CS-TR-77-638.html.

Green:1977:PSK

[654] Cordell C. Green and David R. Barstow. On program synthesis knowledge. Technical Report STAN-CS-77-639 (AIM-306, AD-A053175), Stanford University, Department of Computer Science, Stanford, CA, USA,

November 1977. 63 pp. URL http://i.stanford.edu/TR/CS-TR-77-
639.html.

Barstow:1977:ACA

[655] David Barstow. Automatic construction of algorithms. Technical Report STAN-CS-77-641 (AIM-308, AD-A053184), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1977. 220 pp.

Yao:1977:CMS

[656] Andrew Chi-Chih Yao. On constructing minimum spanning trees in kdimensional spaces and related problems. Technical Report STAN-CS-77-642, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1977. ?? pp. URL http://i.stanford.edu/TR/

CS-TR-77-642.html.

Tanabe:1977:GMN

[657] Kunio Tanabe. A geometric method in nonlinear programming. Technical Report STAN-CS-77-643, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1977. 53 pp.

Mont-Reynaud:1977:HPF

[658] Bernard Mont-Reynaud. Hierarchical properties of flows and the determination of inner loops. Technical Report STAN-CS-77-644, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1977. 164 pp.

Lipton:1977:GND

[659] Richard J. Lipton, Donald J. Rose, and Robert Endre Tarjan. Generalized nested dissection. Technical Report STAN-CS-77-645, Stanford University, Department of Computer Science, Stanford, CA, USA, December

1977. 32 pp. URL http://i.stanford.edu/TR/CS-TR-77-645.html.

Yao:1977:LBP

[660] Andrew Chi-Chih Yao. A lower bound to palindrome recognition by probabilistic Turing machines. Technical Report STAN-CS-77-647, Stanford University, Department of Computer Science, Stanford, CA, USA,

December 1977. ?? pp. URL http://i.stanford.edu/TR/CS-TR-77-
647.html.

Manna:1978:SPR

[661] Zohar Manna and Richard J. Waldinger. Structured programming with recursion. Technical Report STAN-CS-77-640 (AIM-307, AD-A053176), Stanford University, Department of Computer Science, Stanford, CA,

USA, January 1978. ii + 5 pp. URL http://i.stanford.edu/TR/CS-
TR-77-640.html.

Nelson:1978:FDA

[662] Charles Gregory Nelson and Derek C. Oppen. Fast decision algorithms based on congruence closure. Technical Report STAN-CS-77-646 (AIM-309 MU-309), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1978. i + 13 pp. URL http://i.

stanford.edu/TR/CS-TR-77-646.html.

Knuth:1978:MT

[663] Donald E. Knuth. Mathematical typography. Technical Report STAN-CS-78-648 (AD-A054143), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1978. v + 68 pp. URL https:

//apps.dtic.mil/sti/tr/pdf/ADA054143.pdf. Josiah Willard Gibb’s

Lecture, given under the auspices of the American Mathematical Society, January 4, 1978.

Buchanan:1978:DMD

[664] Bruce G. Buchanan and Edward A. Feigenbaum. DENDRAL and Meta-DENDRAL: their applications dimension. Technical Report STAN-CS-78-649 (HPP-78-1), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1978. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-78-649.html.

Lengauer:1978:FAF

[665] Thomas Lengauer and Robert Endre Tarjan. A fast algorithm for finding dominators in a flow graph. Technical Report STAN-CS-78-650 (AD-A054144), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1978. ii + 40 pp.

Dershowitz:1978:PTM

[666] Nachum Dershowitz and Zohar Manna. Proving termination and multiset orderings. Technical Report STAN-CS-78-651 (AIM-310, AD-A058601), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1978. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
78-651.html.

Nelson:1978:SCD

[667] Charles Gregory Nelson and Derek C. Oppen. Simplification by cooperating decision procedures. Technical Report STAN-CS-78-652 (AIM-311), Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1978. ?? pp. URL http://i.stanford.edu/TR/CS-TR-78-
652.html.

Shiloach:1978:MTF

[668] Yossi Shiloach. Multi-terminal 0–1 flow. Technical Report STAN-CS-78-653, Stanford University, Department of Computer Science, Stanford,

CA, USA, April 1978. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-78-653.html.

Shiloach:1978:TPP

[669] Yossi Shiloach. The two paths problem is polynomial. Technical Report STAN-CS-78-654 (PB288520/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1978. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-78-654.html.

Dahlquist:1978:AUS

[670] Germund Dahlquist. On accuracy and unconditional stability of linear multistep methods for second order differential equations. Technical Report STAN-CS-78-655, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1978. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-78-655.html.

Heath:1978:NAN

[671] Michael Thomas Heath. Numerical algorithms for nonlinearly constrained optimization. Technical Report STAN-CS-78-656, Stanford University, Department of Computer Science, Stanford, CA, USA,

April 1978. x + 142 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA057962.pdf.

McCarthy:1978:MTK

[672] John McCarthy, Masahiko Sato, Takeshi Hayashi, and Shigeru Igarashi. On the model theory of knowledge. Technical Report STAN-CS-78-657 (AIM-312), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1978. 12 pp. URL http://i.stanford.edu/TR/

CS-TR-78-657.html.

Shustek:1978:APC

[673] Leonard Jay Shustek. Analysis and performance of computer instruction sets. Technical Report STAN-CS-78-658 (SLAC-205), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1978. x + 167 pp.

Zolnowsky:1978:TCG

[674] John E. Zolnowsky. Topics in computational geometry. Technical Report STAN-CS-78-659 (SLAC-206), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1978. 63 pp.

Shimano:1978:KDF

[675] Bruce E. Shimano. The kinematic design and force control of computer controlled manipulators. Technical Report STAN-CS-78-660 (AIM-313), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1978. 135 pp.

Gilbert:1978:VPG

[676] John R. Gilbert and Robert Endre Tarjan. Variations of a pebble game on graphs. Technical Report STAN-CS-78-661 (AD-A060794), Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1978. ?? pp. URL http://i.stanford.edu/TR/CS-TR-78-
661.html.

Yao:1978:NAP

[677] Andrew Chi-Chih Yao. New algorithms in bin packing. Technical Report STAN-CS-78-662, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1978. ii + 50 pp. URL http://i.

stanford.edu/TR/CS-TR-78-662.html.

Grosse:1978:SRG

[678] Eric H. Grosse. Software restyling in graphics and programming languages. Technical Report STAN-CS-78-663 (AD-A060793), Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1978. ?? pp. URL http://i.stanford.edu/TR/CS-TR-78-
663.html.

Bjorstad:1978:ANA

[679] Petter E. Bjørstad and Jorge Nocedal. Analysis of a new algorithm for one-dimensional minimization. Technical Report STAN-CS-78-664, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1978. 18 pp.

McWilliams:1978:SSC

[680] Thomas M. McWilliams and Lawrence C. Widdoes, Jr. SCALD: Structured Computer-Aided Logic Design. Technical Report STAN-CS-78-665 (CSL-TR-152), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1978. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-78-665.html.

McWilliams:1978:SPD

[681] Thomas M. McWilliams and Lawrence C. Widdoes, Jr. The SCALD physical design subsystem. Technical Report STAN-CS-78-666 (CSL-TR-153), Stanford University, Department of Computer Science, Stanford,

CA, USA, March 1978. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-78-666.html.

Smith:1978:DPS

[682] Reid G. Smith and Randall Davis. Distributed problem solving: The contract net approach. Technical Report STAN-CS-78-667 (HPP-78-7, AD-A060795), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1978. 27 pp.

Bonnet:1978:BPR

[683] Alain Bonnet. BAOBAB, a parser for a rule-based system using a semantic grammar. Technical Report STAN-CS-78-668 (HPP-78-10), Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1978. ?? pp. URL http://i.stanford.edu/TR/CS-TR-78-
668.html.

Stockmeyer:1978:OLM

[684] Paul K. Stockmeyer and F. Frances Yao. On the optimality of linear merge. Technical Report STAN-CS-78-669, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1978. 11 pp.

Graham:1978:IBW

[685] Ronald L. Graham, Andrew C. Yao, and F. Frances Yao. Information bounds are weak in the shortest distance problem. Technical Report STAN-CS-78-670, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1978. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-78-670.html.

Ginsparg:1978:NLP

[686] Jerrold M. Ginsparg. Natural language processing in an automatic programming domain. Technical Report STAN-CS-78-671 (AIM-316, AD-A058531), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1978. vii + 166 pp.

Chan:1978:CNM

[687] Tony F. C. Chan. Comparison of numerical methods for initial value problems. Technical Report STAN-CS-78-672, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1978. 195 pp.

Chan:1978:NLS

[688] Tony F. Chan, William M. Coughran, Jr., Eric H. Grosse, and Michael T. Heath. A numerical library and its support. Technical Report STAN-CS-78-673 (SU326 P30-59), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. iv + 20 pp. URL http:/

/i.stanford.edu/TR/CS-TR-78-673.html.

Chan:1978:FEA

[689] Tony F. Chan and Roland Glowinski. Finite element approximation and iterative solution of a class of mildly non-linear elliptic equations. Technical Report STAN-CS-78-674 (SU326 P30-60), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978.

?? pp. URL http://i.stanford.edu/TR/CS-TR-78-674.html.

Knuth:1978:TEC

[690] Donald E. Knuth. Tau Epsilon Chi, a system for technical text. Technical Report STAN-CS-78-675 (AIM-317), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. 200 pp.

Banning:1978:MDS

[691] John Phineas Banning. A method for determining the side effects of procedure calls. Technical Report STAN-CS-78-676 (SLACR-213), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. 283 pp.

fsSUCSD:1978:CE

[692] The faculty and students of the Stanford University Computer Science Department. Comprehensive examinations 1972–1978. Technical Report STAN-CS-78-677, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ?? pp. URL http://i.stanford.

edu/pub/cstr/reports/cs/tr/78/677/CS-TR-78-677.pdf.

Oppen:1978:RAR

[693] Derek C. Oppen. Reasoning about recursively defined data structures. Technical Report STAN-CS-78-678 (AIM-314), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1978. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-78-678.html.

Murray:1978:SAM

[694] Walter Murray and Michael L. Overton. Steplength algorithms for minimizing a class of nondifferentiable functions. Technical Report STAN-CS-78-679, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ?? pp. URL http://i.stanford.edu/TR/

CS-TR-78-679.html.

Stanley:1978:BSC

[695] Connie J. Stanley. Bibliography of Stanford Computer Science reports, 1963–1978. Technical Report STAN-CS-78-680, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ??

pp. URL http://i.stanford.edu/TR/CS-TR-78-680.html.

TrabbPardo:1978:SRS

[696] Luis Trabb Pardo. Set representation and set intersection. Technical Report STAN-CS-78-681 (AD-A065283), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. viii +

81 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA065283.pdf;
https://dl.acm.org/doi/10.1145/800133.804328.

Valdes:1978:PFS

[697] Jacobo Valdes. Parsing flowcharts and series parallel graphs. Technical Report STAN-CS-78-682 (AD-A065265), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. iv +

228 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA065265.pdf.

Tarjan:1978:SST

[698] Robert Endre Tarjan. Storing a sparse table. Technical Report STAN-CS-78-683 (AD-A065284), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-78-683.html.

Boley:1978:MIE

[699] Daniel L. Boley and Gene H. Golub. The matrix inverse eigenvalue problem for periodic Jacobi matrices. Technical Report STAN-CS-78-684 (SU326 P30-63), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. [5] + 14 pp. URL http://i.

stanford.edu/TR/CS-TR-78-684.html.

Luk:1978:SPM

[700] Franklin Tai-Cheung Luk. Sparse and parallel matrix computations. Technical Report STAN-CS-78-685 (AD-A065285), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. 168 pp.

Lipton:1978:EHS

[701] Richard J. Lipton, Arnold L. Rosenberg, and Andrew C. Yao. External hashing schemes for collections of data structures. Technical Report STAN-CS-78-686, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. 33 pp.

Weyhrauch:1978:PTF

[702] Richard W. Weyhrauch. Prolegomena to a theory of formal reasoning. Technical Report STAN-CS-78-687 (AIM-315, AD-A065698), Stanford University, Department of Computer Science, Stanford, CA, USA, De-

cember 1978. ?? pp. URL http://i.stanford.edu/TR/CS-TR-78-687.
html.

Manna:1978:SLL

[703] Zohar Manna. Six lectures on the logic of computer programming. Technical Report STAN-CS-78-688 (AIM-318), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. 54 pp.

Nelson:1978:ATV

[704] Charles Gregory Nelson. An n^{log n} algorithm for the two-variable-per-constraint linear programming satisfiability problem. Technical Report STAN-CS-78-689 (AIM-319), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-78-689.html.

Manna:1978:DAP

[705] Zohar Manna and Richard J. Waldinger. A deductive approach to program synthesis. Technical Report STAN-CS-78-690 (AIM-320, AD-A065558), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ?? pp. URL http://i.stanford.edu/

TR/CS-TR-78-690.html.

Bube:1978:CID

[706] Kenneth P. Bube. The construction of initial data for hyperbolic systems from nonstandard data. Technical Report STAN-CS-78-691 (AD-A066058), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. 119 pp.

Buchanan:1979:MLS

[707] Bruce G. Buchanan, Tom M. Mitchell, Reid G. Smith, and C. Richard Johnson, Jr. Models of learning systems. Technical Report STAN-CS-79-692 (HPP-77-39, AD-A066147), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. iii + 37 pp.

West:1978:CSG

[708] Douglas B. West. A class of solutions to the gossip problem. Technical Report STAN-CS-78-693 (AD-A066099), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-78-693.html.

King:1978:CSS

[709] Jonathan J. King. Computer science at Stanford, 1977–1978. Technical Report STAN-CS-78-694, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-78-694.html.

McCarthy:1978:RRA

[710] John McCarthy, Tom Binford, Cordell Green, David Luckham, Zohar Manna, and Les Earnest. Recent research in artificial intelligence and foundations of programming. Technical Report STAN-CS-78-695 (AIM-321, AD-A066562), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1978. iv + 92 pp. Edited by Les Earnest.

Smith:1978:CMB

[711] Reid G. Smith and Tom M. Mitchell. Considerations for microprocessorbased terminal design. Technical Report STAN-CS-78-696 (HPP-78-22, AD-A068538), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. iv + 14 pp. URL https://apps.

dtic.mil/sti/tr/pdf/ADA068538.pdf.

Gander:1978:LLS

[712] Walter Gander. On the linear least squares problem with a quadratic constraint. Technical Report STAN-CS-78-697, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. ??

pp. URL http://i.stanford.edu/TR/CS-TR-78-697.html.

Sweet:1978:EEP

[713] Richard E. Sweet. Empirical estimates of program entropy. Technical Report STAN-CS-78-698, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1978. 167 pp.

Bennett:1978:SKB

[714] James Bennett, Lewis Creary, Robert S. Engelmore, and Robert Melosh. SACON: a knowledge-based consultant for structural analysis. Technical Report STAN-CS-78-699 (HPP-78-23, AD-A068539), Stanford University, Department of Computer Science, Stanford, CA, USA, September

1978. ?? pp. URL http://i.stanford.edu/TR/CS-TR-78-699.html.

Smith:1978:FPS

[715] Reid Garfield Smith. A framework for problem solving in a distributed processing environment. Technical Report STAN-CS-78-700 (HPP-78-28, AD-A068230), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. 150 pp.

Shiloach:1978:MFA

[716] Yossi Shiloach. An O(n · I log^{2} I) maximum-flow algorithm. Technical Report STAN-CS-78-702, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. 33 pp. URL http://i.

stanford.edu/TR/CS-TR-78-702.html.

Brown:1978:DAD

[717] Mark R. Brown and Robert E. Tarjan. Design and analysis of a data structure for representing sorted lines. Technical Report STAN-CS-78-709 (AD-A068231), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. 50 pp. URL http://i.

stanford.edu/TR/CS-TR-78-709.html.

Mitchell:1978:VSA

[718] Tom Michael Mitchell. Version spaces: an approach to concept learning. Technical Report STAN-CS-78-711, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1978. 216 pp.

Model:1979:MSB

[719] Mitch L. Model. Monitoring system behavior in a complex computational environment. Technical Report STAN-CS-79-701, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. 189 pp.

Aspvall:1979:PTA

[720] Bengt Aspvall and Yossi Shiloach. A polynomial time algorithm for solving systems of linear inequalities with two variables per inequality. Technical Report STAN-CS-79-703 (AD-A068228), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. 25

pp. URL http://i.stanford.edu/TR/CS-TR-79-703.html.

Sweet:1979:SSS

[721] Roland A. Sweet. A survey of the state of software for partial differential equations. Technical Report STAN-CS-79-704 (AD-A068232), Stanford University, Department of Computer Science, Stanford, CA, USA, Jan-

uary 1979. 31 pp. URL http://i.stanford.edu/TR/CS-TR-79-704.
html.

Drysdale:1979:GVD

[722] Robert Lewis (Scot) Drysdale, III. Generalized Voronoi diagrams and geometric searching. Technical Report STAN-CS-79-705, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. 196 pp.

Yao:1979:GIN

[723] F. Francis Yao. Graph 2-isomorphism is NP-complete. Technical Report STAN-CS-79-706, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. 12 pp. URL http://i.stanford.

edu/TR/CS-TR-79-706.html.

VanWyk:1979:PPS

[724] Christopher J. Van Wyk and Donald E. Knuth. A programming and problem solving seminar. Technical Report STAN-CS-79-707, Stanford University, Department of Computer Science, Stanford, CA, USA, Jan-

uary 1, 1979. 83 pp. URL http://i.stanford.edu/TR/CS-TR-79-707.
html;   http://www-db.stanford.edu/TR/CS-TR-79-707.html;   http:
//www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&
id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-79-707.

Yao:1979:AMA

[725] Andrew C. Yao. An analysis of a memory allocation scheme for implementing stacks. Technical Report STAN-CS-79-708, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. ??

pp. URL http://i.stanford.edu/TR/CS-TR-79-708.html.

Trefethen:1979:NCS

[726] Lloyd N. Trefethen. Numerical computation of the Schwarz–Christoffel transformation. Technical Report STAN-CS-79-710 (AD-A068210), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1979. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
79-710.html.

Knuth:1979:ECP

[727] Donald E. Knuth. The errata of computer programming. Technical Report STAN-CS-79-712 (AD-A068393), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1, 1979.

ii + 56 pp. URL http://i.stanford.edu/TR/CS-TR-79-712.html;
 http://www-db.stanford.edu/TR/CS-TR-79-712.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/search?formname=detail&id=
oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-79-712;   https://apps.
dtic.mil/sti/pdfs/ADA068393.pdf.

Golub:1979:HSMa

[728] Gene H. Golub, Stephen Nash, and Charles F. Van Loan. A Hessenberg– Schur method for the problem AX + XB = C. Technical Report STAN-CS-79-713, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. 50 pp.

Castaneda:1979:PFP

[729] Fernando Castaneda, Frederick C. Chow, Peter Nye, Daniel D. Sleator, and Gio Wiederhold. PCFORT: a Fortran-to-Pcode translator. Technical Report STAN-CS-79-714 (CSL-TN-160), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. vi + 85 pp. URL http://i.stanford.edu/

pub/cstr/reports/cs/tr/79/714/CS-TR-79-714.pdf;   http://i.
stanford.edu/TR/CS-TR-79-714.html;   https://bitsavers.org/
pdf/stanford/cs_techReports/STAN-CS-79-714_PCFORT_Jan79.pdf.

Hailpern:1979:AM

[730] Brent T. Hailpern and Bruce L. Hitson. S-1 architecture manual. Technical Report STAN-CS-79-715 (CSL-TR-79-161), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. iii

+ iv + 366 pp. URL http://i.stanford.edu/TR/CS-TR-79-715.html;
 https://bitsavers.org/pdf/stanford/cs_techReports/STAN-CS-
79-715_S-1_Arch_Man.pdf.

Georgeff:1979:FCP

[731] Michael P. Georgeff. A framework for control in production systems. Technical Report STAN-CS-79-716 (AIM-322), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. ??

pp. URL http://i.stanford.edu/TR/CS-TR-79-716.html.

Cartwright:1979:RPF

[732] Robert Cartwright and John McCarthy. Recursive programs as functions in a first order theory. Technical Report STAN-CS-79-717 (AIM-324, AD-A071421), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1979. 32 pp.

Mujtaba:1979:AUM

[733] Mohamed Shahid Mujtaba and Ron Goldman. AL users’ manual. Technical Report STAN-CS-79-718 (AIM-323), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-79-718.html.

Bjorstad:1979:EAE

[734] Petter Bjørstad, Germund Dahlquist, and Eric H. Grosse. Extrapolation of asymptotic expansions by a modified Aitken δ^{2}-formula. Technical Report STAN-CS-79-719 (SU326 P30-61), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1979. iv + 45 +

7 pp. URL http://i.stanford.edu/TR/CS-TR-79-719.html.

Glowinski:1979:GOB

[735] Roland Glowinski. On grid optimization for boundary value problems. Technical Report STAN-CS-79-720, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1979. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-79-720.html.

Yao:1979:FTN

[736] Andrew C. Yao and F. Frances Yao. On fault-tolerant networks for sorting. Technical Report STAN-CS-79-721, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1979. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-79-721.html.

Wiederhold:1979:SMD

[737] Gio Wiederhold and Ramez A. El-Masri. A structural model for database systems. Technical Report STAN-CS-79-722 (AD-A074077), Stanford University, Department of Computer Science, Stanford, CA, USA, Febru-

ary 1979. ?? pp. URL http://i.stanford.edu/TR/CS-TR-79-722.
html.

Shortliffe:1979:KEM

[738] Edward H. Shortliffe, Bruce G. Buchanan, and Edward A. Feigenbaum. Knowledge engineering for medical decision making: a review of computer-based clinical decision aids. Technical Report STAN-CS-79-723 (AD-A074076), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1979. v + 48 pp.

McCarthy:1979:FOT

[739] John McCarthy. First order theories of individual concepts and propositions. Technical Report STAN-CS-79-724 (AIM-325), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1979. 26 pp.

McCarthy:1979:AMQ

[740] John McCarthy. Ascribing mental qualities to machines. Technical Report STAN-CS-79-725 (AIM-326, AD-A071423), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1979. 25 pp.

Yao:1979:AHK

[741] Andrew Chi-Chih Yao. An analysis of (h, k, l)-shellsort. Technical Report STAN-CS-79-726, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1979. ?? pp. URL http://i.stanford.edu/

TR/CS-TR-79-726.html.

Filman:1979:IOI

[742] Robert Elliot Filman. The interaction of observation and inference. Technical Report STAN-CS-79-727 (AIM-327, AD-A076794), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. vii + 226 pp.

Shiloach:1979:UMA

[743] Yossi Shiloach. Union-member algorithms for non-disjoint sets. Technical Report STAN-CS-79-728, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-79-728.html.

Tarjan:1979:UAP

[744] Robert Endre Tarjan. A unified approach to path problems. Technical Report STAN-CS-79-729 (AD-A068229), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-79-729.html.

Liang:1979:QEC

[745] Frank M. Liang. Qualifying examinations in computer science, 1965– 1978. Technical Report STAN-CS-79-730, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-79-730.html.

Luckham:1979:SPV

[746] David C. Luckham, Steven M. German, Friedrich W. von Henke, Richard A. Karp, P. W. Milne, Derek C. Oppen, Wolfgang Polak, and William L. Scherlis. Stanford Pascal Verifier user manual. Technical Report STAN-CS-79-731 (PVG-11, AD-A071900), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1979. ??

pp. URL http://i.stanford.edu/TR/CS-TR-79-731.html.

Woods:1979:NIC

[747] Donald R. Woods. Notes on introductory combinatorics. Technical Report STAN-CS-79-732, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-79-732.html.

Yao:1979:LBF

[748] Andrew Chi-Chih Yao. A lower bound to finding convex hulls. Technical Report STAN-CS-79-733, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-79-733.html.

Tarjan:1979:FAS

[749] Robert Endre Tarjan. Fast algorithms for solving path problems. Technical Report STAN-CS-79-734 (AD-A074079), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-79-734.html.

Wilkinson:1979:KCF

[750] James Hardy Wilkinson. Kronecker’s canonical form and the QZ algorithm. Technical Report STAN-CS-79-735, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-79-735.html.

Wilkinson:1979:NPS

[751] James Hardy Wilkinson. Note on the practical significance of the Drazin inverse. Technical Report STAN-CS-79-736, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1979. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-79-736.html.

Yao:1979:ACC

[752] Andrew C. Yao and F. Frances Yao. On the average-case complexity of selecting the k-th best. Technical Report STAN-CS-79-737, Stanford University, Department of Computer Science, Stanford, CA, USA, April

1979. ?? pp. URL http://i.stanford.edu/TR/CS-TR-79-737.html.

LeVeque:1979:CRG

[753] Randall J. LeVeque, Germund Dahlquist, and Dan Andree. Computations related to g-stability of linear multistep methods. Technical Report STAN-CS-79-738 (SU326 P30-65), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1979. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-79-738.html.

Quinlan:1979:ILD

[754] J. R. Quinlan. Induction over large data bases. Technical Report STAN-CS-79-739 (HPP-79-14, AD-A074075), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1979. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-79-739.html.

Cartwright:1979:LA

[755] Robert Cartwright and Derek C. Oppen. The logic of aliasing. Technical Report STAN-CS-79-740 (PVG-12), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1979. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-79-740.html.

Ramshaw:1979:FAA

[756] Lyle Harold Ramshaw. Formalizing the analysis of algorithms. Technical Report STAN-CS-79-741, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1979. 123 pp.

Gardner:1979:HAI

[757] Anne Gardner. Handbook of artificial intelligence: Search. Technical Report STAN-CS-79-742 (HPP-79-12, AD-A074078), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1979. ix + 104 pp. A section of the Handbook of Artificial Intelligence edited by Avron Barr and Edward A. Feigenbaum.

Bulnes-Rozas:1979:GGO

[758] Juan Bautista Bulnes-Rozas. GOAL: a goal oriented command language for interactive proof construction. Technical Report STAN-CS-79-743 (AIM-328), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1979. vi + 172 pp.

Garcia-Molina:1979:PUA

[759] Hector Garcia-Molina. Performance of update algorithms for replicated data in a distributed database. Technical Report STAN-CS-79-744 (CSL-TR-172, AD-A075268), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1979. xii + 307 pp.

URL https://apps.dtic.mil/sti/citations/ADA075268;   https:
//apps.dtic.mil/sti/pdfs/ADA075268.pdf.

Lengauer:1979:ULB

[760] Thomas Lengauer. Upper and lower bounds on time-space tradeoffs in a pebble game. Technical Report STAN-CS-79-745 (AD-A076264), Stanford University, Department of Computer Science, Stanford, CA, USA,

July 1979. vii + 75 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA076264.pdf.

Borning:1979:TCO

[761] Alan Borning. ThingLab —- a constraint- oriented simulation laboratory. Technical Report STAN-CS-79-746, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. 109 pp.

Wilkins:1979:UPP

[762] David E. Wilkins. Using patterns and plans to solve problems and control search. Technical Report STAN-CS-79-747 (AIM-329, AD-A076872), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1979. 264 pp.

Yun:1979:FAS

[763] David Y. Y. Yun. Fast algorithms for solving Toeplitz systems of equations and finding rational Hermite interpolants. Technical Report STAN-CS-79-748 (AD-A075376), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-79-748.html.

Clancey:1979:AOA

[764] William J. Clancey, James S. Bennett, and Paul R. Cohen. Applicationsoriented AI research: Education. Technical Report STAN-CS-79-749 (HPP-79-17, AD-A075517), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. ix + 56 pp. URL https:/

/apps.dtic.mil/sti/pdfs/ADA075517.pdf. A section of the Handbook

of Artificial Intelligence edited by Avron Barr and Edward A. Feigenbaum.

Gacs:1979:KAL

[765] Peter Gács and Laszlo Lovász. Khachian’s algorithm for linear programming. Technical Report STAN-CS-79-750 (AD-A075171), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. iv + 12 pp.

Manna:1979:MLP

[766] Zohar Manna and Amir Pnueli. The modal logic of programs. Technical Report STAN-CS-79-751 (AIM-330, AD-A155071), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1979. ii + 35 pp.

Overton:1979:PLA

[767] Michael Lockhart Overton. Projected Lagrangian algorithms for nonlinear minimax and l_{1} optimization. Technical Report STAN-CS-79-752, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. 164 pp.

Yao:1979:STS

[768] Andrew Chi-Chih Yao. Should tables by sorted? Technical Report STAN-CS-79-753, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. ?? pp. URL http://i.stanford.edu/

TR/CS-TR-79-753.html.

Gardner:1979:NLU

[769] Anne Gardner, James Davidson, and Terry A. Winograd. Natural language understanding. Technical Report STAN-CS-79-754 (HPP-79-21, AD-A076873), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. xii + 93 pp.

Kant:1979:ECP

[770] Elaine Kant. Efficiency considerations in program synthesis: a knowledge-based approach. Technical Report STAN-CS-79-755 (AIM-331), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. 160 pp.

Bennett:1979:AOA

[771] James S. Bennett, Bruce G. Buchanan, Paul R. Cohen, and Fritz Fisher. Applications-oriented AI research: Science and mathematics. Technical Report STAN-CS-79-756 (HPP-79-22, AD-A076875), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1979. ix + 100 pp. A section of the Handbook of Artificial Intelligence edited by Avron Barr and Edward A. Feigenbaum.

Clesielski:1979:AOA

[772] Victor B. Clesielski, James S. Bennett, and Paul R. Cohen. Applicationsoriented AI research: Medicine. Technical Report STAN-CS-79-757 (HPP-79-23, AD-A075402), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979. ix + 53 pp. URL https:

//apps.dtic.mil/sti/tr/pdf/ADA075402.pdf.

Elschlager:1979:AP

[773] Robert Elschlager and Jorge Phillips. Automatic programming. Technical Report STAN-CS-79-758 (HPP-79-24, AD-A076874), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1979.

v + 89 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA076874.pdf.

Bonnet:1979:SSS

[774] Alain Bonnet. Schema-shift strategies for understanding structured texts in natural language. Technical Report STAN-CS-79-759 (HPP-79-25), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1979. vi + ii + 40 pp. URL http://i.stanford.edu/TR/

CS-TR-79-759.html.

Graham:1979:SMP

[775] Ronald L. Graham, Andrew C. Yao, and F. Frances Yao. Some monotonicity properties of partial orders. Technical Report STAN-CS-79-760, Stanford University, Department of Computer Science, Stanford, CA,

USA, September 1979. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-79-760.html.

West:1979:GDT

[776] Douglas B. West. Gossiping without duplicate transmissions. Technical Report STAN-CS-79-761 (AD-A084021), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1979. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-79-761.html.

Knuth:1979:MSA

[777] Donald E. Knuth. METAFONT: a system for alphabet design. Technical Report STAN-CS-79-762 (AIM-332, AD-A083229), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1979.

?? pp. URL http://i.stanford.edu/TR/CS-TR-79-762.html.

West:1979:SCD

[778] Douglas B. West. A symmetric chain decomposition of L(4, n). Technical Report STAN-CS-79-763 (AD-A076876), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1979. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-79-763.html.

Yao:1979:TST

[779] Andrew Chi-Chih Yao. On the time-space tradeoff for sorting with linear queries. Technical Report STAN-CS-79-764, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1979. ??

pp. URL http://i.stanford.edu/TR/CS-TR-79-764.html.

Gacs:1979:RBC

[780] Peter Gács. Relation between the complexity and the probability of large numbers. Technical Report STAN-CS-79-765 (AD-A083192), Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1979. ?? pp. URL http://i.stanford.edu/TR/CS-TR-79-
765.html.

Kautsky:1979:EMC

[781] J. Kautsky and N. K. Nichols. Equidistributing meshes with constraints. Technical Report STAN-CS-79-766 (SU326, P30-69), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1979. 27 pp.

VanLoan:1979:SSV

[782] Charles Van Loan. On Stewart’s singular value decomposition for partitioned orthogonal matrices. Technical Report STAN-CS-79-767, Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1979. ?? pp. URL http://i.stanford.edu/TR/CS-TR-79-
767.html.

Clancey:1979:TRB

[783] William John Clancey. Transfer of rule-based expertise through a tutorial dialogue. Technical Report STAN-CS-79-769 (AD-A083432), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1979. xi + 450 + xiii pp. URL https://

apps.dtic.mil/sti/citations/ADA083432;   https://apps.dtic.
mil/sti/tr/pdf/ADA083432.pdf;   https://dl.acm.org/doi/10.5555/
908721;   https://www.semanticscholar.org/paper/Transfer-of-
rule-based-expertise-through-a-tutorial-Clancey/754d8ce1fb8511a780a72b120bfad7b7d73a583e.

Oppen:1979:PP

[784] Derek C. Oppen. Pretty printing. Technical Report STAN-CS-79-770 (PVG-13), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1979. ?? pp. URL http://i.stanford.edu/

TR/CS-TR-79-770.html.

Friedland:1979:KBE

[785] Peter E. Friedland. Knowledge-based experiment design in molecular genetics. Technical Report STAN-CS-79-771 (HPP-79-29), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1979. 137 pp.

McCune:1979:BPM

[786] Brian P. McCune. Building program models incrementally from informal descriptions. Technical Report STAN-CS-79-772 (AIM-333, AD-A086504), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1979. vii + 140 pp. URL https://apps.dtic.

mil/sti/tr/pdf/ADA086504.pdf.

Chan:1979:UFP

[787] Tony F. Chan, Gene H. Golub, and Randall J. LeVeque. Updating formulae and a pairwise algorithm for computing sample variances. Technical Report STAN-CS-79-773 (AD-A083170), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1979. 19 pp.

URL http://i.stanford.edu/TR/CS-TR-79-773.html.

Golub:1979:LSG

[788] Gene H. Golub and Robert J. Plemmons. Large scale geodetic least squares adjustment by dissection and orthogonal decomposition. Technical Report STAN-CS-79-774 (AD-A083193), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1979. 33

pp. URL http://i.stanford.edu/TR/CS-TR-79-774.html.

Diaconis:1979:ASE

[789] Persi Diaconis and Ronald L. Graham. The analysis of sequential experiments with feedback to subjects. Technical Report STAN-CS-79-775 (AD-A083288), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1979. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-79-775.html.

Aspvall:1979:KLP

[790] Bengt Aspvall and Richard E. Stone. Khachiyan’s linear programming algorithm. Technical Report STAN-CS-79-776, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1979. 13 pp. This paper supersedes STAN-CS-79-750 [765].

Graham:1979:CWC

[791] Ronald L. Graham and Neil J. A. Sloane. On constant weight codes and harmonious graphs. Technical Report STAN-CS-79-777 (AD-A083176), Stanford University, Department of Computer Science, Stanford, CA,

USA, December 1979. ?? pp. URL http://i.stanford.edu/TR/CS-
TR-79-777.html.

Shaw:1979:HAA

[792] David Elliot Shaw. A hierarchical associative architecture for the parallel evaluation of relational algebraic database primitives. Technical Report STAN-CS-79-778 (AD-A083573), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1979. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-79-778.html.

King:1979:EUD

[793] Jonathan J. King. Exploring the use of domain knowledge for query processing efficiency. Technical Report STAN-CS-79-781, Stanford University, Department of Computer Science, Stanford, CA, USA, December

1979. iv + 21 pp. URL http://i.stanford.edu/TR/CS-TR-79-781.
html.

Karp:1979:PCS

[794] Richard Alan Karp. Proving concurrent systems correct. Technical Report STAN-CS-80-783 (PVG-14), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1979. 151 pp.

Chow:1980:UFU

[795] Frederick Chow, Peter Nye, and Gio Wiederhold. UFORT: a Fortran-to-Universal PCODE translator. Technical Report CSL-TR-79-168, Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1980. iv + 88 pp. URL http://i.stanford.edu/pub/cstr/
reports/csl/tr/79/168/CSL-TR-79-168.pdf.

Gacs:1980:CNW

[796] Peter Gács and Leonid A. Levin. Causal nets or what is a deterministic computation. Technical Report STAN-CS-80-768, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1980. 20

pp. URL http://www-db.stanford.edu/TR/CS-TR-80-768.html.

Manna:1980:PFP

[797] Zohar Manna and Richard J. Waldinger. Problematic features of programming languages: a situational-calculus approach. Technical Report STAN-CS-80-779, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1980. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-80-779.html.

Knuth:1980:CMF

[798] Donald E. Knuth. The Computer Modern Family of typefaces. Technical Report STAN-CS-80-780, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1, 1980. 406

pp. URL http://www-db.stanford.edu/TR/CS-TR-80-780.html;
 http://www.ncstrl.org:8900/ncstrl/servlet/search?formname=
detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-80-780.

Anonymous:1980:NP

[799] Anonymous. NEVER PRINTED. Technical Report STAN-CS-80-782, Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1980. ???? pp.

Stefik:1980:PC

[800] Mark Jeffrey Stefik. Planning with constraints. Technical Report STAN-CS-80-784 (HPP-80-2), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1980. 230 pp.

Huet:1980:ERR

[801] Gerard Huet and Derek C. Oppen. Equations and rewrite rules: a survey. Technical Report STAN-CS-80-785 (PVG-15), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1980. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-80-785.html.

Knuth:1980:AMM

[802] Donald E. Knuth. Algorithms in modern mathematics and computer science. Technical Report STAN-CS-80-786 (AD-A089912), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1980. iv + 25 pp. URL http://i.stanford.

edu/pub/cstr/reports/cs/tr/80/786/CS-TR-80-786.pdf;   http:
//www-db.stanford.edu/TR/CS-TR-80-786.html.

Symm:1980:REBa

[803] H. J. Symm and James H. Wilkinson. Realistic error bounds for a simple eigenvalue and its associated eigenvector. Technical Report STAN-CS-80-787, Stanford University, Department of Computer Science, Stanford, CA, USA, 1980. ?? pp. Published as [1742].

McCarthy:1980:CFN

[804] John McCarthy. Circumscription — a form of non-monotonic reasoning. Technical Report STAN-CS-80-788 (AIM-334, AD-A086574), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1980. iii + 15 pp. URL http://www-db.

stanford.edu/TR/CS-TR-80-788.html;   https://apps.dtic.mil/
sti/tr/pdf/ADA086574.pdf.

Luckham:1980:AES

[805] David C. Luckham and Wolfgang Polak. ADA exceptions: specification and proof techniques. Technical Report STAN-CS-80-789, Stanford University, Department of Computer Science, Stanford, CA, USA, February

1980. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-80-789.
html.

Wiederhold:1980:DH

[806] Gio Wiederhold. Databases in healthcare. Technical Report STAN-CS-80-790, Stanford University, Department of Computer Science, Stanford,

CA, USA, March 1980. ?? pp. URL http://www-db.stanford.edu/
TR/CS-TR-80-790.html.

Wilcox:1980:MLM

[807] Clark R. Wilcox, Mary L. Dageforde, and Gregory A. Jirak. Mainsail language manual. Technical Report STAN-CS-80-791 (CSL 78-166), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1980. 247 pp. This reported was misnumbered as CS-80-790.

Wilcox:1980:MIO

[808] Clark R. Wilcox, Mary L. Dageforde, and Gregory A. Jirak. MAINSAIL implementation overview. Technical Report STAN-CS-80-792 (CSL 78-167), Stanford University, Department of Computer Science, Stanford,

CA, USA, March 1980. ?? pp. URL http://www-db.stanford.edu/
TR/CS-TR-80-792.html.

Barr:1980:RK

[809] Avron Barr and James Davidson. Representation of knowledge. Technical Report STAN-CS-80-793, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1980. 82 pp.

Tarjan:1980:RDC

[810] Robert Endre Tarjan. Recent developments in the complexity of combinatorial algorithms. Technical Report STAN-CS-80-794, Stanford University, Department of Computer Science, Stanford, CA, USA, June

1980. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-80-794.
html.

Knuth:1980:L

[811] Donald E. Knuth. The letter S. Technical Report STAN-CS-80-795, Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1980. iii + 34 pp. URL https://apps.dtic.mil/sti/tr/
pdf/ADA090470.pdf.

Samuel:1980:E

[812] Arthur L. Samuel. Essential E. Technical Report STAN-CS-80-796 (AIM-335), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1980. ii + 33 pp. URL http://www-db.stanford.

edu/TR/CS-TR-80-796.html.

Garcia-Molina:1980:ROT

[813] Hector Garcia-Molina and Gio Wiederhold. Read-only transactions in a distributed database. Technical Report STAN-CS-80-797, Stanford University, Department of Computer Science, Stanford, CA, USA, April

1980. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-80-797.
html.

Floyd:1980:CRE

[814] Robert W. Floyd and Jeffrey D. Ullman. The compilation of regular expressions into integrated circuits. Technical Report STAN-CS-80-798, Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1980. v + 28 pp. URL https://apps.dtic.mil/sti/tr/
pdf/ADA090507.pdf.

Friedman:1980:MAS

[815] Jerome H. Friedman, Eric Grosse, and Werner Stuetzle. Multidimensional additive spline approximation. Technical Report STAN-CS-80-799, Stanford University, Department of Computer Science, Stanford,

CA, USA, May 1980. ?? pp. URL http://www-db.stanford.edu/TR/
CS-TR-80-799.html.

Blum:1979:ASC

[816] Robert L. Blum. Automating the study of clinical hypotheses on a timeoriented data base: The RX Project. Technical Report STAN-CS-79-816, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1979. 12 pp. URL http://i.stanford.edu/TR/

CS-TR-79-816.html.

Vitter:1980:ACH

[817] Jeffrey Scott Vitter. Analysis of coalesced hashing. Technical Report STAN-CS-79-817, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1980. 111 pp.

Knuth:1980:DLC

[818] Donald E. Knuth. Deciphering a linear congruential encryption. Technical Report STAN-CS-80-800, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1980. 10 pp.

El-Masri:1980:DUI

[819] Ramez Aziz El-Masri. On the design, use, and integration of data models. Technical Report STAN-CS-80-801, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1980. 228 pp.

Polak:1980:TCS

[820] Wolfgang Heinz Polak. Theory of compiler specification and verification. Technical Report STAN-CS-80-802 (PVG-17, AD-A094604), Stanford University, Department of Computer Science, Stanford, CA, USA, May

1980. xiv + 281 pp. URL https://apps.dtic.mil/sti/citations/
ADA094604;   https://apps.dtic.mil/sti/pdfs/ADA094604.pdf.

VanWyk:1980:LTG

[821] Christopher John Van Wyk. A language for typesetting graphics. Technical Report STAN-CS-80-803, Stanford University, Department of Computer Science, Stanford, CA, USA, 1980 1980. may pp.

Brooks:1980:DCT

[822] Martin Brooks. Determining correctness by testing. Technical Report STAN-CS-80-804 (AIM-336), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1980. 59 pp.

Gennery:1980:MEE

[823] Donald B. Gennery. Modelling the environment of an exploring vehicle by means of stereo vision. Technical Report STAN-CS-80-805 (AIM-339, AD-A091081), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1980. viii + 144 pp. URL https://apps.

dtic.mil/sti/tr/pdf/ADA091081.pdf.

Coughran:1980:ASH

[824] William Marvin Coughran. On the approximate solution of hyperbolic initial-boundary value problems. Technical Report STAN-CS-80-806, Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1980. 177 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA091183.pdf.

Matula:1980:PRG

[825] David W. Matula and Danny Dolev. Path-regular graphs. Technical Report STAN-CS-80-807 (AD-A091123), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1980. 39 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-80-807.html.

McCarthy:1980:FRB

[826] John McCarthy, Thomas O. Binford, David C. Luckham, Zohar Manna, Richard W. Weyhrauch, and Les Earnest. Final report: Basic research in artificial intelligence and foundations of programming. Technical Report STAN-CS-80-808 (AIM-337, AD-A091183), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1980. iii +

73 pp. URL http://www-db.stanford.edu/TR/CS-TR-80-808.html;
 https://apps.dtic.mil/sti/tr/pdf/ADA091183.pdf. Edited by Les

Earnest.

Ohwovoriole:1980:EST

[827] Morgan S. Ohwovoriole. An extension of screw theory and its application to the automation of industrial assemblies. Technical Report STAN-CS-80-809 (AIM-338), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1980. 186 pp.

Steele:1980:LBA

[828] J. Michael Steele and Andrew C. Yao. Lower bounds for algebraic decision trees. Technical Report STAN-CS-80-810 (AD-A091124), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1980. 12 pp.

German:1980:ESD

[829] Steven M. German. An extended semantic definition of Pascal for proving the absence of common runtime errors. Technical Report STAN-CS-80-811 (PVG-18, AD-A091313), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1980. ?? pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-80-811.html.

Feigenbaum:1980:KEA

[830] Edward A. Feigenbaum. Knowledge engineering: The applied side of artificial intelligence. Technical Report STAN-CS-80-812 (HPP-80-14), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1980. vi + 14 pp. URL https:/

/apps.dtic.mil/sti/html/tr/ADA092574/index.html;   https://
apps.dtic.mil/sti/tr/pdf/ADA092574.pdf.

Moravec:1980:OAN

[831] Hans Peter Moravec. Obstacle avoidance and navigation in the real world by a seeing robot rover. Technical Report STAN-CS-80-813 (AIM-340, AD-A092604), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1980. 174 pp.

Aikins:1980:PPR

[832] Janice S. Aikins. Prototypes and production rules: a knowledge representation for computer consultations. Technical Report STAN-CS-80-814 (HPP-80-17, AD-A091177), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1980. viii + 204 pp.

URL https://apps.dtic.mil/sti/citations/ADA091177;   https:
//apps.dtic.mil/sti/pdfs/ADA091177.pdf.

Shortliffe:1980:TPM

[833] Edward H. Shortliffe. Two papers on medical computing — (1) medical cybernetics: The challenges of clinical computing, (2) consultation systems for physicians: The role of artificial intelligence techniques. Technical Report STAN-CS-80-815 (HPP-80-16), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1980. 56 pp.

Scherlis:1980:EPP

[834] William Louis Scherlis. Expression procedures and program derivation. Technical Report STAN-CS-80-818 (AIM-341, AD-A091187), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1980. x + 178 pp.

Goad:1980:CUM

[835] Christopher Alan Goad. Computational uses of the manipulation of formal proofs. Technical Report STAN-CS-80-819 (AD-A091180), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1980. vii + 122 pp.

vanMelle:1980:DIS

[836] William van Melle. A domain-independent system that aids in constructing knowledge-based consultation programs. Technical Report STAN-CS-80-820 (HPP-80-22), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1980. 192 pp.

West:1980:SUC

[837] Douglas B. West and Craig A. Tovey. Semiantichains and unichain coverings in direct products of partial orders. Technical Report STAN-CS-80-821, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1980. ?? pp. URL http://www-db.stanford.

edu/TR/CS-TR-80-821.html.

Aspvall:1980:EAC

[838] Bengt Aspvall. Efficient algorithms for certain satisfiability and linear programming problems. Technical Report STAN-CS-80-822, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1980. 59 pp.

Shaw:1980:KBR

[839] David Elliot Shaw. Knowledge-based retrieval on a relational database machine. Technical Report STAN-CS-80-823, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1980. 280 pp.

Mei:1980:LLC

[840] Tung Yun Mei. LCCD, a language for Chinese character design. Technical Report STAN-CS-80-824, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1980. ?? pp. URL http://www-

db.stanford.edu/TR/CS-TR-80-824.html.

Schnorr:1980:RAI

[841] C. P. Schnorr. Refined analysis and improvements on some factoring algorithm. Technical Report STAN-CS-80-825, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1980. 30 pp.

Wiederhold:1980:DAC

[842] Gio Wiederhold, Anne Beetem, and Garrett Short. A database approach to communication in VLSI design. Technical Report STAN-CS-80-826, Stanford University, Department of Computer Science, Stanford, CA,

USA, October 1980. ?? pp. URL http://www-db.stanford.edu/TR/
CS-TR-80-826.html.

Yao:1980:PCK

[843] Andrew Chi-Chih Yao. On the parallel computation for the knapsack problem. Technical Report STAN-CS-80-827, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1980. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-80-827.html.

Knuth:1980:BPL

[844] Donald E. Knuth and Michael F. Plass. Breaking paragraphs into lines. Technical Report STAN-CS-80-828, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1980. 66 pp.

Aspvall:1980:DTP

[845] Bengt Aspvall and Frank M. Liang. The dinner table problem. Technical Report STAN-CS-80-829, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1980. 13 pp. URL http://www-

db.stanford.edu/TR/CS-TR-80-829.html.

Matula:1980:TLT

[846] David W. Matula, Yossi Shiloach, and Robert E. Tarjan. Two linear-time algorithms for five-coloring a planar graph. Technical Report STAN-CS-80-830, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1980. ?? pp. URL http://www-db.stanford.edu/

TR/CS-TR-80-830.html.

Sleator:1980:AMN

[847] Daniel D. K. Sleator. An O(nm log n) algorithm for maximum network flow. Technical Report STAN-CS-80-831, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1980. 81 pp.

Dolev:1980:SWG

[848] Danny Dolev. Scheduling wide graphs. Technical Report STAN-CS-80-832, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1980. ?? pp. URL http://www-db.stanford.

edu/TR/CS-TR-80-832.html.

Gilbert:1980:GST

[849] John Russell Gilbert. Graph separator theorems and sparse Gaussian elimination. Technical Report STAN-CS-80-833, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1980. 104 pp.

Bjorstad:1980:NSB

[850] Petter E. Bjørstad. Numerical solution of the biharmonic equation. Technical Report STAN-CS-80-834 (SU326 P3070), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1980. 139

pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA104084.pdf.

Grosse:1980:AOE

[851] Eric H. Grosse. Approximation and optimization of electron density maps. Technical Report STAN-CS-80-835, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1980. 118 pp.

Spector:1980:PRO

[852] Alfred Z. Spector. Performing remote operations efficiently on a local computer network. Technical Report STAN-CS-80-850 (CSL 81-207), Stanford University, Department of Computer Science, Stanford, CA,

USA, December 1980. 23 pp. URL http://www-db.stanford.edu/TR/
CS-TR-80-850.html.

Genesereth:1980:RPI

[853] Michael R. Genesereth. The role of plans in intelligent teaching systems. Technical Report STAN-CS-81-842, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1980. 19 pp.

Hailpern:1980:VCP

[854] Brent T. Hailpern. Verifying concurrent processes using temporal logic. Technical Report STAN-CS-82-942, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1980. 114 pp.

Manna:1981:TVCa

[855] Zohar Manna and Amir Pnueli. Verification of concurrent programs, Part I: The temporal framework. Technical Report STAN-CS-81-836, Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1981. ii + 62 pp. URL http://www-db.stanford.edu/TR/
CS-TR-81-836.html.

Buchanan:1981:RES

[856] Bruce G. Buchanan. Research on expert systems. Technical Report STAN-CS-81-837, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1981. iv + 38 pp. URL http://www-db.

stanford.edu/TR/CS-TR-81-837.html.

Brown:1981:DPB

[857] Peter Brown. Dynamic program building. Technical Report STAN-CS-81-838, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1981. 13 pp. URL http://www-db.stanford.edu/

TR/CS-TR-81-838.html.

Samuel:1981:SW

[858] Arthur L. Samuel. Short WAITS. Technical Report STAN-CS-81-839, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1981. ix + 33 pp. URL http://i.stanford.

edu/pub/cstr/reports/cs/tr/81/839/CS-TR-81-839.pdf;   http:
//www-db.stanford.edu/TR/CS-TR-81-839.html.

Knuth:1981:VLL

[859] Donald E. Knuth. Verification of link-level protocols. Technical Report STAN-CS-81-840, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1981. iii + 6 pp. URL https://apps.dtic.

mil/sti/pdfs/ADA099140.pdf.

Knuth:1981:HAA

[860] Donald E. Knuth. Huffman’s algorithm via algebra. Technical Report STAN-CS-81-841 (AD-A099103), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1981. ii + 6 pp. URL

https://apps.dtic.mil/sti/tr/pdf/ADA099103.pdf.

Manna:1981:TVCb

[861] Zohar Manna and Amir Pnueli. Temporal verification of concurrent programs, Part II: Proving invariances. Technical Report STAN-CS-81-843, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1981. 30 pp.

Anonymous:1981:NPa

[862] Anonymous. NEVER PRINTED. Technical Report STAN-CS-81-844, Stanford University, Department of Computer Science, Stanford, CA, USA, ??? 1981. ?? pp.

Anonymous:1981:NPb

[863] Anonymous. NEVER PRINTED. Technical Report STAN-CS-81-845, Stanford University, Department of Computer Science, Stanford, CA, USA, ??? 1981. ?? pp.

Dolev:1981:BGS

[864] Danny Dolev. The Byzantine Generals strike again. Technical Report STAN-CS-81-846, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1981. 26 pp. URL http://www-db.

stanford.edu/TR/CS-TR-81-846.html.

Korth:1981:OLP

[865] Henry F. Korth. The optimal locking problem in a directed acyclic graph. Technical Report STAN-CS-81-847, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1981. 6 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-81-847.html.

Tang:1981:PIC

[866] Chih sung Tang. On the problem of inputting Chinese characters. Technical Report STAN-CS-81-848, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1981. 9 pp. URL http://www-

db.stanford.edu/TR/CS-TR-81-848.html.

Nishigaki:1981:EKC

[867] Tohru Nishigaki. Experiments on the Knee Criterion in a multiprogrammed computer system. Technical Report STAN-CS-81-849 (CSL 81-205), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1981. 28 pp. URL http://www-db.stanford.

edu/TR/CS-TR-81-849.html.

Wiederhold:1981:BIP

[868] Gio Wiederhold. Binding in information processing. Technical Report STAN-CS-81-851, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1981. 41 pp. URL http://www-db.stanford.

edu/TR/CS-TR-81-851.html.

Ullman:1981:VDR

[869] Jeffrey D. Ullman. A view of directions in relational database theory. Technical Report STAN-CS-81-852, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1981. 9 pp.

Maier:1981:CAH

[870] David Maier and Jeffrey D. Ullman. Connections in acyclic hypergraphs. Technical Report STAN-CS-81-853, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1981. 10 pp.

Dolev:1981:SPK

[871] Danny Dolev and Andrew C. Yao. On the security of public key protocols. Technical Report STAN-CS-81-854, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1981. 22 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-81-854.html.

Manna:1981:DSU

[872] Zohar Manna and Richard Waldinger. Deductive synthesis of the unification algorithm. Technical Report STAN-CS-81-855, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 51 pp.

Gabriel:1981:OPF

[873] Richard Paul Gabriel. An organization for programs in fluid domains. Technical Report STAN-CS-81-856, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 190 pp.

King:1981:QOS

[874] Jonathan Jay King. Query optimization by semantic reasoning. Technical Report STAN-CS-81-857, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 128 pp.

Pan:1981:BOC

[875] Pan and V. Y. The bit operation complexity of approximate evaluation of matrix and polynomial products using arithmetic. Technical Report STAN-CS-81-858, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 6 pp.

Pan:1981:ALC

[876] V. Y. Pan. The additive and logical complexities of linear and bilinear arithmetic algorithms. Technical Report STAN-CS-81-859, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 21 pp.

Boley:1981:CCO

[877] Daniel Boley. Computing the controllability/observability decomposition of a linear time-invariant dynamic system: a numerical approach.

Technical Report STAN-CS-81-860, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 81 pp.

Brooks:1981:SRA

[878] Rodney A. Brooks. Symbolic reasoning among 3-D models and 2-D images. Technical Report STAN-CS-81-861, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 181 pp.

Pan:1981:LBA

[879] V. Y. Pan. The lower bounds on the additive complexity of bilinear problems in terms of some algebraic quantities. Technical Report STAN-CS-81-862, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 4 pp.

Knuth:1981:PPS

[880] Donald E. Knuth and Allan A. Miller. A programming and problemsolving seminar. Technical Report STAN-CS-81-863, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1, 1981.

84 pp. URL http://www-db.stanford.edu/TR/CS-TR-81-863.html;
 http://www.ncstrl.org:8900/ncstrl/servlet/search?formname=
detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-81-863.

Doyle:1981:TSE

[881] Jon Doyle. Three short essays on decisions, reasons, and logics. Technical Report STAN-CS-81-864, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 19 pp.

Tang:1981:TUL

[882] Chih sung Tang. Toward a unified logical basis for programming languages. Technical Report STAN-CS-81-865, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 21 pp.

URL http://www-db.stanford.edu/TR/CS-TR-81-865.html.

German:1981:VAC

[883] Steven Mark German. Verifying the absence of common runtime errors in computer programs. Technical Report STAN-CS-81-866 (CSL 81-208, CSL-TR-208, PVG-19, AD-A109433), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. v + 179 pp. URL

https://apps.dtic.mil/sti/tr/pdf/ADA109433.pdf.

Luckham:1981:AAB

[884] David C. Luckham, Howard J. Larsen, David R. Stevenson, and Friedrich W. von Henke. ADAM — an Ada based language for multiprocessing. Technical Report STAN-CS-81-867, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1981. v +

71 pp. URL http://www-db.stanford.edu/TR/CS-TR-81-867.html;
 https://apps.dtic.mil/sti/pdfs/ADA110920.pdf.

Knuth:1981:LWE

[885] Donald E. Knuth. The last whole errata catalog. Technical Report STAN-CS-81-868, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1, 1981. 42 pp. URL http://www-db.

stanford.edu/TR/CS-TR-81-868.html;   http://www.ncstrl.org:
8900/ncstrl/servlet/search?formname=detail&id=oai%3Ancstrlh%
3Astan%3ASTAN%2F%2FCS-TR-81-868.

Tajnai:1981:CSC

[886] Carolyn E. Tajnai. Computer Science comprehensive examinations, 1978/79–1980/81. Technical Report STAN-CS-81-869, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1981.

xvi + 216 pp. URL http://www-db.stanford.edu/TR/CS-TR-81-869.
html.

Plass:1981:OPT

[887] Michael F. Plass. Optimal pagination techniques for automatic typesetting systems. Technical Report STAN-CS-81-870, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 77 pp.

Trickey:1981:GLP

[888] Howard W. Trickey. Good layouts for pattern recognizers. Technical Report STAN-CS-81-871, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1981. iv + 14 pp. URL http:/

/i.stanford.edu/pub/cstr/reports/cs/tr/81/871/CS-TR-81-871.
pdf;   http://www-db.stanford.edu/TR/CS-TR-81-871.html.

Manna:1981:SCP

[889] Zohar Manna and Pierre Wolper. Synthesis of communicating processes from temporal logic specifications. Technical Report STAN-CS-81-872, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1981. 28 pp.

Carr:1981:VMM

[890] Richard William Carr. Virtual memory management. Technical Report STAN-CS-81-873, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 238 pp.

Spector:1981:MAL

[891] Alfred Z. Spector. Multiprocessing architectures for local computer networks. Technical Report STAN-CS-81-874, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1981. 125 pp.

Hu:1981:CMC

[892] T. C. Hu and M. T. Shing. Computation of matrix chain products: Part I, part II. Technical Report STAN-CS-81-875, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1981.

124 pp. URL http://www-db.stanford.edu/TR/CS-TR-81-875.html.

Dolev:1981:LAE

[893] Danny Dolev and Howard W. Trickey. On linear area embedding of planar graphs. Technical Report STAN-CS-81-876, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1981. ii +

21 pp. URL ehttp://www-db.stanford.edu/TR/CS-TR-81-876.html;
 http://i.stanford.edu/pub/cstr/reports/cs/tr/81/876/CS-TR-
81-876.pdf.

Manna:1981:VSP

[894] Zohar Manna. Verification of sequential programs: Temporal axiomatization. Technical Report STAN-CS-81-877, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1981. 45 pp.

Maier:1981:MOS

[895] David Maier and Jeffrey D. Ullman. Maximal objects and the semantics of universal relation databases. Technical Report STAN-CS-81-878, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1981. 10 pp.

Masinter:1981:IVR

[896] Larry M. Masinter. Interlisp-VAX: a report. Technical Report STAN-CS-81-879, Stanford University, Department of Computer Science, Stanford,

CA, USA, August 1981. 12 pp. URL http://www-db.stanford.edu/
TR/CS-TR-81-879.html.

Mayr:1981:WSP

[897] Ernst W. Mayr. Well structured parallel programs are not easier to schedule. Technical Report STAN-CS-81-880, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1981. ii

+ 16 pp. URL http://www-db.stanford.edu/TR/CS-TR-81-880.html.

Ullman:1981:URS

[898] Jeffrey D. Ullman. The U.R. strikes back. Technical Report STAN-CS-81-881 (AD-A119145), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1981. iv + 15 pp. URL https://

apps.dtic.mil/sti/pdfs/ADA119145.pdf.

Pan:1981:FMM

[899] V. Pan. Fast matrix multiplication without APP algorithms. Technical Report STAN-CS-81-882, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1981. 30 pp.

Pepper:1981:PTA

[900] P. Pepper. On program transformations for abstract data types and concurrency. Technical Report STAN-CS-81-883, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1981. 37

pp. URL http://www-db.stanford.edu/TR/CS-TR-81-883.html.

Ceri:1981:ODD

[901] Stefano Ceri, Shamkant Navathe, and Gio Wiederhold. Optimal design of distributed databases. Technical Report STAN-CS-81-884, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 48 pp.

vanMelle:1981:EM

[902] William James van Melle, A. C. Scott, J. S. Bennett, and M. Peairs. The EMYCIN manual. Technical Report STAN-CS-81-885 (HPP-81-16), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1981. iii + 138 pp.

Knuth:1981:CMF

[903] Donald E. Knuth. The concept of a Meta-Font. Technical Report STAN-CS-81-886, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1981. 12 pp. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V16N1_1982_E.pdf.

Graham:1981:FCH

[904] Ronald L. Graham and Frances Yao. Finding the convex hull of a simple polygon. Technical Report STAN-CS-81-887, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1981. 9

pp. URL http://www-db.stanford.edu/TR/CS-TR-81-887.html.

Gropp:1981:NST

[905] William D. Gropp. Numerical solution of transport equations. Technical Report STAN-CS-81-888, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 108 pp.

Mujtaba:1981:AUM

[906] Shahid Mujtaba and Ron Goldman. AL users’ manual. Technical Report STAN-CS-81-889, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 168 pp. URL http://www-db.

stanford.edu/TR/CS-TR-81-889.html.

Higdon:1981:BCH

[907] Robert L. Higdon. Boundary conditions for hyperbolic systems of partial differential equations having multiple time scales. Technical Report STAN-CS-81-890, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1981. 136 pp.

Dietterich:1981:RCL

[908] T. G. Dietterich and B. G. Buchanan. The role of the critic in learning systems. Technical Report STAN-CS-81-891 (HPP-81-19, CSL 81-222, AD-A113479), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 23 pp. URL https://apps.dtic.

mil/sti/tr/pdf/ADA113479.pdf.

Clancey:1981:MBI

[909] William J. Clancey. Methodology for building an intelligent tutoring system. Technical Report STAN-CS-81-894 (HPP-81-18), Stanford University, Department of Computer Science, Stanford, CA, USA, October

1981. 55 pp. URL http://www-db.stanford.edu/TR/CS-TR-81-894.
html.

Clancey:1981:ERB

[910] William J. Clancey. The epistemology of a rule-based expert system: a framework for explanation. Technical Report STAN-CS-81-896 (HPP-81-17), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1981. vii + 60 pp. URL http://www-db.

stanford.edu/TR/CS-TR-81-896.html;   https://apps.dtic.mil/
sti/tr/pdf/ADA112672.pdf.

Whang:1981:SPD

[911] Kyu-Young Whang, Gio Wiederhold, and Daniel Sagalowicz. Separability as a physical database design methodology. Technical Report STAN-CS-81-898 (CSL-TR-222), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1981. iv + 60 pp. URL http://

www-db.stanford.edu/TR/CS-TR-81-898.html.

Paulson:1981:CGS

[912] Lawrence Paulson. A compiler generator for semantic grammars. Technical Report STAN-CS-82-893, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 166 pp.

Appelt:1981:PNL

[913] Douglas E. Appelt. Planning natural-language utterances to satisfy multiple goals. Technical Report STAN-CS-82-918, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 188 pp.

Roberts:1981:ETE

[914] Teresa L. Roberts. Evaluation of text editors. Technical Report STAN-CS-82-920, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1981. 188 pp.

Woods:1981:DPG

[915] Donald R. Woods. Drawing planar graphs. Technical Report STAN-CS-82-943, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1981. 58 pp.

Anonymous:1989:NP

[916] Anonymous. NEVER PRINTED. Technical Report STAN-CS-89-1298, Stanford University, Department of Computer Science, Stanford, CA, USA, ??? 1981. ?? pp.

Kuper:1982:ARA

[917] Gabriel M. Kuper. An algorithm for reducing acyclic hypergraphs. Technical Report STAN-CS-82-892, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1982. 9 pp. URL http://

www-db.stanford.edu/TR/CS-TR-82-892.html.

Novak:1982:GUM

[918] Gordon S. Novak, Jr. GLISP users’ manual. Technical Report STAN-CS-82-895 (HPP-82-1), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1982. 38 pp. URL http://www-

db.stanford.edu/TR/CS-TR-82-895.html.

Goad:1982:ACS

[919] Christopher Goad. Automatic construction of special purpose programs. Technical Report STAN-CS-82-897, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1982. 15 pp.

Bolstad:1982:AFD

[920] John H. Bolstad. An adaptive finite difference method for hyperbolic systems in one space dimension. Technical Report STAN-CS-82-899, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1982. 175 pp.

Blum:1982:DRC

[921] Robert L. Blum. Discovery and representation of causal relationships from a large time-oriented clinical database: The RX Project. Technical Report STAN-CS-82-900, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1982. 264 pp.

Fuchs:1982:OFC

[922] David R. Fuchs and Donald E. Knuth. Optimal font caching. Technical Report STAN-CS-82-901 (AD-A119439), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1982. ii +

19 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA119439.pdf;
https://dl.acm.org/doi/pdf/10.1145/2363.2367.

Manna:1982:SRP

[923] Zohar Manna and Richard Waldinger. Special relations in programsynthetic deduction. Technical Report STAN-CS-82-902, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1982. 75 pp.

McCarthy:1982:CMK

[924] John McCarthy. Coloring maps and the Kowalski Doctrine. Technical Report STAN-CS-82-903, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1982. 8 pp. URL http://www-db.

stanford.edu/TR/CS-TR-82-903.html.

LeVeque:1982:TSM

[925] Randall John LeVeque. Time-split methods for partial differential equations. Technical Report STAN-CS-82-904, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1982. 102 pp.

Trefethen:1982:WPS

[926] Lloyd N. Trefethen. Wave propagation and stability for finite difference schemes. Technical Report STAN-CS-82-905, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1982. 207 pp.

Nash:1982:TNM

[927] Stephen G. Nash. Truncated-Newton methods. Technical Report STAN-CS-82-906, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. 120 pp.

Mayr:1982:CA

[928] Ernst W. Mayr. Combinatorial algorithms I. Technical Report STAN-CS-82-907, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. 83 pp.

Clancey:1982:NRR

[929] William J. Clancey and Reed Letsinger. Neomycin: Reconfiguring a rulebased expert system for application to teaching. Technical Report STAN-CS-82-908 (HPP-81-2), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. 12 pp. URL http://www-db.

stanford.edu/TR/CS-TR-82-908.html.

London:1982:PRS

[930] Bob London and William J. Clancey. Plan recognition strategies in student modeling: prediction and description. Technical Report STAN-CS-82-909 (HPP-82-7), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. 13 pp. URL http://www-db.

stanford.edu/TR/CS-TR-82-909.html.

Clancey:1982:ETP

[931] William J. Clancey and Bruce G. Buchanan. Exploration of teaching and problem-solving strategies, 1979–1982. Technical Report STAN-CS-82-910 (HPP-82-8), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. 17 pp. URL http://www-db.stanford.

edu/TR/CS-TR-82-910.html.

Roberts:1982:BSC

[932] Barbara J. Roberts and Irris Marashian. Bibliography of Stanford Computer Science reports, 1963–1982. Technical Report STAN-CS-82-911, Stanford University, Department of Computer Science, Stanford, CA,

USA, May 1982. 59 pp. URL http://www-db.stanford.edu/TR/CS-
TR-82-911.html.

Vardi:1982:IFI

[933] Moshe Y. Vardi. The implication and finite implication problems for typed template dependencies. Technical Report STAN-CS-82-912, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. iv + 34 pp. URL http://i.stanford.

edu/pub/cstr/reports/cs/tr/82/912/CS-TR-82-912.pdf;   http:
//www-db.stanford.edu/TR/CS-TR-82-912.html.

Dietterich:1982:LII

[934] Thomas Dietterich, Bob London, Kenneth Clarkson, and Geoff Dromey. Learning and inductive inference. Technical Report STAN-CS-82-913 (HPP-82-10, AD-A122351), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. [various] pp. URL https:

//apps.dtic.mil/sti/pdfs/ADA122351.pdf. A section of the Hand-

book of Artificial Intelligence, edited by Paul R. Cohen and Edward A. Feigenbaum.

Guoan:1982:USM

[935] Gu Guoan and John Hobby. Using string matching to compress Chinese characters. Technical Report STAN-CS-82-914, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. ii + 15

pp. URL http://i.stanford.edu/pub/cstr/reports/cs/tr/82/914/
CS-TR-82-914.pdf;   http://www-db.stanford.edu/TR/CS-TR-82-
914.html.

Manna:1982:VCP

[936] Zohar Manna and Amir Pnueli. Verification of concurrent programs: proving eventualities by well-founded ranking. Technical Report STAN-CS-82-915, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. iii + 25 pp. URL http://www-db.stanford.

edu/TR/CS-TR-82-915.html.

Bent:1982:DWD

[937] Samuel Watkins Bent. Dynamic weighted data structures. Technical Report STAN-CS-82-916, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1982. vii + 75 + 2 pp. URL https:/

/apps.dtic.mil/sti/tr/pdf/ADA122046.pdf.

Mujtaba:1982:MSM

[938] Mohamed Shahid Mujtaba. Motion sequencing of manipulators. Technical Report STAN-CS-82-917, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1982. 307 pp.

Wall:1982:MBS

[939] David Wall. Mechanisms for broadcast and selective broadcast. Technical Report STAN-CS-82-919 (CSL 80-190), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1982. 122 pp.

Salisbury:1982:KFA

[940] J. Kenneth Salisbury. Kinematic and force analysis of articulated hands. Technical Report STAN-CS-82-921, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1982. 106 pp.

Suwa:1982:AVC

[941] Motoi Suwa, A. Carlisle Scott, and Edward H. Shortliffe. An approach to verifying completeness and consistency in a rule-based expert system. Technical Report STAN-CS-82-922 (HPP-81-5, AD-A120806), Stanford University, Department of Computer Science, Stanford, CA, USA, June

1982. iii + 19 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-
922.html;   https://apps.dtic.mil/sti/pdfs/ADA120806.pdf.

Wallis:1982:EPM

[942] Jerold W. Wallis and Edward H. Shortliffe. Explanatory power for medical expert systems: studies in the representation of causal relationships for clinical consultations. Technical Report STAN-CS-82-923, Stanford University, Department of Computer Science, Stanford, CA, USA, July

1982. 44 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-923.
html.

Berger:1982:AMR

[943] Marsha J. Berger. Adaptive mesh refinement for hyperbolic partial differential equations. Technical Report STAN-CS-82-924, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1982. 123 pp.

Wolper:1982:SCP

[944] Pierre L. Wolper. Synthesis of communicating processes from temporal logic specifications. Technical Report STAN-CS-82-925, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1982. 120 pp.

Buchanan:1982:PRB

[945] Bruce G. Buchanan and Richard O. Duda. Principles of rule-based expert systems. Technical Report STAN-CS-82-926 (HPP-82-14), Stanford University, Department of Computer Science, Stanford, CA, USA, August

1982. 58 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-926.
html.

Ullman:1982:CSM

[946] Jeffrey D. Ullman. Combining state machines and regular expressions for automatic synthesis of VLSI circuits. Technical Report STAN-CS-82-927, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1982. 14 pp. URL http://www-db.stanford.

edu/TR/CS-TR-82-927.html.

Kuhn:1982:AAM

[947] Ingeborg M. Kuhn, Gio Wiederhold, Jonathan E. Rodnick, Diane M. Ramsey-Klee, Sanford Benett, and Donald D. Beck. Automated ambulatory medical record systems in the U.S. Technical Report STAN-CS-82-928, Stanford University, Department of Computer Science, Stanford,

CA, USA, August 1982. 70 pp. URL http://www-db.stanford.edu/
TR/CS-TR-82-928.html.

Maier:1982:FR

[948] David Maier and Jeffrey D. Ullman. Fragments of relations. Technical Report STAN-CS-82-929, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1982. 11 pp.

Baker:1982:DEI

[949] Henry Harlyn Baker. Depth from edge and intensity based stereo. Technical Report STAN-CS-82-930 (AIM-347), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1982. viii + 90 pp.

Aikins:1982:PES

[950] Janice S. Aikins, John C. Kunz, Edward H. Shortliffe, and Robert J. Fallat. PUFF: an expert system for interpretation of pulmonary function data. Technical Report STAN-CS-82-931 (HPP-82-13), Stanford University, Department of Computer Science, Stanford, CA, USA, September

1982. iv + 21 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-
931.html.

Shortliffe:1982:ESR

[951] Edward H. Shortliffe and Lawrence M. Fagan. Expert systems research: modeling the medical decision making process. Technical Report STAN-CS-82-932 (HPP-82-3), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1982. iv + 23 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-82-932.html.

Klein:1982:AMS

[952] Shmuel T. Klein and Eli Shamir. An algorithmic method for studying percolation clusters. Technical Report STAN-CS-82-933, Stanford University, Department of Computer Science, Stanford, CA, USA, Septem-

ber 1982. 13 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-
933.html.

Barr:1982:USL

[953] A. Barr, P. Cohen, and L. Fagan. Understanding spoken language. Technical Report STAN-CS-82-934 (HPP-82-16), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1982. 52 pp.

Tappel:1982:PLA

[954] Tappel, Westfold, and Barr. Programming languages for AI research. Technical Report STAN-CS-82-935 (HPP-82-17), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1982. 90 pp.

Cohen:1982:MC

[955] Paul R. Cohen. Models of cognition. Technical Report STAN-CS-82-936 (HPP-82-18, AD-A222300), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1982. 81 pp.

Ballantyne:1982:AD

[956] Michael Ballantyne, W. W. Bledsoe, Jon Doyle, Robert C. Moore, Richard Pattis, and Stanley J. Rosenschein. Automatic deduction. Technical Report STAN-CS-82-937 (HPP-82-19, PB96-150883), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1982. 77–132 pp. Chapter XII of Volume III of the Handbook of Artificial Intelligence, edited by Paul R. Cohen and Edward A. Feigenbaum.

Kanade:1982:V

[957] Takeo Kanade. Vision. Technical Report STAN-CS-82-938 (HPP-82-20), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1982. 220 pp.

Cohen:1982:PPS

[958] Paul R. Cohen. Planning and problem solving. Technical Report STAN-CS-82-939 (HPP-82-21, AD-A222298), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1982. 515–567 pp.

Chapter XV of Volume III of the Handbook of Artificial Intelligence, edited by Paul R. Cohen and Edward A. Feigenbaum.

Ullman:1982:EUR

[959] Jeffrey D. Ullman, Moshe Y. Vardi, and David Maier. The equivalence of universal relation definitions. Technical Report STAN-CS-82-940, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1982. i + 27 pp. URL https://apps.dtic.mil/

sti/html/tr/ADA324622/;   https://apps.dtic.mil/sti/tr/pdf/
ADA324622.pdf.

Martin:1982:ILI

[960] Paul A. Martin. Integrating local information to understand dialog. Technical Report STAN-CS-82-941 (AIM-348), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1982. 131 pp.

Rowe:1982:MDI

[961] Neil C. Rowe. Modelling degrees of item interest for a general database query system. Technical Report STAN-CS-82-947, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1982. vii +

34 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-947.html.

Rowe:1982:TPR

[962] Neil C. Rowe. Three papers on rule-based estimation of statistics on databases. Technical Report STAN-CS-82-948, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1982. 38 pp.

Broder:1982:SN

[963] Andrei Z. Broder. The r-Stirling numbers. Technical Report STAN-CS-82-949, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1982. ii + 22 pp. URL http://www-db.stanford.

edu/TR/CS-TR-82-949.html.

Pratt:1982:FPS

[964] Vaughan Pratt. Five paradigm shifts in programming language design and their realization in Viron, a dataflow programming environment. Technical Report STAN-CS-82-951, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1982. ii + 9 pp. URL

http://www-db.stanford.edu/TR/CS-TR-82-951.html.

Moszkowski:1982:TLM

[965] Ben Moszkowski. A temporal logic for multi-level reasoning about hardware. Technical Report STAN-CS-82-952, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1982. 25 pp.

Buchanan:1982:PBW

[966] Bruce G. Buchanan. Partial bibliography of work on expert systems. Technical Report STAN-CS-82-953 (HPP-82-30), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1982.

15 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-953.html.

Manna:1982:HCT

[967] Zohar Manna and Amir Pnueli. How to cook a temporal proof system for your pet language. Technical Report STAN-CS-82-954, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1982. 14 pp.

Goldman:1982:DIM

[968] Ron Goldman. Design of an interactive manipulator programming environment. Technical Report STAN-CS-82-955 (AIM-350), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1982. 149 pp.

Barr:1982:AIC

[969] Avron Barr. Artificial intelligence: Cognition as computation. Technical Report STAN-CS-82-956 (HPP-82-29, AD-A131804), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1982. iv + 28 pp.

Zabala-Salelles:1982:IGO

[970] Ignacio Andres Zabala-Salelles. Interfacing with graphic objects. Technical Report STAN-CS-82-960, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1982. 146 pp.

Korth:1983:SUD

[971] Henry Korth, Gabriel Kuper, and Jeffrey Ullman. System/U: a database system based on the universal relation assumption. Technical Report STAN-CS-82-944, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1983. 17 pp.

Gilbert:1983:APT

[972] Erik James Gilbert. Algorithm partitioning tools for a high-performance multiprocessor. Technical Report STAN-CS-82-946, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1983. 133 pp.

Winston:1983:LPD

[973] Patrick H. Winston, Thomas O. Binford, Boris Katz, and Michael Lowry. Learning physical description from functional definitions, examples and precedents. Technical Report STAN-CS-82-950 (AIM-349), Stanford University, Department of Computer Science, Stanford, CA, USA, Jan-

uary 1983. 25 pp. URL http://www-db.stanford.edu/TR/CS-TR-82-
950.html.

Navathe:1983:VPP

[974] Navathe, Ceri, Wiederhold, and Dou. Vertical partitioning for physical and distribution design of databases. Technical Report STAN-CS-82-957, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1983. 33 pp.

Mulsant:1983:KED

[975] Benoit Mulsant and David Servan-Schreiber. Knowledge engineering: a daily activity on a hospital ward. Technical Report STAN-CS-82-998 (HPP-83-40), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1983. iii + 41 pp. URL http://www-

db.stanford.edu/TR/CS-TR-82-998.html.

Zwaenepoel:1983:PRP

[976] Willy Zwaenepoel and Keith A. Lantz. Perseus: Retrospective on a portable operating system. Technical Report STAN-CS-83-945 (CSL-TR-206), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1983. ii + 8 pp. URL http://www-db.

stanford.edu/TR/CS-TR-83-945.html.

Lantz:1983:TGG

[977] Keith A. Lantz, David R. Cheriton, and William I. Nowicki. Third generation graphics for distributed systems. Technical Report STAN-CS-83-958 (CSL Technical Report 235, AD-A126695), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1983. v

+ 36 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA126695.pdf.

Karplus:1983:CIP

[978] Kevin Karplus. CHISEL — an introduction to the programming language C for VLSI layout. Technical Report STAN-CS-83-959, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1983. 137 pp.

Arnold:1983:ASP

[979] R. David Arnold. Automated stereo perception. Technical Report STAN-CS-83-961 (AIM-351, AD-A130645), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1983. vii + 123 pp.

Berg:1983:BSC

[980] Kathryn A. Berg. Bibliography of Stanford Computer Science reports, 1963–1983. Technical Report STAN-CS-83-962, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1983. 65 pp.

URL http://www-db.stanford.edu/TR/CS-TR-83-962.html.

Halpern:1983:HSB

[981] Joseph Halpern, Zohar Manna, and Ben Moszkowski. A hardware semantics based on temporal intervals. Technical Report STAN-CS-83-963, Stanford University, Department of Computer Science, Stanford,

CA, USA, March 1983. 26 pp. URL http://www-db.stanford.edu/
TR/CS-TR-83-963.html.

Manna:1983:PPP

[982] Zohar Manna and Amir Pnueli. Proving precedence properties: The temporal way. Technical Report STAN-CS-83-964, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1983. ii + 38

pp. URL http://i.stanford.edu/pub/cstr/reports/cs/tr/83/964/
CS-TR-83-964.pdf;   http://www-db.stanford.edu/TR/CS-TR-83-
964.html.

Ghosh:1983:ATD

[983] Pijush K. Ghosh. An approach to type design and text composition in Indian scripts. Technical Report STAN-CS-83-965, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1983. 131

pp. URL http://www-db.stanford.edu/TR/CS-TR-83-965.html.

Ghosh:1983:FAL

[984] Pijush K. Ghosh and Charles A. Bigelow. A formal approach to lettershape description for type design. Technical Report STAN-CS-83-966, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1983. vi + 51 pp. URL http://i.stanford.

edu/pub/cstr/reports/cs/tr/83/966/CS-TR-83-966.pdf;   http:
//www-db.stanford.edu/TR/CS-TR-83-966.html.

Manna:1983:VCP

[985] Zohar Manna and Amir Pnueli. Verification of concurrent programs: a temporal proof system. Technical Report STAN-CS-83-967, Stanford University, Department of Computer Science, Stanford, CA, USA, June

1983. 92 pp. URL http://www-db.stanford.edu/TR/CS-TR-83-967.
html.

Whang:1983:PDD

[986] Kyu-Young Whang. A physical database design methodology using the property of separability. Technical Report STAN-CS-83-968, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1983. 271 pp.

Moszkowski:1983:RIT

[987] Ben Moszkowski and Zohar Manna. Reasoning in interval temporal logic. Technical Report STAN-CS-83-969, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1983. 18 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-83-969.html.

Moszkowski:1983:RAD

[988] Ben Moszkowski. Reasoning about digital circuits. Technical Report STAN-CS-83-970, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1983. 146 pp.

Ruggles:1983:LDS

[989] Lynn Ruggles. Letterform design systems. Technical Report STAN-CS-83-971, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1983. iv + 24 pp. URL http://i.stanford.

edu/pub/cstr/reports/cs/tr/83/971/CS-TR-83-971.pdf;   http:
//www-db.stanford.edu/TR/CS-TR-83-971.html.

Karlin:1983:ERE

[990] Anna R. Karlin, Howard W. Trickey, and Jeffrey D. Ullman. Experience with a regular expression compiler. Technical Report STAN-CS-83-972, Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1983. ii + 23 pp. URL http://www-db.stanford.edu/TR/
CS-TR-83-972.html.

Cheriton:1983:DVK

[991] David R. Cheriton and Willy Zwaenepoel. The distributed V kernel and its performance for diskless workstations. Technical Report STAN-CS-83-973, Stanford University, Department of Computer Science, Stanford,

CA, USA, July 1983. 12 pp. URL http://www-db.stanford.edu/TR/
CS-TR-83-973.html.

Hobby:1983:CMF

[992] John Hobby and Gu Guoan. A Chinese Meta-Font. Technical Report STAN-CS-83-974, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1983. ii + 22 pp. URL http://i.stanford.

edu/pub/cstr/reports/cs/tr/83/974/CS-TR-83-974.pdf;   http:
//www-db.stanford.edu/TR/CS-TR-83-974.html.

Rowe:1983:RBS

[993] Neil C. Rowe. Rule-based statistical calculations on a database abstract. Technical Report STAN-CS-83-975, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1983. 167 pp.

Beeri:1983:ADD

[994] Catriel Beeri and Moshe Vardi. On acyclic database decompositions. Technical Report STAN-CS-83-976, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1983. 11 pp.

Liang:1983:WHP

[995] Franklin Mark Liang. Word Hy-phen-a-tion by Com-put-er. Technical Report STAN-CS-83-977, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1983. v + 85 pp. URL https://

www.tug.org/docs/liang/liang-thesis.pdf.

Knuth:1983:LLM

[996] Donald E. Knuth. Lessons learned from METAFONT. Technical Report STAN-CS-83-978, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1983. 22 pp. URL https://journals.uc.

edu/index.php/vl/article/view/5401/4265.

Graham:1983:NDS

[997] M. H. Graham, A. O. Mendelzon, and M. Y. Vardi. Notions of dependency satisfaction. Technical Report STAN-CS-83-979, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1983. 35 pp.

Knuth:1983:WSS

[998] Donald E. Knuth. The WEB system of structured documentation. Technical Report STAN-CS-83-980, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1, 1983. iv + 206

pp. URL http://i.stanford.edu/pub/cstr/reports/cs/tr/83/980/
CS-TR-83-980.pdf;   http://www-db.stanford.edu/TR/CS-TR-83-
980.html;   http://www.ncstrl.org:8900/ncstrl/servlet/search?

formname=detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-83-

980. Version 2.3.

Knuth:1983:LP

[999] Donald E. Knuth. Literate programming. Technical Report STAN-CS-83-981, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1983. 15 pp.

Greene:1983:LFL

[1000] Daniel H. Greene. Labelled formal languages and their uses. Technical Report STAN-CS-83-982, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1983. 150 pp.

Ullman:1983:CTT

[1001] Jeffrey D. Ullman. A communication-time tradeoff. Technical Report STAN-CS-83-983, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1983. 11 pp.

Brown:1983:PEE

[1002] Harold Brown, Christopher Tong, and Gordon Foyster. Palladio: an exploratory environment for circuit design. Technical Report STAN-CS-83-984 (HPP-83-31, AD-A266696), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1983. 48 pp. URL https:

//apps.dtic.mil/sti/tr/pdf/ADA266696.pdf.

Samuel:1983:FGT

[1003] Arthur L. Samuel. First grade TEX: a beginner’s TEX manual. Technical Report STAN-CS-83-985, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1983. iv + 34 pp. URL http:/

/i.stanford.edu/pub/cstr/reports/cs/tr/83/985/CS-TR-83-985.
pdf;   http://www-db.stanford.edu/TR/CS-TR-83-985.html.

Cohen:1983:HRA

[1004] Paul R. Cohen. Heuristic reasoning about uncertainty: an artificial intelligence approach. Technical Report STAN-CS-83-986, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1983. 200 pp.

Ullman:1983:STA

[1005] Jeffrey D. Ullman. Some thoughts about supercomputer organization. Technical Report STAN-CS-83-987, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1983. 17 pp.

Mairson:1983:PCS

[1006] Harry George Mairson. The program complexity of searching a table. Technical Report STAN-CS-83-988, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1983. 85 pp.

Knuth:1983:PPS

[1007] Donald E. Knuth and Joseph S. Weening. A programming and problem-solving seminar. Technical Report STAN-CS-83-989, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1983. vi + 91 pp. URL http://www-db.

stanford.edu/TR/CS-TR-83-989.html;   http://www.ncstrl.org:
8900/ncstrl/servlet/search?formname=detail&id=oai%3Ancstrlh%
3Astan%3ASTAN%2F%2FCS-TR-83-989.

Hobby:1983:PPS

[1008] John D. Hobby and Donald E. Knuth. A programming and problem-solving seminar. Technical Report STAN-CS-83-990, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1983. vii + 61 pp. URL http://www-db.

stanford.edu/TR/CS-TR-83-990.html;   http://www.ncstrl.org:
8900/ncstrl/servlet/search?formname=detail&id=oai%3Ancstrlh%
3Astan%3ASTAN%2F%2FCS-TR-83-990.

Morgensteren:1983:PAA

[1009] Moshe Morgensteren and Eli Shamir. Parallel algorithms for arithmetics, irreducibility and factoring of GFq-polynomials. Technical Report STAN-CS-83-991, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1983. 10 pp. URL http://www-

db.stanford.edu/TR/CS-TR-83-991.html.

Ketonen:1983:LIP

[1010] Jussi Ketonen and Joseph S. Weening. The language of an interactive proof checker. Technical Report STAN-CS-83-992, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1983.

34 pp. URL http://www-db.stanford.edu/TR/CS-TR-83-992.html.

Lansky:1983:SAC

[1011] Amy Lansky. Specification and analysis of concurrency. Technical Report STAN-CS-83-993, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1983. 290 pp.

Chapiro:1983:SRP

[1012] Daniel M. Chapiro. Sorting by recursive partitioning. Technical Report STAN-CS-83-994, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1983. 32 pp. URL http://www-db.

stanford.edu/TR/CS-TR-83-994.html.

Clancey:1983:AAC

[1013] William J. Clancey. The advantages of abstract control knowledge in expert system design. Technical Report STAN-CS-83-995 (HPP-83-17, AD-A139978), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1983. v + 14 pp. URL http://www-db.

stanford.edu/TR/CS-TR-83-995.html;   https://apps.dtic.mil/
sti/tr/pdf/ADA139978.pdf.

Hasling:1983:SED

[1014] Diane Warner Hasling, William J. Clancey, and Glenn Rennels. Strategic explanations for a diagnostic consultation system. Technical Report STAN-CS-83-996 (HPP-83-41, AD-A140014), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1983. ii

+ 23 pp. URL http://www-db.stanford.edu/TR/CS-TR-83-996.html;
 https://apps.dtic.mil/sti/tr/pdf/ADA140014.pdf.

Clancey:1983:G

[1015] William J. Clancey. GUIDON. Technical Report STAN-CS-83-997 (HPP-83-42), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1983. 13 pp.

Phillips:1983:SDP

[1016] Jorge Phillips. Self-described programming environments — an application of a theory of design to programming systems. Technical Report STAN-CS-84-1008, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1983. 262 pp.

Ullman:1984:ILQ

[1017] Jeffrey D. Ullman. Implementation of logical query languages for databases. Technical Report STAN-CS-84-1000, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1984. 41 pp.

Brinkley:1984:UTD

[1018] James F. Brinkley. Ultrasonic three-dimensional organ modelling. Technical Report STAN-CS-84-1001, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1984. 141 pp.

VanGelder:1984:STN

[1019] Allen Van Gelder. A satisfiability tester for non-clausal propositional calculus. Technical Report STAN-CS-84-1002, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1984. 14 pp.

Anderson:1984:PGA

[1020] Richard Anderson and Ernst Mayr. Parallelism and greedy algorithms. Technical Report STAN-CS-84-1003, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1984. 17 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-84-1003.html.

Goldschlager:1984:CTH

[1021] Leslie M. Goldschlager. A computational theory of higher brain function. Technical Report STAN-CS-84-1004, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1984. 29 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-84-1004.html.

Manna:1984:APP

[1022] Zohar Manna and Amir Pnueli. Adequate proof principles for invariance and liveness properties of concurrent programs. Technical Report STAN-CS-84-1005, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1984. 35 pp. URL http://www-db.

stanford.edu/TR/CS-TR-84-1005.html.

Ketonen:1984:EIP

[1023] Jussi Ketonen and Joseph S. Weening. EKL — an interactive proof checker user’s reference manual. Technical Report STAN-CS-84-1006, Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1984. 55 pp. URL http://www-db.stanford.edu/TR/CS-
TR-84-1006.html.

Gabriel:1984:QBM

[1024] Richard P. Gabriel and John McCarthy. Queue-based multi-processing Lisp. Technical Report STAN-CS-84-1007, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1984. 35 pp. URL

http://www-db.stanford.edu/TR/CS-TR-84-1007.html.

Sagiv:1984:CTC

[1025] Yehoshua Sagiv and Jeffrey D. Ullman. Complexity of a top-down capture rule. Technical Report STAN-CS-84-1009, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1984. ii +

35 pp. URL http://i.stanford.edu/pub/cstr/reports/cs/tr/84/
1009/CS-TR-84-1009.pdf;   http://www-db.stanford.edu/TR/CS-
TR-84-1009.html.

Lazowska:1984:FAP

[1026] Edward D. Lazowska, John Zahorjan, David R. Cheriton, and Willy Zwaenepoel. File access performance of diskless workstations. Technical Report STAN-CS-84-1010, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1984. 29 pp.

Cheriton:1984:OMI

[1027] David R. Cheriton and Willy Zwaenepoel. One-to-many interprocess communication in the V-System. Technical Report STAN-CS-84-1011, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1984. 8 pp.

Malachi:1984:TDT

[1028] Yonathan Malachi, Zohar Manna, and Richard Waldinger. TABLOG: The deductive-tableau programming language. Technical Report STAN-CS-84-1012, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1984. 16 pp. URL http://www-db.stanford.

edu/TR/CS-TR-84-1012.html.

Desarmenien:1984:HRT

[1029] Jacques Desarmenien. How to run TEX in French. Technical Report STAN-CS-84-1013, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1984. 42 pp.

Anderson:1984:PCP

[1030] Richard Anderson and Ernst W. Mayr. A P-complete problem and approximations to it. Technical Report STAN-CS-84-1014, Stanford University, Department of Computer Science, Stanford, CA, USA, Septem-

ber 1984. 12 pp. URL http://www-db.stanford.edu/TR/CS-TR-84-
1014.html.

Kirousis:1984:CRP

[1031] Lefteris Kirousis and Christos Papadimitriou. The complexity of recognizing polyhedral scenes. Technical Report STAN-CS-84-1015, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1984. 33 pp.

Papadimitriou:1984:IPC

[1032] Christos Papadimitriou and John Tsitsiklis. Intractable problems in control theory. Technical Report STAN-CS-84-1016, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1984. 23 pp.

Papadimitriou:1984:TPB

[1033] Christos Papadimitriou and John Tsitsiklis. The throughput of a precedence-based queuing discipline. Technical Report STAN-CS-84-1017, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1984. 9 pp.

Clancey:1984:CPS

[1034] William J. Clancey. Classification problem solving. Technical Report STAN-CS-84-1018, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1984. iv + 23 pp. URL http://www-db.

stanford.edu/TR/CS-TR-84-1018.html.

Lifschitz:1984:SRC

[1035] Vladimir Lifschitz. Some results on circumscription. Technical Report STAN-CS-84-1019, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1984. 14 pp.

Lowe:1984:POV

[1036] David Lowe. Perceptual organization and visual recognition. Technical Report STAN-CS-84-1020, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1984. 152 pp.

Kunz:1984:UAI

[1037] John C. Kunz. Use of artificial intelligence and simple mathematics. Technical Report STAN-CS-84-1021, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1984. 182 pp.

Konolige:1984:DMB

[1038] Kurt Konolige. A deduction model of belief and its logics. Technical Report STAN-CS-84-1022, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1984. 312 pp.

Gordon:1984:MME

[1039] Jean Gordon and Edward H. Shortliffe. A method for managing evidential reasoning in a hierarchical hypothesis space. Technical Report STAN-CS-84-1023, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1984. 41 pp. URL http://www-

db.stanford.edu/TR/CS-TR-84-1023.html.

Upfal:1984:HSM

[1040] Eli Upfal and Avi Wigderson. How to share memory in a distributed system. Technical Report STAN-CS-84-1024, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1984. 19

pp. URL http://www-db.stanford.edu/TR/CS-TR-84-1024.html.

Helmbold:1984:FSA

[1041] David Helmbold and Ernst Mayr. Fast scheduling algorithms on parallel computers. Technical Report STAN-CS-84-1025, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1984.

31 pp. URL http://www-db.stanford.edu/TR/CS-TR-84-1025.html.

Chapiro:1984:GAL

[1042] Daniel M. Chapiro. Globally-asynchronous locally-synchronous systems. Technical Report STAN-CS-84-1026, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1984. 136 pp.

Knuth:1984:TTT

[1043] Donald E. Knuth. A torture test for TEX. Technical Report STAN-CS-84-1027, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1984. iv + 142

pp. URL http://www-db.stanford.edu/TR/CS-TR-84-1027.html;
 http://www.ncstrl.org:8900/ncstrl/servlet/search?formname=
detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-84-1027.

Hochschild:1984:PGA

[1044] Peter H. Hochschild, Ernst W. Mayr, and Alan R. Siegel. Parallel graph algorithms. Technical Report STAN-CS-84-1028, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1984.

57 pp. URL http://www-db.stanford.edu/TR/CS-TR-84-1028.html.

Dietterich:1984:CPT

[1045] Thomas G. Dietterich. Constraint propagation techniques for theorydriven data interpretation. Technical Report STAN-CS-84-1030, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1984. 180 pp.

Cooper:1984:NCB

[1046] Gregory Floyd Cooper. NESTOR: a computer-based medical diagnostic aid that integrates causal and probabilistic knowledge. Technical Report STAN-CS-84-1031 (HPP-84-48, AD-A152046), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1984. xiii

+ 239 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA152046.pdf.

Genesereth:1984:SPD

[1047] Michael R. Genesereth, Matthew L. Ginsberg, and Jeffrey S. Rosenschein. Solving the Prisoner’s Dilemma. Technical Report STAN-CS-84-1032, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1984. 12 pp. URL http://www-db.stanford.edu/

TR/CS-TR-84-1032.html.

Gischer:1984:POA

[1048] Jay L. Gischer. Partial orders and the axiomatic theory of shuffle. Technical Report STAN-CS-84-1033, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1984. 46 pp.

Hayes-Roth:1984:BAB

[1049] Barbara Hayes-Roth. BB1: an architecture for blackboard systems that control, explain, and learn about their own behavior. Technical Report STAN-CS-84-1034, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1984. ?? pp. URL http://www-

db.stanford.edu/TR/CS-TR-84-1034.html.

Ullman:1984:CFW

[1050] Jeffrey Ullman, Harry Mairson, Danny Dolev, and David Maier. Correcting faults in write-once memory. Technical Report STAN-CS-84-999, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1984. 5 pp.

Ginsberg:1984:C

[1051] Matthew L. Ginsberg. Counterfactuals. Technical Report STAN-CS-85-1029, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1984. 13 pp.

Gardner:1984:AIA

[1052] Anne von der Leith Gardner. An artificial intelligence approach to legal reasoning. Technical Report STAN-CS-85-1045, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1984. 205 pp.

Apers:1983:TCS

[1053] Peter M. G. Apers and Gio Wiederhold. Transaction classification to survive a network partition. Technical Report STAN-CS-85-1053, Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1983. ii + 24 pp. URL http://i.stanford.edu/pub/
cstr/reports/cs/tr/85/1053/CS-TR-85-1053.pdf;   http://www-
db.stanford.edu/TR/CS-TR-85-1053.html.

Ossher:1984:NPS

[1054] Harold L. Ossher. A new program structuring mechanism based on layered graphs. Technical Report STAN-CS-85-1078, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1984. 247 pp.

Hayes-Roth:1985:BAB

[1055] Barbara Hayes-Roth. BB1: an architecture for blackboard systems that control and explain, and learn about their own behavior. Technical Report STAN-CS-85-1034, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1985. 22 pp.

Finger:1985:RDA

[1056] J. J. Finger and Michael R. Genesereth. RESIDUE: a deductive approach to design synthesis. Technical Report STAN-CS-85-1035, Stanford University, Department of Computer Science, Stanford, CA, USA, January

1985. 20 pp. URL http://www-db.stanford.edu/TR/CS-TR-85-1035.
html.

Hayes-Roth:1985:LCH

[1057] Barbara Hayes-Roth and Michael Hewett. Learning control heuristics in BB1. Technical Report STAN-CS-85-1036, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1985. 12 pp.

URL http://www-db.stanford.edu/TR/CS-TR-85-1036.html.

MacKinlay:1985:ELC

[1058] Jock MacKinlay and Michael R. Genesereth. Expressiveness and language choice. Technical Report STAN-CS-85-1037, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1985. 17

pp. URL http://www-db.stanford.edu/TR/CS-TR-85-1037.html.

Yao:1985:UHO

[1059] Andrew C. Yao. Uniform hashing is optimal. Technical Report STAN-CS-85-1038, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1985. 9 pp. URL http://www-db.

stanford.edu/TR/CS-TR-85-1038.html.

Feigenbaum:1985:PTA

[1060] Joan Feigenbaum, John Hershberger, and Alejandro A. Schaffer. A polynomial time algorithm for finding the prime factors of cartesian product graphs. Technical Report STAN-CS-85-1039, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1985. 27 pp.

Keller:1985:URD

[1061] Arthur M. Keller. Updating relational databases through views. Technical Report STAN-CS-85-1040, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1985. 119 pp.

Blicher:1985:EDG

[1062] A. Peter Blicher. Edge detection and geometric methods in computer vision. Technical Report STAN-CS-85-1041, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1985. 266 pp.

Rosenschein:1985:DAR

[1063] Jeffrey S. Rosenschein and Michael R. Genesereth. Deals among rational agents. Technical Report STAN-CS-85-1042, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1985. 25 pp.

Karp:1985:CPM

[1064] Richard M. Karp, Eli Upfal, and Avi Wigderson. Constructing a perfect matching is in random NC. Technical Report STAN-CS-85-1043, Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1985. 15 pp. URL http://www-db.stanford.edu/TR/CS-TR-85-
1043.html.

Manna:1985:OBS

[1065] Zohar Manna and Richard Waldinger. The origins of the binary-search paradigm. Technical Report STAN-CS-85-1044 (AD-A154367), Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1985. iii + 23 pp. URL https://apps.dtic.mil/sti/tr/pdf/

ADA154367.pdf. See revised version [1066].

Manna:1986:OBS

[1066] Zohar Manna and Richard Waldinger. The origins of the binary-search paradigm. SRI Technical Note 351R (AD-A193996), SRI International, Menlo Park, CA 94025, USA, October 1986. i + 50 pp. URL https://

apps.dtic.mil/sti/tr/pdf/ADA193996.pdf. Revision of [1065].

Ullman:1985:TAT

[1067] Jeffrey D. Ullman and Allen Van Gelder. Testing applicability of topdown capture rules. Technical Report STAN-CS-85-1046, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1985. 13 pp.

Hobby:1985:SEC

[1068] John D. Hobby. Smooth, easy to compute interpolating splines. Technical Report STAN-CS-85-1047, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1985. ii + 14 pp. URL

http://i.stanford.edu/pub/cstr/reports/cs/tr/85/1047/CS-TR-
85-1047.pdf;   http://www-db.stanford.edu/TR/CS-TR-85-1047.
html.

Pratt:1985:SCO

[1069] Vaughan Pratt. Some constructions for order-theoretic models of concurrency. Technical Report STAN-CS-85-1048, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1985. 17 pp.

URL http://www-db.stanford.edu/TR/CS-TR-85-1048.html.

Pratt:1985:PMP

[1070] Vaughan Pratt. The Pomset model of parallel processes: Unifying the temporal and the special. Technical Report STAN-CS-85-1049, Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1985. 20 pp. URL http://www-db.stanford.edu/TR/CS-TR-
85-1049.html.

Hershberger:1985:FSA

[1071] John Hershberger and Ernst Mayr. Fast sequential algorithms to find shuffle-minimizing and shortest paths in a shuffle-exchange network. Technical Report STAN-CS-85-1050, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1985. 18 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-85-1050.html.

Manna:1985:SRA

[1072] Zohar Manna and Richard Waldinger. Special relations in automated deduction. Technical Report STAN-CS-85-1051, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1985. 63 pp.

URL http://www-db.stanford.edu/TR/CS-TR-85-1051.html.

Tajnai:1985:FSV

[1073] Carolyn E. Tajnai and Fred Terman. the father of silicon valley. Technical Report STAN-CS-85-1052, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1985. 19 pp.

Broder:1985:WRM

[1074] Andrei Zary Broder. Weighted random mappings; properties and applications. Technical Report STAN-CS-85-1054, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1985. 77 pp.

Haddad:1985:PPS

[1075] Ramsey W. Haddad and Donald E. Knuth. A programming and problem-solving seminar. Technical Report STAN-CS-85-1055, Stanford University, Department of Computer Science, Stanford, CA, USA,

June 1, 1985. iv + 103 pp. URL http://www-db.stanford.edu/
TR/CS-TR-85-1055.html;   http://www.ncstrl.org:8900/ncstrl/
servlet/search?formname=detail&id=oai%3Ancstrlh%3Astan%3ASTAN%
2F%2FCS-TR-85-1055.

Abadi:1985:NTD

[1076] Martin Abadi and Zohar Manna. Nonclausal temporal deduction. Technical Report STAN-CS-85-1056, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1985. ii + 17 pp. URL

http://www-db.stanford.edu/TR/CS-TR-85-1056.html.

Mason:1985:MEP

[1077] Ian A. Mason and Carolyn L. Talcott. Memories of S-expressions proving properties of Lisp-like programs that destructively alter memory. Technical Report STAN-CS-85-1057, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1985. 46 pp.

Cheriton:1985:HGM

[1078] David R. Cheriton and Stephen E. Deering. Host groups: a multicast extension for datagram internetworks. Technical Report STAN-CS-85-1058, Stanford University, Department of Computer Science, Stanford,

CA, USA, July 1985. i + 8 pp. URL http://www-db.stanford.edu/
TR/CS-TR-85-1058.html.

Trickey:1985:CPP

[1079] Howard Wellington Trickey. Compiling Pascal programs into silicon. Technical Report STAN-CS-85-1059, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1985. x + 184 pp.

Talcott:1985:ERT

[1080] Carolyn L. Talcott. The essence of Rum: a theory of the intensional and extensional aspects of Lisp-like computation. Technical Report STAN-CS-85-1060 (AD-A327435), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1985. viii + 237 pp. URL

https://apps.dtic.mil/sti/tr/pdf/ADA327435.pdf.

Knuth:1985:TP

[1081] Donald E. Knuth. Theory and practice. Technical Report STAN-CS-85-1061, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1985. 10 pp.

Keller:1985:CSC

[1082] Arthur M. Keller. Computer science comprehensive examinations, 1981/ 82–1984/85. Technical Report STAN-CS-85-1062, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1985. 294

pp. URL http://www-db.stanford.edu/TR/CS-TR-85-1062.html.

Smith:1985:CRI

[1083] David Smith, Michael Genesereth, and Matthew Ginsberg. Controlling recursive inference. Technical Report STAN-CS-85-1063, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1985. 58 pp.

Ginsberg:1985:DP

[1084] Matthew L. Ginsberg. Decision procedures. Technical Report STAN-CS-85-1064, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1985. 21 pp.

Clancey:1985:RSC

[1085] William J. Clancey. Review of Sowa’s “Conceptual Structures”. Technical Report STAN-CS-85-1065, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1985. 22 pp. URL http://

www-db.stanford.edu/TR/CS-TR-85-1065.html.

Clancey:1985:HC

[1086] William J. Clancey. Heuristic classification. Technical Report STAN-CS-85-1066, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1985. 86 pp. URL http://www-db.stanford.

edu/TR/CS-TR-85-1066.html.

Clancey:1985:ARE

[1087] William J. Clancey. Acquiring, representing, and evaluating a competence model of diagnostic strategy. Technical Report STAN-CS-85-1067, Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1985. 94 pp. URL http://www-db.stanford.edu/TR/CS-
TR-85-1067.html.

Richer:1985:GWG

[1088] Mark H. Richer and William J. Clancey. GUIDON-WATCH: a graphic interface for viewing a knowledge-based system. Technical Report STAN-CS-85-1068, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1985. v + 33 pp. URL http://www-

db.stanford.edu/TR/CS-TR-85-1068.html.

Kuper:1985:LDM

[1089] Gabriel M. Kuper. The logical data model: a new approach to database logic. Technical Report STAN-CS-85-1069, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. 107 pp.

Hobby:1985:DBT

[1090] John D. Hobby. Digitized brush trajectories. Technical Report STAN-CS-85-1070, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. 125 pp. URL https://

searchworks.stanford.edu/view/1181669.

Greiner:1985:LUA

[1091] Russell Greiner. Learning by understanding analogies. Technical Report STAN-CS-85-1071, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. 423 pp.

Mayr:1985:PPS

[1092] Ernst W. Mayr, Richard J. Anderson, and Peter H. Hochschild. A programming and problem-solving seminar. Technical Report STAN-CS-85-1072, Stanford University, Department of Computer Science, Stanford,

CA, USA, October 1985. 68 pp. URL http://www-db.stanford.edu/
TR/CS-TR-85-1072.html.

Hochschild:1985:REP

[1093] Peter H. Hochschild. Resource-efficient parallel algorithms. Technical Report STAN-CS-85-1073, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. 90 pp.

Southall:1985:DNT

[1094] Richard Southall. Designing new typefaces with Metafont. Technical Report STAN-CS-85-1074, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. ii + 37 pp. URL

http://i.stanford.edu/pub/cstr/reports/cs/tr/85/1074/CS-TR-
85-1074.pdf;   http://www-db.stanford.edu/TR/CS-TR-85-1074.
html.

Buchanan:1985:ESWa

[1095] Bruce G. Buchanan. Expert systems: Working systems and the research literature. Technical Report STAN-CS-85-1075, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1985. 26

pp. URL http://www-db.stanford.edu/TR/CS-TR-85-1075.html.

Buchanan:1985:SAK

[1096] Bruce G. Buchanan. Some approaches to knowledge acquisition. Technical Report STAN-CS-85-1076, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1985. 25 pp. URL http://www-

db.stanford.edu/TR/CS-TR-85-1076.html. This paper is superseded

by STAN-CS-86-1094 [1112].

McCarthy:1985:ACF

[1097] John McCarthy. Applications of circumscription to formalizing common sense knowledge. Technical Report STAN-CS-85-1077, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. 29 pp.

Helmbold:1985:TPS

[1098] David Helmbold and Ernst Mayr. Two-processor scheduling is in NC. Technical Report STAN-CS-85-1079, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1985. 18 pp. URL

http://www-db.stanford.edu/TR/CS-TR-85-1079.html.

Russell:1985:CGM

[1099] Stuart Russell. The compleat guide to MRS. Technical Report STAN-CS-85-1080, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1985. 126 pp. URL http://www-db.stanford.

edu/TR/CS-TR-85-1080.html.

Rosenschein:1985:RIC

[1100] Jeffrey Solomon Rosenschein. Rational interaction: Cooperation among intelligent agents. Technical Report STAN-CS-85-1081, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1985. 133 pp.

Nowicki:1985:PFD

[1101] William I. Nowicki. Partitioning of function in a distributed graphics system. Technical Report STAN-CS-85-1082 (CSL-85-282, AD-A16634), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1985. vii + 136 pp. URL https://apps.dtic.mil/sti/

tr/pdf/ADA166935.pdf.

Zwaenepoel:1985:MPL

[1102] Willy Zwaenepoel. Message passing on a local network. Technical Report STAN-CS-85-1083, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1985. 112 pp.

Edighoffer:1985:TDB

[1103] Judy L. Edighoffer and Keith A. Lantz. Taliesin: a distributed bulletin board system. Technical Report STAN-CS-85-1084, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985.

11 pp. URL http://www-db.stanford.edu/TR/CS-TR-85-1084.html.

Lantz:1985:TUD

[1104] Keith A. Lantz, Judy L. Edighoffer, and Bruce L. Hitson. Towards a universal directory service. Technical Report STAN-CS-85-1086, Stanford University, Department of Computer Science, Stanford, CA, USA,

August 1985. 19 pp. URL http://www-db.stanford.edu/TR/CS-TR-
85-1086.html.

Theimer:1985:PRE

[1105] Marvin M. Theimer, Keith A. Lantz, and David R. Cheriton. Preemptable remote execution facilities for the V-system. Technical Report STAN-CS-85-1087, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. 15 pp. URL http://www-

db.stanford.edu/TR/CS-TR-85-1087.html.

VanGelder:1985:MPF

[1106] Allen Van Gelder. A message passing framework for logical query evaluation. Technical Report STAN-CS-85-1088, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1985. 18 pp.

Ullman:1985:PCL

[1107] Jeffrey D. Ullman and Allen Van Gelder. Parallel complexity of logical query programs. Technical Report STAN-CS-85-1089, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1985. 42 pp.

Yue:1985:CAS

[1108] Kaizhi Yue. Constructing and analyzing specifications of real world systems. Technical Report STAN-CS-86-1090, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1985. 290 pp.

Fu:1985:LOL

[1109] Li-Min Fu. Learning object-level and meta-level knowledge in expert systems. Technical Report STAN-CS-86-1091, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1985. 229 pp.

Anderson:1985:CPA

[1110] Richard Anderson. The complexity of parallel algorithms. Technical Report STAN-CS-86-1092, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1985. 72 pp.

Subramanian:1985:GRL

[1111] Devika Subramanian and Bruce G. Buchanan. A general reading list for artificial intelligence. Technical Report STAN-CS-86-1093, Stanford University, Department of Computer Science, Stanford, CA, USA, December

1985. 66 pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1093.
html.

Buchanan:1985:ESWb

[1112] Bruce G. Buchanan. Expert systems: Working systems and the research literature. Technical Report STAN-CS-86-1094, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1985. 57

pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1094.html.

Malik:1985:ILD

[1113] Jitendra Malik. Interpreting line drawings of curved objects. Technical Report STAN-CS-86-1099, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1985. 138 pp.

Malachi:1985:TNA

[1114] Yonathan Malachi, Zohar Manna, and Richard Waldinger. Tablog — a new approach to logic programming. Technical Report STAN-CS-86-1110, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1985. 22 pp.

Fu:1985:IKA

[1115] Li-Min Fu and Bruce G. Buchanan. Inductive knowledge acquisition for rule-based expert systems. Technical Report STAN-CS-86-1116 (KSL-85-42), Stanford University, Department of Computer Science, Stanford,

CA, USA, October 1985. 36 pp. URL http://www-db.stanford.edu/
TR/CS-TR-86-1116.html.

Lantz:1985:ESD

[1116] Keith A. Lantz, William I. Nowicki, and Marvin M. Theimer. An empirical study of distributed application performance. Technical Report STAN-CS-86-1117 (CSL-85-287, AD-A221759), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1985. iv +

20 pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1117.html.

Clancey:1985:RCK

[1117] William J. Clancey and Conrad Bock. Representing control knowledge as abstract task and metarules. Technical Report STAN-CS-87-1168 (KSL-85-16, AD-A187021), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1985. iii + 56

pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1168.html;
https://apps.dtic.mil/sti/tr/pdf/ADA187021.pdf.

Berg:1986:BCS

[1118] Kathryn A. Berg and Taleen Marashian. Bibliography of Computer Science reports, 1963–1986. Technical Report STAN-CS-86-1085, Stanford University, Department of Computer Science, Stanford, CA, USA, June

1986. iv + 71 pp. URL http://www-db.stanford.edu/TR/CS-TR-86-
1085.html.

Knuth:1986:TTM

[1119] Donald E. Knuth. A torture test for . Technical Report STAN-CS-86-1095, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1, 1986. 79

pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1095.html;
 http://www.ncstrl.org:8900/ncstrl/servlet/search?formname=
detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-86-1095.

Wilkins:1986:MTA

[1120] Marianne Winslett Wilkins. A model-theoretic approach to updating logical databases. Technical Report STAN-CS-86-1096, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1986. 23

pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1096.html.

Knuth:1986:T

[1121] Donald E. Knuth. TEXware. Technical Report STAN-CS-86-1097, Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1, 1986. 146 pp. URL http://www-db.stanford.edu/
TR/CS-TR-86-1097.html;   http://www.ncstrl.org:8900/ncstrl/
servlet/search?formname=detail&id=oai%3Ancstrlh%3Astan%3ASTAN%
2F%2FCS-TR-86-1097.

Cheriton:1986:DNF

[1122] David Cheriton and Timothy Mann. A decentralized naming facility. Technical Report STAN-CS-86-1098, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1986. 26 pp.

Abadi:1986:MTP

[1123] Martin Abadi and Zohar Manna. Modal theorem proving. Technical Report STAN-CS-86-1100, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1986. ii + 20 pp. URL

http://i.stanford.edu/pub/cstr/reports/cs/tr/86/1100/CS-TR-
86-1100.pdf;   http://www-db.stanford.edu/TR/CS-TR-86-1100.
html.

Foulser:1986:RSS

[1124] David E. Foulser. On random strings and sequence comparisons. Technical Report STAN-CS-86-1101, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1986. 138 pp.

Naughton:1986:DIR

[1125] Jeffrey F. Naughton. Data independent recursion in deductive databases. Technical Report STAN-CS-86-1102, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1986. 34 pp. URL

http://www-db.stanford.edu/TR/CS-TR-86-1102.html.

Mogul:1986:RIA

[1126] Jeffrey C. Mogul. Representing information about files. Technical Report STAN-CS-86-1103, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1986. 204 pp.

Subramanian:1986:CSA

[1127] Devika Subramanian. CS229b: a survey of AI classnotes for winter 84–85. Technical Report STAN-CS-86-1104, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1986. 201 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-86-1104.html.

Cheriton:1986:SCC

[1128] David R. Cheriton, Gert A. Slavenburg, and Patrick D. Boyle. Softwarecontrolled caches in the VMP multiprocessor. Technical Report STAN-CS-86-1105, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 12 pp. URL http://www-db.

stanford.edu/TR/CS-TR-86-1105.html.

Abadi:1986:TR

[1129] Martin Abadi and Zohar Manna. A timely resolution. Technical Report STAN-CS-86-1106, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1986. 21 pp. URL http://www-db.

stanford.edu/TR/CS-TR-86-1106.html.

Smith:1986:CI

[1130] David E. Smith. Controlling inference. Technical Report STAN-CS-86-1107, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1986. 199 pp.

Morris:1986:DON

[1131] Katherine Morris, Jeffrey D. Ullman, and Allen Van Gelder. Design overview of the NAIL! system. Technical Report STAN-CS-86-1108, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1986. 13 pp.

Casley:1986:PEP

[1132] Ross Casley. A proof editor for propositional temporal logic. Technical Report STAN-CS-86-1109, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1986. 26 pp. URL http://www-

db.stanford.edu/TR/CS-TR-86-1109.html.

Rosenbloom:1986:MEB

[1133] Paul S. Rosenbloom and John E. Laird. Mapping explanation-based generalization onto Soar. Technical Report STAN-CS-86-1111, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 18 pp.

Demetrescu:1986:SLA

[1134] Stefan G. Demetrescu. Scan line access memories for high speed image rasterization. Technical Report STAN-CS-86-1112, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 137 pp.

Pratt:1986:MCP

[1135] Vaughan Pratt. Modelling concurrency with partial orders. Technical Report STAN-CS-86-1113, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 33 pp.

Naughton:1986:OFF

[1136] Jeffrey F. Naughton. Optimizing function-free recursive inference rules. Technical Report STAN-CS-86-1114, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1986. 27 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-86-1114.html.

Buchanan:1986:HRM

[1137] Bruce G. Buchanan, Barbara Hayes-Roth, Olivier Lichtarge, Russ Altman, James Brinkley, Michael Hewett, Craig Cornelius, Bruce Duncan, and Oleg Jardetzky. The heuristic refinement method for deriving solution structures of proteins. Technical Report STAN-CS-86-1115, Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1986. 25 pp. URL http://www-db.stanford.edu/TR/CS-TR-86-
1115.html.

Helmbold:1986:APS

[1138] David Helmbold and Ernst Mayr. Applications of parallel scheduling to perfect graphs. Technical Report STAN-CS-86-1118, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 20 pp.

URL http://www-db.stanford.edu/TR/CS-TR-86-1118.html.

Rosenblum:1986:SUS

[1139] David S. Rosenblum and Ernst W. Mayr. Simulation of an Ultracomputer with several ’hot spots’. Technical Report STAN-CS-86-1119, Stanford University, Department of Computer Science, Stanford, CA, USA, June

1986. 37 pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1119.
html.

Moses:1986:KDE

[1140] Yoram Moses. Knowledge in distributed environment. Technical Report STAN-CS-86-1120, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 113 pp.

Feigenbaum:1986:PGS

[1141] Joan Feigenbaum. Product graphs: Some algorithmic and combinatorial results. Technical Report STAN-CS-86-1121, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 69 pp.

Rennels:1986:CMR

[1142] Glenn D. Rennels. A computational model of reasoning from the clinical literature. Technical Report STAN-CS-86-1122, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 263 pp.

Nii:1986:BS

[1143] H. Penny Nii. Blackboard systems. Technical Report STAN-CS-86-1123 (KSL-86-18), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 95 pp. URL http://www-db.

stanford.edu/TR/CS-TR-86-1123.html.

Scales:1986:EMA

[1144] Daniel J. Scales. Efficient matching algorithms for the SOARIOPSS production system. Technical Report STAN-CS-86-1124 (KSL-86-47), Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1986. 59 pp. URL http://www-db.stanford.edu/TR/CS-
TR-86-1124.html.

Schoen:1986:CS

[1145] Eric Schoen. The CAOS system. Technical Report STAN-CS-86-1125 (KSL-86-22), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 75 pp. URL http://www-db.

stanford.edu/TR/CS-TR-86-1125.html.

Davies:1986:CVD

[1146] Byron Davies. CAREL: a visible distributed Lisp. Technical Report STAN-CS-86-1126 (KSL-86-14), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 23 pp. URL

http://www-db.stanford.edu/TR/CS-TR-86-1126.html.

Malachi:1986:NLP

[1147] Yonathan Malachi. Nonclausal logic programming. Technical Report STAN-CS-86-1127, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 167 pp.

Theimer:1986:PRE

[1148] Marvin M. Theimer. Preemptable remote execution facilities for looselycoupled distributed systems. Technical Report STAN-CS-86-1128 (CSL-86-302), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 139 pp.

Cohn:1986:BOE

[1149] Evan R. Cohn and Ramsey W. Haddad. Beta operations: Efficient implementation of a primitive parallel operation. Technical Report STAN-CS-86-1129, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1986. 18 pp. URL http://www-db.stanford.

edu/TR/CS-TR-86-1129.html.

Nalwa:1986:DE

[1150] Vishvjit S. Nalwa and Thomas O. Binford. On detecting edges. Technical Report STAN-CS-86-1130, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 50 pp. URL http://

www-db.stanford.edu/TR/CS-TR-86-1130.html.

Bar-Noy:1986:PRA

[1151] Amotz Bar-Noy and David Peleg. Processor renaming in asynchronous environments. Technical Report STAN-CS-86-1131, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1986.

15 pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1131.html.

Sagiv:1986:ODP

[1152] Yehoshua Sagiv. Optimizing datalog programs. Technical Report STAN-CS-86-1132, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 29 pp. URL http://www-db.

stanford.edu/TR/CS-TR-86-1132.html.

Edighoffer:1986:DRC

[1153] Judy Lynn Edighoffer. Distributed, replicated computer bulletin board service. Technical Report STAN-CS-86-1133, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 152 pp.

Cheriton:1986:UUS

[1154] David R. Cheriton. UIO: a uniform I/O system interface for distributed systems. Technical Report STAN-CS-86-1134, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1986. 30

pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1134.html.

Treitel:1986:SLP

[1155] Richard Treitel. Sequentialization of logic programs. Technical Report STAN-CS-86-1135, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1986. 167 pp.

Brown:1986:EKB

[1156] Harold D. Brown, Eric Schoen, and Bruce A. Delagi. An experiment in knowledge-based signal understanding using parallel architectures. Technical Report STAN-CS-86-1136 (KSL-86-69), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1986. ix +

36 pp. URL http://www-db.stanford.edu/TR/CS-TR-86-1136.html.

Mogul:1986:LFA

[1157] Jeffrey Mogul. The Leaf file access protocol. Technical Report STAN-CS-86-1137, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1986. 25 pp. URL http://www-db.stanford.

edu/TR/CS-TR-86-1137.html.

Mackinlay:1986:ADG

[1158] Jock Mackinlay. Automatic design of graphical presentations. Technical Report STAN-CS-86-1138, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1986. 200 pp.

Healey:1986:LSS

[1159] Glenn Healey and Thomas O. Binford. Local shape from specularity. Technical Report STAN-CS-86-1139, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1986. 29 pp. URL

http://www-db.stanford.edu/TR/CS-TR-86-1139.html.

Newell:1987:SAG

[1160] John E. Laird, Allen Newell, and Paul S. Rosenbloom. Soar: an architecture for general intelligence. Technical Report STAN-CS-86-1140 (KSL-86-70, AD-A179583), Stanford University, Department of Computer Science, Stanford, CA, USA, December 3, 1986. vii + 63 pp. URL

https://apps.dtic.mil/sti/tr/pdf/ADA179583.pdf.

Manna:1986:HCB

[1161] Zohar Manna and Richard Waldinger. How to clear a block: a theory of plans. Technical Report STAN-CS-87-1141, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1986. 39 pp.

Hayes-Roth:1986:LER

[1162] Barbara Hayes-Roth, Alan Garvey, M. Vaughan Johnson, Jr., and Michael Hewett. A layered environment for reasoning about action. Technical Report STAN-CS-87-1147 (KSL-86-38), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1986.

83 pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1147.html.

Russell:1986:AIR

[1163] Stuart J. Russell. Analogical and inductive reasoning. Technical Report STAN-CS-87-1150, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1986. 230 pp.

Bellin:1986:EAT

[1164] Gianluigi Bellin and Jussi Ketonen. Experiments in automatic theorem proving. Technical Report STAN-CS-87-1155, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1986. 265

pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1155.html.

Thompson:1986:AQM

[1165] Timothy F. Thompson and William J. Clancey. Applying a qualitative modeling shell to process diagnosis: The Caster system. Technical Report STAN-CS-87-1169 (KSL-85-32, AD-A186994), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1986. 35 + 9 pp.

Clancey:1986:VKB

[1166] William J. Clancey. Viewing knowledge bases as qualitative models. Technical Report STAN-CS-87-1170 (KSL-86-27, AD-A187091), Stanford University, Department of Computer Science, Stanford, CA, USA,

May 1986. v + 20 pp. URL http://www-db.stanford.edu/TR/CS-TR-
87-1170.html;   https://apps.dtic.mil/sti/tr/pdf/ADA187091.
pdf.

Clancey:1986:QSM

[1167] William J. Clancey. Qualitative student models. Technical Report STAN-CS-87-1171 (KSL-86-15), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1986. 50 pp.

Clancey:1986:GNH

[1168] William J. Clancey. From GUIDON to NEOMYCIN and HERACLES in twenty short lessons. Technical Report STAN-CS-87-1172 (KSL-86-11), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1986. 40–60, 187–187 pp. URL https://apps.dtic.mil/

sti/tr/pdf/ADA186953.pdf.

Clancey:1986:RWF

[1169] William J. Clancey. Review of Winograd and Flores’ Understanding Computers and Cognition. Technical Report STAN-CS-87-1173 (KSL-86-48), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1986. vi + 13 pp. URL http://www-

db.stanford.edu/TR/CS-TR-87-1173.html;   https://apps.dtic.
mil/sti/tr/pdf/ADA187092.pdf.

Clancey:1986:ITS

[1170] William J. Clancey. Intelligent tutoring systems: a tutorial survey. Technical Report STAN-CS-87-1174 (KSL-86-58), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1986.

ii + 43 pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1174.
html;   https://apps.dtic.mil/sti/tr/pdf/ADA187066.pdf.

Brinkley:1987:HRS

[1171] James Brinkley, Bruce Buchanan, Russ Altman, Bruce Duncan, and Craig Cornelius. A heuristic refinement for spacial constraint satisfaction problems. Technical Report STAN-CS-87-1142 (KSL-87-05), Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1987. vi + 15 pp. URL http://www-db.stanford.edu/TR/CS-
TR-87-1142.html.

Winslett:1987:UDI

[1172] Marianne S. Winslett. Updating databases with incomplete information. Technical Report STAN-CS-87-1143, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1987. 164 pp.

Byrd:1987:CMT

[1173] Gregory T. Byrd and Bruce A. Delagi. Considerations for multiprocessor topologies. Technical Report STAN-CS-87-1144 (KSL-87-07 PB96-148416), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1987. 6 pp. URL http://www-db.stanford.

edu/TR/CS-TR-87-1144.html.

Cheriton:1987:NMV

[1174] David R. Cheriton and Cary L. Williamson. Network measurement of the VMTP request-response protocol in the V Distributed System. Technical Report STAN-CS-87-1145, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1987. 23 pp.

Byrd:1987:PPM

[1175] Gregory Byrd, Russell Nakano, and Bruce Delagi. A point-to-point multicast communications protocol. Technical Report STAN-CS-87-1146 (KSL-87-02), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1987. viii + 30 pp. URL http://www-

db.stanford.edu/TR/CS-TR-87-1146.html.

Delagi:1987:IASa

[1176] Bruce Delagi, Nakul Saraiya, Sayuri Nishimura, and Greg Byrd. An instrumented architectural simulation system. Technical Report STAN-CS-87-1148 (KSL-86-36), Stanford University, Department of Computer Science, Stanford, CA, USA, January 29, 1987. vii + 21 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-87-1148.html.

Millen:1987:PNA

[1177] Katie Mac Millen, Ann Diaz-Barriga, and Carolyn Tajnai. Proceedings from the Nineteenth Annual Meeting of the Stanford Computer Forum. Technical Report STAN-CS-87-1149, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1987. 24 pp. URL

http://www-db.stanford.edu/TR/CS-TR-87-1149.html.

Abadi:1987:TLT

[1178] Martin Abadi. Temporal-logic theorem proving. Technical Report STAN-CS-87-1151, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1987. 179 pp.

Davidson:1987:INL

[1179] James E. Davidson. Interpreting natural language database updates. Technical Report STAN-CS-87-1152, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1987. 119 pp.

Markenscoff:1987:OGP

[1180] Xanthippi Markenscoff and Christos Papadimitriou. Optimum grip of a polygon. Technical Report STAN-CS-87-1153, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1987. 21

pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1153.html.

Rokicki:1987:PPS

[1181] Tomas G. Rokicki and Donald E. Knuth. A programming and problemsolving seminar. Technical Report STAN-CS-87-1154, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1987.

vi + 89 pp. URL http://i.stanford.edu/pub/cstr/reports/cs/
tr/87/1154/CS-TR-87-1154.pdf;   http://www-db.stanford.edu/
TR/CS-TR-87-1154.html;   http://www.ncstrl.org:8900/ncstrl/
servlet/search?formname=detail&id=oai%3Ancstrlh%3Astan%3ASTAN%
2F%2FCS-TR-87-1154.

Mayr:1987:DTE

[1182] Ernst W. Mayr. The dynamic tree expression problem. Technical Report STAN-CS-87-1156, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1987. 19 pp. URL http://www-db.

stanford.edu/TR/CS-TR-87-1156.html.

Mayr:1987:NID

[1183] Ernst W. Mayr and C. Greg Plaxton. Network implementation of the DTEP algorithm. Technical Report STAN-CS-87-1157, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1987.

22 pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1157.html.

Winograd:1987:LAP

[1184] Terry A. Winograd. A language/action perspective on the design of cooperative work. Technical Report STAN-CS-87-1158, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1987. 34 pp.

Winograd:1987:MTL

[1185] Terry A. Winograd. Muir: a tool for language design. Technical Report STAN-CS-87-1159, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1987. 30 pp. URL http://www-db.

stanford.edu/TR/CS-TR-87-1159.html.

Winograd:1987:SCR

[1186] Terry A. Winograd. Strategic computing research and the universities. Technical Report STAN-CS-87-1160, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1987. 17 pp. URL

http://www-db.stanford.edu/TR/CS-TR-87-1160.html.

Winogad:1987:TMC

[1187] Terry A. Winogad. Thinking machines: Can there be? Are we? Technical Report STAN-CS-87-1161, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1987. 34 pp.

Spencer:1987:WMA

[1188] Thomas Spencer. Weighted matching algorithms. Technical Report STAN-CS-87-1162, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1987. 130 pp.

Hershberger:1987:EAS

[1189] John E. Hershberger. Efficient algorithms for shortest path and visibility problems. Technical Report STAN-CS-87-1163, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1987. 193 pp.

Karlin:1987:SMD

[1190] Anna R. Karlin. Sharing memory in distributed systems — methods and applications. Technical Report STAN-CS-87-1164, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1987. 76 pp.

Fraley:1987:SNL

[1191] Christina Fraley. Solution of nonlinear least-squares problems. Technical Report STAN-CS-87-1165, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1987. 325 pp.

Okuno:1987:PEO

[1192] Hiroshi G. Okuno and Anoop Gupta. Parallel execution of OPS5 in QLISP. Technical Report STAN-CS-87-1166 (KSL-87-43, AD-A323588), Stanford University, Department of Computer Science, Stanford, CA,

USA, June 1987. 19 pp. URL http://www-db.stanford.edu/TR/CS-
TR-87-1166.html.

Tang:1987:SST

[1193] Wei-Pai Tang. Schwarz splitting and template operators. Technical Report STAN-CS-87-1167 (Classic-87-03, AD-A187956), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1987. x

+ 129 pp. URL https://apps.dtic.mil/sti/tr/pdf/ADA187956.pdf.

Wilkin:1987:UED

[1194] David C. Wilkins, William J. Clancey, and Bruce G. Buchanan. Using and evaluating differential modeling in intelligent tutoring and apprentice learning systems. Technical Report STAN-CS-87-1175 (KSL-86-62, AD-A187503), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1987. v + 26 pp. URL http://www-

db.stanford.edu/TR/CS-TR-87-1175.html;   https://apps.dtic.

mil/sti/tr/pdf/ADA187503.pdf. To appear in: Intelligent Tutoring Systems: Lessons Learned, J. Psotka, D. Massey and S. Mutter, editors, Lawrence Erlbaum Publishers, 1987.

Flajolet:1987:BPC

[1195] Philippe Flajolet, Danièle Gardy, and Loÿs Thimonier. Birthday paradox, coupon collectors, caching algorithms, and self-organizing search. Technical Report STAN-CS-87-1176 (RR-0720), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1987. 18

pp. URL https://inria.hal.science/inria-00075832v1/file/RR-
0720.pdf.

Finlayson:1987:LFE

[1196] Ross S. Finlayson and David R. Cheriton. Log files: an extended file service exploiting write-once storage. Technical Report STAN-CS-87-1177, Stanford University, Department of Computer Science, Stanford,

CA, USA, August 1987. 14 pp. URL http://www-db.stanford.edu/
TR/CS-TR-87-1177.html.

Byrd:1987:DCT

[1197] Gregory T. Byrd, Russell Nakano, and Bruce A. Delagi. A dynamic, cut-through communications protocol with multicast. Technical Report STAN-CS-87-1178 (KSL-87-44, AD-A197006), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1987. ii +

23 pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1178.html.

Mann:1987:DND

[1198] Timothy Paul Mann. Decentralized naming in distributed computer systems. Technical Report STAN-CS-87-1179, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1987. 101 pp.

Wilkins:1987:DRS

[1199] David C. Wilkins and Bruce G. Buchanan. On debugging rule sets when reasoning under uncertainty. Technical Report STAN-CS-87-1181 (KSL-86-30), Stanford University, Department of Computer Science, Stanford,

CA, USA, May 1987. 21 pp. URL http://www-db.stanford.edu/TR/
CS-TR-87-1181.html.

Wilkins:1987:KBR

[1200] David C. Wilkins and Bruce G. Buchanan. Knowledge base refinement by monitoring abstract control knowledge. Technical Report STAN-CS-87-1182 (KSL-87-01), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1987. 21 pp. URL http://www-db.

stanford.edu/TR/CS-TR-87-1182.html.

Clancey:1987:KES

[1201] William J. Clancey. The knowledge engineer as student: Metacognitive bases for asking good questions. Technical Report STAN-CS-87-1183 (KSL-87-12, AD-A186995), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1987. iv + 34

pp. URL http://www-db.stanford.edu/TR/CS-TR-87-1183.html;
https://apps.dtic.mil/sti/tr/pdf/ADA186995.pdf.

Okuno:1987:FAF

[1202] Hiroshi G. Okuno, Nobuyasu Osato, and Ikuo Takeuchi. Firmware approach to fast Lisp interpreter. Technical Report STAN-CS-87-1184 (KSL-87-57), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1987. v + 22 pp. URL http://www-

db.stanford.edu/TR/CS-TR-87-1184.html.

Haas:1987:BPI

[1203] Zygmunt Haas and David R. Cheriton. Blazenet: a photonic implementable wide-area network. Technical Report STAN-CS-87-1185 (CSL-TR-87-346, AD-A197005), Stanford University, Department of Computer Science, Stanford, CA, USA, October 5, 1987. ii + 21 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-87-1185.html.

Manna:1987:HTP

[1204] Zohar Manna and Amir Pnueli. A hierarchy of temporal properties. Technical Report STAN-CS-87-1186, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1987. 44 pp. URL

http://www-db.stanford.edu/TR/CS-TR-87-1186.html.

Morris:1987:YYA

[1205] K. Morris, J. F. Naughton, Y. Saraiya, J. D. Ullman, and A. Van Gelder. YAWN! (YET ANOTHER WINDOW ON NAIL!). Technical Report STAN-CS-87-1187, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1987. 16 pp.

Nakano:1987:EKB

[1206] Russell Nakano and Masafumi Minami. Experiments with a knowledgebased system on a multiprocessor. Technical Report STAN-CS-87-1188 (KSL-87-61, AD-A198708), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1987. vii + 45 pp. URL

http://www-db.stanford.edu/TR/CS-TR-87-1188.html.

Delagi:1987:IASb

[1207] Bruce A. Delagi, Nakul Saraiya, Sayuri Nishimura, and Greg Byrd. Instrumented architectural simulation. Technical Report STAN-CS-87-1189 (KSL-87-65), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1987. iv + 6 pp. URL http://

www-db.stanford.edu/TR/CS-TR-87-1189.html.

Finger:1987:ECD

[1208] Joseph Jeffrey Finger. Exploiting constraints in design synthesis. Technical Report STAN-CS-88-1204, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1987. 128 pp.

Nazarian:1988:BDC

[1209] Taleen Marashian Nazarian. Bibliography; Department of Computer Science technical reports, 1963–1988. Technical Report CS-TR-87-1180, Stanford University, Department of Computer Science, Stanford, CA,

USA, January 1988. 92 pp. URL http://www-db.stanford.edu/TR/
CS-TR-87-1180.html.

Kim:1988:VTC

[1210] Scott Edward Kim. Viewpoint: Toward a computer for visual thinkers. Technical Report STAN-CS-88-1190, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1988. 137 pp.

Cai:1988:IRM

[1211] Chunsheng Cai. Instantaneous robot motion with contact between surfaces. Technical Report STAN-CS-88-1191 (AD-A198709), Stanford University, Department of Computer Science, Stanford, CA, USA, Jan-

uary 1988. x + 173 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA198709.pdf.

Cheriton:1988:UMM

[1212] David R. Cheriton. The unified management of memory in the V distributed system. Technical Report STAN-CS-88-1192 (CSL-TR-88-359), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1988. 24 pp.

Knuth:1988:MW

[1213] Donald E. Knuth, Tracy Larrabee, and Paul M. Roberts. Mathematical writing. Technical Report STAN-CS-88-1193, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1988. 115

pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1193.html.

Musen:1988:GMB

[1214] Mark A. Musen. Generation of model-based knowledge-acquisition tools for clinical-trial advice systems. Technical Report STAN-CS-88-1194, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1988. 294 pp.

Bar-Noy:1988:LBR

[1215] A. Bar-Noy, N. Linial, and D. Peleg. A lower bound for radio broadcast. Technical Report STAN-CS-88-1195, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1988. 15 pp. URL

http://www-db.stanford.edu/TR/CS-TR-88-1195.html.

Latombe:1988:MPU

[1216] Jean-Claude Latombe. Motion planning with uncertainty: The preimage backchaining approach. Technical Report STAN-CS-88-1196, Stanford University, Department of Computer Science, Stanford, CA, USA, March

1988. i + 62 pp. URL http://www-db.stanford.edu/TR/CS-TR-88-
1196.html.

Cheriton:1988:VMI

[1217] D. R. Cheriton, Anoop Gupta, P. D. Boyle, and H. A. Goosen. The VMP multiprocessor: Initial experience, refinements and performance evaluation. Technical Report STAN-CS-88-1197 (CSL-TR-88-354), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1988. 24 pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1197.
html.

Lamping:1988:USP

[1218] John Lamping. A unified system of parameterization for programming languages. Technical Report STAN-CS-88-1198, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1988. 141 pp.

Buning:1988:PVA

[1219] H. K. Buning, T. Lettman, and E. W. Mayr. Projections of vector addition system reachability sets are semilinear. Technical Report STAN-CS-88-1199, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1988. 9 pp. URL http://www-db.stanford.

edu/TR/CS-TR-88-1199.html.

Anderson:1988:PAA

[1220] R. J. Anderson, E. W. Mayr, and M. K. Warmuth. Parallel approximation algorithms for bin packing. Technical Report STAN-CS-88-1200, Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1988. 16 pp. URL http://www-db.stanford.edu/TR/CS-
TR-88-1200.html.

Flajolet:1988:SAG

[1221] Philippe Flajolet and Andrew Odlyzko. Singularity analysis of generating functions. Technical Report STAN-CS-88-1201, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1988. 25 pp.

Baudinet:1988:PTP

[1222] Marianne Baudinet. Proving termination properties of PROLOG programs: a semantic approach. Technical Report STAN-CS-88-1202, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1988. 14 pp.

Baudinet:1988:STL

[1223] Marianne Baudinet. On the semantics of temporal logic programming. Technical Report STAN-CS-88-1203, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 30 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-88-1203.html.

Armstrong:1988:DRC

[1224] Brian Stewart Randall Armstrong. Dynamics for robot control: Friction modeling and ensuring excitation during parameter identification. Technical Report STAN-CS-88-1205, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1988. 198 pp.

Weening:1988:PLS

[1225] Joseph S. Weening. A parallel Lisp simulator. Technical Report STAN-CS-88-1206, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1988. 27 pp. URL http://www-db.

stanford.edu/TR/CS-TR-88-1206.html.

Burdick:1988:KAD

[1226] Joel W. Burdick. Kinematic analysis and design of redundant robot manipulators. Technical Report STAN-CS-88-1207, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1988. 267 pp.

Feder:1988:T

[1227] Tomas Feder. Toetjes. Technical Report STAN-CS-88-1208, Stanford University, Department of Computer Science, Stanford, CA, USA, June

1988. 13 pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1208.
html.

Goldberg:1988:CAG

[1228] A. V. Goldberg, S. A. Plotkin, and E. Tardos. Combinatorial algorithms for the generalized circulation problem. Technical Report STAN-CS-88-1209, Stanford University, Department of Computer Science, Stanford,

CA, USA, June 1988. 36 pp. URL http://www-db.stanford.edu/TR/
CS-TR-88-1209.html.

Bronstein:1988:SFS

[1229] Alexandre Bronstein and Carolyn L. Talcott. String-functional semantics for formal verification of synchronous circuits. Technical Report STAN-CS-88-1210, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 62 pp. URL http://www-db.

stanford.edu/TR/CS-TR-88-1210.html.

Goldberg:1988:STP

[1230] A. V. Goldberg, S. A. Plotkin, and P. M. Vaidya. Sublinear-time parallel algorithms. Technical Report STAN-CS-88-1211, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 27

pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1211.html.

Worley:1988:IRI

[1231] Patrick Haven Worley. Information requirements and the implications for parallel computation. Technical Report STAN-CS-88-1212, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 148 pp.

Cheriton:1988:ERS

[1232] David R. Cheriton. Exploiting recursion to simplify RPC communication architectures. Technical Report STAN-CS-88-1213 (CSL-TR-88-360, AD-A198711), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 14 pp.

Deering:1988:MRI

[1233] Stephen E. Deering. Multicast routing in internetworks and extended LANs. Technical Report STAN-CS-88-1214 (CSL-TR-88-361, AD-A200911), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1988. 14 pp. URL http://www-db.stanford.edu/

TR/CS-TR-88-1214.html.

Stolfi:1988:PCG

[1234] Jorge Stolfi. Primitives for computational geometry. Technical Report STAN-CS-88-1215, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 236 pp.

Schaffer:1988:API

[1235] Alejandro A. Schaffer. Algorithmic problems on intersection graphs. Technical Report STAN-CS-88-1216, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1988. 139 pp.

Brooks:1988:TVD

[1236] Kenneth P. Brooks. A two-view document editor with user-definable document structure. Technical Report STAN-CS-88-1217, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 203 pp.

Bar-Noy:1988:SMA

[1237] Amotz Bar-Noy and David Peleg. Square meshes are not always optimal. Technical Report STAN-CS-88-1218, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1988. 28 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-88-1218.html.

Flajolet:1988:FCE

[1238] Philippe Flajolet, Donald E. Knuth, and Boris Pittel. The first cycles in an evolving graph. Technical Report STAN-CS-88-1219, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1988. 49 pp.

Vistnes:1988:CTA

[1239] Richard L. Vistnes. Computer texture analysis and segmentation. Technical Report STAN-CS-88-1220, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1988. 186 pp.

Beigel:1988:QLR

[1240] Richard Beigel. Query-limited reducibilities. Technical Report STAN-CS-88-1221, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1988. 144 pp.

Hailpern:1988:LBM

[1241] Max Hailpern. Load balancing for massively-parallel soft-real-time systems. Technical Report STAN-CS-88-1222 (KSL-88-62, AD-A200912), Stanford University, Department of Computer Science, Stanford, CA, USA, August 30, 1988. ii + 18 pp.

Knuth:1988:ET

[1242] Donald E. Knuth. The errors of TEX. Technical Report STAN-CS-88-1223, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1988. 75 pp.

Singh:1988:DBC

[1243] Vineet Singh. Distributing backward-chaining deductions to multiple processors. Technical Report STAN-CS-88-1224, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1988. 222 pp.

Mayr:1988:PAA

[1244] Ernst W. Mayr. Parallel approximation algorithms. Technical Report STAN-CS-88-1225, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1988. 17 pp. URL http://www-

db.stanford.edu/TR/CS-TR-88-1225.html.

Hayes-Roth:1988:MIS

[1245] Barbara Hayes-Roth. Making intelligent systems adaptive. Technical Report STAN-CS-88-1226 (KSL-88-64, AD-A200913), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1988. iv +

24 pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1226.html;
 https://apps.dtic.mil/sti/tr/pdf/ADA200913.pdf. To appear as

a chapter in the book, Architectures for Intelligence, edited by K. Van Lehn and published by Lawrence Erlbaum, 1988.

Ahuja:1988:FMC

[1246] Ravindra K. Ahuja, Andrew V. Goldberg, James B. Orlin, and Robert E. Tarjan. Finding minimum-cost flows by double scaling. Technical Report STAN-CS-88-1227, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1988. iv + 27 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-88-1227.html.

Goldberg:1988:PAF

[1247] A. V. Goldberg and R. E. Tarjan. A parallel algorithm for finding a blocking flow in an acyclic network. Technical Report STAN-CS-88-1228, Stanford University, Department of Computer Science, Stanford,

CA, USA, November 1988. 9 pp. URL http://www-db.stanford.edu/
TR/CS-TR-88-1228.html.

Hayes-Roth:1988:DIWa

[1248] Barbara Hayes-Roth, Michael Hewett, Richard Washington, Rattikorn Hewett, and Adam Seiver. Distributing intelligence within an individual. Technical Report STAN-CS-88-1229 (KSL-88-50), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1988.

23 pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1229.html.

Manna:1988:SVC

[1249] Zohar Manna and Amir Pnueli. Specification and verification of concurrent programs by for-all automata. Technical Report STAN-CS-88-1230, Stanford University, Department of Computer Science, Stanford, CA,

USA, November 1988. 41 pp. URL http://www-db.stanford.edu/TR/
CS-TR-88-1230.html.

Cohn:1988:BOP

[1250] Evan R. Cohn. The beta operation: a parallel primitive. Technical Report STAN-CS-88-1231, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1988. 69 pp.

Baudinet:1988:TLP

[1251] Marianne Baudinet. Temporal logic programming is complete and expressive. Technical Report STAN-CS-88-1232, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1988. 15 pp.

Ross:1988:PSW

[1252] Kenneth A. Ross. A procedural semantics for well founded negation in logic programs. Technical Report STAN-CS-88-1233, Stanford University, Department of Computer Science, Stanford, CA, USA,

December 1988. ii + 29 pp. URL http://i.stanford.edu/pub/
cstr/reports/cs/tr/88/1233/CS-TR-88-1233.pdf;   http://www-
db.stanford.edu/TR/CS-TR-88-1233.html.

Pittel:1988:ANS

[1253] Boris Pittel. The average number of stable matchings. Technical Report STAN-CS-88-1234, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988. 21 pp. URL http://www-

db.stanford.edu/TR/CS-TR-88-1234.html.

Knuth:1988:CME

[1254] Donald E. Knuth. Concrete mathematics errata. Technical Report STAN-CS-88-1235, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988. 30 pp.

Shoham:1988:TAR

[1255] Yoav Shoham. Time for action: On the relation between time, knowledge, and action. Technical Report STAN-CS-88-1236, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988.

16 pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1236.html.

Shoham:1988:BDK

[1256] Yoav Shoham and Yoram Moses. Belief as defeasible knowledge. Technical Report STAN-CS-88-1237, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988. 11 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-88-1237.html.

Bothner:1988:ECL

[1257] Per Magnus Bothner. Efficiently combining logical constraints with functions. Technical Report STAN-CS-88-1238, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988. 191 pp.

Bar-Noy:1988:SMF

[1258] Amotz Bar-Noy and Joseph Naor. Sorting, minimal feedback sets and Hamilton paths in tournaments. Technical Report STAN-CS-88-1239, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988. iv + 20 pp. URL http://www-db.stanford.

edu/TR/CS-TR-88-1239.html.

Gafni:1988:SEC

[1259] E. Gafni, J. Naor, and P. Ragde. On separating the EREW and CREW PRAM models. Technical Report STAN-CS-88-1240, Stanford University, Department of Computer Science, Stanford, CA, USA, December

1988. 6 pp. URL http://www-db.stanford.edu/TR/CS-TR-88-1240.
html.

Knuth:1988:SH

[1260] Donald Knuth, Rajeev Motwani, and Boris Pittel. Stable husbands. Technical Report STAN-CS-88-1241, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988. 16 pp.

Wilkins:1988:ALT

[1261] David C. Wilkins. Apprenticeship learning techniques for knowledge based systems. Technical Report STAN-CS-88-1242, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1988. 153 pp.

Hayes-Roth:1988:DIWb

[1262] Barbara Hayes-Roth, Michael Hewett, Richard Washington, Rattikorn Hewett, and Adam Seiver. Distributing intelligence within an individual. Technical Report STAN-CS-88-1992 (KSL-88-50), Stanford University, Department of Computer Science, Stanford, CA, USA, October

1988. iv + 21 pp. URL https://bitsavers.org/pdf/stanford/cs_
techReports/CS-TR-88-1229_Distributing_Intelligence_Within_
An_Individual.pdf.

DeMoor:1989:RSV

[1263] Bart L. R. De Moor and Gene H. Golub. The restricted singular value decomposition: properties and applications. Technical Report STAN-CS-2001, Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1989. 69 pp.

DeMoor:1989:GSV

[1264] Bart L. R. De Moor and Gene H. Golub. Generalized singular value decompositions: a proposal for a standardized nomenclature. Technical Report STAN-CS-2002, Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1989. 12 pp.

Elman:1989:IMC

[1265] Howard C. Elman and Gene H. Golub. Iterative methods for cyclically reduced non-self-adjoint linear systems II. Technical Report STAN-CS-2238 (UMIACS-TR-89-45), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1989. ii + 26

pp. URL https://dl.acm.org/doi/10.5555/891608;   https://
www.jstor.org/stable/2008506.

Elhay:1989:UDO

[1266] Sylvan Elhay, Gene H. Golub, and Jaroslav Kautský. Updating and downdating of orthogonal polynomials with data fitting applications. Technical Report STAN-CS-89-04 (NA-89-04), Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1989. 32 pp.

URL https://dl.acm.org/doi/10.1137/0612024.

Golub:1989:MMI

[1267] Gene H. Golub and Martin H. Gutknecht. Modified moments for indefinite weight functions. Technical Report STAN-CS-89-08 (NA-89-08), Stanford University, Department of Computer Science, Stanford, CA,

USA, ???? 1989. 19 pp. URL https://people.math.ethz.ch/~mhg/
talks/GolGreview/GolGrev.pdf   (slides).

Lin:1989:ASU

[1268] Fangzhen Lin and Yoav Shoham. Argument systems: a uniform basis for nonmonotonic reasoning. Technical Report STAN-CS-89-1243, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1989. 28 pp.

Fraley:1989:SPN

[1269] Christina Fraley. Software performance on nonlinear least-squares problems. Technical Report STAN-CS-89-1244, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1989. 128 pp.

URL http://www-db.stanford.edu/TR/CS-TR-89-1244.html.

Wiederhold:1989:MOR

[1270] Gio Wiederhold, Thierry Barsalou, and Surajit Chaoudhuri. Managing objects in a relational framework. Technical Report STAN-CS-89-1245, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1989. 103 pp.

Byrd:1989:MCM

[1271] Gregory Byrd, Nakul Saraiya, and Bruce Delagi. Multicast communication in multiprocessor systems. Technical Report STAN-CS-89-1246, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1989. 19 pp.

Rabinov:1989:PQC

[1272] Arkady Rabinov and Igor Rivin. Programming in Qlisp — a case study. Technical Report STAN-CS-89-1247, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1989. 17 pp.

Goldberg:1989:ENS

[1273] Andrew V. Goldberg, Michael D. Grigoriadis, and Robert E. Tarjan. Efficiency of the network simplex algorithm for the maximum flow problem. Technical Report STAN-CS-89-1248, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1989. 18 pp. URL

http://www-db.stanford.edu/TR/CS-TR-89-1248.html.

Baudinet:1989:LPS

[1274] Marianne Baudinet. Logic programming semantics: Techniques and applications. Technical Report STAN-CS-89-1249, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1989. 172 pp.

Mason:1989:SCA

[1275] Ian Mason and Carolyn Talcott. A sound and complete axiomatization of operational equivalence between programs with memory. Technical Report STAN-CS-89-1250, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1989. 27 pp. URL http://www-db.

stanford.edu/TR/CS-TR-89-1250.html.

Huggins:1988:EHM

[1276] K. Cleo R. Huggins. Egyptian hieroglyphs for modern printing devices: an outline font of Egyptian hieroglyphs for PostScript(R) printers. Technical Report STAN-CS-89-1251, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1, 1988. ii + 60 pp. URL

https://apps.dtic.mil/sti/tr/pdf/ADA326695.pdf.

Goldberg:1989:NFA

[1277] Andrew Goldberg, Eva Tardos, and Robert Tarjan. Network flow algorithms. Technical Report STAN-CS-89-1252, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1989. 82 pp.

Vavasis:1989:CFP

[1278] Stephen A. Vavasis. Complexity of fixed point computations. Technical Report STAN-CS-89-1253, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1989. 102 pp.

Knuth:1989:M

[1279] Donald E. Knuth, Tomas G. Rokicki, and Arthur L. Samuel. META-FONTware. Technical Report STAN-CS-89-1255, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1,

1989. 209 pp. URL http://i.stanford.edu/pub/cstr/reports/cs/
tr/89/1255/CS-TR-89-1255.pdf;   http://www-db.stanford.edu/
TR/CS-TR-89-1255.html;   http://www.ncstrl.org:8900/ncstrl/
servlet/search?formname=detail&id=oai%3Ancstrlh%3Astan%3ASTAN%
2F%2FCS-TR-89-1255.

Billawala:1989:MPS

[1280] Neenie Billawala. Metamarks: Preliminary studies for a Pandora’s box of shapes. Technical Report STAN-CS-89-1256, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989. 131 pp.

Barraquand:1989:RMP

[1281] Jerome Barraquand and Jean-Claude Latombe. Robot motion planning: a distributed representation approach. Technical Report STAN-CS-89-1257, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989. 56 pp.

Chen:1989:HSB

[1282] Pang-Chieh Chen. Heuristic sampling on backtrack trees. Technical Report STAN-CS-89-1258, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989. 172 pp.

Goldberg:1989:IPM

[1283] Andrew V. Goldberg, Serge A. Plotkin, David B. Shmoys, and Eva Tardos. Interior-point methods in parallel computation. Technical Report STAN-CS-89-1259, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989. 15 pp. URL http://www-db.

stanford.edu/TR/CS-TR-89-1259.html.

Subramanian:1989:TJR

[1284] Devika Subramanian. A theory of justified reformulations. Technical Report STAN-CS-89-1260, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989. 134 pp.

Mayr:1989:PPP

[1285] Ernst W. Mayr and C. Greg Plaxton. Pipelined parallel computations, and sorting on a pipelined hypercube. Technical Report STAN-CS-89-1261, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989. iv + 15 pp. URL http://i.stanford.edu/

pub/cstr/reports/cs/tr/89/1261/CS-TR-89-1261.pdf;   http://
www-db.stanford.edu/TR/CS-TR-89-1261.html.

Swami:1989:OLJ

[1286] Arun N. Swami. Optimization of large join queries. Technical Report STAN-CS-89-1262, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1989. 145 pp.

Karp:1989:HFQ

[1287] Peter D. Karp. Hypothesis formation and qualitative reasoning in molecular biology. Technical Report STAN-CS-89-1263, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1989. 339 pp.

Fischer:1989:CPA

[1288] Bernd Fischer and Roland Freund. Chebyshev polynomials are not always optimal. Technical Report STAN-CS-89-1264, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1989. 14

pp. URL http://www-db.stanford.edu/TR/CS-TR-89-1264.html.

Weening:1989:PEL

[1289] Joseph S. Weening. Parallel execution of Lisp programs. Technical Report STAN-CS-89-1265, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1989. 94 pp.

Cheriton:1989:MLS

[1290] David R. Cheriton, Hendrik A. Goosen, and Patrick D. Boyle. Multi-level shared caching techniques for scalability in VMP-MC. Technical Report STAN-CS-89-1266, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989. ii + 18 pp. URL http://www-db.

stanford.edu/TR/CS-TR-89-1266.html.

Alur:1989:RTL

[1291] Rajeev Alur and Thomas A. Henzinger. A really temporal logic. Technical Report STAN-CS-89-1267, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1989. 27 pp. URL http://www-

db.stanford.edu/TR/CS-TR-89-1267.html.

Floyd:1989:AMa

[1292] Robert W. Floyd and Donald E. Knuth. Addition machines. Technical Report STAN-CS-89-1254, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1989. 16 pp.

Floyd:1989:AMb

[1293] Robert W. Floyd and Donald E. Knuth. Addition machines. Technical Report STAN-CS-89-1268, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1, 1989. 16

pp. URL http://www-db.stanford.edu/TR/CS-TR-89-1268.html;
 http://www.ncstrl.org:8900/ncstrl/servlet/search?formname=
detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-89-1268.

Ross:1989:PPS

[1294] Kenneth A. Ross and Donald E. Knuth. A programming and problem solving seminar. Technical Report STAN-CS-89-1269, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1, 1989.

88 pp. URL http://www-db.stanford.edu/TR/CS-TR-89-1269.html;
 http://www.ncstrl.org:8900/ncstrl/servlet/search?formname=
detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-89-1269.

Holstege:1989:MDN

[1295] Mary A. Holstege. Marking and the design of notations. Technical Report STAN-CS-89-1270, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1989. 255 pp.

Kent:1989:CKL

[1296] Mark David Kent. Chebyshev, Krylov, Lanczos: Matrix relationships and computations. Technical Report STAN-CS-89-1271, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1989.

x + 126 pp. URL https://www.proquest.com/pqdtglobal/docview/
303724638.

Finlayson:1989:LFS

[1297] Ross S. Finlayson. A log file service exploiting write-once storage. Technical Report STAN-CS-89-1272, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1989. 98 pp.

Cheriton:1989:SHP

[1298] David R. Cheriton. Sirpent^{TM}: a high-performance internetworking approach. Technical Report STAN-CS-89-1273, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1989. ii + 12

pp. URL http://www-db.stanford.edu/TR/CS-TR-89-1273.html.

Hirsh:1989:IVS

[1299] Haym Hirsh. Incremental version-space merging: a general framework for concept learning. Technical Report STAN-CS-89-1274, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 103 pp.

Subramanian:1989:NAS

[1300] Ashok Subramanian. A new approach to stable matching problems. Technical Report STAN-CS-89-1275, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 35 pp. URL

http://www-db.stanford.edu/TR/CS-TR-89-1275.html.

Plaxton:1989:NCS

[1301] C. Greg Plaxton. On the network complexity of selection. Technical Report STAN-CS-89-1276, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 20 pp. URL http://www-

db.stanford.edu/TR/CS-TR-89-1276.html.

Mitchell:1989:TSP

[1302] John C. Mitchell. Type systems for programming languages. Technical Report STAN-CS-89-1277, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 90 pp.

Mayr:1989:CCV

[1303] Ernst W. Mayr and Ashok Subramanian. The complexity of circuit value and network stability. Technical Report STAN-CS-89-1278, Stanford University, Department of Computer Science, Stanford, CA, USA, Au-

gust 1989. 19 pp. URL http://www-db.stanford.edu/TR/CS-TR-89-
1278.html.

Zhu:1989:HAE

[1304] David Zhu and Jean-Claude Latombe. The heuristic algorithms for efficient hierarchical path planning. Technical Report STAN-CS-89-1279, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 44 pp.

Plotkin:1989:SBU

[1305] Serge A. Plotkin. Sticky bits and universality of consensus. Technical Report STAN-CS-89-1280, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 19 pp. URL http://www-

db.stanford.edu/TR/CS-TR-89-1280.html.

Plaxton:1989:LBH

[1306] C. Greg Plaxton. Load balancing on the hypercube and shuffle-exchange. Technical Report STAN-CS-89-1281, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 19 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-89-1281.html.

Sankar:1989:ARC

[1307] Sriram Sankar. Automatic runtime consistency checking and debugging of formally specified programs. Technical Report STAN-CS-89-1282, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1989. 210 pp.

Plaxton:1989:ECS

[1308] C. Gregory Plaxton. Efficient computation on sparse interconnection networks. Technical Report STAN-CS-89-1283, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1989. 122 pp.

Knuth:1989:TP

[1309] Donald E. Knuth. Theory and practice. Technical Report STAN-CS-89-1284, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1989. 25 pp. URL https://arxiv.org/pdf/cs/

9301114.

Barraquand:1989:NPF

[1310] J. Barraquand, B. Langlois, and J. Latombe. Numerical potential field techniques for robot path planning. Technical Report STAN-CS-89-1285, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1989. 39 pp.

Rothberg:1989:FSM

[1311] Edward Rothberg and Anoop Gupta. Fast sparse matrix factorization on modern workstations. Technical Report STAN-CS-89-1286, Stanford University, Department of Computer Science, Stanford, CA, USA, Octo-

ber 1989. ii + 15 pp. URL http://www-db.stanford.edu/TR/CS-TR-
89-1286.html.

DeMichiel:1989:PDO

[1312] Linda Gail DeMichiel. Performing database operations over mismatched domains. Technical Report STAN-CS-89-1287, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1989. 172 pp.

Talcott:1989:PPF

[1313] Carolyn Talcott. Programming and proving with function and control abstractions. Technical Report STAN-CS-89-1288, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1989.

121 pp. URL http://www-db.stanford.edu/TR/CS-TR-89-1288.html.

Healey:1989:UOI

[1314] Glenn Healey. The use of optical info in a machine vision system. Technical Report STAN-CS-89-1289, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1989. 131 pp.

Myers:1989:RLQ

[1315] Karen Myers, Devika Subramanian, and Ramin Zabih. Reading list for the qualifying examination in artificial intelligence. Technical Report STAN-CS-89-1290, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1989. 16 pp. URL http://www-

db.stanford.edu/TR/CS-TR-89-1290.html.

Quian:1989:DSD

[1316] Xiaolei Quian. The deductive synthesis of database transactions. Technical Report STAN-CS-89-1291, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1989. 194 pp.

Latombe:1989:RMP

[1317] Jean-Claude Latombe, Anthony Lazanas, and Shashank Shekhar. Robot motion planning with uncertainty in control and sensing. Technical Report STAN-CS-89-1292, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1989. iv + 46 pp.

URL https://apps.dtic.mil/sti/citations/ADA323613;   https:
//apps.dtic.mil/sti/pdfs/ADA323613.pdf.

Bronstein:1989:SFS

[1318] Alexandre Bronstein. MLP: String-functional semantics and Boyer– Moore mechanization for the formal verification of synchronous circuits. Technical Report STAN-CS-89-1293, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1989. viii + 279 pp.

URL https://www.proquest.com/pqdtglobal/docview/303871822.

Rice:1989:DIP

[1319] James Rice. The design and implementation of Poligon and a highperformance concurrent blackboard system shell. Technical Report STAN-CS-89-1294, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1989. 88 pp.

Langlotz:1989:DTA

[1320] Curtis Philip Langlotz. A decision-theoretic approach to heuristic planning. Technical Report STAN-CS-89-1295, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1989. 381 pp.

Manna:1989:CTP

[1321] Zohar Manna and Amir Pnueli. Completing the temporal picture. Technical Report STAN-CS-89-1296, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1989. 28 pp. URL

http://www-db.stanford.edu/TR/CS-TR-89-1296.html.

Casley:1989:TS

[1322] Ross Casley, Roger F. Crew, José Meseguer, and Vaughan Pratt. Temporal structures. Technical Report STAN-CS-89-1297, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1989.

43 pp. URL http://boole.stanford.edu/pub/man.pdf.

Tuminaro:1989:MAP

[1323] Ray Tuminaro. Multigrid algorithms on parallel processing systems. Technical Report STAN-CS-90-1299, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1989. 143 pp.

Pallas:1989:MSI

[1324] Joseph I. Pallas. Multiprocessor Smalltalk: Implementation, performance, and analysis. Technical Report STAN-CS-90-1315, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1989. 136 pp.

Gray:1990:LEF

[1325] Cary G. Gray and David R. Cheriton. Leases: an efficient fault-tolerant mechanism for distributed file cache consistency. Technical Report STAN-CS-90-1298, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1990. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-90-1298.html.

Guibas:1990:RIC

[1326] Leo Guibas, Donald E. Knuth, and Micha Sharir. Randomized incremental construction of Delaunay and Voronoi diagrams. Technical Report STAN-CS-90-1300, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1990. 36 pp.

Goldberg:1990:PEI

[1327] Andrew Goldberg. Processor-efficient implementation of a maximum flow algorithm. Technical Report STAN-CS-90-1301, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1990. 11 pp.

Larrabee:1990:EGT

[1328] Tracy Larrabee. Efficient generation of test patterns using Boolean satisfiability. Technical Report STAN-CS-90-1302, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1990.

68 pp. URL https://dl.acm.org/doi/10.5555/100195;   https://
ieeexplore.ieee.org/document/82368.

Wiederhold:1990:MAA

[1329] Gio Wiederhold, Tore Risch, Peter Rathmann, Linda DeMichiel, Surajit Chaudhuri, Byung Suk Lee, Kincho H. Law, Thierry Barsalou, and Dallan Quass. A mediator architecture for abstract data access. Technical Report STAN-CS-90-1303, Stanford University, Department of Computer Science, Stanford, CA, USA, February 23, 1990. vii + 164 pp.

URL https://apps.dtic.mil/sti/tr/pdf/ADA227362.pdf.

Matsushima:1990:MOI

[1330] Toshiyuki Matsushima and Gio Wiederhold. A model of object identities and values. Technical Report STAN-CS-90-1304, Stanford University, Department of Computer Science, Stanford, CA, USA, Febru-

ary 23, 1990. ii + 64 pp. URL http://i.stanford.edu/pub/
cstr/reports/cs/tr/90/1304/CS-TR-90-1304.pdf;   http://www-
db.stanford.edu/TR/CS-TR-90-1304.html.

Rothberg:1990:CEN

[1331] Edward Rothberg and Anoop Gupta. A comparative evaluation of nodal and supernodal parallel sparse matrix factorization: detailed simulation results. Technical Report STAN-CS-90-1305, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1990. 28

pp. URL http://www-db.stanford.edu/TR/CS-TR-90-1305.html.

Altman:1990:EMD

[1332] Russ Biagio Altman. Exclusion methods for the determination of protein structure from experimental data. Technical Report STAN-CS-90-1306, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1990. 205 pp.

Alur:1990:RTL

[1333] Rajeev Alur and Thomas A. Henzinger. Real-time logics: Complexity and expressiveness. Technical Report STAN-CS-90-1307, Stanford University, Department of Computer Science, Stanford, CA, USA, March

1990. 35 pp. URL http://www-db.stanford.edu/TR/CS-TR-90-1307.
html.

Patashnik:1990:OCS

[1334] Oren Patashnik. Optimal circuit segmentation for pseudo-exhaustive testing. Technical Report STAN-CS-90-1308, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1990. 110 pp.

Pratt:1990:DAW

[1335] Vaughan Pratt. Dynamic algebras as a well-behaved fragment of relation algebras. Technical Report STAN-CS-90-1309, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1990. 30 pp.

Barsalou:1990:VOR

[1336] Thierry Barsalou. View objects for relational databases. Technical Report STAN-CS-90-1310, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1990. 350 pp.

Subramanian:1990:CCC

[1337] Ashok Subramanian. The computational complexity of the circuit value and network stability problems. Technical Report STAN-CS-90-1311, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1990. 171 pp.

Kambhampati:1990:VSB

[1338] Subbarao Kambhampati and James A. Hendler. A validation structure based theory of plan modification and reuse. Technical Report STAN-CS-90-1312, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 56 pp. URL http://www-db.stanford.

edu/TR/CS-TR-90-1312.html.

Goldberg:1990:BRP

[1339] Andrew V. Goldberg and Dan Gusfield. Book review: Potokovye Algoritmy (Flow Algorithms) by G. M. Adel’son-Vel’ski, E. A. Dinic, and A. V. Karzanov. Technical Report STAN-CS-90-1313, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. ii +

14 pp. URL http://i.stanford.edu/pub/cstr/reports/cs/tr/90/
1313/CS-TR-90-1313.pdf;   http://www-db.stanford.edu/TR/CS-
TR-90-1313.html.

Koza:1990:GPP

[1340] John R. Koza. Genetic programming: a paradigm for genetically breeding populations of computer programs to solve problems. Technical Report STAN-CS-90-1314, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 130 pp. URL http://www-db.

stanford.edu/TR/CS-TR-90-1314.html.

Heckerman:1990:PSN

[1341] David Earl Heckerman. Probabilistic similarity networks. Technical Report STAN-CS-90-1316, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 285 pp.

Barraquand:1990:CMR

[1342] Jerome Barraquand and Jean-Claude Latombe. Controllability of mobile robots with kinematic constraints. Technical Report STAN-CS-90-1317, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 26 pp.

Rothberg:1990:TIP

[1343] Edward Rothberg and Anoop Gupta. Techniques for improving the performance of sparse matrix factorization on multiprocessor workstations. Technical Report STAN-CS-90-1318, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 14 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-90-1318.html.

Bellin:1990:MPT

[1344] Gianluigi Bellin. Mechanizing proof theory: Resource-aware logics and proof transformations to extract implicit information. Technical Report STAN-CS-90-1319, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 237 pp.

Jimison:1990:RGI

[1345] Holly Brugge Jimison. A representation for gaining insight into clinical decision models. Technical Report STAN-CS-90-1320, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 202 pp.

Manna:1990:TRP

[1346] Zohar Manna and Amir Pnueli. Tools and rules for the practicing verifier. Technical Report STAN-CS-90-1321, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1990. 35 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-90-1321.html.

Shoham:1990:ESF

[1347] Yoav Shoham and Fangzhen Lin. Epistemic semantics for fixed-point nonmonotonic logics. Technical Report STAN-CS-90-1322, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1990. 11 pp.

Moses:1990:P

[1348] Eyal Mozes and Yoav Shoham. Protograms. Technical Report STAN-CS-90-1323, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1990. i + 18 pp. URL http://i.stanford.edu/

pub/cstr/reports/cs/tr/90/1323/CS-TR-90-1323.pdf;   http://
www-db.stanford.edu/TR/CS-TR-90-1323.html.

Guerreira:1990:CMIb

[1349] Ramiro A. de T. Guerreira, Andrea S. Hemerly, and Yoav Shoham. On the complexity of monotonic inheritance with roles. Technical Report STAN-CS-90-1324, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1990. ii + 6 pp. URL

http://i.stanford.edu/pub/cstr/reports/cs/tr/90/1324/CS-TR-
90-1324.pdf;   http://www-db.stanford.edu/TR/CS-TR-90-1324.
html.

Lavignon:1990:TA

[1350] Jean-François Lavignon and Yoav Shoham. Temporal automata. Technical Report STAN-CS-90-1325, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1990. 41 pp.

Kaelbling:1990:LES

[1351] Leslie Pack Kaelbling. Learning in embedded systems. Technical Report STAN-CS-90-1326, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1990. 199 pp.

Plambeck:1990:STC

[1352] Thane E. Plambeck. Semigroups and transitive closure. Technical Report STAN-CS-90-1327, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1990. 123 pp.

Radzik:1990:TBN

[1353] Tomasz Radzik and Andrew V. Goldberg. Tight bounds on the number of minimum-mean cycle cancellations. Technical Report STAN-CS-90-1328, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1990. 19 pp.

Henzinger:1990:IMR

[1354] Thomas A. Henzinger, Zohar Manna, and Amir Pnueli. An interleaving model for real time. Technical Report STAN-CS-90-1329, Stanford University, Department of Computer Science, Stanford, CA, USA, Septem-

ber 1990. 36 pp. URL http://www-db.stanford.edu/TR/CS-TR-90-
1329.html.

Rothberg:1990:PIH

[1355] Edward Rothberg and Anoop Gupta. Parallel ICCG on a hierarchical memory multiprocessor — addressing the triangular solve bottleneck. Technical Report STAN-CS-90-1330, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1990. i + 20 pp.

URL http://www-db.stanford.edu/TR/CS-TR-90-1330.html.

Snoeyink:1990:TAC

[1356] Jack Scott Snoeyink. Topological approaches in computational geometry. Technical Report STAN-CS-90-1331, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1990. 104 pp.

Schoen:1990:IAD

[1357] Eric Jonathan Schoen. Intelligent assistance for the design of knowledgebased systems. Technical Report STAN-CS-90-1332, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1990. 258 pp.

Spreitzer:1990:CSD

[1358] Michael J. Spreitzer. Comparing structurally different views of a VLSI design. Technical Report STAN-CS-90-1333, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1990. 161 pp.

Blatt:1990:SCW

[1359] Miriam Greta Blatt. Soft configurable wafer scale integration design, implementation, and yield analysis. Technical Report STAN-CS-90-1334, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1990. vii + 116 pp. URL https://apps.dtic.mil/sti/

tr/pdf/ADA326034.pdf.

Shoham:1990:AOP

[1360] Yoav Shoham. Agent-oriented programming. Technical Report STAN-CS-90-1335, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1990. 52 pp.

Knuth:1990:CUM

[1361] Donald E. Knuth. CWEB user manual: The CWEB system of structured documentation. Technical Report STAN-CS-90-1336, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1990. 214 pp.

Galbiati:1990:SUL

[1362] Louis Galbiati and Carolyn Talcott. A simplifier for untyped lambda expressions. Technical Report STAN-CS-90-1337, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1990.

ii + 26 pp. URL http://www-db.stanford.edu/TR/CS-TR-90-1337.
html.

Haddad:1990:TTP

[1363] Ramsey W. Haddad. Triangularization: a two-processor schedule problem. Technical Report STAN-CS-90-1338, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1990. 126 pp.

Nazarian:1990:BDC

[1364] Taleen Nazarian. Bibliography, Department of Computer Science technical reports, 1963–1990. Technical Report STAN-CS-90-1339, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1990. 86 pp.

Mason:1990:PQ

[1365] Ian A. Mason, Joseph D. Pehoushek, Carolyn L. Talcott, and Joseph S. Weening. Programming in QLisp. Technical Report STAN-CS-90-1340, Stanford University, Department of Computer Science, Stanford, CA, USA, October 24, 1990. iv + 56 pp. URL http://www-db.stanford.

edu/TR/CS-TR-90-1340.html.

Wiederhold:1990:TM

[1366] G. Wiederhold, P. Wagner, and Stefano Ceri. Towards megaprogramming. Technical Report STAN-CS-90-1341, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1990. 45 pp.

Pratt:1990:MCG

[1367] Vaughan Pratt. Modeling concurrency with geometry. Technical Report STAN-CS-90-1342, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1990. 13 pp. URL http://www-

db.stanford.edu/TR/CS-TR-90-1342.html.

Pratt:1990:ALP

[1368] Vaughan Pratt. Action logic and pure induction. Technical Report STAN-CS-90-1343, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1990. 23 pp. URL http://www-

db.stanford.edu/TR/CS-TR-90-1343.html.

Cheriton:1990:PHS

[1369] David R. Cheriton, Hendrik A. Goosen, and Patrick D. Boyle. ParaDiGM: a highly scalable shared-memory multi-computer architecture. Technical Report STAN-CS-90-1344, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1990. 19 pp.

URL http://www-db.stanford.edu/TR/CS-TR-90-1344.html.

Laumond:1990:NMP

[1370] Jean-Paul Laumond. Nonholonomic motion planning versus controllability via the multibody car system example. Technical Report STAN-CS-90-1345, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1990. i + 52 pp. URL http://www-db.

stanford.edu/TR/CS-TR-90-1345.html.

Lee:1990:EIO

[1371] Byung S. Lee. Efficiency in instantiating objects from relational databases through views. Technical Report STAN-CS-90-1346, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1990. 147 pp.

Merchant:1990:STB

[1372] Arif Merchant. Settling time bounds for M—G—1 queues. Technical Report STAN-CS-91-1349, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1990. 6 pp.

Horvitz:1990:CAU

[1373] Eric Horvitz. Computation and action under bounded resources. Technical Report STAN-CS-92-1430 (KSL-90-76), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1990. 320 pp.

Merchant:1991:AMP

[1374] Arif Merchant. Analytical models for the performance analysis of banyan networks. Technical Report STAN-CS-91-1347, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1991. 120 pp.

Merchant:1991:MCA

[1375] Arif Merchant. A Markov chain approximation for the analysis of Banyan networks. Technical Report STAN-CS-91-1348, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1991. 16 pp.

Chang:1991:PPS

[1376] Edward Chang, Steven J. Phillips, and Jeffrey D. Ullman. A programming and problem solving seminar. Technical Report STAN-CS-91-1350, Stanford University, Department of Computer Science, Stanford, CA,

USA, February 1991. 99 pp. URL http://www-db.stanford.edu/TR/
CS-TR-91-1350.html.

Zhu:1991:SVP

[1377] Liping Zhu, Arthur M. Keller, and Gio Wiederhold. Sequence vs. pipeline parallel multiple joins in Paradata. Technical Report STAN-CS-91-1351, Stanford University, Department of Computer Science, Stanford, CA,

USA, February 1991. 79 pp. URL http://www-db.stanford.edu/TR/
CS-TR-91-1351.html.

Chavez:1991:AAA

[1378] R. Martin Chavez. Architectures and approximation algorithms for probabilistic systems. Technical Report STAN-CS-91-1352, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1991. 214 pp.

Phipps:1991:GMV

[1379] Geoffrey Phipps. Glue manual: Version 1.0. Technical Report STAN-CS-91-1353, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1991. 32 pp.

Winograd:1991:IPP

[1380] Terry Winograd. Introduction to the Project on People, Computers, and Design. Technical Report STAN-CS-91-1354 (CSLI-91-150, PCD-1), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1991. 31 pp.

Casley:1991:SCS

[1381] Ross Casley. On the specification of concurrent systems. Technical Report STAN-CS-91-1355, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1991. 101 pp.

Saraiya:1991:SEA

[1382] Yatin Saraiya. Subtree elimination algorithms in deductive databases. Technical Report STAN-CS-91-1356, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1991. 159 pp.

Christensen:1991:AAP

[1383] Jens Christensen. Automatic abstraction planning. Technical Report STAN-CS-91-1357, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1991. 140 pp.

Goldberg:1991:COL

[1384] Andrew V. Goldberg. Combinatorial optimization lecture notes for CS363/OR349 winter 1991. Technical Report STAN-CS-91-1358, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1991. 78 pp.

Alur:1991:BRP

[1385] Rajeev Alur, Tomas Feder, and Thomas A. Henzinger. The benefits of relaxing punctuality. Technical Report STAN-CS-91-1359, Stanford University, Department of Computer Science, Stanford, CA, USA, May

1991. 38 pp. URL http://www-db.stanford.edu/TR/CS-TR-91-1359.
html.

Henzinger:1991:SST

[1386] Thomas A. Henzinger. Sooner is safer than later. Technical Report STAN-CS-91-1360, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1991. 9 pp. URL http://www-db.

stanford.edu/TR/CS-TR-91-1360.html.

Gangolli:1991:CBM

[1387] Anil Ramesh Gangolli. Convergence bounds for Markov chains and applications to sampling. Technical Report STAN-CS-91-1361, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1991. 153 pp.

Feder:1991:SNP

[1388] Tomas Feder. Stable networks and product graphs. Technical Report STAN-CS-91-1362, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1991. 220 pp.

Gray:1991:PFT

[1389] Cary G. Gray. Performance and fault-tolerance in a cache for distributed file service. Technical Report STAN-CS-91-1363, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1991. 140 pp.

Winograd:1991:LAA

[1390] Terry Winograd and Finn Kensing. The language/action approach to the design of computer-support for cooperative work: a preliminary study in work mapping. Technical Report STAN-CS-91-1364 (CSLI-91-152, PCD-2), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1991. 31 pp.

Paek:1991:JBT

[1391] Eunok Paek. A justification-based theory of explanation. Technical Report STAN-CS-91-1365, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1991. 103 pp.

Cohen:1991:CAO

[1392] Edith Cohen. Combinatorial algorithms for optimization problems. Technical Report STAN-CS-91-1366, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1991. 168 pp.

Herskovits:1991:CBP

[1393] Edward Herskovits. Computer-based probabilistic-network construction. Technical Report STAN-CS-91-1367, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1991. 215 pp.

Friedman:1991:CAC

[1394] Joseph Friedman. Computational aspects of compliant motion planning. Technical Report STAN-CS-91-1368, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1991. 122 pp.

Fischer:1991:AMP

[1395] Ted Fischer, Andrew V. Goldberg, and Serge Plotkin. Approximating matchings in parallel. Technical Report STAN-CS-91-1369, Stanford University, Department of Computer Science, Stanford, CA, USA, June

1991. 5 pp. URL http://www-db.stanford.edu/TR/CS-TR-91-1369.
html.

Nagayama:1991:NME

[1396] Misao Nagayama and Carolyn Talcott. An NQTHM mechanization of “An Exercise in the Verification of Multi-Process Programs”. Technical Report STAN-CS-91-1370, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1991. ii + 84 pp. URL

http://i.stanford.edu/pub/cstr/reports/cs/tr/91/1370/CS-TR-
91-1370.pdf;   http://www-db.stanford.edu/TR/CS-TR-91-1370.
html.

Rathmann:1991:NSP

[1397] Peter Rathmann. Nonmonotonic semantics for partitioned knowledge bases. Technical Report STAN-CS-91-1371, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1991. 116 pp.

Goldberg:1991:NRS

[1398] A. V. Goldberg. A natural randomization strategy for multicommodity flow and related algorithms. Technical Report STAN-CS-91-1372, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1991. 11 pp.

Kanakia:1991:HPH

[1399] Hemant Ratubhai Kanakia. High-performance host interfacing for packet-switched networks. Technical Report STAN-CS-91-1373, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1991. 53 pp.

Orlin:1991:PDN

[1400] James B. Orlin, Serge A. Plotkin, and Eva Tardos. Polynomial dual network simplex algorithms. Technical Report STAN-CS-91-1374 (AD-A254340), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. 28 pp. URL http://www-

db.stanford.edu/TR/CS-TR-91-1374.html;   https://apps.dtic.
mil/sti/tr/pdf/ADA254340.pdf.

Leighton:1991:FAA

[1401] Tom Leighton, Fillia Makedon, Serge Plotkin, Clifford Stein, Eva Tardos, and Spyros Tragoudas. Fast approximation algorithms for multicommodity flow problems. Technical Report STAN-CS-91-1375, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991.

25 pp. URL http://www-db.stanford.edu/TR/CS-TR-91-1375.html.

Strat:1991:NOR

[1402] Thomas M. Strat. Natural object recognition. Technical Report STAN-CS-91-1376, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. xv + 156 pp. URL https://apps.

dtic.mil/sti/tr/pdf/ADA325974.pdf.

Rothberg:1991:ELL

[1403] Edward Rothberg and Anoop Gupta. An evaluation of left-looking, right-looking and multifrontal approaches to sparse Cholesky factorization and hierarchical-memory machines. Technical Report STAN-CS-91-1377 (CSL-TR-91-487), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. ii + 47 pp. URL http://

www-db.stanford.edu/TR/CS-TR-91-1377.html.

Alur:1991:TAV

[1404] Rajeev Alur. Techniques for automatic verification of real-time systems. Technical Report STAN-CS-91-1378, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. 188 pp.

Hall:1991:FCM

[1405] Keith Hall. A framework for change management in a design database. Technical Report STAN-CS-91-1379, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. 185 pp.

Henzinger:1991:TSV

[1406] Thomas Henzinger. The temporal specification and verification of realtime systems. Technical Report STAN-CS-91-1380, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. 300 pp.

Hara:1991:IHD

[1407] Yoshinori Hara, Arthur M. Keller, Peter K. Rathmann, and Gio Wiederhold. Implementing hypertext database relationships through aggregations and exceptions. Technical Report STAN-CS-91-1381, Stanford University, Department of Computer Science, Stanford, CA, USA, Septem-

ber 1991. 36 pp. URL http://www-db.stanford.edu/TR/CS-TR-91-
1381.html.

Morris:1991:SOQ

[1408] Katherine A. Morris. Subgoal order for query optimization in logic databases. Technical Report STAN-CS-91-1382, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1991. 116 pp.

Henzinger:1991:TPM

[1409] Thomas A. Henzinger, Zohar Manna, and Amir Pnueli. Temporal proof methodologies for real-time systems. Technical Report STAN-CS-91-1383, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1991. 49 pp. URL http://www-db.stanford.

edu/TR/CS-TR-91-1383.html.

Myers:1991:UAI

[1410] Karen L. Myers. Universal attachment: an integration method for logic hybrids. Technical Report STAN-CS-91-1384, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1991. 188 pp.

Lin:1991:SNR

[1411] Fangzhen Lin. A study of nonmonotonic reasoning. Technical Report STAN-CS-91-1385, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. 106 pp.

Ross:1991:SDD

[1412] Kenneth A. Ross. The semantics of deductive databases. Technical Report STAN-CS-91-1386, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. 168 pp.

Wilson:1991:APS

[1413] Randall Wilson and Achim Schweikard. Assembling polyhedra with single translations. Technical Report STAN-CS-91-1387, Stanford University, Department of Computer Science, Stanford, CA, USA, October

1991. 16 pp. URL http://www-db.stanford.edu/TR/CS-TR-91-1387.
html.

Whang:1991:DMD

[1414] Kyu-Young Whang, Sang-Wook Kim, and Gio Wiederhold. Dynamic maintenance of data distribution for selectivity estimation. Technical Report STAN-CS-91-1388, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1991. 30 pp.

Torrance:1991:AM

[1415] Mark C. Torrance and Paul A. Viola. The AGENT0 manual. Technical Report STAN-CS-91-1389, Stanford University, Department of Computer Science, Stanford, CA, USA, April 9, 1991. ii + 8 pp. URL

http://www-db.stanford.edu/TR/CS-TR-91-1389.html.

Shoham:1991:VC

[1416] Yoav Shoham. Varieties of context. Technical Report STAN-CS-91-1390, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1991. 15 pp.

Shoham:1991:LPB

[1417] Yoav Shoham and Alvaro del Val. A logic for perception and belief. Technical Report STAN-CS-91-1391, Stanford University, Department of Computer Science, Stanford, CA, USA, September 24, 1991. ii + 23

pp. URL http://www-db.stanford.edu/TR/CS-TR-91-1391.html.

Ceri:1991:CUM

[1418] Stefano Ceri, Maurice A. W. Houtsma, Arthur M. Keller, and Pierangela Samarati. A classification of update methods for replicated databases. Technical Report STAN-CS-91-1392, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1991. 17 pp. URL

http://www-db.stanford.edu/TR/CS-TR-91-1392.html.

Goosen:1991:SMC

[1419] Hendrik A. Goosen. Shared multilevel caches for scalable multiprocessors. Technical Report STAN-CS-91-1393, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1991. 76 pp.

Harty:1991:ACP

[1420] Kieran Harty and David R. Cheriton. Application-controlled physical memory using external page-cache management. Technical Report STAN-CS-91-1394, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1991. 16 pp. URL http://www-db.

stanford.edu/TR/CS-TR-91-1394.html.

Hartfield:1991:LHD

[1421] Brad Hartfield, Terry Winograd, and John Bennett. Learning HCI design: Mentoring project groups in a course on human–computer interaction. Technical Report STAN-CS-91-1395 (CSLI-91-161, PCD-3), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1991. 20 pp.

Scales:1991:POM

[1422] Daniel J. Scales. Parallelizing the OPS5 matching algorithm in Qlisp. Technical Report STAN-CS-91-1396 (AD-A323588PB96-148556), Stanford University, Department of Computer Science, Stanford, CA, USA,

October 1991. iii + 20 pp. URL https://apps.dtic.mil/sti/tr/pdf/
ADA323588.pdf.

Roy:1991:AMP

[1423] Shaibal Roy. Adaptive methods in parallel databases. Technical Report STAN-CS-91-1397, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1991. 160 pp.

Salesin:1991:EGB

[1424] David Salesin. Epsilon geometry: Building robust algorithms from imprecise computations. Technical Report STAN-CS-91-1398, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1991. 130 pp.

Guha:1991:CFS

[1425] Ramanathan Guha. Contexts: a formalization and some applications. Technical Report STAN-CS-91-1399 (KSL-91-74), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1991. 280 pp.

Mumick:1991:QOD

[1426] Inderpal Mumick. Query optimization in deductive and relational databases. Technical Report STAN-CS-91-1400, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1991. 190 pp.

Lehmann:1991:BCB

[1427] Harold P. Lehmann. A Bayesian computer-based approach to the physician’s use of the clinical research literature. Technical Report STAN-CS-92-1402 (KSL-91-76), Stanford University, Department of Computer Science, Stanford, CA, USA, December 1991. xxii + 277 pp.

Golding:1991:PNC

[1428] Andrew Golding. Pronouncing names by a combination of rule-based and case-based reasoning. Technical Report STAN-CS-92-1403, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1991. 380 pp.

Cha:1991:KMB

[1429] Sang Kyun Cha. Kaleidoscope: a model-based grammar-driven menu interface for databases. Technical Report STAN-CS-92-1405, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1991. 141 pp.

Williamson:1991:MRT

[1430] Carey Lee Williamson. Minimizing round-trip times for highperformance transport communication. Technical Report STAN-CS-92-1409, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1991. 129 pp.

Deering:1991:MRD

[1431] Stephen Deering. Multicast routing in a datagram internetwork. Technical Report STAN-CS-92-1415, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1991. 141 pp.

Rothberg:1992:PID

[1432] Edward Rothberg and Anoop Gupta. The performance impact of data reuse in parallel dense Cholesky factorization. Technical Report STAN-CS-92-1401 (CSL-TR-92-503), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1992. 30 pp. URL http://

www-db.stanford.edu/TR/CS-TR-92-1401.html.

Manna:1992:FDP

[1433] Zohar Manna and Richard Waldinger. Fundamentals of deductive program synthesis. Technical Report STAN-CS-92-1404, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1992. 62 pp.

Radzik:1992:NMF

[1434] Tomasz Radzik. Newton’s method for fractional combinatorial optimization. Technical Report STAN-CS-92-1406, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1992. 22 pp.

Maler:1992:THS

[1435] Oded Maler, Zohar Manna, and Amir Pnueli. From timed to hybrid systems. Technical Report STAN-CS-92-1407, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1992. 38 pp.

Chang:1992:SPC

[1436] Edward Chang, Zohar Manna, and Amir Pnueli. The safety-progress classification. Technical Report STAN-CS-92-1408, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1992. 60 pp.

Alizadeh:1992:IPR

[1437] Farid Alizadeh and Andrew Goldberg. Implementing the push-relabel method for the maximum flow problem on the connection machine. Technical Report STAN-CS-92-1410, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1992. 17 pp.

Wong-Toi:1992:CDR

[1438] Howard Wong-Toi and Gerard Hoffmann. The control of dense real-time discrete event systems. Technical Report STAN-CS-92-1411, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1992. 50 pp.

Nilsson:1992:TAP

[1439] Nils J. Nilsson. Toward agent programs with circuit semantics. Technical Report STAN-CS-92-1412, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1992. 33 pp. URL http://

www-db.stanford.edu/TR/CS-TR-92-1412.html.

Zhu:1992:EIG

[1440] David Zhu. Exploring the interaction of geometry and search in path planning. Technical Report STAN-CS-92-1413, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1992. 203 pp.

Miller:1992:PTF

[1441] Robert Miller. Proceedings from the Twenty-Fourth Annual Meeting of the Stanford Computer Forum. Technical Report STAN-CS-92-1414, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1992. 186 pp.

Wilson:1992:GAP

[1442] Randall Wilson. On geometric assembly planning. Technical Report STAN-CS-92-1416, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1992. 154 pp.

Suermondt:1992:EBB

[1443] Henri Jacques Suermondt. Explanation in Bayesian belief networks. Technical Report STAN-CS-92-1417, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1992. 263 pp.

Goldberg:1992:ISP

[1444] Andrew Goldberg and Michael Kharitonov. On implementing scaling push–relabel algorithms for the minimum cost flow problems. Technical Report STAN-CS-92-1418, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1992. 37 pp.

Plotkin:1992:FAA

[1445] Serge A. Plotkin, David B. Shmoys, and Éva Tardos. Fast approximation algorithms for fractional packing and covering problems. Technical Report STAN-CS-92-1419, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1992. ii + 52 pp. URL http://

www-db.stanford.edu/TR/CS-TR-92-1419.html.

Chambers:1992:DIS

[1446] Craig Chambers. The design and implementation of the self compiler, an optimizing compiler for object-oriented programming languages. Technical Report STAN-CS-92-1420, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1992. 246 pp.

Wang:1992:FTC

[1447] Alexander Wang. Fault-tolerant computation on hypercubes. Technical Report STAN-CS-92-1421, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1992. 148 pp.

Walker:1992:PEC

[1448] Michael Walker. Probability estimation for classification trees and DNA sequence analysis. Technical Report STAN-CS-92-1422 (KSL-92-27), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1992. 213 pp.

Dwork:1992:TLS

[1449] Cynthia Dwork, Maurice Herlihy, Serge A. Plotkin, and Orli Waarts. Time-lapse snapshots. Technical Report STAN-CS-92-1423, Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1992. ii + 19 pp. URL http://i.stanford.edu/pub/
cstr/reports/cs/tr/92/1423/CS-TR-92-1423.pdf;   http://www-
db.stanford.edu/TR/CS-TR-92-1423.html.

Takeda:1992:PMM

[1450] Haruo Takeda and Jean-Claude Latombe. Planning the motions of a mobile robot in a sensory uncertainty field. Technical Report STAN-CS-92-1424, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1992. 31 pp.

Kavraki:1992:CCS

[1451] Lydia Kavraki. Computation of configuration-space obstacles using the Fast Fourier Transform. Technical Report STAN-CS-92-1425, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1992. 21 pp.

Danvy:1992:PAS

[1452] Olivier Danvy and Carolyn Talcott. Proceedings of the ACM SIGPLAN Workshop on Continuations CW92. Technical Report STAN-CS-92-1426, Stanford University, Department of Computer Science, Stanford, CA, USA, June 21, 1992. ii + 103 pp. URL http://i.stanford.edu/

pub/cstr/reports/cs/tr/92/1426/CS-TR-92-1426.pdf;   http://
www-db.stanford.edu/TR/CS-TR-92-1426.html.

Goldberg:1992:PAR

[1453] Andrew Goldberg, Bruce Maggs, and Serge Plotkin. A parallel algorithm for reconfiguring a multibutterfly network with faulty switches. Technical Report STAN-CS-92-1427, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1992. 15 pp.

Lazanas:1992:LBR

[1454] Anthony Lazanas and Jean-Claude Latombe. Landmark-based robot navigation. Technical Report STAN-CS-92-1428, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1992. 54 pp.

Goldberg:1992:SAS

[1455] Andrew Goldberg. Scaling algorithms for the shortest paths problem. Technical Report STAN-CS-92-1429, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1992. 11 pp.

Garcia-Molina:1992:ATR

[1456] Hector Garcia-Molina, Ben Kao, and Daniel Barbará. Aggressive transmissions over redundant paths for time critical messages. Technical Report STAN-CS-92-1431, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1992. ii + 43 pp. URL

http://i.stanford.edu/pub/cstr/reports/cs/tr/92/1431/CS-TR-
92-1431.pdf;   http://www-db.stanford.edu/TR/CS-TR-92-1431.
html.

Chaudhuri:1992:DRT

[1457] Surajit Chaudhuri. Detecting redundant tuples during query evaluation. Technical Report STAN-CS-92-1433, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1992. 230 pp.

Tomasic:1992:PII

[1458] Anthony Tomasic and Hector Garcia-Molina. Performance of inverted indices in distributed text document retrieval systems. Technical Report STAN-CS-92-1434, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1992. 25 pp.

Motwani:1993:LNA

[1459] Rajeev Motwani. Lecture notes on approximation algorithms — Volume I. Technical Report STAN-CS-92-1435, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1992. ii + 132 pp. URL

http://www-db.stanford.edu/TR/CS-TR-92-1435.html.

Winograd:1992:CES

[1460] Terry Winograd. Computers, ethics, and social responsibility. Technical Report STAN-CS-92-1436, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1992. 30 pp.

Phipps:1992:GDD

[1461] Geoffrey Phipps. Glue: a deductive database programming language. Technical Report STAN-CS-92-1437, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1992. 139 pp.

Rothberg:1992:EBO

[1462] Edward Rothberg and Anoop Gupta. An efficient block-oriented approach to parallel sparse Cholesky factorization. Technical Report STAN-CS-92-1438 (CSL-TR-92-533, PB96-151543), Stanford University, Department of Computer Science, Stanford, CA, USA, July 13, 1992. iv + 25 pp.

Goldberg:1992:EIS

[1463] Andrew Goldberg. An efficient implementation of a scaling Minimum-Cost flow algorithm. Technical Report STAN-CS-92-1439, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1992. 21 pp.

Woodfill:1992:MVT

[1464] John Woodfill. Motion vision and tracking for robots in dynamic, unstructured environments. Technical Report STAN-CS-92-1440, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1992. 163 pp.

Schweikard:1992:MPS

[1465] Achim Schweikard, John R. Adler, and Jean-Claude Latombe. Motion planning in stereotaxic radiosurgery. Technical Report STAN-CS-92-1441, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1992. 20 pp. URL http://www-db.stanford.

edu/TR/CS-TR-92-1441.html.

Wilson:1992:RAM

[1466] Randall H. Wilson and Jean-Claude Latombe. Reasoning about mechanical assembly. Technical Report STAN-CS-92-1442, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1992. 31 pp.

Nayak:1992:AMP

[1467] P. Pandurang Nayak. Automated modeling of physical systems. Technical Report STAN-CS-92-1443 (KSL-92-69, AD-A263755), Stanford University, Department of Computer Science, Stanford, CA, USA, Septem-

ber 1992. xix + 295 pp. URL https://apps.dtic.mil/sti/citations/
ADA263755;   https://apps.dtic.mil/sti/pdfs/ADA263755.pdf.

Goldberg:1992:MPD

[1468] Aaron J. Goldberg. Multiprocessor performance debugging and memory bottlenecks. Technical Report STAN-CS-92-1444 (CSL-TR-92-542, AD-A268387), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1992. xi + 124 pp. URL https://apps.dtic.

mil/sti/tr/pdf/ADA268387.pdf.

Kharitonov:1992:CHD

[1469] Michael Kharitonov. Cryptographic hardness of distribution-specific learning. Technical Report STAN-CS-92-1445, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1992. 24

pp. URL https://www.cis.upenn.edu/~mkearns/teaching/Crypto/
kharitonov.pdf.

Ceri:1992:IUI

[1470] Stefano Ceri, Maurice A. W. Houtsma, Arthur M. Keller, and Pierangela Samarati. Independent updates and incremental agreement in replicated databases. Technical Report STAN-CS-92-1446, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1992. 13

pp. URL http://www-db.stanford.edu/TR/CS-TR-92-1446.html.

Goldberg:1992:LNT

[1471] Andrew V. Goldberg and Serge Plotkin. Lecture notes: Topics in combinatorial optimization. Technical Report STAN-CS-92-1447, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1992. 85 pp.

Crew:1992:MPM

[1472] Roger F. Crew. Metric process models. Technical Report STAN-CS-92-1448, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1992. 87 pp.

Maydan:1992:AAA

[1473] Dror Eliezer Maydan. Accurate analysis of array references. Technical Report STAN-CS-92-1449 (CSL-TR-92-547, AD-A268069), Stanford University, Department of Computer Science, Stanford, CA, USA, September 1992. viii + 134 pp. URL https://apps.dtic.mil/sti/

pdfs/ADA268069.pdf.

Waarts:1992:NAP

[1474] Orli Waarts. New algorithms and primitives for multi processor coordination. Technical Report STAN-CS-92-1450, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1992. 165 pp.

Radzik:1992:ASL

[1475] Tomasz Radzik. Algorithms for some linear and fractional combinatorial optimization problems. Technical Report STAN-CS-92-1451, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1992. 100 pp.

Kao:1992:DAD

[1476] Ben Kao and Hector Garcia-Molina. Deadline assignment in a distributed soft real-time system. Technical Report STAN-CS-92-1452, Stanford University, Department of Computer Science, Stanford, CA, USA, October

1992. 25 pp. URL http://www-db.stanford.edu/TR/CS-TR-92-1452.
html.

Kao:1992:RTC

[1477] Ben Kao and Hector Garcia-Molina. Real-time communication over multiple standard networks. Technical Report STAN-CS-92-1453, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1992. 23 pp.

Yan:1992:ISS

[1478] Tak W. Yan and Hector Garcia-Molina. Index structures for selective dissemination of information. Technical Report STAN-CS-92-1454, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1992. 48 pp.

Howard:1992:FPE

[1479] Brian Howard. Fixed points and extensionality in typed functional programming languages. Technical Report STAN-CS-92-1455, Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1992. 66 pp.

Tomasic:1992:CDS

[1480] Anthony Tomasic and Hector Garcia-Molina. Caching and database scaling in distributed shared-nothing information retrieval systems. Technical Report STAN-CS-92-1456, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1992. 21 pp.

Khatib:1992:IPR

[1481] Oussama Khatib. Inertial properties in robotics manipulation. Technical Report STAN-CS-92-1457, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1992. 34 pp.

Wilson:1992:CPA

[1482] Randall Wilson, Jean-Claude Latombe, and Tomas Lozano-Perez. On the complexity of partitioning an assembly. Technical Report STAN-CS-92-1458, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1992. 14 pp.

Rothberg:1992:EMH

[1483] Edward Rothberg. Exploiting the memory hierarchy in sequential and parallel sparse Cholesky factorization. Technical Report STAN-CS-92-1459 (AD-A262849CSL-TR-92-555), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1992. xii + 154 pp.

URL https://apps.dtic.mil/sti/tr/pdf/ADA262849.pdf.

Derr:1992:AOD

[1484] Marcia A. Derr. Adaptive optimization in a database programming language. Technical Report STAN-CS-92-1460, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1992. 143 pp.

Breitbart:1992:OMT

[1485] Yuri Breitbart, Hector Garcia-Molina, and Avi Silberschatz. Overview of multidatabase transaction management. Technical Report STAN-CS-92-1432, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1992. 38 pp. URL http://i.stanford.edu/

pub/cstr/reports/cs/tr/92/1432/CS-TR-92-1432.pdf;   http://
www-db.stanford.edu/TR/CS-TR-92-1432.html.

Manna:1993:MR

[1486] Zohar Manna and Amir Pnueli. Models for reactivity. Technical Report STAN-CS-93-1461, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1993. 73 pp.

Gunawardena:1993:PBT

[1487] Jeremy Gunawardena. Periodic behaviour in timed systems with {And/ OR} causality. Part I: Systems of dimension 1 and 2. Technical Report STAN-CS-93-1462, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1993. 25 pp.

Manna:1993:TPM

[1488] Zohar Manna and Amir Pnueli. A temporal PROOF methodology for reactive systems. Technical Report STAN-CS-93-1463, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1993. 37 pp.

Kharitonov:1993:CHM

[1489] Michael Kharitonov. Cryptographic hardness of machine learning. Technical Report STAN-CS-93-1464, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1993. 131 pp. URL https:

//books.google.com/books?id=QvREAQAAIAAJ;   https://dl.acm.
org/doi/book/10.5555/164350.

Gupta:1993:IPA

[1490] Ashish Gupta and Inderpal Singh Mumick. Improvement to the PF algorithm. Technical Report STAN-CS-93-1473, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1993. 4 pp.

Gunawardena:1993:MMF

[1491] Jeremy Gunawardena. Min-max functions. Technical Report STAN-CS-93-1474, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1993. 25 pp.

Pieper:1993:PCI

[1492] Karen Pieper. Parallelizing compilers: Implementation and effectiveness. Technical Report STAN-CS-93-1475, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1993. 151 pp.

Goldberg:1993:TFE

[1493] Andrew Goldberg and Alexander Karzanov. Transitive fork environments and minimum cost flows. Technical Report STAN-CS-93-1476, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1993. 44 pp.

Choi:1993:CTR

[1494] Wonyun Choi. Contingency-tolerant robot motion planning and control. Technical Report STAN-CS-93-1477, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1993. 130 pp.

Wilson:1993:THA

[1495] Randall Wilson, Lydia Kavraki, Tomas Lozano-Perez, and Jean-Claude Latombe. Two-handed assembly sequencing. Technical Report STAN-CS-93-1478, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1993. 26 pp.

Shekhar:1993:CUF

[1496] Shashank Shekhar. Control uncertainty in fine motion planning. Technical Report STAN-CS-93-1479, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1993. 110 pp.

Cherkassky:1993:SPA

[1497] Boris Cherkassky, Andrew Goldberg, and Tomas Radzik. Shortest paths algorithms: Theory and experimental evaluation. Technical Report STAN-CS-93-1480, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1993. 46 pp.

Goldberg:1993:ECS

[1498] Andrew V. Goldberg and Robert Kennedy. Efficient cost scaling algorithm for the assignment problem. Technical Report STAN-CS-93-1481, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1993. 24 pp.

Levy:1993:IRK

[1499] Alon Levy. Irrelevance reasoning in knowledge based systems. Technical Report STAN-CS-93-1482, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1993. 201 pp.

Avrahami:1993:IAC

[1500] Gideon Avrahami. Identification and analysis of curves in digital images. Technical Report STAN-CS-93-1483, Stanford University, Department of Computer Science, Stanford, CA, USA, ???? 1993. ?? pp.

Phillips:1993:TAO

[1501] Steven J. Phillips. Theory and applications of online algorithms. Technical Report STAN-CS-93-1484, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1993. 93 pp.

Cheriton:1993:ULC

[1502] David Cheriton and Dale Skeen. Understanding the limitations of causally and totally ordered communication. Technical Report STAN-CS-93-1485, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1993. 15 pp.

vanGlabbeek:1993:WBT

[1503] Robert van Glabbeek. What is branching time semantics and why to use it. Technical Report STAN-CS-93-1486, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1993. 8 pp.

Thomas:1993:PAO

[1504] Sarah Rebecca Thomas. PLACA: An agent oriented programming language. Technical Report STAN-CS-93-1487, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1993. 157 pp.

Takeda:1993:PMM

[1505] Haruo Takeda, Cladio Facchinetti, and Jean-Claude Latombe. Planning the motions of a mobile robot in a sensory uncertainty field. Technical Report STAN-CS-93-1488, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1993. 42 pp.

Goldberg:1993:PPS

[1506] Andrew V. Goldberg and Alexander V. Karzanov. Path problems in skew-symmetric graphs. Technical Report STAN-CS-93-1489, Stanford University, Department of Computer Science, Stanford, CA, USA, Oc-

tober 1993. i + 31 pp. URL https://alexander-karzanov.net/
Publications/93_path_probl_skew.pdf.

Kavraki:1993:RPC

[1507] Lydia Kavraki and Jean-Claude Latombe. Randomized preprocessing of configuration. Technical Report STAN-CS-93-1490, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1993. 18 pp.

Kao:1993:SDA

[1508] Ben Kao and Hector Garcia-Molina. Subtask deadline assignment for complex distributed soft real-time tasks. Technical Report STAN-CS-93-1491, Stanford University, Department of Computer Science, Stanford,

CA, USA, October 1993. 25 pp. URL http://www-db.stanford.edu/
TR/CS-TR-93-1491.html.

Jakobsson:1993:TBT

[1509] Hakan Jakobsson. Tree-based techniques for query evaluation. Technical Report STAN-CS-93-1492, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1993. 59 pp.

Davis:1993:MSA

[1510] Helen Davis. Multiprocessor simulation: Achieving accuracy, efficiency and flexibility. Technical Report STAN-CS-93-1493, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1993. 175 pp.

Yan:1993:ISI

[1511] Tak W. Yan and Hector Garcia-Molina. Index structures for information filtering under the vector space model. Technical Report STAN-CS-93-1494, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1993. i + 33 pp. URL http://www-db.stanford.

edu/TR/CS-TR-93-1494.html.

Baralis:1993:UDR

[1512] Elena Baralis and Jennifer Widom. Using delta relations to optimize condition evaluation in active databases. Technical Report STAN-CS-93-1495, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1993. 25 pp.

Hayes-Roth:1993:AAI

[1513] Barbara Hayes-Roth. An architecture for adaptive intelligent systems. Technical Report STAN-CS-93-1496, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1993. 51 pp.

Kutter:1993:OMB

[1514] Robert A. Kutter. Optimizing memory-based messaging for scalable shared-memory multiprocessor. Technical Report STAN-CS-93-1497, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1993. 82 pp.

Ash:1993:DUA

[1515] David Ash. Diagnosis using action-based hierarchies for optimal real-time performance. Technical Report STAN-CS-93-1498, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1993. 150 pp.

Knuth:1993:ST

[1516] Donald E. Knuth. The Sandwich Theorem. Technical Report STAN-CS-93-1499, Stanford University, Department of Computer Science, Stanford, CA, USA, December 6, 1993. ii + 49 pp. URL

http://i.stanford.edu/pub/cstr/reports/cs/tr/93/1499/CS-TR-
93-1499.pdf;   http://www-db.stanford.edu/TR/CS-TR-93-1499.
html;   http://www.ncstrl.org:8900/ncstrl/servlet/search?formname=
detail&id=oai%3Ancstrlh%3Astan%3ASTAN%2F%2FCS-TR-93-1499.

Chang:1993:CVR

[1517] Edward Chang. Compositional verification of reactive and real-time systems. Technical Report STAN-CS-94-1522, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1993. ?? pp.

URL http://www-db.stanford.edu/TR/CS-TR-94-1522.html.

Siroker:1994:PSS

[1518] Marianne Siroker. 1993 publications summary for the Stanford Database Group. Technical Report STAN-CS-94-1500, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1994. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1500.html.

delVal:1994:DPB

[1519] Alvaro del Val and Yoav Shoham. Deriving properties of belief update from theories of action. Technical Report STAN-CS-94-1501, Stanford University, Department of Computer Science, Stanford, CA, USA, Febru-

ary 1994. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-94-
1501.html.

Magerman:1994:NLP

[1520] David M. Magerman. Natural language parsing as statistical pattern recognition. Technical Report STAN-CS-94-1502, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1994.

161 pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1502.html.

Dabija:1994:DWP

[1521] Vlad G. Dabija. Deciding whether to plan to react. Technical Report STAN-CS-94-1503, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1994. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1503.html.

Baralis:1994:AAR

[1522] Elena Baralis and Jennifer Widom. An algebraic approach to rule analysis in expert database systems. Technical Report STAN-CS-94-1504, Stanford University, Department of Computer Science, Stanford, CA,

USA, February 1994. ?? pp. URL http://www-db.stanford.edu/TR/
CS-TR-94-1504.html.

Singhal:1994:UPH

[1523] Sandeep K. Singhal and David R. Cheriton. Using a position historybased protocol for distributed object visualization. Technical Report STAN-CS-94-1505, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1994. 27 pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1505.html.

Cheriton:1994:OMB

[1524] David R. Cheriton and Robert A. Kutter. Optimized memory-based messaging: Leveraging the memory system for high-performance communication. Technical Report STAN-CS-94-1506, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1994.

26 pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1506.html.

Mashack:1994:BDC

[1525] Thea Mashack. Bibliography Department of Computer Science technical reports, 1963–1993. Technical Report STAN-CS-94-1507, Stanford University, Department of Computer Science, Stanford, CA, USA, March

1994. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1507.
html.

Kondo:1994:IKH

[1526] Koichi Kondo. Inverse kinematics of a human arm. Technical Report STAN-CS-94-1508, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1994. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1508.html.

Goldberg:1994:GPU

[1527] Andrew V. Goldberg and Robert Kennedy. Global price updates help. Technical Report STAN-CS-94-1509, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1994. 20 pp. URL

http://www-db.stanford.edu/TR/CS-TR-94-1509.html.

Hayes:1994:KOG

[1528] Barry Hayes. Key objects in garbage collection. Technical Report STAN-CS-94-1510, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1994. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1510.html.

Shoham:1994:CLE

[1529] Yoav Shoham and Moshe Tennenholtz. Co-learning and the evolution of social activity. Technical Report STAN-CS-94-1511, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1994. 38

pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1511.html.

Washington:1994:APR

[1530] Richard Washington. Abstraction planning in real time. Technical Report STAN-CS-94-1512, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1994. 130 pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1512.html.

Egar:1994:CND

[1531] John W. Egar. Construction of normative decision models using abstract graph grammars. Technical Report STAN-CS-94-1513, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1994.

247 pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1513.html.

Hailperin:1994:LBU

[1532] Max Hailperin. Load balancing using time series analysis for soft real time systems with statistically periodic loads. Technical Report STAN-CS-94-1514, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1994. 147 pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1514.html.

Wolverton:1994:RSD

[1533] Michael Wolverton. Retrieving semantically distant analogies. Technical Report STAN-CS-94-1515, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1994. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-94-1515.html.

Goyal:1994:FRP

[1534] Nita Goyal. A framework for reasoning precisely with vague concepts. Technical Report STAN-CS-94-1516, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1994. ?? pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-94-1516.html.

vanGlabbeek:1994:RGS

[1535] Rob J. van Glabbeek, Scott A. Smolka, and Bernhard Steffen. Reactive, generative and stratified models of probabilistic processes. Technical Report STAN-CS-94-1517, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1994. 39 pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1517.html.

Manna:1994:SST

[1536] Zohar Manna, Anuchit Anuchitanukul, Nikolaj Bjorner, Anca Browne, Edward Chang, Michael Colon, Luca de Alfaro, Harish Devarajan, Henny Sipma, and Tomas Uribe. STeP: The Stanford Temporal Prover. Technical Report STAN-CS-94-1518, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1994. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-94-1518.html.

Kavraki:1994:PRP

[1537] Lydia Kavraki, Petr Svestka, Jean-Claude Latombe, and Mark Overmars. Probabilistic roadmaps for path planning in high-dimensional configuration spaces. Technical Report STAN-CS-94-1519, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1994.

37 pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1519.html.

Holzle:1994:AOS

[1538] Urs Holzle. Adaptive optimization for SELF: Reconciling high performance with exploratory programming. Technical Report STAN-CS-94-1520, Stanford University, Department of Computer Science, Stanford,

CA, USA, August 1994. ?? pp. URL http://www-db.stanford.edu/
TR/CS-TR-94-1520.html.

Gupta:1994:CSM

[1539] Vineet Gupta. Chu spaces: a model for concurrency. Technical Report STAN-CS-94-1521, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1994. 101 pp. URL http://www-

db.stanford.edu/TR/CS-TR-94-1521.html.

Cherkassky:1994:IPR

[1540] Boris V. Cherkassky and Andrew V. Goldberg. On implementing pushrelabel method for the maximum flow problem. Technical Report STAN-CS-94-1523, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1994. 21 pp. URL http://www-

db.stanford.edu/TR/CS-TR-94-1523.html.

deAlfaro:1994:CVD

[1541] Luca de Alfaro and Zohar Manna. Continuous verification by discrete reasoning. Technical Report STAN-CS-94-1524, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1994. 106

pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1524.html.

Anuchitanukul:1994:DB

[1542] Anuchit Anuchitanukul and Zohar Manna. Differential BDDs. Technical Report STAN-CS-94-1525, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1994. 12 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-94-1525.html.

Mohammed:1994:CET

[1543] John Llewelyn Mohammed. Combining experiential and theoretical knowledge in the domain of semiconductor manufacturing. Technical Report STAN-CS-94-1526, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1994. 265 pp. URL http://

www-db.stanford.edu/TR/CS-TR-94-1526.html.

Koller:1994:KB

[1544] Daphne Koller. From knowledge to belief. Technical Report STAN-CS-94-1527, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1994. 219 pp. URL http://www-db.stanford.

edu/TR/CS-TR-94-1527.html.

Koza:1994:AAO

[1545] John R. Koza. Architecture-altering operations for evolving the architecture of a multi-part program in genetic programming. Technical Report STAN-CS-94-1528, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1994. ?? pp. URL http://www-

db.stanford.edu/TR/CS-TR-94-1528.html.

Shahar:1994:KBM

[1546] Yuval Shahar. A knowledge-based method for temporal abstraction of clinical data. Technical Report STAN-CS-94-1529, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1994.

330 pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1529.html.

Koga:1994:CMA

[1547] Yoshihito Koga. On computing multi-arm manipulation trajectories. Technical Report STAN-CS-94-1530, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1994. 143 pp. URL

http://www-db.stanford.edu/TR/CS-TR-94-1530.html.

Li:1994:LMP

[1548] Tsai-Yen Li and Jean-Claude Latombe. On-line manipulation planning for two robot arms in a dynamic environment. Technical Report STAN-CS-94-1531, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1994. 34 pp. URL http://www-db.

stanford.edu/TR/CS-TR-94-1531.html.

Banon:1994:PCF

[1549] Jose Banon. Planning the collision-free paths of an actively flexible manipulator. Technical Report STAN-CS-94-1532, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1994. 29

pp. URL http://www-db.stanford.edu/TR/CS-TR-94-1532.html.

Raghavan:1994:RQP

[1550] L. Raghavan, J-C. Kavraki, R. Latombe, and P. Motwani. Randomized query processing in robot motion planning. Technical Report STAN-CS-94-1533, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1994. 18 pp. URL http://www-db.stanford.

edu/TR/CS-TR-94-1533.html.

Lazanas:1994:RAU

[1551] Anthony Lazanas. Reasoning about uncertainty in robot motion planning. Technical Report STAN-CS-95-1539, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1994. 255 pp.

URL http://www-db.stanford.edu/TR/CS-TR-95-1539.html.

Wong-Toi:1994:SAV

[1552] Howard Wong-Toi. Symbolic approximations for verifying real-time systems. Technical Report STAN-CS-95-1546, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1994. 207 pp.

URL http://www-db.stanford.edu/TR/CS-TR-95-1546.html.

Gupta:1995:PIB

[1553] Ashish Gupta. Partial information based integrity constraint checking. Technical Report STAN-CS-95-1534, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1995. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-95-1534.html.

Kavraki:1995:RNC

[1554] Lydia E. Kavraki. Random networks in configuration space for fast path planning. Technical Report STAN-CS-95-1535, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1995. 150

pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1535.html.

Yim:1995:LUM

[1555] Mark Yim. Locomotion with a unit-modular reconfigurable robot. Technical Report STAN-CS-95-1536, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1995. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-95-1536.html.

Quinlan:1995:RTM

[1556] Sean Quinlan. Real-time modification of collision-free paths. Technical Report STAN-CS-95-1537, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1995. 144 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-95-1537.html.

Hammer:1995:PSS

[1557] Joachim Hammer. 1994 publications summary of the Stanford Database Group. Technical Report STAN-CS-95-1538, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1995. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1538.html.

Murdock:1995:MMI

[1558] Janet L. Murdock. Model-matching and individuation for model-based diagnosis. Technical Report STAN-CS-95-1540, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1995. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1540.html.

Karger:1995:RSG

[1559] David R. Karger. Random sampling in graph optimization problems. Technical Report STAN-CS-95-1541, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1995. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-95-1541.html.

Koza:1995:PGP

[1560] John R. Koza and David Andre. Parallel genetic programming on a network of transputers. Technical Report STAN-CS-95-1542, Stanford University, Department of Computer Science, Stanford, CA, USA, January

1995. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1542.
html.

Tomasi:1995:SS

[1561] Carlo Tomasi and Roberto Manduchi. Stereo without search. Technical Report STAN-CS-95-1543, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1995. 27 pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-95-1543.html.

Basch:1995:DVB

[1562] Julien Basch, Sanjeev Khanna, and Rajeev Motwani. On diameter verification and Boolean matrix multiplication. Technical Report STAN-CS-95-1544, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1995. 5 pp. URL http://www-db.stanford.

edu/TR/CS-TR-95-1544.html.

Khanna:1995:AAL

[1563] Sanjeev Khanna, Rajeev Motwani, and Frances F. Yao. Approximation algorithms for the largest common subtree problem. Technical Report STAN-CS-95-1545, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1995. 8 pp. URL http://www-db.

stanford.edu/TR/CS-TR-95-1545.html.

Roy:1995:SRP

[1564] H. Scott Roy. Sharp, reliable predictions using supervised mixture models. Technical Report STAN-CS-95-1547, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1995. ?? pp.

URL http://www-db.stanford.edu/TR/CS-TR-95-1547.html.

Gawlick:1995:RAC

[1565] Rainer Gawlick, Anil Kamath, Serge Plotkin, and K. G. Ramakrishnan. Routing and admission control in general topology networks. Technical Report STAN-CS-95-1548, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1995. 19 pp. URL http://www-

db.stanford.edu/TR/CS-TR-95-1548.html.

Rutledge:1995:DSM

[1566] Geoffrey William Rutledge. Dynamic selection of models. Technical Report STAN-CS-95-1549, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1995. 188 pp. URL http://www-

db.stanford.edu/TR/CS-TR-95-1549.html.

Drakopoulos:1995:TDH

[1567] John A. Drakopoulos. Theory and design of a hybrid pattern recognition system. Technical Report STAN-CS-95-1550, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1995. 109

pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1550.html.

McGuire:1995:TMC

[1568] Hugh W. McGuire. Two methods for checking formulas of temporal logic. Technical Report STAN-CS-95-1551, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1995. 145 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-95-1551.html.

Brafman:1995:ETR

[1569] Ronen I. Brafman and Moshe Tennenholtz. Embedded teaching of reinforcement learners. Technical Report STAN-CS-95-1552, Stanford University, Department of Computer Science, Stanford, CA, USA, June

1995. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1552.
html.

Srinivas:1995:MTA

[1570] Sampath Srinivas. Modeling techniques and algorithms for probabilistic model-based diagnosis and repair. Technical Report STAN-CS-95-1553, Stanford University, Department of Computer Science, Stanford, CA,

USA, July 1995. ?? pp. URL http://www-db.stanford.edu/TR/CS-
TR-95-1553.html.

Lasher:1995:CST

[1571] Rebecca Lasher, Vicky Reich, and Greg Anderson. The Computer Science Technical Report (CS-TR) Project: Considerations from the library perspective. Technical Report STAN-CS-95-1554, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1995. 27 pp.

URL http://www-db.stanford.edu/TR/CS-TR-95-1554.html.

Kiriha:1995:RTD

[1572] Yoshiaki Kiriha. Real-time database experiences in network management application. Technical Report STAN-CS-95-1555, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1995.

?? pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1555.html.

Kennedy:1995:SUW

[1573] J. Robert Kennedy, Jr. Solving unweighted and weighted bipartite matching problems in theory and practice. Technical Report STAN-CS-95-1556, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1995. ?? pp. URL http://www-db.stanford.

edu/TR/CS-TR-95-1556.html.

Wolf:1995:HMS

[1574] Elizabeth Susan Wolf. Hierarchical models of synchronous circuits for formal verification and substitution. Technical Report STAN-CS-95-1557, Stanford University, Department of Computer Science, Stanford, CA,

USA, October 1995. 210 pp. URL http://www-db.stanford.edu/TR/
CS-TR-95-1557.html.

Greenwald:1995:DAF

[1575] Michael B. Greenwald, Sandeep K. Singhal, Jonathan R. Stone, and David R. Cheriton. Designing an academic firewall: Policy, practice and experience with SURF. Technical Report STAN-CS-95-1558, Stanford University, Department of Computer Science, Stanford, CA, USA,

December 1995. ?? pp. URL http://www-db.stanford.edu/TR/CS-
TR-95-1558.html.

Kavraki:1995:NEP

[1576] Lydia E. Kavraki. On the number of equilibrium placements of mass distributions in elliptic potential fields. Technical Report STAN-CS-95-1559, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1995. ?? pp. URL http://www-db.stanford.edu/

TR/CS-TR-95-1559.html.

Kohavi:1995:WPE

[1577] Ron Kohavi. Wrappers for performance enhancements and oblivious decision graphs. Technical Report STAN-CS-95-1560, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1995.

?? pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1560.html.

Hu:1995:TEF

[1578] Alan John Hu. Techniques for efficient formal verification using binary decision diagrams. Technical Report STAN-CS-95-1561, Stanford University, Department of Computer Science, Stanford, CA, USA, December

1995. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-95-1561.
html.

Bjorner:1995:SST

[1579] Nikolaj Bjorner, Anca Browne, Eddie Chang, Michael Colon, Arjun Kapur, Zohar Manna, Henny B. Sipma, and Tomas E. Uribe. STeP: The Stanford Temporal Prover (educational release) user’s manual. Technical Report STAN-CS-95-1562, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1995. 46 pp. URL http:

//www-db.stanford.edu/TR/CS-TR-95-1562.html.

Silberschatz:1996:DRA

[1580] Avi Silberschatz, Michael Stonebraker, and Jeffrey D. Ullman. Database research: Achievements and opportunities into the 21st century. Technical Report STAN-CS-96-1563, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1996. ?? pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-96-1563.html.

Ohno-Machado:1996:MAN

[1581] Lucila Ohno-Machado. Medical applications of neural networks: Connectionist models of survival. Technical Report STAN-CS-96-1564, Stanford University, Department of Computer Science, Stanford, CA, USA, March

1996. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-96-1564.
html.

Sujansky:1996:FMB

[1582] Walter Sujansky. A formal model for bridging heterogeneous relational databases in clinical medicine. Technical Report STAN-CS-96-1565, Stanford University, Department of Computer Science, Stanford, CA,

USA, April 1996. ?? pp. URL http://www-db.stanford.edu/TR/CS-
TR-96-1565.html.

Manna:1996:CTS

[1583] Zohar Manna and Amir Pnueli. Clocked transition systems. Technical Report STAN-CS-96-1566, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1996. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-96-1566.html.

Anuchitanukul:1996:SRP

[1584] Anuchit Anuchitanukul. Synthesis of reactive programs. Technical Report STAN-CS-96-1567, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1996. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-96-1567.html.

Cazals:1996:ACI

[1585] Frederic Cazals and G. D. S. Ramkumar. Algorithms for computing intersection and union of toleranced polygons with applications. Technical Report STAN-CS-96-1568, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1996. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-96-1568.html.

Unruh:1996:UAA

[1586] Amy Unruh. Using automatic abstraction for problem-solving and learning. Technical Report STAN-CS-96-1569, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1996. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-96-1569.html.

Hasan:1996:OSQ

[1587] Waqar Hasan. Optimization of SQL queries for parallel machines. Technical Report STAN-CS-96-1570, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1996. ?? pp. URL http://www-

db.stanford.edu/TR/CS-TR-96-1570.html.

DeAlfaro:1996:FVP

[1588] Luca DeAlfaro. Formal verification of performance and reliability of realtime systems. Technical Report STAN-CS-96-1571, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1996. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-96-1571.html.

Geddis:1996:CNH

[1589] Donald F. Geddis. Caching and non-Horn inference in model elimination theorem provers. Technical Report STAN-CS-96-1572, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1996.

?? pp. URL http://www-db.stanford.edu/TR/CS-TR-96-1572.html.

Birchfield:1996:DDP

[1590] Stan Birchfield and Carlo Tomasi. Depth discontinuities by pixel-to-pixel stereo. Technical Report STAN-CS-96-1573, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1996. ?? pp.

URL http://www-db.stanford.edu/TR/CS-TR-96-1573.html.

Singhal:1996:ERM

[1591] Sandeep K. Singhal. Effective remote modeling in large-scale distributed simulation and visualization environments. Technical Report STAN-CS-96-1574, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1996. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-96-1574.html.

Kamath:1996:RAC

[1592] Anil Kamath, Omri Palmon, and Serge Plotkin. Routing and admission control in general topology networks with Poisson arrivals. Technical Report STAN-CS-96-1575, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1996. ?? pp. URL http://www-

db.stanford.edu/TR/CS-TR-96-1575.html.

Huyn:1996:QRU

[1593] Nam Huyn. Query reformulation under incomplete mappings. Technical Report STAN-CS-96-1576, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1996. ?? pp. URL http:

//www-db.stanford.edu/TR/CS-TR-96-1576.html.

Huyn:1996:MAU

[1594] Nam Huyn. A more aggressive use of views to extract information. Technical Report STAN-CS-96-1577, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1996. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-96-1577.html.

Ip:1996:SRM

[1595] C. Norris Ip. State reduction methods for automatic formal verification. Technical Report STAN-CS-96-1578, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1996. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-96-1578.html.

Tajnai:1997:VHD

[1596] Carolyn Tajnai. From the Valley of Heart’s Delight to Silicon Valley: a study of Stanford University’s role in the transformation. Technical Report STAN-CS-97-1579, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1997. ?? pp. URL http://www-

db.stanford.edu/TR/CS-TR-97-1579.html.

Gravano:1997:SSP

[1597] Luis Gravano, Kevin Chang, Hector Garcia-Molina, and Andreas Paepcke. STARTS: Stanford Protocol Proposal for Internet Retrieval and Search. Technical Report STAN-CS-97-1580, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1997. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-97-1580.html.

Paepcke:1997:TID

[1598] Andreas Paepcke, Steve B. Cousins, Hector Garcia-Molina, Scott W. Hassan, Steven K. Ketchpel, Martin Roscheisen, and Terry Winograd. Towards interoperability in digital libraries: Overview and selected highlights of the Stanford Digital Library Project. Technical Report STAN-CS-97-1581, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1997. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-97-1581.html.

Roscheisen:1997:SWA

[1599] Martin Roscheisen, Christian Mogensen, and Terry Winograd. Shared web annotations as a platform for third-party value-added, information providers: Architecture, protocols, and usage examples. Technical Report STAN-CS-97-1582, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1997. ?? pp. URL http://www-

db.stanford.edu/TR/CS-TR-97-1582.html.

Chang:1997:BQM

[1600] Kevin Chen-Chuan Chang, Hector Garcia-Molina, and Andreas Paepcke. Boolean query mapping across heterogeneous information sources (extended version). Technical Report STAN-CS-97-1583, Stanford University, Department of Computer Science, Stanford, CA, USA, January

1997. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-97-1583.
html.

Kamiya:1997:GSP

[1601] Kenichi Kamiya, Martin Roscheisen, and Terry Winograd. Grassroots: a system providing a uniform framework for communicating, structuring, sharing information, and organizing people. Technical Report STAN-CS-97-1584, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1997. ?? pp. URL http://www-db.stanford.

edu/TR/CS-TR-97-1584.html.

Baldonado:1997:TTM

[1602] Michelle Q. Wang Baldonado and Terry Winograd. Techniques and tools for making sense out of heterogeneous search service results. Technical Report STAN-CS-97-1585, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1997. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-97-1585.html.

Pichumani:1997:CTD

[1603] Ramani Pichumani. Construction of a three-dimensional geometric model for segmentation and visualization of cervical spine images. Technical Report STAN-CS-97-1586, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1997. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-97-1586.html.

Matan:1997:ESC

[1604] Ofer Matan. Ensembles for supervised classification learning. Technical Report STAN-CS-97-1587, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1997. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-97-1587.html.

Cohen:1997:SBE

[1605] Scott Cohen and Carlo Tomasi. Systems of bilinear equations. Technical Report STAN-CS-97-1588, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1997. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-97-1588.html.

Benson:1997:LAM

[1606] Scott Sherwood Benson. Learning action models for reactive autonomous agents. Technical Report STAN-CS-97-1589, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1997. ?? pp.

URL http://www-db.stanford.edu/TR/CS-TR-97-1589.html.

Goldwasser:1997:CMA

[1607] Michael Goldwasser. Complexity measures for assembly sequences. Technical Report STAN-CS-97-1590, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1997. ?? pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-97-1590.html.

Friedman:1997:MBD

[1608] Nir Friedman. Modeling beliefs in dynamic systems. Technical Report STAN-CS-97-1591, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1997.

Goel:1997:OTC

[1609] Ashish Goel, Monika R. Henzinger, and Serge Plotkin. Online throughput-competitive algorithm for multicast routing and admission control. Technical Report STAN-CS-97-1592, Stanford University, Department of Computer Science, Stanford, CA, USA, July 1997. ?? pp.

URL http://www-db.stanford.edu/TR/CS-TR-97-1592.html.

Nourbakhsh:1997:IPEa

[1610] Illah Reza Nourbakhsh. Interleaving planning and execution for autonomous robots. Technical Report STAN-CS-97-1593, Stanford University, Department of Computer Science, Stanford, CA, USA, 1997. ISBN 0-7923-9828-9. 168 pp. Published as [1743].

Kapur:1997:IPB

[1611] Arjun Kapur. Interval and point-based approaches to hybrid system verification. Technical Report STAN-CS-97-1594, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1997.

?? pp. URL http://www-db.stanford.edu/TR/CS-TR-97-1594.html.

Huyn:1997:MDW

[1612] Nam Huyn. Maintaining data warehouses under limited source access. Technical Report STAN-CS-97-1595, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1997. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-97-1595.html.

Campbell:1997:DDL

[1613] Keith Eugene Campbell. Distributed development of a logic-based controlled medical terminology. Technical Report STAN-CS-97-1596, Stanford University, Department of Computer Science, Stanford, CA, USA,

October 1997. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-
97-1596.html.

Cohen:1997:EMD

[1614] Scott Cohen and Leonidas Guibas. The Earth Mover’s Distance: Lower bounds and invariance under translation. Technical Report STAN-CS-97-1597, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1997. 44 pp. URL http://www-db.stanford.edu/

TR/CS-TR-97-1597.html.

Duschka:1997:QPO

[1615] Oliver M. Duschka. Query planning and optimization in information integration. Technical Report STAN-CS-97-1598, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1997.

?? pp. URL http://www-db.stanford.edu/TR/CS-TR-97-1598.html.

Sim:1997:TBI

[1616] Ida Sim. Trial banks: an informatics foundation for evidence-based medicine. Technical Report STAN-CS-97-1599, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1997. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-97-1599.html.

Goldberg:1997:ICA

[1617] Andrew Goldberg, Jeffrey D. Oldham, Serge Plotkin, and Cliff Stein. An implementation of a combinatorial approximation algorithm for minimum-cost multicommodity flow. Technical Report STAN-CS-97-1600, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1997. ?? pp. URL http://www-db.stanford.edu/

TR/CS-TR-97-1600.html.

deAlfaro:1998:FVP

[1618] Luca de Alfaro. Formal verification of probabilistic systems. Technical Report STAN-CS-98-1601, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1998. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-98-1601.html.

Fisher:1998:TSO

[1619] Kathleen Fisher. Type systems for object-oriented programming languages. Technical Report STAN-CS-98-1602, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1998. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-98-1602.html.

Herrod:1998:UCM

[1620] Stephen Alan Herrod. Using complete machine simulation to understand computer system behavior. Technical Report STAN-CS-98-1603, Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1998. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-
98-1603.html.

Teo:1998:TAS

[1621] Patrick C. Teo. Theory and applications of steerable functions. Technical Report STAN-CS-98-1604, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1998. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-98-1604.html.

Balabanovic:1998:LSM

[1622] Marko Balabanovic. Learning to surf: Multiagent systems for adaptive web page recommendation. Technical Report STAN-CS-98-1605, Stanford University, Department of Computer Science, Stanford, CA, USA,

March 1998. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-
98-1605.html.

Basu:1998:ACC

[1623] Julie Basu. Associative caching in client–server databases. Technical Report STAN-CS-98-1606, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1998. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-98-1606.html.

Roscheisen:1998:NCD

[1624] Martin Roscheisen. A network-centric design for relationship-based rights management. Technical Report STAN-CS-98-1607, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1998. ?? pp.

URL http://www-db.stanford.edu/TR/CS-TR-98-1607.html.

Katz:1998:NPP

[1625] Morris J. Katz. A new perspective on partial evaluation and use analysis. Technical Report STAN-CS-98-1608, Stanford University, Department of Computer Science, Stanford, CA, USA, June 1998. ?? pp. URL http:/

/www-db.stanford.edu/TR/CS-TR-98-1608.html.

Sanders:1998:ACC

[1626] Gillian D. Sanders. Automated creation of clinical-practice guidelines from decision models. Technical Report STAN-CS-98-1609, Stanford University, Department of Computer Science, Stanford, CA, USA, July

1998. xxiv + 244 pp. URL http://www-db.stanford.edu/TR/CS-TR-
98-1609.html.

Veach:1998:RMC

[1627] Eric Veach. Robust Monte Carlo methods for light transport simulation. Technical Report STAN-CS-98-1610, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1998. xxv + 406 pp.

URL https://searchworks.stanford.edu/view/10386091.

Chekuri:1998:AAS

[1628] Chandra Chekuri. Approximation algorithms for scheduling problems. Technical Report STAN-CS-98-1611, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1998. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-98-1611.html.

Iliano:1998:PPC

[1629] Iliano Cervesato and John C. Mitchell. Pleiades Project: Collected work 1997–1998. Technical Report STAN-CS-98-1612, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1998. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-98-1612.html.

Prabhakar:1998:SPP

[1630] Balaji Prabhakar, Nicholas Bambos, and Tom Mountford. On the synchronization of Poisson processes and queueing networks with service and synchronization nodes. Technical Report STAN-CS-98-1613, Stanford University, Department of Computer Science, Stanford, CA, USA,

October 1998. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-
98-1613.html.

deAlfaro:1998:DTC

[1631] Luca de Alfaro, Zohar Manna, and Henny Sipma. Decomposing, transforming and composing diagrams: The joys of modular verification. Technical Report STAN-CS-98-1614, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1998. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-98-1614.html.

Sahami:1998:UML

[1632] Mehran Sahami. Using machine learning to improve information access. Technical Report STAN-CS-98-1615, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1998. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-98-1615.html.

Bjorner:1998:DVR

[1633] Nikolaj S. Bjørner, Zohar Manna, Henny B. Sipma, and Tomás E. Uribe. Deductive verification of real-time systems using STeP. Technical Report STAN-CS-98-1616, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1998. Preliminary version appeared in 4th Intl. AMAST Workshop on Real-Time Systems, LNCS, Vol. 1231, Springer-Verlag, May 1997, pp. 484–498. Published in Published in Theoretical Computer Science 253(1) 27–60, 17 February 2001.

Shiffman:1999:SMI

[1634] Smadar Shiffman. Segmentation of medical image volumes using intrinsic shape information. Technical Report STAN-CS-99-1617, Stanford University, Department of Computer Science, Stanford, CA, USA, February

1999. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-99-1617.
html.

Uribe:1999:ABD

[1635] Tomas E. Uribe. Abstraction-based deductive-algorithmic verification of reactive systems. Technical Report STAN-CS-99-1618, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1999.

?? pp. URL http://www-db.stanford.edu/TR/CS-TR-99-1618.html.

Huang:1999:IAA

[1636] Cecil Huang. Intelligent alarms: Allocating attention among concurrent processes. Technical Report STAN-CS-99-1619, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1999. ?? pp.

URL http://www-db.stanford.edu/TR/CS-TR-99-1619.html.

Cohen:1999:FCS

[1637] Scott Cohen. Finding color and shape patterns in images. Technical Report STAN-CS-99-1620, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1999. ?? pp. URL http://www-db.

stanford.edu/TR/CS-TR-99-1620.html.

Rubner:1999:PMI

[1638] Yossi Rubner. Perceptual metrics for image database navigation. Technical Report STAN-CS-99-1621, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1999. ?? pp. URL http://

www-db.stanford.edu/TR/CS-TR-99-1621.html.

Oldham:1999:MGF

[1639] Jeffrey David Oldham. Multicommodity and generalized flow algorithms: Theory and practice. Technical Report STAN-CS-99-1622, Stanford University, Department of Computer Science, Stanford, CA, USA, August

1999. ?? pp. URL http://www-db.stanford.edu/TR/CS-TR-99-1622.
html.

Labio:1999:EMR

[1640] Wilburt Juan Labio. Efficient maintenance and recovery of data warehouses. Technical Report STAN-CS-99-1623, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1999. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-99-1623.html.

Greenwald:1999:NBS

[1641] Michael Greenwald. Non-blocking synchronization and system design. Technical Report STAN-CS-99-1624, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1999. ?? pp. URL

http://www-db.stanford.edu/TR/CS-TR-99-1624.html.

Iliano:1999:PPC

[1642] Iliano Cervesato and John C. Mitchell. Pleiades Project: Collected work 1998–1999. Technical Report STAN-CS-99-1625, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1999. ??

pp. URL http://www-db.stanford.edu/TR/CS-TR-99-1625.html.

Guimbretiere:2000:FCC

[1643] François Guimbretière and Terry Winograd. FlowMenu: Combining command, text, and data entry. Technical Report CSTR 2000-01, Stanford University, Department of Computer Science, Stanford, CA, USA,

???? 2000. 4 pp. URL http://graphics.stanford.edu/papers/
FlowMenu/flowmenu.pdf.

Johanson:2000:EHE

[1644] Brad Johanson, Armando Fox, Pat Hanrahan, and Terry Winograd. The Event Heap: an enabling infrastructure for interactive workspaces. Technical Report CSTR 2000-02, Stanford University, Department of Computer Science, Stanford, CA, USA, May 2000. ?? pp. URL https://

graphics.stanford.edu/papers/eheap/.

Johanson:2000:PSP

[1645] Brad Johanson, Greg Hutchins, and Terry Winograd. PointRight: a system for pointer/keyboard redirection among, multiple displays and machines. Technical Report CSTR 2000-03, Stanford University, Department of Computer Science, Stanford, CA, USA, May 2000. ?? pp.

URL http://graphics.stanford.edu/papers/pointright/.

Ringel:2000:BIF

[1646] Meredith Ringel, Henry Berg, Yuhui Jin, and Terry Winograd. Barehands: Implement-free interaction with a wall-mounted display. Technical Report CSTR 2000-04, Stanford University, Department of Computer Science, Stanford, CA, USA, May 2000. 7 pp. URL https://hci.

stanford.edu/cstr/reports/bhlong.pdf.

Guimbretiere:2000:GWE

[1647] François Guimbretière, Terry Winograd, and Sha Xin Wei. The Geometer’s Workbench: an experiment in interacting with a large, high resolution display. Technical Report CSTR 2000-05, Stanford University, Department of Computer Science, Stanford, CA, USA, June 2000.

7 pp. URL http://graphics.stanford.edu/%7Efrancois/Papers/
UIST2000/geometerworkbench.pdf.

Chen:2000:LMU

[1648] Xing Chen and James Davis. LumiPoint: Multi-user laser-based interaction on large tiled displays. Technical Report CSTR 2000-06, Stanford University, Department of Computer Science, Stanford, CA, USA,

June 2000. ?? pp. URL http://graphics.stanford.edu/papers/
multiuser.

Chen:2000:CPC

[1649] Xing Chen and James Davis. Camera placement considering occlusion for robust motion capture. Technical Report CSTR 2000-07, Stanford University, Department of Computer Science, Stanford, CA, USA, De-

cember 2000. ?? pp. URL http://graphics.stanford.edu/papers/
OcclusionMetric/.

Davis:2000:MSM

[1650] James Davis and Xing Chen. Mixed scale motion recovery using guidable cameras. Technical Report CSTR 2000-08, Stanford University, Department of Computer Science, Stanford, CA, USA, December 2000. ?? pp.

URL http://graphics.stanford.edu/papers/GuidableCameras/.

Johanson:2001:SSI

[1651] Brad Johanson, Shankar Ponnekanti, Emre Kiciman, Caesar Sengupta, and Armando Fox. System support for interactive workspaces. Technical Report CSTR 2001-01, Stanford University, Department of Computer Science, Stanford, CA, USA, March 2001. ?? pp. URL http://

graphics.stanford.edu/papers/iwork-sosp18/.

Marschner:2001:FHC

[1652] Stephen R. Marschner, James Davis, Matt Garr, and Marc Levoy. Filling holes in complex surfaces using volumetric diffusion. Technical Report CSTR 2001-07, Stanford University, Department of Computer Science, Stanford, CA, USA, January 2001. ?? pp. URL https://graphics.

stanford.edu/papers/holefill-tr-2001-07/.

Davis:2001:FHC

[1653] James Davis, Stephen R. Marschner, Matt Garr, and Marc Levoy. Filling holes in complex surfaces using volumetric diffusion. Technical Report CSTR 2001-08, Stanford University, Department of Computer Science, Stanford, CA, USA, December 2001. ?? pp. URL https://graphics.

stanford.edu/papers/holefill-tr-2001-08/.

Wei:2002:OIT

[1654] Li-Yi Wei and Marc Levoy. Order-independent texture synthesis. Technical Report CSTR 2002-01, Stanford University, Department of Computer Science, Stanford, CA, USA, January 2002. ?? pp. URL http://

graphics.stanford.edu/papers/texture-synthesis-tr-2002-01/.

Chuang:2002:PDF

[1655] Erica Chuang and Chris Bregler. Performance driven facial animation using blendshape interpolation. Technical Report CSTR 2002-02, Stanford University, Department of Computer Science, Stanford, CA, USA, April

2002. ?? pp. URL http://graphics.stanford.edu/%7Eechuang/face.

Ionescu:2002:WCR

[1656] Arna Ionescu, Maureen Stone, and Terry Winograd. WorkspaceNavigator: Capture, recall and reuse using spatial cues in an interactive workspace. Technical Report CSTR 2002-04, Stanford University, Department of Computer Science, Stanford, CA, USA, May 2002. 16 pp.

URL http://hci.stanford.edu/research/wkspcNavTR.pdf.

Ikemoto:2003:HMA

[1657] Leslie Ikemoto. A hierarchical method for aligning warped meshes. Technical Report CSTR 2003-01, Stanford University, Department of Computer Science, Stanford, CA, USA, June 2003. 23 pp. URL https://hci.

stanford.edu/cstr/reports/2003-01.pdf. Master’s with Distinction

in Research Report.

Chang:2003:HE

[1658] Erica Chang and Chris Bregler. Head emotion. Technical Report CSTR 2003-02, Stanford University, Department of Computer Science, Stanford, CA, USA, April 7, 2003. ?? pp. URL http://graphics.stanford.

edu/%7Eechuang/heademotion/index.html.

Buck:2003:DPC

[1659] Ian Buck and Pat Hanrahan. Data parallel computation on graphics hardware. Technical Report CSTR 2003-03, Stanford University, Department of Computer Science, Stanford, CA, USA, January 22, 2003.

?? pp. URL https://hci.stanford.edu/cstr/abstracts/2003-03.
html.

Abstract:2003:BSV

[1660] Ian Buck. Brook specification v0.2. Technical Report CSTR 2003-04, Stanford University, Department of Computer Science, Stanford, CA, USA, October 31, 2003. ?? pp. URL https://hci.stanford.edu/

cstr/abstracts/2003-04.html.

Grant:2003:BSB

[1661] Karen D. Grant, Adrian Graham, Tom Nguyen, Andreas Paepcke, and Terry Winograd. Beyond the shoe box: Foundations for flexibly organizing photographs on a computer. Technical Report CSTR 2003-05, Stanford University, Department of Computer Science, Stanford, CA, USA, January 15, 2003. ?? pp. URL https://hci.stanford.edu/

cstr/reports/2003-05.pdf.

Song:2003:ISF

[1662] Yee Jiun Song, Wendy Tobagus, Der Yao Leong, Brad Johanson, and Armando Fox. iSecurity: a security framework for interactive workspaces. Technical Report CSTR 2004-03, Stanford University, Department of Computer Science, Stanford, CA, USA, September 3, 2003. ?? pp. URL

https://hci.stanford.edu/cstr/abstracts/2004-03.html.

Wilburn:2004:SSI

[1663] Bennett Wilburn, Neel Joshi, Katherine Chou, Marc Levoy, and Mark Horowitz. Spatiotemporal sampling and interpolation for dense camera arrays. Technical Report CSTR 2004-01, Stanford University, Department of Computer Science, Stanford, CA, USA, January 21, 2004. ?? pp.

URL https://hci.stanford.edu/cstr/abstracts/2004-01.html.

and:2004:CCA

[1664] Neel Joshi. Color calibration for arrays of inexpensive image sensors. Technical Report CSTR 2004-02, Stanford University, Department of Computer Science, Stanford, CA, USA, March 31, 2004. iv + 26 pp.

URL https://hci.stanford.edu/cstr/reports/2004-02.pdf. Mas-

ter’s with Distinction in Research Report.

Yeh:2004:FNF

[1665] Ron B. Yeh and Scott Klemmer. Field notes on field notes: Informing technology support for biologists. Technical Report CSTR 2004-04, Stanford University, Department of Computer Science, Stanford, CA, USA, December 13, 2004. ?? pp. URL http://hci.stanford.edu/

publications/techreports/CHI2005-Biology.pdf.

Johanson:2004:SIW

[1666] Brad Johanson, Armando Fox, and Terry Winograd. The Stanford Interactive Workspaces Project. Technical Report CSTR 2004-05, Stanford University, Department of Computer Science, Stanford, CA, USA, Au-

gust 9, 2004. 30 pp. URL https://hci.stanford.edu/cstr/reports/
2004-05.pdf.

and:2005:SBX

[1667] Jan Chong. Social behaviors on XP teams and non-XP teams: a comparative study. Technical Report CSTR 2005-01, Stanford University, Department of Computer Science, Stanford, CA, USA, March 15, 2005.

10 pp. URL https://hci.stanford.edu/cstr/reports/2005-01.pdf.

Ng:2005:LFP

[1668] Ren Ng, Marc Levoy, Mathieu Brédif, Gene Duval, Mark Horowitz, and Pat Hanrahan. Light field photography with a hand-held plenoptic camera. Technical Report CSTR 2005-02, Stanford University, Department of Computer Science, Stanford, CA, USA, April 20, 2005. 11 pp. URL

http://graphics.stanford.edu/papers/lfcamera/.

Pauly:2005:EBS

[1669] Mark Pauly, Niloy J. Mitra, Joachim Giesen, Leonidas Guibas, and Markus Gross. Example-based 3D scan completion. Technical Report CSTR 2005-03, Stanford University, Department of Computer Science, Stanford, CA, USA, April 29, 2005. ?? pp. URL https://hci.

stanford.edu/cstr/reports/2005-03.pdf.

Cadar:2005:EGT

[1670] Cristian Cadar and Dawson Engler. Execution generated test cases: How to make systems code crash itself. Technical Report CSTR 2005-04, Stanford University, Department of Computer Science, Stanford, CA,

USA, March 25, 2005. 14 pp. URL https://hci.stanford.edu/cstr/
reports/2005-04.pdf.

Houston:2005:HFB

[1671] Mike Houston, Arcot Preetham, and Mark Segal. A hardware F-buffer implementation. Technical Report CSTR 2005-05, Stanford University, Department of Computer Science, Stanford, CA, USA, April 11, 2005. 6

pp. URL https://hci.stanford.edu/cstr/reports/2005-05.pdf.

Hartmann:2005:DTI

[1672] Björn Hartmann, Scott R. Klemmer, and Michael Bernstein. d.tools: Integrated prototyping for physical interaction design. Technical Report CSTR 2005-06, Stanford University, Department of Computer Science, Stanford, CA, USA, September 22, 2005. 10 pp. URL https://hci.

stanford.edu/cstr/reports/2005-06.pdf.

Piper:2005:SCT

[1673] Anne Marie Piper, Eileen O’Brien, Meredith Ringel Morris, and Terry Winograd. SIDES: a cooperative tabletop computer game for social skills development. Technical Report CSTR 2005-07, Stanford University, Department of Computer Science, Stanford, CA, USA, September 22, 2005.

?? pp. URL https://hci.stanford.edu/cstr/reports/2005-07.pdf.

RingelMorris:2005:SCL

[1674] Meredith Ringel Morris, Anne Marie Piper, Anthony Cassanego, and Terry Winograd. Supporting cooperative language learning: Issues in interface design for an interactive table. Technical Report CSTR 2005-08, Stanford University, Department of Computer Science, Stanford, CA, USA, September 22, 2005. 9 pp. URL https://hci.stanford.edu/

cstr/reports/2005-08.pdf.

Wang:2005:MUP

[1675] QianYing Wang, Tony Hsieh, Meredith Ringel Morris, and Andreas Paepcke. Multi-user piles across space. Technical Report CSTR 2005-09, Stanford University, Department of Computer Science, Stanford, CA, USA, September 22, 2005. 4 pp. URL https://hci.stanford.edu/

cstr/reports/2005-09.pdf.

Cadar:2006:ESA

[1676] Cristian Cadar, Paul Twohey, Vijay Ganesh, and Dawson Engler. EXE: a system for automatically generating inputs of death using symbolic execution. Technical Report CSTR 2006-01, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1, 2006.

20 pp. URL https://hci.stanford.edu/cstr/reports/2006-01.pdf.

Hartmann:2006:RPP

[1677] Björn Hartmann, Scott R. Klemmer, Michael Bernstein, Leith Abdulla, Brandon Burr, Avi Robinson-Mosher, and Jennifer Gee. Reflective physical prototyping through integrated design, test, and analysis. Technical Report CSTR 2006-02, Stanford University, Department of Computer Science, Stanford, CA, USA, April 7, 2006. 10 pp. URL https://hci.

stanford.edu/cstr/reports/2006-02.pdf.

Yeh:2006:IGP

[1678] Ron B. Yeh, Joel Brandt, Jonas Boli, and Scott R. Klemmer. Interactive gigapixel prints: Large, paper-based interfaces for visual context and collaboration. Technical Report CSTR 2006-03, Stanford University, Department of Computer Science, Stanford, CA, USA, April 7, 2006. 10

pp. URL https://hci.stanford.edu/cstr/reports/2006-03.pdf.

Morris:2006:HRI

[1679] Dan Morris and Neel Joshi. Hybrid rendering for interactive virtual scenes. Technical Report CSTR 2006-04, Stanford University, Department of Computer Science, Stanford, CA, USA, May 23, 2006. 7 pp.

URL https://hci.stanford.edu/cstr/reports/2006-04.pdf.

RingelMorris:2006:SEI

[1680] Meredith Ringel Morris. Supporting effective interaction with tabletop groupware. Technical Report CSTR 2006-05, Stanford University, Department of Computer Science, Stanford, CA, USA, April 7, 2006. xviii

+ 233 pp. URL https://hci.stanford.edu/cstr/reports/2006-05.
pdf.

and:2006:ADS

[1681] Dan Morris. Algorithms and data structures for haptic rendering: Curve constraints, distance maps, and data logging. Technical Report CSTR 2006-06, Stanford University, Department of Computer Science, Stanford, CA, USA, June 11, 2006. 12 pp. URL https://hci.stanford.

edu/cstr/reports/2006-06.pdf.

and:2006:APC

[1682] Dan Morris. Automatic preparation, calibration, and simulation of deformable objects. Technical Report CSTR 2006-07, Stanford University, Department of Computer Science, Stanford, CA, USA, July 24, 2006. 15

pp. URL https://hci.stanford.edu/cstr/reports/2006-07.pdf.

and:2006:RCE

[1683] Manu Kumar. Reducing the cost of eye tracking systems. Technical Report CSTR 2006-08, Stanford University, Department of Computer Science, Stanford, CA, USA, April 4, 2006. 4 pp. URL https://hci.

stanford.edu/cstr/reports/2006-08.pdf.

Lee:2006:EAI

[1684] Brian Lee, Heidy Maldonado, and Scott R. Klemmer. Evaluating augmented idea logs for design education. Technical Report CSTR 2006-09, Stanford University, Department of Computer Science, Stanford, CA,

USA, May 28, 2006. 10 pp. URL https://hci.stanford.edu/cstr/
reports/2006-09.pdf.

Lee:2006:AID

[1685] Brian Lee, Scott R. Klemmer, and Ronen Brafman. Adaptive interfaces for declarative presentation of heterogeneous content. Technical Report CSTR 2006-10, Stanford University, Department of Computer Science, Stanford, CA, USA, September 29, 2006. 10 pp. URL https://hci.

stanford.edu/cstr/reports/2006-10.pdf.

Lee:2006:LSA

[1686] Brian Lee, Heidy Maldonado, Scott R. Klemmer, Isabelle Kim, and Paz Hilfinger-Pardo. Longitudinal studies of augmented notebook usage informing the design of sharing mechanisms. Technical Report CSTR 2006-11, Stanford University, Department of Computer Science, Stanford, CA, USA, September 29, 2006. 11 pp. URL https://hci.stanford.edu/

cstr/reports/2006-11.pdf.

Ju:2006:REI

[1687] Wendy Ju, Brian Lee, and Scott R. Klemmer. Range: Exploring implicit interaction through electronic whiteboard design. Technical Report CSTR 2006-12, Stanford University, Department of Computer Science, Stanford, CA, USA, September 29, 2006. 10 pp. URL https://hci.

stanford.edu/cstr/reports/2006-12.pdf.

Morris:2006:VSB

[1688] Dan Morris, Christopher Sewell, Federico Barbagli, Nikolas Blevins, Sabine Girod, and Kenneth Salisbury. Visuohaptic simulation of bone surgery. Technical Report CSTR 2006-13, Stanford University, Department of Computer Science, Stanford, CA, USA, November 7, 2006. 12

pp. URL https://hci.stanford.edu/cstr/reports/2006-13.pdf.

Hartmann:2006:HMG

[1689] Björn Hartmann, Scott Doorley, and Scott R. Klemmer. Hacking, mashing, gluing: a study of opportunistic design. Technical Report CSTR 2006-14, Stanford University, Department of Computer Science, Stanford, CA, USA, September 29, 2006. 10 pp. URL https://hci.

stanford.edu/cstr/reports/2006-14.pdf.

Brandt:2007:TLL

[1690] Joel Brandt, Noah Weiss, and Scott R. Klemmer. txt 4 l8r: Lowering the burden for diary studies under mobile conditions. Technical Report CSTR 2007-01, Stanford University, Department of Computer Science, Stanford, CA, USA, January 12, 2007. 6 pp. URL https://

hci.stanford.edu/cstr/reports/2007-01.pdf.

Kumar:2007:ESA

[1691] Manu Kumar, Andreas Paepcke, and Terry Winograd. EyeExposé: Switching applications with your eyes. Technical Report CSTR 2007-02, Stanford University, Department of Computer Science, Stanford, CA,

USA, February 15, 2007. 7 pp. URL https://hci.stanford.edu/cstr/
reports/2007-02.pdf.

and:2007:GSD

[1692] Manu Kumar. GUIDe saccade detection and smoothing algorithm. Technical Report CSTR 2007-03, Stanford University, Department of Computer Science, Stanford, CA, USA, February 15, 2007. 2 pp. URL https:

//hci.stanford.edu/cstr/reports/2007-03.pdf.

Ganesh:2007:CRM

[1693] Vijay Ganesh, Sergey Berezin, Cesare Tinelli, and David L. Dill. Combination results for many-sorted theories with overlapping signatures. Technical Report CSTR 2007-04, Stanford University, Department of Computer Science, Stanford, CA, USA, March 12, 2007. 58 pp. URL

https://hci.stanford.edu/cstr/reports/2007-04.pdf.

Kumar:2007:RSS

[1694] Manu Kumar, Tal Garfinkel, Dan Boneh, and Terry Winograd. Reducing shoulder-surfing by using gaze-based password entry. Technical Report CSTR 2007-05, Stanford University, Department of Computer Science, Stanford, CA, USA, March 19, 2007. 7 pp. URL https://hci.

stanford.edu/cstr/reports/2007-05.pdf.

Ganesh:2007:DPF

[1695] Vijay Ganesh, Sergey Berezin, and David L. Dill. A decision procedure for fixed-width bit-vectors. Technical Report CSTR 2007-06, Stanford University, Department of Computer Science, Stanford, CA, USA, March

24, 2007. 27 pp. URL https://hci.stanford.edu/cstr/reports/
2007-06.pdf.

Berezin:2007:OPP

[1696] Sergey Berezin, Vijay Ganesh, and David L. Dill. Online proof-producing decision procedure for mixed-integer linear arithmetic. Technical Report CSTR 2007-07, Stanford University, Department of Computer Science, Stanford, CA, USA, March 24, 2007. 16 pp. URL https://hci.

stanford.edu/cstr/reports/2007-07.pdf.

Ganesh:2007:SS

[1697] Vijay Ganesh, Hassan Saidi, and Natarajan Shankar. Slicing SAL. Technical Report CSTR 2007-08, Stanford University, Department of Computer Science, Stanford, CA, USA, March 24, 2007. 16 pp. URL https:/

/hci.stanford.edu/cstr/reports/2007-08.pdf. The report is dated

7 October 1999.

Hartmann:2007:PSR

[1698] Björn Hartmann, Leslie Wu, Kevin Collins, and Scott R. Klemmer. Programming by a sample: Rapidly prototyping web applications with d.mix. Technical Report CSTR 2007-09, Stanford University, Department of Computer Science, Stanford, CA, USA, March 30, 2007. 10 pp.

URL https://hci.stanford.edu/cstr/reports/2007-09.pdf.

Yeh:2007:IDP

[1699] Ron B. Yeh, Scott R. Klemmer, Andreas Paepcke, Marcello Bastéa-Forte, Joel Brandt, and Jonas Boli. Iterative design of a paper + digital toolkit: Supporting designing, developing, and debugging. Technical Report CSTR 2007-10, Stanford University, Department of Computer Science, Stanford, CA, USA, March 30, 2007. 10 pp. URL https://

hci.stanford.edu/cstr/reports/2007-10.pdf.

Kumar:2007:GES

[1700] Manu Kumar and Terry Winograd. Gaze-enhanced scrolling techniques. Technical Report CSTR 2007-11, Stanford University, Department of Computer Science, Stanford, CA, USA, March 30, 2007. 4 pp. URL

https://hci.stanford.edu/cstr/reports/2007-11.pdf.

Kumar:2007:IAG

[1701] Manu Kumar, Jeff Klingner, Rohan Puranik, Terry Winograd, and Andreas Paepcke. Improving the accuracy of gaze input. Technical Report CSTR 2007-12, Stanford University, Department of Computer Science, Stanford, CA, USA, March 30, 2007. 4 pp. URL https://hci.

stanford.edu/cstr/reports/2007-12.pdf.

Brandt:2007:DLA

[1702] Joel Brandt, Noah Weiss, and Scott R. Klemmer. Designing for limited attention. Technical Report CSTR 2007-13, Stanford University, Department of Computer Science, Stanford, CA, USA, October 3, 2007. 8 pp.

URL https://hci.stanford.edu/cstr/reports/2007-13.pdf.

and:2007:SVI

[1703] Leslie Wu. Social values at the interface: Toward “Just” humancomputer ranking designs at scale. Technical Report CSTR 2007-14, Stanford University, Department of Computer Science, Stanford, CA,

USA, October 3, 2007. 3 pp. URL https://hci.stanford.edu/cstr/
reports/2007-14.pdf.

Wu:2007:RWT

[1704] Leslie Wu, Joel Brandt, and Scott Klemmer. Remixing the Web: Tailoring applications using programmable proxies inside Web browsers. Technical Report CSTR 2007-15, Stanford University, Department of Computer Science, Stanford, CA, USA, October 3, 2007. 4 pp. URL https:

//hci.stanford.edu/cstr/reports/2007-15.pdf.

Lee:2007:AIS

[1705] Brian Lee, Scott R. Klemmer, Savil Srivastava, and Ronen Brafman. Adaptive interfaces for supporting design by example. Technical Report CSTR 2007-16, Stanford University, Department of Computer Science, Stanford, CA, USA, October 9, 2007. 10 pp. URL https://hci.

stanford.edu/cstr/reports/2007-16.pdf.

Lee:2007:CSA

[1706] Brian Lee, Heidy Maldonado, Isabelle Kim, Paz Hilfinger-Pardo, and Scott R. Klemmer. Classroom studies of augmented notebook usage informing the design of sharing mechanisms. Technical Report CSTR 2007-17, Stanford University, Department of Computer Science, Stanford, CA,

USA, October 9, 2007. 8 pp. URL https://hci.stanford.edu/cstr/
reports/2007-17.pdf.

Chen:2008:VCE

[1707] Xing Chen, Lucas Pereira, and Pat Hanrahan. Viewing complex environments using hierarchical light fields. Technical Report CSTR 2002-03, Stanford University, Department of Computer Science, Stanford, CA,

USA, January 2008. URL http://graphics.stanford.edu/papers/
hierarchical_lf/.

Talton:2008:CMP

[1708] Jerry Talton, Daniel Gibson, Pat Hanrahan, and Vladlen Koltun. Collaborative mapping of a parametric design space. Technical Report CSTR 2008-01, Stanford University, Department of Computer Science, Stanford, CA, USA, January 28, 2008. 10 pp. URL https://hci.stanford.

edu/cstr/reports/2008-01.pdf.

Chaudhuri:2008:DRV

[1709] Siddhartha Chaudhuri, Daniel Horn, Pat Hanrahan, and Vladlen Koltun. Distributed rendering of virtual worlds. Technical Report CSTR 2008-02, Stanford University, Department of Computer Science, Stanford, CA,

USA, January 28, 2008. 9 pp. URL https://hci.stanford.edu/cstr/
reports/2008-02.pdf.

Cadar:2008:KUA

[1710] Cristian Cadar, Daniel Dunbar, and Dawson Engler. KLEE: Unassisted and automatic generation of high-coverage tests for complex systems programs. Technical Report CSTR 2008-03, Stanford University, Department of Computer Science, Stanford, CA, USA, May 8, 2008. 14 pp.

URL https://hci.stanford.edu/cstr/reports/2008-03.pdf.

and:2008:ULR

[1711] Philip J. Guo. Using logistic regression to predict developer responses to Coverity Scan bug reports. Technical Report CSTR 2008-04, Stanford University, Department of Computer Science, Stanford, CA, USA, July

6, 2008. 52 pp. URL https://hci.stanford.edu/cstr/reports/2008-
04.pdf.

Brandt:2008:EIO

[1712] Joel Brandt, Philip J. Guo, Joel Lewenstein, Mira Dontcheva, and Scott R. Klemmer. An empirical investigation of opportunistic programming: Interleaving web foraging, learning, and writing code. Technical Report CSTR 2008-05, Stanford University, Department of Computer Science, Stanford, CA, USA, September 19¿, 2008. 10 pp. URL https:

//hci.stanford.edu/cstr/reports/2008-05.pdf.

Brandt:2009:ECP

[1713] Joel Brandt, Mira Dontcheva, Marcos Weskamp, and Scott R. Klemmer. Example-centric programming: Integrating web search into the development environment. Technical Report CSTR 2009-01, Stanford University, Department of Computer Science, Stanford, CA, USA, March 31, 2009.

9 pp. URL https://hci.stanford.edu/cstr/reports/2009-01.pdf.

Chaudhuri:2009:IBE

[1714] Siddhartha Chaudhuri, Daniel Horn, Pat Hanrahan, and Vladlen Koltun. Image-based exploration of massive online environments. Technical Report CSTR 2009-02, Stanford University, Department of Computer Science, Stanford, CA, USA, December 15, 2009. 8 pp. URL https://hci.

stanford.edu/cstr/reports/2009-02.pdf.

Dow:2009:EPP

[1715] Steven P. Dow, Alana Glassco, Jonathan Kass, Melissa Schwarz, and Scott R. Klemmer. The effect of parallel prototyping on design performance, learning, and self-efficacy. Technical Report CSTR 2009-03, Stanford University, Department of Computer Science, Stanford, CA, USA,

December 16, 2009. URL http://hci.stanford.edu/publications/
2009/EffectOfParallelPrototyping.pdf.

Horn:2009:IBS

[1716] Daniel Horn, Ewen Cheslack-Postava, Behram F. T. Mistree, Tahir Azim, Jeff Terrace, Michael J. Freedman, and Philip Levis. To infinity and not beyond: Scaling communication in virtual worlds with meru. Technical Report CSTR 2010-01, Stanford University, Department of Computer Science, Stanford, CA, USA, May 11, 2009. 13 pp. URL https://hci.

stanford.edu/cstr/reports/2010-01.pdf.

Roy:2010:EMM

[1717] Arjun Roy, Stephen M. Rumble, Ryan Stutsman, Philip Levis, David Mazières, and Nickolai Zeldovich. Energy management in mobile devices with the cinder operating system. Technical Report CSTR 2010-02, Stanford University, Department of Computer Science, Stanford, CA, USA,

June 3, 2010. 15 pp. URL https://hci.stanford.edu/cstr/reports/
2010-02.pdf.

Kazandjieva:2010:IEW

[1718] Maria Kazandjieva, Omprakash Gnawali, Brandon Heller, Philip Levis, and Christos Kozyrakis. Identifying energy waste through dense power sensing and utilization monitoring. Technical Report CSTR 2010-03, Stanford University, Department of Computer Science, Stanford, CA,

USA, August 25, 2010. 14 pp. URL https://hci.stanford.edu/cstr/
reports/2010-03.pdf.

Kumar:2010:BSP

[1719] Ranjitha Kumar, Jerry O. Talton, Salman Ahmad, and Scott R. Klemmer. Bricolage: a structured-prediction algorithm for example-based web design. Technical Report CSTR 2010-04, Stanford University, Department of Computer Science, Stanford, CA, USA, September 29, 2010. 9

pp. URL https://hci.stanford.edu/cstr/reports/2010-04.pdf.

Brandt:2010:RHP

[1720] Joel Brandt, Vignan Pattamatta, William Choi, Ben Hsieh, and Scott R. Klemmer. Rehearse: Helping programmers adapt examples by visualizing execution and highlighting related code. Technical Report CSTR 2010-05, Stanford University, Department of Computer Science, Stanford, CA, USA, October 7, 2010. 4 pp. URL https://hci.stanford.

edu/cstr/reports/2010-05.pdf.

Guo:2011:CUS

[1721] Philip J. Guo and Dawson Engler. CDE: Using system call interposition to automatically create portable software packages. Technical Report CSTR 2011-01, Stanford University, Department of Computer Science, Stanford, CA, USA, April 5, 2011. 15 pp. URL https://hci.stanford.

edu/cstr/reports/2011-01.pdf.

Kazandjieva:2011:SHF

[1722] Maria Kazandjieva, Brandon Heller, Omprakash Gnawali, Wanja Hofer, Philip Levis, and Christos Kozyrakis. Software or hardware: The future of green enterprise computing. Technical Report CSTR 2011-02, Stanford University, Department of Computer Science, Stanford, CA, USA, July

7, 2011. 13 pp. URL https://hci.stanford.edu/cstr/reports/2011-
02.pdf.

Karpenko:2011:DVS

[1723] Alexandre Karpenko, David Jacobs, Jongmin Baek, and Marc Levoy. Digital video stabilization and rolling shutter correction using gyroscopes. Technical Report CSTR 2011-03, Stanford University, Department of Computer Science, Stanford, CA, USA, October 1, 2011. 7 pp.

URL https://hci.stanford.edu/cstr/reports/2011-03.pdf.

Kokkalis:2011:PAP

[1724] Nicolas Kokkalis, Johannes Huebner, Moontae Lee, Steven Diamond, Michael Chang, Dominic Becker, Binna Kim, and Scott Klemmer. Providing action plans helps people complete tasks. Technical Report CSTR 2011-09, Stanford University, Department of Computer Science, Stanford, CA, USA, September 24, 2011. 1 pp. URL https://hci.stanford.

edu/cstr/reports/2011-04.pdf.

Kim:2012:AIE

[1725] Young Min Kim, Niloy Mitra, Dongming Yan, and Leonidas Guibas. Acquisition of 3D indoor environments with variability and repetition. Technical Report CSTR 2012-01, Stanford University, Department of Computer Science, Stanford, CA, USA, April 3, 2012. 8 pp. URL https:

//hci.stanford.edu/cstr/reports/2012-01.pdf.

Handigol:2012:MPF

[1726] Nikhil Handigol, Brandon Heller, Vimalkumar Jeyakumar, Bob Lantz, and Nick McKeown. Mininet performance fidelity benchmarks. Technical Report CSTR 2012-02, Stanford University, Department of Computer Science, Stanford, CA, USA, October 21, 2012. 11 pp. URL https://

hci.stanford.edu/cstr/reports/2012-02.pdf.

Lim:2012:LSS

[1727] Maxine Lim, Ranjitha Kumar, Arvind Satyanarayan, Cesar Torres, Jerry O. Talton, and Scott R. Klemmer. Learning structural semantics for the Web. Technical Report CSTR 2012-03, Stanford University, Department of Computer Science, Stanford, CA, USA, November 15, 2012.

8 pp. URL https://hci.stanford.edu/cstr/reports/2012-03.pdf.

Kulkarni:2012:LDW

[1728] Chinmay Kulkarni and Scott R. Klemmer. Learning design wisdom by augmenting physical studio critique with online self-assessment. Technical Report CSTR 2012-04, Stanford University, Department of Computer Science, Stanford, CA, USA, July 2, 2012. 10 pp. URL https://hci.

stanford.edu/cstr/reports/2012-04.pdf.

Azim:2013:DLU

[1729] Tahir Azim, Ewen Cheslack-Postava, and Philip Levis. Displaying large user-generated virtual worlds from the cloud. Technical Report CSTR 2013-01, Stanford University, Department of Computer Science, Stan-

ford, CA, USA, May 25, 2013. URL https://hci.stanford.edu/cstr/
reports/2013-01.pdf.

Kim:2013:JNM

[1730] Jae Young Kim, Omid Mashayekhi, Hang Qu, Maria Kazandjieva, and Philip Levis. Janus: a novel MAC protocol for full duplex radio. Technical Report CSTR 2013-02, Stanford University, Department of Computer Science, Stanford, CA, USA, July 23, 2013. 12 pp. URL https://hci.

stanford.edu/cstr/reports/2013-02.pdf.

Wu:2013:SCR

[1731] Leslie Wu, Jesse Cirimele, Kristen Leach, Stuart Card, Larry Chu, Kyle Harrison, and Scott Klemmer. Supporting crisis response with dynamic procedure aids. Technical Report CSTR 2013-03, Stanford University, Department of Computer Science, Stanford, CA, USA, September 17,

2013. 29 pp. URL https://hci.stanford.edu/cstr/reports/2013-
03.pdf.

Mistree:2014:RAN

[1732] Behram F. T. Mistree, Jay Thomason, Gabriel Kho, Harrison Ho, Edric Kyauk, and Philip Levis. Rethinking application networking as transactional scripting. Technical Report CSTR 2014-01, Stanford University, Department of Computer Science, Stanford, CA, USA, February 4, 2014.

14 pp. URL https://hci.stanford.edu/cstr/reports/2014-01.pdf.

Chang:2014:TDC

[1733] Angel X. Chang and Christopher D. Manning. TokensRegex: Defining cascaded regular expressions over tokens. Technical Report CSTR 2014-02, Stanford University, Department of Computer Science, Stanford, CA,

USA, April 9, 2014. 4 pp. URL https://hci.stanford.edu/cstr/
reports/2014-02.pdf.

Chang:2014:LSK

[1734] Angel X. Chang, Manolis Savva, and Christopher D. Manning. Learning spatial knowledge for text to 3D scene generation. Technical Report CSTR 2014-03, Stanford University, Department of Computer Science, Stanford, CA, USA, April 9, 2014. 11 pp. URL https:

//hci.stanford.edu/cstr/reports/2014-03.pdf;   https://nlp.
stanford.edu/pubs/spatial-emnlp2014.pdf.

Nayak:2014:DNA

[1735] Neha Nayak, Mark Kowarsky, Gabor Angeli, and Christopher D. Manning. A dictionary of nonsubsective adjectives. Technical Report CSTR 2014-04, Stanford University, Department of Computer Science, Stanford, CA, USA, October 22, 2014. 7 pp. URL https://hci.stanford.

edu/cstr/reports/2014-04.pdf.

Mashayekhi:2015:DGS

[1736] Omid Mashayekhi, Chinmayee Shah, Hang Qu, Andrew Lim, and Philip Levis. Distributed graphical simulation in the cloud. Technical Report CSTR 2015-01, Stanford University, Department of Computer Science, Stanford, CA, USA, April 24, 2015. 7 pp. URL https://hci.stanford.

edu/cstr/reports/2015-01.pdf.

Qu:2016:CSA

[1737] Hang Qu, Omid Mashayekhi, David Terei, and Philip Levis. Canary: a scheduling architecture for high performance cloud computing. Technical Report CSTR 2016-01, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1, 2016. 13 pp. URL https://

hci.stanford.edu/cstr/reports/2016-01.pdf.

Mashayekhi:2016:SFC

[1738] Omid Mashayekhi, Hang Qu, Chinmayee Shah, and Philip Levis. Scalable, fast cloud computing with execution templates. Technical Report CSTR 2016-02, Stanford University, Department of Computer Science, Stanford, CA, USA, May 10, 2016. 14 pp. URL https://hci.stanford.

edu/cstr/reports/2016-02.pdf.

Subhraveti:2017:ARA

[1739] Dinesh Subhraveti, Sri Goli, Serge Hallyn, Ravi Chamarthy, and Christos Kozyrakis. AppSwitch: Resolving the application identity crisis. Technical Report CSTR 2017-07, Stanford University, Department of Computer Science, Stanford, CA, USA, November 5, 2017. 7 pp. URL https://

hci.stanford.edu/cstr/reports/2017-01.pdf.

Wilkinson:1968:ADM

[1740] James H. Wilkinson. Almost diagonal matrices with multiple or close eigenvalues. Linear Algebra and its Applications, 1:1–12, 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Golub:1976:ICE

[1741] Gene H. Golub and James H. Wilkinson. Ill-conditioned eigensystems and the computation of the Jordan canonical form. SIAM Review, 18 (4):578–619, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Reprinted in [1744].

Symm:1980:REBb

[1742] H. J. Symm and James H. Wilkinson. Realistic error bounds for a simple eigenvalue and its associated eigenvector. Numerische Mathematik, 35 (2):113–126, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nourbakhsh:1997:IPEb

[1743] Illah Reza Nourbakhsh. Interleaving Planning and Execution for Autonomous Robots, volume SECS 385 of Kluwer international series in engineering and computer science; Robotics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-9828-9. xvi + 145 pp. LCCN TJ211.415 .N68 1997.

Chan:2007:MMC

[1744] Raymond H. Chan, Chen Greif, and Dianne P. O’Leary, editors. Milestones in Matrix Computation: the Selected Works of Gene H. Golub with Commentaries. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920681-3. xi + 565 + 3 pp. LCCN

QA188 .G67 2007. URL http://www.loc.gov/catdir/enhancements/
fy0737/2007276086-d.html.

Knuth:1979:TMN

[1745] Donald E. Knuth. TEX and : New Directions in Typesetting. Digital Press and American Mathematical Society, 12 Crosby Drive, Bedford, MA 01730, USA and Providence, RI, USA, 1979. ISBN 0-932376-02-9, 0-8218-0209-7. xi + 45 + 201 + 105 pp. LCCN Z253.3 .K58 1979. Revised version of Stanford computer science report number STAN-CS-78-675, originally published in September, 1978 [690].